A Complete Bibliography of Publications in Technology and Culture: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 September 2023 Version 1.00

Title word cross-reference

$5 [Bee79]. ^{e} [Rob71]. ème [Day74].

-Webmaschine [Fre73].

Żelazo [Pia71]. Żelazo-Wegiel [Pia71].

1 [Sha78a]. 10 [Mar76b]. 1100 [Bla74a]. 1600 [Hod79]. 1640 [Had76c]. 16th [Kel71b, Str71a]. 16th-Century [Kel71b, Str71a]. 1708 [Tha72]. 1775 [Fli72]. 1786 [Bal73b]. 1794 [Hin71a]. 1797 [Dot75]. 1798 [Fri79a]. 1800 [Gre72, Mid78]. 1803 [Bri72]. 1820 [Fus74b, Lay77c]. 1823 [Web76]. 1824 [Hen76b]. 1827 [Car72b]. 1830 [Gil73b]. 1830s [Lie79b]. 1831 [Jer70a, Ste70]. 1840 [Hyd73, Rae73c, Sch78c]. 1841 [May75a]. 1844 [All79a]. 1845 [Hil75b]. 1850 [Buc70b, Fer72, Gha70a, Hai79]. 1854 [Sin72]. 1860 [Lew76, Ros76b]. 1861 [Rob76b]. 1862 [Sin70a, Smi72c]. 1863 [Min79b]. 1865 [Tha76]. 1870 [Koh71a, Sta78a]. 1875 [Lay71b, Sal77]. 1876 [Con73d, Sta78b]. 1879 [Whi72b]. 1880 [Bee76]. 1883 [vS72]. 1888 [Whi72b, Pos76a]. 1890 [Bry73b]. 1893 [Jac75b, Seg77a]. 1898 [Syl72]. 18th [May71a]. 18th-Century [May71a]. 1900 [Seg77a, Don77, Eis75a, Mer76b, Ste76c, Sti77, Wik72, Tha73b]. 1905 [Fra76, Sau72]. 1906 [Whi72b]. 1910 [Bat70a, Bay75, Bry70, Bry71a, Mat75, Par74, Seg78]. 1911 [Hig79a]. 1913 [Har75a, Lie73b]. 1914 [Pur76a, Bak72, Cow75b, Gai75, Kui78]. 1915 [Car75a, Noe77]. 1917 [Bla72, Con74e, Mar76a, May70, Sto73]. 1919 [vD75]. 1920 [Ben78, Hou77, Jam76, Mat78]. 1923 [Mur79]. 1925 [Cha77]. 1926 [Jac79]. 1929 [Hac73b]. 1930 [Bal75a, Cra76a, Fel76a, Lie73d, Ros72, Top72]. 1930s [Bai76]. 1935 [Con71e]. 1936 [Gie78]. 1939 [Abb72a, Fre76, HS78, Rae79a]. 1940 [Pos75a, Con76a, Fuk75a, Gor71b, Led77, San73c]. 1941 [Bal75b, Per74, Seg77b, Süs76a]. 1945 [Lie73d, Edw72, Emm72, She78b]. 1950 [Cal71b]. 1960 [De 73b, De 74a, Hol70a, Rob76a]. 1962 [Rey75]. 1965 [Hac71, Lie74f]. 1966 [Koh76b, Rob75a]. 1967 [Kel70a]. 1968 [Hil76]. 1969 [Ano70k, Bon71, HS70, Ord70, Slo70, Bat71, Sca72]. 1970 [Bal71, Emm71, MBB^{+}71, O’B71, Swe71, Cow75a, DeV75b, Fra74a, Goi73, Lie73c, Lie74a, Por74, Pri75, Rob75b]. 1970/71 [Har73a]. 1970’s [Hun71]. 1971 [Hal72d, Hal74, LDS^{+}72, Ung72, dLM73, Rop76d]. 1972 [RHTH73, Kob77, Ras76]. 1973 [Jud74a, Kel74, LLC^{+}74, Bri79]. 1974 [Bat74, BMP^{+}75, Dev75a]. 1975 [Buc76b, Fle77, HRM^{+}76, Kra77a, McG76, Rah77, All79c]. 1976 [Bra77d, Mor78, Pic77, PPH^{+}77, SB78, Tar77, Zup77, Bry77a, Cai78, Had78c, Jac77, Kop77, Mor77c, Sta77a]. 1977 [Abr78, BPV^{+}78, Byr78b, FS78, Web79, Chu79]. 1978 [AGS^{+}79]. 1980 [Bil75a, Laf72]. 19th [Bro79, Fer76a, Fri79b, Gla78, Jer73, Mei74, Per70b, Wel77, Hac77, Lay71d, O’B71, Par72b, Pos72, Rez70]. 19th-Century [Fri79b, Wel77, Hac77, Lay71d, O’B71, Rez70].

2 [Sha78a]. 200 [Nes77]. 20th [Cow76b, Eis75b]. 20th-Century [Eis75b].

32/33 [Har73a].

6 [Gar72b]. 60 [Swi78].

7 [Wel73b]. 75 [Ste73]. 751 [Gil73c]. ’77 [FS78, Had79].

Ôkura [Siv77].

Aaron [Ces70]. Abba [Sol76b]. Abbey [Con72c]. Abbott [Ano70a, Ano71a, Ano72a, Ano73a, Ano74a, Ano75a, Ano76a, Ano77a, Ano78a, Ano79a, Par74, Pos75a]. Abraham [Siv75]. Abroad [Cah76].

Absolutely [dC72a]. Abstract [HS78]. Academy [Bri72, Cah76, Car78d, Hac78, Kar76, Lei70, Nut79c, Ols76]. Acceleration [Kob74]. Accomplishment [Uch75]. According [De 75b]. Account [Fer70a, Gai75, Jer70b, Lew74, Siv72, dC70, Lie75c]. Accoutrements [Sto70]. Achieving [DeG71]. Acid [Bee70]. Acquisition [Bal71]. Acragas [Bur74a]. Across [vD75]. Act [Bar71a]. Acta [Koh70a]. Acting [Ret70]. Action [Ste76a, Zin79]. Activity [SG75]. AD [Bla74a]. Adam [May71a]. Adams [Wol78]. Adaptation [Jer74, Wal76b, Lew74]. Addresses [Fai71]. Adelman [Hog75]. Adie [SB72]. Adjustment [Wik72]. Adkins [Vac72b]. Adler [Syl72]. Administration [Bra77b]. administrativen [Bry70]. Administrator [Con72b]. Adopts [Bry71a]. Advanced [Bel78]. Advancement [Ros76b, Swe71]. Adventure [Pau71]. Adventures [Bag79, Sta78d, Hac71]. Advertising [Bur79, Nes77]. Advisors [San77]. Aegean [Wal73b, Ste76d]. Aerial [Per74]. Aero [Slo70]. Aeronautics [Cha70, Cha71a, Han73, Swe71]. Aerospace [Bar70b, Sca72]. Aesthetics [Qui74]. Aether [Top72]. Aether-Drift [Top72]. Afanasyev [Lie77d]. Affairs [Wik77]. Affect [Die73]. Affluent [Wei76]. Africa [De 76, WK71, De 73a, De 73b, De 74b]. African [Bar71b, De 74d, De 75b, DeG76, Hop75]. after [Fri75, Hyd73, Lie79b, Whi75a]. Against [Had79, CZ73, Sha76c]. Age [Cra76a, Fen78, Fit73b, Fre71b, Gor71b, Lew76, McM76a, Mei71b, Por74, Pos75b, Row72b, Sal77, She78b, Web76, Woo79c, Cra78, Dav70, Jud74a, Koh76a, Dau72, Kle76, McN78, Ste75b, Ste76d]. Agencies [Hel71, Smi70c]. Agenda [Gar78, Bay78a]. Agent [Ben75, Fer70b, Mar79b]. Ages [De 75b, Kre77, Mas77a, Sha78c, Web74]. Agostino [Gnu74, Hal70]. Agrarian [Fuk75a, Fus73, McM75]. Agreements [Dan70b]. Agricola [Dib72]. Agriculturae [Koh70a]. Agricultural [Bak72, Bee76, Bra78, Koh76b, Koh78, Ras75, RH73, Arn70, Dal70, Hun71, tZS73a]. Agriculture [Fel76a, Koh70b, Koh71a, Ove75, Koh74, Wil78, Fus72, Koh72, Ras74, Koh77]. Agrimensores [Sky73]. Agroclimatological [Ras71]. Agustín [Gli74]. Ahead [Gol70a]. AIAA [HS70]. Aid [Bla75]. Air [Bra75, Bri79, Bry76a, Ell77, Emm79, Goo74a, HS70, Hal73, Hig79a, Vin79, Bil73, Emm74]. Air-Propeller [Vin79]. Aircraft [Bil77, Bra72, Bri79, Hol70a, Sim71]. Airplane [Han73]. Aitken [Nor77a]. Akin [Seg77b]. al [Con74h, Fer74b, dSP75, Sol76a, Bel78]. al-Jazari [dSP75]. al-Razzaz [dSP75]. al. [Aga77, Bal75a, Bay78a, Bil79, Bos71, Bri71a, Buc74, Ces70, Cha73b, Chu79, Con72c, Cot70, Fer74b, Fra74b, Fuk77, Gla78, Gli70a, Hil78, Hud76, Koh78, Lie75d, Lie78c, Maj77, Man73, Mon71, Nob71, Par79, Per73, PS79, Pur74, Rae73b, Rah79, Row72a, Sal73, Sko75, Spi74, Spi78, Sto71, Swa70, Wil72, Win71, Woo77b]. Alamos [Bra77d]. Alan [Bay76a, Con71c, Dub77, Ger76, Jer78a, Par74]. Alaska [She78b]. Albany [Bri74]. Albert [All79b, Bla74a, Con75b, Gie76b, Hei77a, Par71a, Sha72b, Top74, Whi71b, Rae73c]. Alberti [Kel71a]. Albertson [Lei73]. Albro [Kle77, Sto73]. Album [Pos76a]. Alchemists [Dob78]. Alchemy [Gor79, Dob78]. Alderson [Rol73]. Alekséi [Gli74]. Alex [Bla76a, Woo79c]. Alexander [Bla72, Buc71a, Car79b, Sha74, Sin70b, Rei70]. Alexandria [Kre74]. Alfred [Bed71, Emm77, Har73b, dLM73, Mos79, Sko75]. Ali [Con74b]. Alicante [Gli78]. Alienation [Seg78, Tru71]. Alison [Con71c, Cra78, Pac76]. All-American [Emm78]. Allan [Pen79, Mos75a]. Allen [Con72e, Dru76, Mul79b, Had78d]. Alliance [Mee78]. Allinson [Kob77]. Allison [Spa70]. Allocation [Koh78]. alloys [Pia70]. Allvine [Joh78]. Alma [Bed71, Smi77c]. Almarin [Sim71]. Alonso [Gli78]. Alonza [Sau72]. Alperin [Sma78]. Alps [Str71a]. Alt [Bed71]. Alt-Augsburg [Bed71]. Alternative [For79b, Rah77]. Alternatives [Ols76]. Alvan [Jen70]. Alvin [Bor71b]. am [Hal72d]. Amann [Lie78c]. Amateur [Jen75]. Ambroise [Gnu74]. America [Abb72b, Abb72a, Ben78, Con77b, Cra76a, Cra78, Fer72, Hen76b, Lay71a, Lay77c, Ols76, San73c, Seg77a, Seg78, Sha78a, Sti77, Woo77b, Car78c, Eis75b, Hug79a, Jud74a, Lay71d, Pos72, Rez70, Aus74, Bak76, Bri76a, Con74d, Con75a, Con77c, Con79a, Eat75, Ehr73, Fit73a, Fri78b, Jer74, Lub78, Mas73, Min77a, PS71a, Pur74, Rei71, Rey74, Sal71, Sch78b, Smi75b, Smi77b, War73, Wel77, Wik77, Wri76, Boz77, Bry71a, Cha78b, Cra78, McM76a, Pur70, Rob76c]. American [Bat77, Bra79c, Bri77a, Cha77, Cla78, Con71c, Con71d, Con73a, Con74i, Con79f, Coo77, Cow75a, Emm78, Eze77, Fli72, Gol74a, Gol77b, Gor72, Gro72, Hab72, Hal73, Hea75, Hec73, Hin71a, Hin79, Hok79a, Hol70a, How72, Ino76, Jac75b, Jac79, Jen77a, Kni77, Kra75, Kul79, Lay70a, Lay71b, Lay76a, Lay77a, Lon78, Mar76a, Mor77a, Mul78, Nes77, Nor78, O’B71, Pau71, Por74, Ras74, Ras76, Ros71a, Ros72, Ros76b, Seg77b, Sin72, Smi72b, Smi74a, Sto70, Sto73, Str76, Syl72, Use78, Vol78, Wal73b, Wes77, Wik73, Wil73b, Wil71b, Win78, Woo77a, You72, Cha72a, Cow79, Dan70a, Fel74, Fer79a, Fri75, Hil72, How78, Jer73, Lay76b, Lay79, Ove75, Pic77, Pur76b, Rei70, Rez70, Rol77b, Ros79b, SA77, Sin79a, Swe71, Tob76, Use74, Bay78a, Bee77, Bri77a, Con73c]. American [Con73d, Con74a, DeV75b, Fer73a, Fer79b, Gor71b, Led77, Pen77, Wel73a, Woo79b]. American-ness [Fer79a]. Americans [Bee76, Ell76]. Americas [Cra76a]. Ames [Hac71]. among [Cro71, Mee78, Smi73a]. Amory [Wil76]. Anaheim [HS70]. analisa [Lie75a]. Analysis [Har79b, Ino76, MCA74, Lie75a, Mag76, Mon71]. Analytical [Dru76]. Anand [Eme78]. Anatole [Con71e]. Anatomy [Bri72, Hil78]. Anchor [Alt70]. Anchor-Works [Alt70]. Ancient [Bed71, Bla76b, Bri76a, Bur74a, Con74i, Con78e, Gil79, Sky70, Sky75, Smi77a, Wal71b, dC78, dC79, dc72b, Bul76, Pac78, Per70a]. Anderson [Kop77, Qui74, Gli79, Jer76a, Pos79]. Andes [Gad71]. Andover [Hil76]. André [Her75]. Andrei [Par71b]. Andrew [Nut79a]. Andrews [Hei77a, Chu78b, Con76c]. Angel [Wal73b]. Angelina [Fri79a]. Angelo [Rae73b]. Angles [Shr72]. Angus [Con71b]. Animal [Bla79b]. Animal-powered [Bla79b]. Animals [Arn70]. Anita [Had75b, Had76b, Had77b, Had78b]. Ann [De 74f, San78]. Anne [Rae71a, Ros71a, Whi73c]. Annotated [De 74d, Jac73, Le 75, Ros79d, Siv75]. Announcements [Ano71b, Ano71c, Ano72b, Ano72c, Ano72d, Ano73b, Ano73c, Ano73d, Ano73e, Ano74b, Ano74c, Ano74d, Ano74e, Ano75b, Ano75c, Ano75d, Ano75e, Ano76b, Ano76c, Ano76d, Ano77b, Ano77c, Ano77d, Ano77e, Ano78b, Ano78c, Ano78d, Ano79b]. Annual [Ano70e, Ano78h, Ano79f, Bry77a, Dug79, Fie77, Mei74, Mor79, Spi74]. Anomie [Tru71]. Ansgar [Hal71]. Ante [Mor77a]. Ante-Bellum [Mor77a]. Antebellum [Pos76d, Smi73a]. Antecedents [DeW78, Sko75]. Anthology [Buc73]. Anthony [Aga77, Bar70a, Ber70a, Fri78a, Hin79, Rob76a]. Anthropology [Spi74, Fri79b]. Anti [Sha76c]. Anti-Tank [Sha76c]. antiken [Sch71a]. Antikythera [Dra76a]. Antipollution [Cok72]. antique [Ste76e]. Antiquity [Nut79b]. Antoinette [Con73a]. Anton [Bat77]. Antonio [GD73, Gli75b]. Antony [Lie73d, Lie74f]. Aphoristica [Han79b]. Apocalyptic [Whi71a]. Apollo [Van71]. Apparatus [Mul74c, Whi73c]. Appearance [Vac71]. Appendix [Ano73j, Ano73q, Ano73p]. Appleton [Gil72b]. Applewhite [Fer76c]. Appliances [Dor70]. Application [Hin71b]. Applications [Lol78, Qui71, Tri72]. Applied [Con73g, Lan73, Mar79a, De 74c]. Appreciation [HS78]. Approach [Gil70, Ras75, Row72a]. Approaching [Sto71]. Appropriate [DeW78, Jam79a]. April [Byr78b, McG76, O’B71, Pic77, Zup77, Ano70h, Ano71h, Ano72i, Ano73s, Ano73t, Ano74k, Ano75k, Ano76i, Ano77j, Ano78j, Ano79h, Jud74a]. Arab [Ham78]. Arabic [De 75b]. Arango [Tag72]. árboles [Gli75a]. Arch [Con71c]. Archaeological [Cas76b, Wer78a, Row77]. Archaeology [Alt73a, Buc74, Bur74b, Gol78, Hud76, Lew73, Rae76, Rol70a, Rol70b, Row72a, San73a, Sta78c, Fit71, Gla74, Wal73a, Spi75a, Ung78]. Archaic [Web74]. Archeology [Bat77]. Archimedean [Kel71b]. Archimedes [Sim77]. Architect [Con73f, Cra75, Cra79c, Cra79a, Hal73, Buc78, Coh74]. Architect-in-Chief [Coh74]. Architects [Pac78]. Architectural [Smi72a, Woo79b]. Architecture [Con71e, Con72a, Con72d, Con75b, Con76a, Con78d, Con78b, Con79a, Cra76a, Cra78, Cra79a, Cra79b, Gha70a, Kou70, Mor71, Qui78, Vac71, Can74, Con74b, Cra78, Gra72, Hun75, Rot72, Wat78a, Con74c, Con74d, Hay79]. Archival [Wil70a]. Archives [Tha73b]. Arctic [Kap78]. Ard [Fen78]. Ard-Plough [Fen78]. Area [Har74a]. Areas [Gai75]. Argentina [Ord70]. Argles [Buc70b]. Argolid [Wal73b]. Argument [MN73]. Arid [Gli79, Sal71]. Aristocracy [Kra75]. Armacost [Nie70b]. Armament [Hin71b]. Armory [Eze78, Sta78c, Woo77a]. Arms [Cha78a, Fri75, How78, Smi73a, Dew76]. Armstrong [Bat71, Sta77a]. Army [Con72b, Pri75, Rop76d, Sma78]. Armytage [Lak70]. Arndt [Koh78]. Arnold [Bur77, Con75c, Hal73, Lay75, Mor74, Ber70b]. Art [Con76d, De 75a, Fow78, Kah70, Le 71, Mar77b, Mee78, Nee70, Nes77, Smi70a, Spi75b, Sta78d, Vac72b, Wei73, Cra79a, HS78, Dob78, Le 72, Con76a, Nob71, Spr76, Vac74]. Arthur [Bis76b, Con72b, Con75e, Con76e, Con78d, Cra75, De 79b, Fus74b, Gli79, Har78b, Lew73, Mar77b, Fra73]. Article [Kel70c]. Articles [De 74d, Fer73b]. Articulated [Whi71c]. Artifacts [Ger76, Jen77b]. Artificial [Mic70b]. Artigues [Gli70b]. Artillery [Dra70, Dra72]. Artisan [Cha71c, Cra79a, Sha76a]. Artist [Cra79a, Vac77a]. Artistic [Fer78a]. Artists [Jen70]. Arts [Bat74, Buc70a, Eat73, Eat75, Eis75a, Ger76, Mei74, Mos75a, Nee79, Nes77, Qui78, Sky70, Smi71a, Con78d, Mar79a]. Artz [Buc70b]. Arzneimittelforschung [Geb72]. Aschenbrenner [Wer79]. Ashtor [Mas77a]. Asia [Bra79d, Hun71, tZS78, Nee79]. Asian [Bra79d]. Asit [Car71b]. Aspects [Cah76, Han73, Jac75a, Jer78a, Lei71, McM76a, Per73, ZIL77, Rah79, Mei71a]. Asquith [Fei78]. Assembly [Abr78]. Assessment [Cai78, Fer74b, Fla75, IS77, Tar77, Tri72, Fer74b, Ros79c, Sin79b]. Assessments [Mor74]. Associates [Bil75a]. Association [Abr79, Con72a, Ros76b, Con75f, Fei78]. Astrolabe [Ber77, Ber77]. Astronautics [Cha70, Emm71, Hac78, Hal74, Ord70, Swe71, Hac78, Cha71a]. Astronomical [War71]. Astronomy [Chu78a, Han79b, Chu78a]. Atlantic [Hen76b]. Atlas [Bay76b, BR74]. Atom [San72]. Atomic [Dun76, San74, San74]. Atoms [Gor79]. Atti [Kel70a, Kel70a]. Attitudes [Gue73, Tav72b]. Attlmayr [Tro72]. Aubrey [Bat70b]. Auger [Con73f]. Augsburg [Bed71]. August [Abr78, Fle77, Web79]. Augusta [Cas76a]. Augustin [Gol74b]. Aurel [Wer78a]. Aurelio [Gol70a]. Ausgang [Rob76d]. Austin [SB78]. Australia [Bed76]. Australian [Bed76, Hig79a]. Author [Nut78]. Authorship [GD76]. Autobiography [Con75d, Hol70b, Rol77a]. Automation [Cos71, May70, Mar72a, Tru71]. automatique [KF76]. Automobile [Bry71a, Edw72, Ell77, Fli72]. Automobiles [Win78]. Automotive [Rae79b]. Autonomous [Car78a]. Autoworkers [Gue77]. Aviation [Bai76, Bil75a, Emm79, Rae73b, Cla78, Rae71b]. Aviator [Hal73]. Awakening [Kni77]. Axel [Fen78]. Axes [Wel73a].

B [Aga70, All79b, Ash71, Bai74, Bri78a, Buc70b, Byr78a, Com74, Con79a, De 73b, Dra78, Fer78b, Gie76b, Hal72a, Hol70a, Hol79, Lei71, Nie70a, Ras75, Rou71, Seg78, Sha72b, Sha78a, She74, Smi77b, Wes76, Wik70, Wik73, Wil76, Woo79b, vS72, Ano73i, Ano73l, Ano73q, Cas76b]. B.C. [Smi77a, Dra76a, Wal73b]. Bache [Rei70]. Bachot [Gnu74]. Bachrach [Gil73c]. Background [GD73, Ano73q]. Badash [Hei77a, Hei70]. Baden [Wei78a]. Bailey [Hok79a]. Baillie [Eat73]. Bailyn [Lay77a]. Bainbridge [Swe78b]. Baker [Fra74b, Eys74, She77]. Balance [SB72]. Ball [Bil75b, Had74b]. Baltimore [Whi70]. Balz [Mag76]. BAM [Bra79b]. Bank [Boz77, Bra79d, Sol76b]. Banking [Clo70a]. Banneker [Str73b]. Baranson [Mei70]. Barbara [Gro72, Wil78]. Barclay [Kel71c]. Barker [Had78c]. Barn [Con70b]. Barnett [Lei73]. Barnouw [Gol77b]. Baroque [Wat77]. Barraclough [Aus77]. Barrett [Gli73b]. Barrie [Hau76]. Barry [Kur73]. Bart [Bri71b]. Barton [Sha79]. Bases [Gli70b]. Basic [Swe75, Wei78b, Lie73a, Zin79]. Basil [Rey74, Ung78]. Bateson [Lew74].

Bath [Rol70b]. Bathe [Maa78]. Battery [Fri76]. Battison [Smi74a]. Battista [Kel71a]. Battleship [Rol78]. Baxter [Sew75]. Bay [Cra77]. Be [Bat79]. Beal [Wil72]. Bearse [Bay78a]. Beatty [Jen76]. Beaux [Con78d]. Beaux-Arts [Con78d]. Beckerman [Wei76]. Becoming [Bla76a]. Bedini [Str73b]. Beedell [Kel77]. Beer [Cas74, Smi76b]. before [Car72e, Koh74, Lie79b, Mas77b, May70]. Beginnings [Hew76, Mat78]. Behavioral [Cli71]. Behind [Vac71]. Being [Hou75, Ihd77, Sha76b]. Beispiele [Har71c]. Beiträge [Bry70, Tro72]. Bela [Fre74, Ros76a, Geb72]. Belgrade [Sol76c]. Belief [San73c]. Bell [Bra77a, Bra79a, Sha74, Fer76a, Sha74]. Bellagio [Fle77]. Bellum [Mor77a]. Belt [Goi71]. Ben [Rie70, Sol76b]. Ben-Shahar [Sol76b]. Bender [Wil78]. Benefit [Ell77]. Bengt [Hud76]. Benign [Sto71]. Benin [De 75a]. Benjamin [Fri79a, How74, Str73b, Wik79, Nor75, Nut78, See78]. Bennett [Rob75b]. Benno [Mar77a]. Benthall [Wei73]. Bergaust [Tri77]. Bergbau [Buc76b]. Bergbau-Museum [Buc76b]. Berkebile [Sum75]. Berlin [Hac73b]. Berliner [Lie77c]. Berlins [Son78]. Berman [All79a, Cat76]. Bern [See78, Wil70b]. Bernal [Rav72]. Bernard [DeG75, Gil73c, Lay77a, Lie74d, Sel77, Sha78b, Smi77a, Spi74, Cra79a]. Bernardo [Gli75a]. Berriz [Gli75a]. Bert [Str78]. Bertrand [Hun75]. Berzelius [Mul72a]. Bessey [Ove75]. Besson [Kel73]. Best [Ste73, vR78]. Betancourt [Gli74]. Betjeman [Con73e]. Betrachtungsweise [Sch71a]. between [Fel76a, Gai75, McG76, Vac71]. Beynon [Ehr74]. Beyond [Hel77, Lew70, Vac70]. Beziehungen [Hal72c, Rob76d]. bibliografia [Gli75b, Gli70a]. Bibliographies [Bri78b, Jen73a, Win70]. Bibliography [Bri78b, Can74, De 74d, Fer71a, Fus72, Goi73, Goo70c, Goo71, Goo72, Goo73, Goo74b, Goo75, Goo76, Goo77, Goo78, Goo79, Jac73, Jen73a, Ros79d, Siv75, Cha70, Le 75, Siv75, Hug70]. Biblische [Hal71]. Bicentennial [Sha78a, HRM^{+}76]. Bicycling [Rol73]. Biennial [Fei78]. Big [Dan70a]. Bill [Hey74]. Binder [Lew76]. Biographical [Con73a, Fer76a, Fer79b, Kar76]. Biography [McM78, Süs71, Top74]. Biology [Mor72b]. Biomedical [Kel71d, Per73]. Bird [Ber70a, Hud79]. Birth [Car72c, Rol72, Tho75a]. bis [Bry70, Hal72c, Rob76d]. Biswas [Car71b]. Black [De 73a, Gal75, Law71, Sch77b]. Blackburn [Ehr74]. Blaine [Con74e, Sha78a]. Blake [Cra79b, Lei71]. Blätter [Har71b, Har73a, Har74b, Har76b, Har77]. Bletter [Con76d]. Bloomfield [Rae79b]. Blue [Gue77]. Bluestone [Wat79a]. Bo [Hor79]. Board [Kev75]. Boas [Kre74]. Boat [Ung78]. Bobinger [Bed71]. Bochum [Buc76b]. Bockstoce [Kap78]. Boczar [Cas76a]. Bodey [Buc76a]. Body [Pos75a]. Boeri [Kap77]. Boffey [Boz77]. Bogdan [Ros74]. Bogoliúbov [Gli74]. Bökönyi [Isa76]. Boltwood [Hei70]. Bolz [Mar72a]. Bonanza [Sha71]. Bonanzas [Bil75b]. Bonelli [Sch76]. Book [Bry77c, Nee79, dSP75, dC76]. Bookcase [Hal70]. Booklets [Han79a]. Books [De 74d, Fer73b, She75a, Wil70b, Ada74]. Boorstin [Ell76]. Boot [You72]. Borden [Bat71]. Bordes [Row73]. Borghetty [Adr70]. Borodin [Mar77c].

Bosch [Bry77c]. Boston [Esp70]. Bostons [Bru73]. Botany [Ove75]. Bound [Woo79a]. Bourke [Jen73b]. Bourke-White [Jen73b]. Boutros [Mul74a]. Bow [Goo70a]. Bow-Drill [Goo70a]. Bowden [Goo70b]. Bowls [Ols74]. Boyd [Con73f]. Boyer [Spo72, She78a]. Boys [Emm78]. Bracegirdle [Buc74, Hud75b]. Brain [Boz77, Cas74]. Bramah [May75a]. Brand [Fow79]. Brandes [Led77]. Branigan [Ste76d]. Brass [Smi76b, Dav70]. Braun [Sha78c, Rob76d, Tri77]. Brazil [Mat74]. Brazilian [Mat78]. Breast [Gil79]. Breda [Whi72c]. Breechloaders [Cha73a]. Bremen [Bec78]. Breslow [Van78b]. Brian [Bru73, Buc74, Hud75b, Pos75a, Pos76b, Rop76d, Ste71, dC79]. Brichford [Wil70a]. Bridge [Wat73]. Bridges [She78a, Wat72a, Bri77a, Con75a]. Brief [Pen79]. Brier [Bay76a]. Brierley [Per72]. Brigade [You72]. Bright [Dre76, Qui71]. Brill [Wal73a]. Brinckerhoff [Con73d]. Bring [Sew72]. Brion [Web76]. Bristol [Rol70b]. Britain [Bru76, Dun76, Har75b, Hau76, Hyd73, Rey74, Ste73, Tha73b, Lee78, May71a, Bru76, Car78d, Chu78a, Cro71, For79a, Hud71, San73a]. British [Bal77, Buc76a, Buc79, Con79c, Fer76a, Gre72, Har75a, Har78a, Har79a, Hud79, Kap74a, Pos75a, Rae79a, Rop72, Sta78a, Str76, Ung72, Fri75, Sin79a, Fus74a, Nut79b, Pos75a, Syl72]. Brittain [Nor78]. Broadcasting [Mar77a]. Broadcloth [Bee75]. Brock [Rey74]. Bronze [Fen78, Fit73b, Row72b, Ste75b, Ste76d]. Bronzes [Nob71]. Brooke [Jen77a, McM76a, Bri76b]. Brooklyn [Per70a]. Brooks [Bra77c, Con72e]. Brow [Har76a]. Brown [dC72a, Hin71b, Hun71, Jen73b]. Bruce [Dal70, Hac79, Hou77, Tha76, Sha74]. Brunel [Emm77]. Brunelleschi [Con71a]. Brussels [Hal74]. Bruxellensis [Abr79]. Bruxelles [Abr79]. Bryant [Fra73]. Bryce [Ros79a]. Bryden [Don77]. Buch [Har71c]. Buchanan [Gol78, Rol70b, Rol70b, San73a]. Buchdrucker [Hal72b]. Buckley [Whi77b]. Buckminster [Fer76c, Sto71]. Budapest [Zup77]. Budget [Nic71]. Bugge [Art70]. Builder [Con75d, Whi72b, Woo72b]. Builders [Con78e, Hug79a]. Building [Bay74a, Bay74b, Con72a, Con73c, CZ73, Con74d, Con77a, Con78e, Lie76b, Pac74, Pur74, Car78b, Con77b, Cra77, McK79]. Buildings [Con74g, Con73e, Gar72a]. Built [Whi72c]. Bulferetti [Clo70b]. Bulfinch [Mor71]. Bulliet [Min78]. Bulmer [Ste76b]. Bureaucracy [Ham71]. Burford [Pac76]. Burg [Seg77a]. Burials [Mei71b]. Burke [Wol78]. Burndy [Wil70b]. Burner [Gol71]. Burnham [Vac70]. Burning [Sim77]. Burnley [Hea72]. Burrs [Win70]. Burstall [Bat70b]. Burt [All74, Min79b]. Bush [Cra76b, Jac73]. Bushnell [Rol77b]. Business [Fer73a, Hit76, Ino76, May78a, Por74, Rae71a, Rob75a, Ros75a, Hea71, Ino76, Lam71b]. Busted [Hil70b]. bye [Kap77]. Byers [Hac79]. Byron [Fri77]. Bythell [Lea70]. Byzantium [Dra76b].

C [Adr70, Aus77, Bal73a, Bat79, Bla74a, Bra79c, Bri76a, Bry77b, Bul76, Car73, Car72d, Car72b, Car78d, Car79c, Cha78a, Chu78b, Cra76a, De 71, De 74d, De 75a, Dru76, Emm72, Had78c, Had78d, Hei77b, Hel71, Hey74, Hig73, Hig79a, Hil76, Hin71b, Hin79, Joh78, Lew71, Lie73d, Lie74f, Maj74, Mar76b, Mar77c, McC70a, McK73, McL70, McM78, Mos75a, Mul74c, Rob70, Ros75d, Ros72, Sha78a, Sha79, Sin70b, Siv75, Sta77b, Str78, Vac71, Wat76a, You72, Ano73j]. C-141 [Hig73]. ca [Dra76a, Jer78b]. Cabinet [Bul76]. Cabinetwork [Smi71b]. Cable [Mas73, Pos76a]. Cain [Con79e]. Calculation [Aga76]. Calculator [Bra79c]. Calendar [Dra76a]. Calgary [Spi75b]. Calhoun [Sok75]. California [All79b, Bar73a, HS70, Hei77a, Sha72b, Pat75]. Caliphate [DeG77]. Calkins [De 74f]. Call [Sin77]. Camel [Min78]. Camera [Mar76a]. Cameron [Clo70a]. Campbell [Bri70, Mar79a]. Can [Mul73a]. Canada [Lei70]. Canadian [Bal75a, Con74f, Lei70, Sin79a]. Canal [Hey74, Hil70a]. Canals [Wil71a]. Canan [Smi76d]. Capital [dLM76, Sal77, San70, Rob71]. Capitalism [Bla73a, Led77, Ran76, Cha78b]. Capitalist [Hay79, May78a]. Capitols [Con78c]. Capron [Ros73a]. Car [Bry77c, Cla72, Fli72, Mas73, Whi72b, Whi76a, Whi77a, Wik73, Con79f]. carbon [Pia71]. Carder [Had72]. Cardew [Wal71a]. Cardwell [Cra79a, Dau72, Fer73c, Sha76a]. Career [Cra75, May78a, She74]. Careers [Kap71]. Carel [Smi76b]. Cargill [Hac78, Swe78a]. Carl [Aga70, Arn70, Bay74a, Bay74b, Bay78b, Bez79, Hal72b, Hec70, Lon78, Sky70, Wil77]. Carleton [Fer71c]. Carlisle [Cot70]. Carlo [Cam74, Clo78, Con74h, Day74, Kel70a]. Carlson [Ste70]. Carmaux [Ste76a]. Carnegie [Tar77]. Carnot [Car73]. Carol [Bay75, Con79d]. Carolina [Sha71]. Carovillano [Tav72a]. Carpenter [Bal73b, Bul76]. Carriage [Ber74, Pur76b]. Carriages [Hea75]. Carroll [Hal73, Lay70a, Lay71a, Jud74b]. Cars [Pos76a]. Carter [Fri79a, Har71a, Ros71a]. Carts [Nel72]. Cartwright [Ris78]. Carver [Gol77a]. Caryl [Mor70]. Case [Con79e, Fel76a, Fri75, Ham71, Hug76b, JR78, Kev75, Rey79b, Sew75, Tay75, Top79b, Vin79, Wil76]. Caslon [Had74b]. Cassatt [Car79b]. Cassimere [Bee75]. Casson [dc72b]. Casting [Par72a]. Caswall [Buc75b]. catalana [Gli73a]. Catalog [Pur75, Wil71b, Hei77a, Bis70, Goo70b]. Cataloging [Kle76]. Catalogue [Hit76, Hok79b, Swe75]. Catch [Mat76]. Catchers [Wes77]. Catherine [Lew74]. Catling [Hil75a]. Causality [MN73]. Causes [Owe73]. Cavendish [Dev75a]. Celestial [Sle78]. Cement [Con75d]. Cementing [ZIL77]. Cent [Gai75]. Centenarian [Bir76]. Centennial [Coh74, Con73d]. Center [All79b, Hac71, Sha72b, Con79d]. Central [Fra71, Gad71, Isa76, Row72b]. Centre [KF76]. Centro [Kel70a, Rap75]. Centuries [Bis76a, Buc76a, Qui77, Con78e, Mul79b, Whi75a]. Century [Bro79, Con72a, Con72b, Cow76b, Fer79b, Fri79b, Han79a, Ihd77, Jer73, Jer74, Jer76a, May75a, McK73, Mok77, Mul78, Nee79, Per72, Per70b, Rei71, Sch74, Seg79a, Smi76b, Ste76a, Tha72, Tho75a, War73, Wel77, Eis75b, Hac77, Kel71b, Lay71d, May71a, O’B71, Par72b, Pos72, Rez70, Str71a, Vac71, Bul76, CZ73, Con78e, Con79c, Fer76a, Gha73a, Gla78, Hen76b, Hit76, Kob74, Le 71, Nut79b, Str76, Tho75b, Tho76, Vac70, Wel71, Bla74b, Koh70b]. Ceramics [Aus74]. Cerchione [Rae73b]. Cerinda [Hin72]. Cervera [Gli75a]. Cervin [Con76d]. Cetron [Gil70]. Chafe [Cow75a]. Challenge [Bra79d, Eze78, Uch75, McL70]. Chalmers [Bla71]. Chamber [Nut79c]. Chance [Mor72b]. Chandler [Ros77]. Change [All79a, Bla76a, Car71a, Con73g, Cow76b, Cro71, DeG71, De 76, Esp72, Fee79, Fel74, Han73, Hel73, Hug76a, Hun71, Jer70a, Jer78b, Mar79b, Mat75, McM75, PS79, Ros73a, Ros71a, Ros76a, SA77, Siv75, Smi75b, Sta78a, tZS73a, Tar74, Tay75, Wal76a, War73, Win70, Ces70, De 79a, Eze78, Fer70b, Fre71a, Lam71b, Lay75, Lie74d, Man72, Mor72a, Mul74a, Top79a, Bla71, Bru73, Cha72b, Koh71a]. Changed [She77]. Changing [Cow75a, Hol79, Laf72, Nev79, Ros75a, Tho76, vR78, Rie70]. Chaos [Hug77]. Chapman [Jer79, Lea76, Lea79, Mal78]. Chapter [GD76, Smi76b]. Chapters [Cra79c]. Chaput [Abb72b]. Characteristics [Sal70]. Charcoal [SA77]. Chariots [Web74]. Charles [Bal73b, Con77b, Dib79, Dru74, Hig79a, Jud74b, McM78, McM76b, Mor71, Mul74a, Pen79, Per78b, PS79, Pos75b, Qui78, Sta78a, Win78, Wis79, tzS75, Les79, May75b, Ove75, Ros77, Sin70a]. Charlotte [Jer74]. Chasm [Gol70a]. Cheers [Wei76]. Chelovek [Lie75a]. Chelovek-nauka-tekhnika [Lie75a]. Chemical [Bee77, Ihd77, Maj75, Maj77, Mul78, Mul79b, Ron78, Mul74c]. Chemist [Ihd77, Rez70]. Chemistry [Maj75, Maj77, Mul76, Mul78, Ron78, Sch77b]. Chemists [Bee77, Mul78]. Chenhall [Kle76]. Cherns [Gie76b]. Chesbrough [Cai72]. Chiasso [MCA74]. Chicago [Con77f, BMP^{+}75, Byr78b, Cai72, Cai74, MBB^{+}71, Con79e, Bay74a, Bay74b, Seg77a]. Chief [Coh74]. Children [Sch71b]. Chilinder [Kre77]. Chilton [Bee70]. China [All79b, Bla75, Fai71, Maj75, Maj77, Sha72b, Siv72, Siv75, Smi77a, tZS73b, Bra78, Bro79, Hac77, Vol78, Bri73, Smi77a, tZS73a, Vol78, Siv72, tZS76]. Chinese [MM74, Siv72, tzS75]. Chitnis [Eme78]. Chittenden [She74]. Choice [Cra76a, Str76]. Choices [Bra77b]. Choose [Ols76]. Chris [Hok79a]. Christiansen [Car75b]. Christine [Fer79b]. Christman [All79b, Sha72b]. Christopher [Bri76b, Con79b]. Chronology [Cha71a, Sca72]. Chumash [Spi78]. Church [Cat76, Rob75a]. Churchman [McL70]. Chymist [Ihd77]. ciencia [Gli70a, Gli70b, Gli70b]. Cincinnati [Bay78b]. Cinema [Gia72]. Cinnabar [Maj77]. Cipolla [Cam74, Day74]. circa [Bac79, Gre72]. Cities [Con76c, Gre72, Ros79a, Goi71, Lie76b, Sew72, Bri76a, War73]. Citizen [Del76]. City [Bal73b, Bay75, Bay78b, Bri73, Bud78, Bur74a, Cai72, Con71e, Con72a, Con74i, Fra71, Fri78b, Joh73, Rob75b, Seg77a, Smi71b, Con74i, Bri71a, Bri76b, Con71d, Ste76a]. Civic [Mon75]. Civil [Bri77a, Con73a, Gil73a, How78, Jon78, tZS73b]. Civilian [Jon78, Str73a]. Civilisation [Maj75, Maj77, tZS73b, Wal73b]. Civility [Bou78]. Civilization [Gor71b, Nee70, Zup77, Car75c, Row76b, Siv72, Vac72c, SR73]. Civilizational [De 70]. Civilizations [De 73a]. Civilizing [Seg77a]. Clara [Whi71b]. Clare [Woo79a]. Clarence [Con72b, Hac73a, Koh71a, Man72, Nes77]. Clark [Hig79b, Jen70, You72, Bla75, Eat75, Gil72b, Mei71a, Sha73]. Clarke [Car78b]. Class [Bal75a, Cow75b, Kra75, Van78b, Dub77, Ran76, Ste76b]. Classic [Con75e]. Classical [Bla73b, Nob71, Wal73b, Ano73p]. Classics [Wat79b]. Classified [De 74d]. Claude [Maq70]. Claudia [Coo77]. claus [Gli73a]. Clausius [Dau72]. Clavicula [Dre75]. Clay [Mer76b, Con74i]. Clean [Ell77]. Clerk [Hei77b]. Clerks [Fai71]. Cleveland [Wat79a]. Cliff [Abb72b, Bag79]. clima [Gli70b]. Climates [Chu78b]. Climb [Hol70a]. Clines [Hal73]. Clinical [Lol78]. Cliometrics [Cro76]. Clock [Cat76, Smi72c, Bla77]. Clocks [Bal78, Bed71, Cat76, Hok79a, Bed71]. Clockwork [Bed71]. Close [Con74i]. Close-Up [Con74i]. Clouds [She78b]. Clyde [Goo70b]. Co [Gli70b]. Co-operation [Gli70b]. Coach [Pos76c]. Coal [Lew76, Sch78c, Lie75b, Lew76, McM77]. Coal-mining [McM77]. Coalmining [Har78a]. Coast [Pur76a, Sma78, Ste76c]. Coastal [Bag79]. Codex [Mar72c, GD76]. Coe [McC70b]. Cohen [Wil78]. Coil [Bri70]. Coke [Jen73c]. Cold [Hac73a]. Cole [Kap74b, Kap74b]. Coles [Spi75a]. Colin [Fit71, Rae73a, Wal73b, Wal76a]. Collaboration [SR73]. Collar [Gue77]. Collected [Hin71b]. Collection [De 73b, Gor79, Hei77a, Jer77, Maa78, Ron78, Ros77, Fei77, Sum75, Swe75]. Collections [Bis76b, Nut79c, Wil71b, Woo72a]. Collectivisim [San73c]. College [How72, Koh72, Mor77a]. Colleges [Sto74]. Collins [Fus71]. Collis [Hin72]. Colombia [Svi77]. Colonial [Bak76, De 76, Ehr73, Lub78, Nes77, Wes76, Cot70, Fri78a]. Colonialism [De 73b]. Colonies [Fow77, Fow79]. Color [Alt73b]. Colt [Fri76]. Colum [Tav74]. Columbia [Ung72, Bal77]. Columbian [Pau71]. Combine [Fus71]. Combing [Hea72]. Combustion [Pos75a]. Comes [Kah70]. Coming [McN78]. Commemorate [Sha76a]. Comment [Bur70, Lay70b, Owe73, Wil73a]. Commentaries [Gre72]. Commentary [De 74e, Süs76b]. Commerce [Koh71b]. commercio [Rap75]. Commission [San74]. Committee [Bri77a, Cah76, Sin72]. Common [Hec73, Mul76, Tha72]. Commons [Wei78a]. Communication [Bla74a, Bos71, Le 75, Rop76b, Swe73]. Communications [Mar79b, Mul73b]. Communist [Bla71, Bla79a, Bri71a, Bri73, Fle77, Fee79]. Communities [Lay71d, PS79, Sto74]. Community [Dub77, Fri78b, Ros76b, Sch78c, Kap74a]. comp [And74, Con72a, Fer76a, Gar72a, Goo70b, Goo74a, Hei77a]. Company [Bal73b, dLM73, Par74, Pos76a, Rob75b, Sal73, Con72a, Bag79]. Comparative [Clo70a, Ino76, Mag76]. Comparato [Ada74]. Comparison [Eze77]. Compass [Kre73]. Competition [Kap74a]. Compilation [Dug79, Fie77, Mor79, Smi75a]. Complex [Bur70]. Component [Fer78a]. Composition [Smi75a]. Comprehensive [Buc70a, Ano73g]. Compton [Dev76, Wei71]. Computer [Ash71, Bar70a, Dra76a, Fer70b, Kle76, Le 72, Rog78, Tro75, Cas73, Con73f, Vac77a, Aga76, Le 72]. Computers [Die73, Bay76a]. Computing [Bra77d, Mar77c]. Concept [May71a]. Conceptual [Hug77]. Concise [Emm74]. Concorde [Smi76d]. Concrete [Jac79]. Condenser [Fer70a]. Condit [Bay74a, Bay74b, Bay78b, Bez79]. Condition [Gha70b, Sel77]. Conducted [Wal73b]. Confederate [Slo72]. Conference [Bra77d, Dra78, Lew74, Mei74, Pos79]. Conflict [Gli79, Ino76, SR73]. Conflicts [Jam79b]. Congress [Buc76b, Ros79c, Ano70f, Ano75n, Cha71a]. Connecticut [Bla77, Smi72c]. Connections [Mee78]. conquête [Her75]. Conquer [Hou77]. Conquest [Sal71, Sha74, Swe72]. Conscience [Gol70b, Ste71]. Conscious [Lew74]. Consensus [Bra77b]. Consequences [DeW78, Lei73, Spo73, Tri72, Oka78]. Conservation [Buc76b, Pen77, Ray77, Coo74]. Consiglio [Rap75]. Conspiracy [Smi76d]. Constance [Emm71]. Constants [Lid76]. Construction [Ber77, Bra72, Cal74, Con72a]. Consumer [Pal79]. Consumption [Bry76a]. contemporanea [Clo70b]. Contemporaries [Con72e, Koh70c]. Contemporary [ZIL77, Lie74c, PS79]. Context [Mor77b]. Continental [Fla75]. Continuity [Cha72b]. Contracting [Man72]. Contribució [Gli73a]. Contribución [Gli75b]. Contribution [Cla78, Con79c, Fen78, Nut79a]. Contributions [Con77b, Kre73, Mas77b, Smi72c]. Control [Car78a, Har79b, Jon78, Kev75, Per78a, Lei71]. Controls [Ell77]. Controversy [Del76, Wri76, Nie70b]. convegno [Kel70a]. Convergence [Bla78a]. Conyngham [Lie74a]. Cooke [Sin70b, Vac72a]. Cooked [Maq70]. cooled [Les79]. Coolidge [Bir76]. Cooper [Cal70b, Dor79, Hou78]. Cooperation [Sha78a]. Cope [Cha72a]. Copernicus [Ros74]. Coppa [Bri76a]. Copper [Abb72b, Les79, Abb72a, Ste75b]. Copper-cooled [Les79]. Copyright [Jac73]. Corbusier [Seg79a]. Cordell [Zin79]. Corliss [San72]. Corning [Had70, Had71, Had76a]. Cornish [All74, Fow78, All74]. Corporate [Cha78b]. Corporation [Gol77c, Man73, Mar72b]. Corporations [Por74]. Corps [Cal74, Con72b, Rop76d]. Correspondence [Hei77a]. cortar [Gli75a]. COSMAT [Cah76, Cah76]. Cossons [Rol70b]. Cost [Ell77, Gol77c]. Costa [Bos71]. Costello [Smi76d]. Costs [Dal72]. Cotton [Bee75, Fel74, Lea70, Sau72]. Cotton-Textile [Fel74]. Cottrell [Car71a]. Could [Bat79]. Coulomb [Gil73a, McK73]. Coulton [Pac78]. Council [Com74]. Count [Hin71b]. Counterrevolution [Shr73]. Countries [Bal71, Sol76c, Bla76a, Gue77, Hog75, Jam79a, Jun70, San78]. Country [Gal75, Koh70c, Nut78, Smi71b]. County [Bal73b, Had77c]. Coupler [Cla72]. Coutts [Le 71]. Coutts-Smith [Le 71]. covering [Lie75c]. Cowan [Con78e, Con78e]. Cox [Ste76c]. CPSU [Lie74a]. Cradle [Smi77a]. Craft [Tyl77]. Crafts [Eat73, Eat75, Sky70]. Craftsmen [Fai71, Kra75, Ste76a, Pac76]. Craig [Bry73b]. Crane [Sto74]. Crash [Mul71]. creado [Gli75a]. Create [Pur74]. Creating [Bra77b]. Creativity [Con72b]. Creed [BM77]. Creveld [Rey79a]. Crises [Bra75]. Crisis [Sko76, Wil78]. Criteria [Gol70b]. Critical [Bri78b, Gar78, HV78, Jen73a, Mor74]. Critics [Dur79]. Critiques [Ano73h, Ano73i, Ano73j, Ano73k, Ano73l]. Crompton [Gue73]. Cromwell [Dew76]. Crop [Koh74]. Crosland [Art70]. Croton [Rou71]. Crouch [Sca72]. Crowther [Cal70a, Dev75a]. Cruz [Mat78, Cat76]. CT [Woo77a]. CT-2 [Woo77a]. Cubs [Sha75]. Cudahy [Bru73, Pos76b]. Culbert [Wik77]. Cullen [Sch77b]. Culp [Joh73]. Cultural [Bla79a, Fee79, Mar79b, McL78, WK71, Wes78, DeG75, DeG75, Gol74b]. Culture [Bla72, Bla79a, Buc73, DeG75, Fle77, Fri78a, Fri79b, Gie77, Kel76, Koh71b, Kul79, Smi70c, Wal76a, Wel70, Flo72, Gor72, Qui78, Rav70, Wei72, Whi76a, Ano70c, Ano71e, Ano72f, Ano73m, Ano74h, Ano75h, Ano76f, Ano77g, Ano78f, Ano79d, Lev79, Whe73]. Cultures [Row72b]. Cultures. [Tur75]. Cummins [Bry77b]. Cumulative [Bri78b, Jen73a]. Cunningham [Emm78, Hil78]. Cunynghame [Bed71]. Current [Bis76b, Goo70c, Goo71, Goo72, Goo73, Goo74b, Goo75, Goo76, Goo77, Goo78, Goo79, Rür74, Con72d]. Curriculum [Lev75]. Curtin [De 76]. Curtis [Lev79]. Custodians [Sma78]. Cybernetic [Bee73]. Cybernetics [Cas74]. Cyclades [Wal73b]. Cycle [Sum75]. Cynthia [Hog75]. Cyril [Dre75]. cznym [Pia70].

D [All79b, Bay76b, Bee77, Bil77, Bla77, Bra79a, Cal71b, Con71a, Con75e, Con75g, Cra78, De 76, Don77, Eat73, Fei78, Fer73c, Fer72, Fla75, Fre71a, Fri78a, Fus72, Fus76a, Gol76, Gre72, Hil78, Hil70a, Jac71, Jer79, Kob77, Koh77, Lea76, Lea79, Led77, Lid76, Mal78, dLM76, Mon71, Mon75, Mos75a, Mur79, Par79, Ras75, Rav72, Rol77a, Rol70b, Row77, Sca72, Sha76a, Sha73, Sha78a, Sin70a, Smi71b, Smi72c, Spa70, Swa70, Tha73a, Wel71, Wel77, Whi72c, Woo79a]. Daguerreian [Jen72]. Daguerreotype [Gor72]. Dahrendorf [Rah79]. Dakin [Cra75]. Dale [Coo77, Wei71]. Dalhousie [Gor71a]. Dallas [Rei70]. Dalton [Mor74]. Daly [Whi73c]. Damals [Bry71b]. Damasceńskiej [Smi76c]. Damascus [Smi76c]. Dampfkesselwesens [Bry70]. Dams [Smi70b, Con72b, Con73b, Dob75]. Dan [Bee79]. Dane [Pen75]. Danhof [Koh71a, Man72]. Daniel [Ben78, Ell76, Gli70b, Jam76, May78a, Sok75, Ste75a, Wat79a]. Daniels [Smi72b]. dans [Ste76e]. Danube [JR78]. Darbys [Hud75b]. Darcy [De 70]. Dare [Cow79]. Daredevil [Hal73]. Dark [McN78, De 75a]. Darkly [Wil73a]. Darrah [Joh73]. Darwin [San73c]. Data [Tar74]. Davenport [Buc73, Flo72]. David [Bir76, Bla76a, Bru76, Byr74, Cha78b, Chu78a, Con75a, De 78, Dra76b, Ell77, Gie77, Gol77c, Hin71a, Hod79, Hol79, Kap77, Koh70c, Lev75, Rae71b, Rop76b, Ros75a, Sch78b, Seg77a, Spa70, Ste75b, Swi78, War73, Wat73, Web76, Wik77, Wri76, Str76]. Davidson [Bel78, Con79a]. Davies [Hen76b, Hou77]. Davis [Gie76b, Car79b, Kel76, Vac74, Web76]. Dawley [Dub77]. Day [Bai74, Bry77c, Fus71]. Days [Joh73]. Deadly [Kra76]. Dean [Kap72, Siv75]. Death [Rol72]. Debate [Sin79b]. Deborah [Jen70]. Debus [Mul79b]. Decade [Bar71a, Cra76b]. December [Ano70k, Bri74, BMP^{+}75, Bri77b, LDS^{+}72, LLC^{+}74, MBB^{+}71, Mer76a, Mor78, PPH^{+}77, RHTH73, Tar77]. Dechen [Car72b]. Decipherment [Cas76b]. Decision [Lie74a, Lie77c, Van71]. Decision-Making [Lie74a]. Decisions [Lam71b, Zin79, Mon75]. Decline [Con71c, Fre76, Har75a, Jer76b, Kap77, Sto73, Vac72c, Ros79a]. Deco [Con76a, Con76d]. Decoration [Cra76a]. Decorative [Bat74, Mei74]. Dedication [The74]. Dee [Han79b, Dra76c]. Defended [Dor76]. Defense [Kev75]. Deficiencies [MGM78]. Definition [Rol77c]. Definitions [Hug77]. Definitive [Hey74]. Dejiny [Win71]. Del [Ord70, Gli73b, Gli74, Gli75b, Kel70a, Rap75, Wat78b]. Delacroix [Jen73c]. Delaunay [HP76]. Delaware [Bay75, Kra77a, Hey74]. Delegation [Vol78]. della [Clo70b, Clo75, Kel70a, Rap75]. Delmo [Str78]. dels [Gli73a]. Democracy [Bla78c, Con78c, Zin79]. Democratic [Ell76]. Demonstration [Bra77b]. Denied [Sto73]. Denis [Jam79b]. Denmark [Esp70]. Dennis [Bla74c, Gil72a, Rop76c]. Deo [Rav70]. Dependency [Hel77]. Depletion [Sha72c]. Depression [Fus74a, Pur79a, Rap77, Con77c]. Derbyshire [Rol70a]. Derek [Dra76a, War76]. Deren [Jen73c]. Described [Art70, Spi78]. Describing [Tyl77]. Description [Bed71]. Desert [Gli79]. Design [Bla78a, Cha78b, Con75f, Con78e, Han79a, Hay79, Qui74, Rot72]. Designing [Con75b, San73c]. Designs [Tar79, Cra79a]. Deslandes [HP76]. Destination [Hac79]. Destroy [Pur74]. Destroyed [Rol78]. Détente [Dor79]. Deterrence [Dun76]. Deutsche [Hal72b]. Deutschen [Hal72c, May78b]. Deutschland [Bry70, Rob76d]. developed [Lie77b, Lie79a]. Developing [Bla76a, Hel77, Hog75, Jam79a, Jun70, Mei70, Sol76c, Bar73b]. Development [Bal75b, Bal78, Bar71b, Bos71, Bra77b, Bra79d, BM77, Bry76b, Buc70b, Bur77, Cla72, Dal70, Dob75, Dru76, Ehr73, Fel78, Fri78a, Gli70b, Gol70b, Gol74b, Had76c, Hai79, Ham78, Ham71, Hew76, Hug76a, Hun71, JR78, Koh72, Lay76a, Lie73d, Lie74f, Mea71, Mil76, Mon75, Mos79, Pal79, Pat75, PS71a, Rae71b, Rob76a, Ros70, Rot72, Rür74, RH73, Sch74, Siv75, Smi73b, Smi76e, Wel73a, Gli70b, Sol76a, Bry77c, Chu78b, Con78b, De 71, DeG76, Dra70, Fel77, Hay79, Hou78, Nic71, Sol76a, Wil72, Wol78, Bil77, Bla75, Con71b, Per70b]. Developments [Car78c, Lei73, dC70, Con76b]. développement [Day74]. Devices [Hin71b, dSP75]. Devon [Sha75]. Dewdney [Car79c]. Dexter [Ano70b, Ano71d, Ano72e, Ano73f, Ano74f, Ano75f, Ano76e, Ano77f, Ano78e, Ano79c]. Dharampal [Gha73a]. d’histoire [KF76]. diagram [Pia71]. Dial [Kre77]. Dialectic [McL79]. Diana [Sto74]. Diane [Hai79]. Diaries [Ait79]. Diary [Har75b]. Dibner [See78, Wil70b]. Dickey [Fer74b]. Dickinson [Gia72]. Dicks [dC79]. Dickson [Cha70, Gor73]. Dictionary [Con73a, Ihd77, Whi72b, Whi73a, Whi73b]. Did [Fri77]. Diebold [Fer70b, Lam71b]. Diesel [Bry76b, Lie75c, Tho78]. Dieter [Fer74b]. Differences [Dan70b, Van78c]. Diffusion [Blo77, Fel74, Fel76b, Rog78, Sto74]. Dignity [Nel79]. Dilemma [Gol70b]. Dilke [Sky73]. Dimension [Ols77]. Diplomes [Abr79]. Direct [De 79a, Mar77a]. Directory [Ano74g]. Dirkzwager [vS72]. Disadvantages [Hou75]. Disasters [Con72b]. Discipline [Fer74a, Bur74b]. Disciplines [Nee79]. Discourses [Sha76b]. Discoverable [Ger76]. Discoveries [Sch77b]. Discovery [Gia72, Maj75, Maj77]. Dispersals [Arn70]. Distance [Tuc78]. Distiller [Kel73]. Distinguishing [Gru71]. Distribution [Ell77, Ste75b]. District [Rop76d]. Diversion [Cai74]. Division [Cha71a]. Dixon [Bil75a, Con71d, Fri78b]. Doctors [Hel71]. Doctrines [Sch77b]. Documentary [DeV75b, Goo74a, Koh77, Pur71, Smi75a]. Documentation [Wil70a, KF76]. Documents [Ano73p, De 73b]. Dodds [She74]. Doeringer [Nob71]. Doers [Hok77]. Dofny [Aga78]. Dolce [Bri76a]. Dollar [Hig79a]. Dollars [Rei71]. Domestic [Isa76, Wil72]. Domestication [Arn70]. Domination [Wag73]. Domus [Kel70a]. Don [Bra77b, Con76a]. Donaghy [Hil75b]. Donald [Abb72b, Chu78b, Con79a, Cra76b, Dat71, Dau72, Fra74a, Har75a, Hel71, Hil78, Hyd73, Lay77a, Nee79, dSP75, Sum75]. Donovan [Sch77b]. Doolittle [Hal73]. Doris [Le 75]. Dörner [Bed71]. Dorothy [Gre73, Gue73, Str73a, Syl72, Top74]. Dossin [Hil70c]. Double [Ret70]. Double-Acting [Ret70]. Doud [Gar72b]. Doughty [Lub78]. Douglas [How72, Law71, Vac74]. Dr [Alt73b, Hal72d, Sha78a]. Dr. [vS72]. Dragon [Smi77d]. Draining [Hil71]. Drama [Pau71]. Dream [Kni77, Le 71, Seg77b]. Dreamers [Hok77]. Dreher [Lon78]. Dressier [Hal72d]. Drexler [Con78d]. Driakhlov [Lie75d]. Drift [Top72]. Drill [Goo70a]. Drilling [Goo70a]. Dritte [Lie75b]. Dritten [Lie76c]. driven [HP72]. Droege [Con70a]. Drug [Lei71]. Dry [Tar79]. Dryden [Swe75]. Duane [Per74, Tha73a]. Dubin [Con75e]. Dublin [Tha72]. Dudley [Gal75, Per72]. Duke [Pur74]. Dumas [Ray77]. Duncan [Lea70, Rog77, Rol72, Rey75, San74]. DuPre [Hal73]. Durand [Vin79]. Durbin [Mor79]. during [All79b, Buc70a, Buc75a, Con72a, Jer73, Lea70, Lie78a]. Dutch [Bal74, Con70b]. Dwight [tZS76]. Dyed [Bor71a]. Dying [Bee75]. Dynamics [Hil72, May75b, Par79, Ros79a, Goi70]. Dynamische [Sch71a]. Dynamism [Rav70]. Dynasties [Nee70]. Dyos [Con74i]. dziejów [Mos78].

E. [Vin79]. Eagle [Sha78a]. Earl [Gol74a, Wik72]. Earle [Qui74]. Earliest [Gre72]. Early [Bac75, Bla74a, Bra77a, BM77, Car77b, Clo70a, Con77b, Dau72, Fen78, Har71a, HP72, Hin79, Jer70b, Jer73, Joh73, Kel76, Mar79b, McM76a, Min79a, Pal79, Ran76, SB78, Smi71a, Smi77a, Ste76d, War76, Bra78, And74, Gie78, Jen77a, Lay71a, Ram70, Web74]. Early-Modern [Mar79b]. Earth [Bee73, Pau71, Tag72]. Easily [Bat70b]. East [Mas77a, Smi77a, tZS78, Ano73j, Ret70]. Eastern [Isa76, Row72b, Wel70]. Eastman [Jen75]. Eastwood [Sha76b]. Ebbin [Del76]. Eben [Rez70]. Ebenezer [Seg79a]. Eberhard [Har71c]. Eckaus [Jam79a]. Eclecticism [Cra76a]. Ecole [Con78d]. Ecological [Str75]. Ecology [Tob76]. Economic [Bai74, Bar71b, Bil75a, Buc71a, Cam74, Clo70a, Com74, Cow75a, Dal70, Dal72, De 74a, De 74b, De 75c, De 79b, Dev78, Dru76, Dug79, Ehr73, Fel78, Fer73a, Fie77, Fra71, Gal75, Gli70b, Hal72a, Han73, Kob74, Lie73d, Lie74f, Mas77a, May71a, PS71b, Ros70, Ros71b, Sin79b, Str76, Tar74, Teb74, Use78, Wal74, Wil73b, Lie76a, Lie79a, De 76, Hai79, Har75a, Hog75, Pal79, Spo73]. Economics [Aga77, Con71f, Fei77, Hil77, Ros76a, Sha78b, Sol76b, Lie77b, Fre74, Lam71a]. Economie [Jun70]. Economies [Mei70]. économique [Day74]. Economy [De 73b, Hyd73, JR78, Jud74b, McM75, McM77, Mon72, Ros79b, tZS73a, tZS76, Lie74b, Lie75b, All74, Joh78, Ros71a, Sky75]. ed [Adr70, Aga77, Aga78, Ait79, All74, Art70, Ash71, Aus74, Aus76, Bal73a, Bal73b, Bar73b, Bay76b, Bee77, Bel78, Bil79, Bla71, Bla79a, Blo77, Bra77a, Bra79a, Bri73, Bri76a, Bri71b, Bri74, Bri77a, Bri77b, Bri78b, Bry76a, Bry77a, Buc70a, Buc73, Byr78a, Byr74, Cai78, Cam74, Can74, Car72c, Car78d, Caw77, Ces70, Cha78a, Cha73b, Chu79, Clo78, Con71d, Con73e, Con74d, Con74i, Con75c, Con77b, Con78d, Coo71, Cra79c, Dal70, Day74, De 71, De 73b, DeV75b, Dra78, Dre75, Dug79, Eat75, Emm72, Emm79, Emm77, Fei78, Fel77, Fer72, Fer74b, Fie77, Fit71, Fit73b, Fow79, Fri79a, Fuk77, Fus73, Gar72b, Gha70b, Gie76a, Gie76b, Gli70a, Gli75b, Goi73, Gol77a, Gol70b, Gre72, Hac78, Had75a, Had75b, Had76b, Had77b, Had77c, Had78b, Hal72b, Hal72d, Han79b, Har74a, Har71b, Har71c]. ed [Har73a, Har74b, Har76b, Har77, Hau76, Hei70, Hel71, Hel77, Hin71b, Hin71a, Hol79, Hol70b, Hor79, Hud76, Hun71, Hyd73, Jac73, Jac79, Jen73a, Jen77a, Jer70a, Jer79, Kel70a, Kel71d, Koh70c, Koh77, Koh78, Kre76, Kul79, Law71, Lay70a, Lay75, Lay77a, Lay77b, Lay77c, Lea76, Lea79, Lei71, Lei73, Lew70, Lie74d, Lie75a, Lie75d, Lie77a, Lie78c, Lie79a, Low76, Mal78, Mar72a, Mar76b, Mat76, May71b, May77, McM76a, Min77b, Mok77, Mol78, Mor70, Mor77c, Mor79, Mul73a, Mul74a, Mul78, Mul79a, Nel79, Nob71, Nor78, Oka78, Par71a, Par79, Pos79, Pur70, Pur74, Qui74, Qui71, Qui77, Rae73b, Rae74, Ran72, Rie70, Rog75, Rol77a, Rop72, Rop76a, Ros79a, Ros73a, Ros75a, Ros74, Ros76a, Ros73b, Ros75c, Ros75d, Rou71, Row77, Row73, Row76a, Row79]. ed [San78, Sch78c, Sch76, Sha76a, Sha76b, Sha78a, Sin70b, Sin72, Siv75, Sky70, Smi70c, Smi71b, Smi75b, Spa70, Spi74, Spi78, Spo73, Sta77a, Sta77b, Ste76b, Ste75a, Str78, tZS76, tZS78, Swe75, Tav72a, Tha73a, Tri72, Tro72, Tru71, Tuc73, Tur76, Vac77a, Van78a, Wal73a, Wal74, Wal76a, Wat79a, Wat79b, Wei71, Wei78a, Wel73b, Wer79, Whe78, Whi75a, Wik70, Wil72, Win78, Win70, Win71, Wol78]. Edelstein [Adr70, Gor79, Ron78, Gor74]. Edgar [Hil70b]. Edge [Chu78a, Wel71]. Edinburgh [Abr78, Nut77, SB72]. Edison [Sch77a]. Edith [Rog75]. Edition [Kre74]. Editors [Fli72]. educación [Gli70b]. Education [Bar70b, Buc70b, De 74f, DeV75b, Dra78, Mor78, Rez70, Bar70b, Lie77d, Smi76a]. Educational [Bar70a]. educativa [Gli70b]. Educator [Con72b]. Edward [Cha78a, Emm72, Fer78b, Fow78, Fri79a, Gil72b, Gol70b, HS78, Nut79b, Pos79]. Edwin [Hab72, Hac71, Man73, Smi74a, Smi75b]. Effect [Dev76, Per78a]. Effects [Lew74, Owe73, Vac70]. Efficiency [Sol76b]. Efremtsev [Lie75c]. Egon [Wat78b]. Egypt [Con78e, Con74c]. Egyptian [Cas76b, Nee70, Per70a]. Ehrenreich [Gro72]. Ehrensvärd [Alt73b]. Eichberg [Bec78]. Eighteenth [Bul76, Gha73a, Ihd77, McK73, Tho75a, Koh70b, Smi76b].

Eighteenth-Century [McK73, Koh70b]. Eiho [Aus76]. Einfluss [Fuk75b]. Eisen [Lie75b]. Eisenberg [Mat75]. Eisenhower [Dob75, Bri78a]. Eisenstein [Mar79b]. Eklund [Ihd77]. ekonomicheskie [Lie79a]. Ekonomicheskoe [Lie76a]. ekonomika [Lie74b, Lie77b, Lie74b]. Ekundare [De 74a]. Elder [Par71a]. Eleanor [Had76c]. Electric [Bal75b, Com74, Mas77b, MN73, Teb74, Whi73a]. electrica [Clo75]. Electrical [Hou74, Nor78]. Electrician [See78]. Electricity [Dib76, Sha72a]. Electrification [Hug79a, Bez79, Con77d]. Electrified [Con75g]. Electron [Ram70]. Electronics [Süs70]. Electrophysiology [Lol78]. Elegant [Fer78a]. Elektropolis [Son78]. Eli [Sta78c, Bla77, Woo77a]. Eliahu [Mas77a]. Elisha [Hou75, Use74]. Elite [Svi77]. Elixirs [Maj77]. Elizabeth [Cra77, Cra78, Mar79b, Per70a]. Ellet [Sin70a]. Elliott [Kra74]. Ellipse [Wes78]. Ellis [Mar72a, Sta77a, Cai72, Cla77]. Ellsworth [Hok77, Smi70c]. Elmer [Goi73, Dan72]. Elmo [Dob75]. Elms [Bat71]. Else [Ros79d]. Elting [Lay76a]. Elvin [MM74]. Emergence [Bee76, Cam74, Chu78a, Gal75, San70, Wal73b, Wer78b]. Emerging [Hug79b]. Emilio [Kle71]. Emission [Ell77]. Emme [Sha78a]. Emmerson [Buc78]. Emphasis [Ros79d, Smi75a]. Empire [Buc79, Lew76, Gro72, Mic70a, Mul73b]. Empires [Hen76b]. Empirical [Kap74c]. Employed [Jac71]. Employment [Mon72]. Encyclopedia [Spr76, Bri79, Con75c]. End [Bry71c, Sha78a, Car79b]. Endrei [Fre71b]. énergétiques [Her75]. énergies [Her75]. Energy [Bra77b, Dun76, Kap74a, Ols76, San74, Wil76, Spo73, Fla75, Ols76]. Engelsberg [Hud76]. Engine [Bry76b, Had78d, Les79, May75b, Whi73e, BR74, Car72a, Con71b, Gol78]. Engineer [Ben75, Buc78, Cla73, Con72b, Mar72b, Min79b, Rae74, Rol74, Rou71, Sin70a, Dan72, Kel70b]. Engineering [Bra77a, Bra79a, Bry77c, Buc79, Cah76, Hab72, Kel71c, Kev71, Lam71a, Lay76b, May75b, McC70b, McK73, Per78b, Ros79d, Sin77, tZS73b, Wat79a, Woo77a, Bri77a, Gil73a, Rob70, Smi73b, Wat76a, Lay71b, Wat79b, dC78]. Engineers [BM77, Cal74, Con70d, Fer76a, Fer79b, Hab72, Pri75, Rae75, Rop76d, Sma78, Wis79, You72, Lie76c, Con72b, Bee77, Con73a, Kel71c]. Engines [Bil77, Bry73a, Bry76a, Lay71a, She75a, Bla79b, Hea75]. England [Ber70b, Buc75a, Car72b, Cow75b, Fus73, Kap77, Rae73c, Rob76d, San73b, Bee75, Jud74b, Lev75, Pal79]. Engle [Had75b, Had76b, Had77b, Had78b]. English [Car75a, Had76c, Jer74, Lea70, Ran76, Jer70b, Han79b, Hea75]. Enhancement [Por76]. Eniac [Bra76b]. Enigma [Fri76]. Enlarger [Hec73]. Enlightenment [Eme78, How72, Sch77b]. Enquiry [Smi77a]. Enrico [Wat77, Kle71]. Enseignement [Abr79]. Enserfment [Esp72]. Enterprise [Coh74, Sto73]. Enterprises [Cha77, Wel71]. Entomological [Per78a]. Entrepreneur [Min79b]. Entrepreneurial [Har75a]. Entrepreneurship [Jer76b, Lie73b]. Entwicklung [Bry70, Fuk75b]. Entwisle [Le 75]. Environment [Ces70, Hea71, Pur74, Ros71b, Sto71, Com74, Fit73b, Fri78a, Nev79, Rae73b, Ros76a, Lei73]. Environmental [Con73c, Con78e, Kou70, Lie74e, Pen77, Fer74b]. Environments [Gli79]. Environs [Wer79]. Episode [Kev71]. Epistemological [Ano73n]. Epochal [Wil70b]. Epoche [Fuk75b]. Epstein [Cro71]. Equipment [Fus76a, Con72a]. Equity [Hog75]. Era [Art70, De 76, McM76b, Sha78a, Dob75, Smi71a, Wat77]. Erasmus [PS79]. Erb [Hel77]. Eric [Buc71b, Dra78, Fuk75b, Har75b, Law71, Pur71, Qui74, Sha76b, Tho75a, Wal74]. Erickson [Jer74]. Ericsson [Con75d]. Erik [Gol77b, Mul72a, Tri77]. Ernest [Sal70]. Ernst [Sha78a, Tro72]. Ervin [Ros73b]. Erwin [Ros75b]. España [Gli70b, Gli70b, Gli70a]. español [Gli74]. Espinos [Sko75]. Essay [Cat76, Gil73a, Ihd77, Koh71b, Lev79, McM75, Mor72b, Bla74b, Con72d]. Essays [Aga77, All74, Bal73a, Fer71c, Hau76, Ras74, Rav70, Sha76a, Str76, Tar74, Hac78, Hyd73]. essence [Lie73a]. Essential [Top79a]. Essers [Ord78]. Establishment [Gor71a, Wer78b]. Estate [Con72a, Gal75]. estestvoznnniia [Gra75]. Esther [Emm72]. estudi [Gli73a]. Estudios [Gli75b]. età [Clo70b]. Etat [Abr79]. Ethereal [Top72]. Ethic [Ben78]. Ethical [Wil76, Ano73h, Ano73i, Ano73j]. Ethics [McG76]. Ethnographic [Spi78]. Ethnology [Cha73b]. Etudes [KF76]. Euan [For79a]. Euclid [Dra76c]. Eudald [Gli73a]. Eugen [Kui78]. Eugene [Fel77, Hol79, Hug70, Sha78a, Vac72c, Wat77]. Eugeniusz [Pia70]. Eulau [Bou78]. Europe [Bla74a, Bla74b, Cam74, Con74d, Dot75, Nee79, Ros75a, Alt73a, Isa76, Mar79b, Mil76, Pos78, Row72b, Buc71a]. European [Bla73b, Fus76b, Zin79, Ano73j, Car72e, Rez70, Rol77b, Gha70a]. Euston [Con71c]. Evaluation [Wil70a]. Evans [Had79, Hin72, Spr76]. Eve [How78, Lie78a]. Evelyn [tZS73a]. Evening [Sha76b]. Events [Sca72]. Everett [Chu79, Lay75, Per73]. Everitt [Hei77b]. Every [Kah70]. Everyman [Wik77]. Evidence [Mea71]. Evolution [Bry73a, BR74, Bry77c, Con70c, De 79b, Gol77b, Maa78, McK73, Fre71b, Aga77, Row79]. Evolutionary [Ras75]. Evry [Bel78]. Ewing [Ros75a]. Ewo [Bro79]. ex [Rav70]. Examination [Gla74]. Example [Cha71b]. Excavating [Mos79]. Excavation [Hun75]. Excavations [Nee70, Wal73b, Wer79]. Excellence [Smi74c]. Excursion [Par72b]. Execution [Dun76]. Exhibit [Lew77]. Exhibition [Coh74]. Exhibitions [Car72e]. Existential [Wat76a]. Expansion [Hen76b, dC74b, Mon72]. Expedition [Fit73b]. Experience [Bar73b, Gol74b, Meu71, Str76, Bar71a, Ell76, Rog78]. Experiment [All79b, Con74c, Kev71, Koh76b, Sch76, Spi75a]. Experimental [Hil78]. Experiments [Top72, Zin79, Bic73]. Expert [Hou75]. Explanation [Wal76a, Row72a]. Explanations [Bur70]. Explicitly [Row72a]. Exploitation [Pen77]. Exploration [Ros71b, Sha78a, Sha78b]. Explorations [Fer76c, Fre74]. Exploratory [Fer71c, Tay75]. Explorer [Wer78a]. Exploring [Kur73]. Exports [Jun70]. Expressionism [HS78]. Expropriation [Kra77b]. Extension [Bak72]. Extracting [Pau71]. Eye [Mar76a]. Eyewitness [Vac72a]. Ezell [Cha78a].

F [Bat70b, Bel78, Bra70b, Cha78b, Con71c, Con76c, Cot72, Dal70, Ehr73, GD73, Hei77a, Hei77b, Hel77, Hin79, Hok79b, Hud71, Jud74b, Les79, Mar77b, Mei72, Nee79, Par78, Par71b, PS71b, Qui74, Ris78, Ros77, Row73, SG75, Seg77a, Sew75, Smi70c, Spr76, Sta77b, Str71b, Vin79, dC74a]. Fabricius [Dob78]. Facsimile [Koh71b, Kre74]. Factor [Coo71]. Factories [Con74g]. Factors [Hel71]. Factory [De 73b, Jam76, Van78a, Ehr74, Wei79, Bat71]. Fagen [Bra77a, Bra79a]. Failed [Mok77]. Fairbairn [Bri71b]. Fairfield [Gie76a]. Fairlie [Pos75a]. Fairservis [Bla76b, Row76b]. Faith [Bee79]. Fall [Süs76a]. Fallingwater [Con79a]. Fallout [Per70b]. Falls [Gol77a]. Family [Abb72a, Gar78, Rob75a, Van78b, Jer76a]. Farago [Vac77b]. farga [Gli73a]. Farm [Ber70b, Fel76a, Fus71, Hil72, Swi78, Fuk75a, Fus76a]. Farmer [Bak72, De 78, Wik72]. Farming [Fus74a, Ras76, Bla73b, Wil78]. Fascinating [Woo77c]. Fashionable [Wat72b]. Fate [Str73a]. Father [Tho78]. Fatherland [Lie74b]. Fathers [Koh70c]. Fathy [Con74c]. Federal [DeG75, Ell77, Fel76b, Hel71, Kev71, Pos72, Nic71]. Feedback [May71a, Woo72a]. Feldman [Mur79]. Ferguson [Rae73a, De 74e, Hug70]. Ferkiss [Car75c, Lew71]. Fermi [Kle71]. Ferraria [Tyl77]. ferrous [Par72a]. ferroviaria [Clo75]. Ferry [Eze78]. Fertility [Fre76]. Fiasco [Cra79b]. Fiction [dC76]. Field [Wei79]. Fields [Koh72]. Fifteenth [Han79a, Sch74]. Fifteenth-Century [Han79a]. Fifth [Hal74, Bry73b]. Fifty [HS70, Süs70, Bra72, Con74e]. Fighters [Bri79]. Fighting [Dor70, Whi73c]. filage [Fre71b]. Filature [Bro79]. Finance [May78a]. Finberg [Fus73]. Finch [Wat79b]. Finding [Pau71, Tho75a]. Fine [Cal74, Cra79a]. Finley [Sky75]. Fire [Whi73c, Whi73e, Whi73c, Bry77b, Dor70]. Firm [Cas74]. First [Abb72b, Bra77c, Bra79c, Bry77a, Cai72, Cok72, CZ73, Con77f, Dib76, Dor79, Gor71b, IS77, Lie76a, Nee70, Rae71a, Rea78, Süs70, Tho75b, Min77a, vD75]. Fisher [Blo77]. Fishermen [Cha73b]. Fishery [Kap77]. Fishman [Seg79a]. Fitch [Con73c, Rol77a]. Fitchen [Con70b]. FitzSimons [Rou71, Wat79b]. Five [Gai75, Qui77, Wis79]. Fixed [dLM76]. Flame [Gol76]. Flaws [Ven76]. Fleming [Lay77a]. Fleron [Bla79a]. Fletcher [Hea75]. Flight [Cla78, Gie79, Sha78a, Swe73, vD75]. Flink [Bry71a, Whi76a]. Flint [Hal73]. Florence [Bon71]. Florence-Vinci [Bon71]. Florman [Wat76a]. Flows [Hey74]. Fluidmechanics [Rou72]. Flying [Hig73, Cha71b]. Focus [Bla78b]. Folk [Wel70]. Follows [Cra79b]. Foner [Van78a]. Fontana [Cam74]. fonti [Kel70a]. Food [De 75b, Wil78, Wil78]. Food-Producer [Wil78]. Foodstuffs [Arn70]. Forbes [Tho75a, Whi74]. Force [Bri79, Goo74a, Hig79a, Ols77]. Forces [Con73c]. Ford [Fit73a, Hea71, Ols76, Rae71a, Wik70]. Forecast [Par72a]. Forecasting [Gil70, Qui71]. Foreign [Hou77, Lie73b, Wik77, De 79a]. Forest [Bal77, Buc75a]. Forests [Bal75b]. Foreword [Kra73]. Forging [Use74]. Forgotten [Cra78]. Form [Con70c, Pos79, Svi77, Woo79b, Gue77, Con76b, McL78, Cra79b]. Formation [BM77, Ros76b, dLM76]. Formed [Bri78b, Jen73a]. Forms [WK71]. Forney [Whi72b]. Forrestal [Bee79]. Forrester [Goi70]. Forschung [Maj74]. Fortifications [Bac75, Bac79]. Forts [Woo79b]. Fortschritt [Hal71]. Forum [Ols76]. Foster [Ran76]. Foundation [Sha76a, Bar70b, Ols76]. Founding [Sch78c]. Foundry [Smi76b]. Four [Buc71a, Gue77, Shr72]. Fourier [Gil73b]. Fowler [Whi73b]. Framework [Cah76, Zin79]. France [Art70, Bel78, Buc70b, Kui78, Rob71, Bac75, Kel76, McK73, McK79]. Frances [Jam79a, Sta78b]. Francine [Dal72]. Francis [Pei72, Rey75, San74, Tho76, Sha75, War76]. Francisco [Bry73b, Con78a, Cra77, Gli78, LLC^{+}74]. Francois [Tuc78]. Frank [Ada74, Bri76a, Cha72a, Chu78b, Con71a, Con72e, Dra76b, Dre76, Fus76b, Kop77, Rol77a, Rol70a, Rop76d, Seg79a, Svi77, Wil71a, Con79a]. Franke [Le 72]. Frankel [Dal72]. Frankenstein [Sch71b]. Franklin [Nor75, See78, Sto71, Tha76]. Franz [dC76]. Fraser [Rae73a]. Frazier [Bis76b]. Fred [Car71a, Fra76, Joh78, Tof72]. Frederic [Bla79a, Goi71]. Frederick [Buc70b, Cra76a, Fei78, Had72, Lew76, Lid76, Mok77, Pur74, Ris78, Rol73, Sha78c, Lay71c]. Frederik [Jac71]. Freight [Cla72]. Freight-Car [Cla72]. French [She78a, Ste76a, Mul71]. Frenchmen [Kui78]. frères [Rob71]. Freudenheim [Cra77]. Friday [Sha76b]. Fridolf [Nes77]. Fridolin [Hal72d]. Friedman [Row79]. Fries [Lub78]. Frisinger [Mid78]. Fritz [Sch71a]. front [Lie78a]. fronta [Lie78a]. Frontier [Fow77, Bar71b]. Frost [Pat75]. Frozen [Mei71b]. frühen [Osp79]. Frühgeschichte [Har70]. Fuel [Bra75, Bry76a]. Fueling [Kra78]. Fuels [Bil77]. fuentes [Gli75b]. Fuhrmann [Bla73a]. Fulk [Bac79]. Fuller [Fer76c, Sto71, Hec73]. Fulton [Pur78]. Function [Rog77, Woo79b]. Fundamental [Lid76]. fünf [Bed71]. fünfzehnten [Hal72b]. Furnaces [Rol70b]. Furniss [Bag79]. Furniture [Cra79a, Smi71b]. Fussell [Koh74, Bla73b, Koh74]. Future [Bra77b, Car75c, Fow78, HV78, Lak70, McL79, Mul73a, Ols76, Pic77, Ros79c, Sch70, Tav72a, Whi73f, Mul74a, Ols76, Bra77b, Gol75, Ros75b, Sch70, Smi77c, Vac74, Bor71b]. Futures [Wil76]. Futurism [HV78].

G [Aus74, Bal75a, Bar70a, Bla73b, Blo77, Cal70a, Car75a, Cas73, Con74f, Con79a, Cot72, De 73b, De 74b, Dev75a, Dev78, Dib75, Dre75, Dru76, Emm72, Emm76, Fer71c, Fus74b, Gar72b, Gli78, Hac73a, Har78b, Hen76b, Hig79b, Jer76b, Jer79, Kap71, Kle76, Koh70c, Koh71b, Koh74, Kul79, Lak70, Lea76, Lea79, Lie73a, Lie74b, Lie75c, Lie77a, Lie77d, Mal78, Mei72, Mos75a, Mul73a, Mul79b, Nor77a, Par74, Rae74, Rey75, Rob76b, Rol78, Rou72, San74, Smi76a, Swe72, Tho75a, Tri72, Wel73b, dC78]. Gabor [Bla74c, Gil72a]. Gabriel [Ros75c]. Gadol [Kel71a]. Gaëtan [HS78]. Gains [Dal72]. Gale [Hud79, Wer70]. Galeona [Cas76a]. Galician [OO73]. Galilaeana [Kel70a]. Galleys [Tho76]. Galveston [Sma78]. Game [Spa70]. Ganzevoort [Bal74]. Gardner [Bla75, Had72]. Garrett [Wei78a]. Gary [Ehr73, Kob77]. Gas [Fla75, Gol71, Jon78, Whi77a, vO78]. Gaston [Kap74a]. Gateway [Bra79b]. Gatherer [Wil78]. Gatovskii [Lie77b]. Gavin [Kra75]. Gazette [Whi73d]. Gazley [Fus74b]. GE [Con77d]. Gears [Dra76a].

Geburtstag [Hal72d]. Gelder [Bal74]. Gellner [Sal70]. Gemini [Sha79]. Gendron [Sel77]. Gene [Sin70a]. General [Abr78, Bil75a, Ros73b, dC70]. Genesis [Bra76b]. Genevieve [Siv75]. genovese [Rap75]. Geoffrey [Bal77, Jen76, Pos75a, Ste73]. Geography [Bed76, Car79c, War73, Fus74a]. Geology [Mul76]. Geometry [Fer76c]. George [Bar71a, Bri78a, Buc78, Byr74, Car72a, Con76c, Fit73b, Gol78, Jac73, Mas73, Mic70a, Qui74, Sin72, Smi72b, Wer79, Whi73c, Whi73b, Wil72, Bra79c, Bri70, Jen75, Rol74]. Georgescu [Aga77]. Georgescu-Roegen [Aga77]. Georgia [FS78, BM77]. Georgii [Lie78a]. Gerald [Blo77, Cal71a, Cut74, Laf72, Mur79, Rae79b, Top79b]. Gerard [Fow77]. Geräte [Fuk75b]. Gerber [Kel71d]. Gerd [Hal72d]. Gerhard [Bry70, Fer74b, Rap77]. German [Hac73a, Mar79a, Mur79, Qui78, Süs76a, Tho75b]. Germany [Buc76b, Emm71, Fre76, DeG75, Rop76c]. Gerrard [All79b]. Gerrard-Gough [All79b]. Gerschenkron [Buc71a]. Geschichte [Bry71b, Fre73, Har71c, Geb72]. Gesellschaft [Fuk75b]. gesellschaftlicher [Fuk75b]. Getreidemühlen [Str72]. gg [Lie74b]. Ghali [Mul74a]. Ghana [De 73b]. Ghanaian [De 73b]. Ghita [Sal70]. Ghost [Cha71c]. Giant [Por74, Sca72]. Gibbon [Whi75a]. Gibbs [Dib79, Nor77b]. Gibbs-Smith [Dib79]. Gibson [Kap77]. Gideon [Cat76]. Gies [Har76a, Sta78b]. Gilbert [Con79c, San73c, Dor70, Hol76, Seg78, Woo72b]. Gilded [Cra76a]. Giles [Dew76]. Gilfillan [Tav74]. Gill [Wes76]. Gillispie [Car73]. Gillmor [McK73]. Gilman [Gor71b]. Gimbutas [Row72b]. Gioanventura [Adr70]. Giovanni [Con78b, Kel70c]. Girard [Gol74b]. Girls [Van78a]. Glare [Sha78c]. Glasgow [Rae76]. Glass [Had70, Had71, Had72, Had73, Had74a, Had75a, Had75b, Had76a, Had76b, Had77a, Had77b, Had78a, Had78b, Per70a, Per72, Pet75, Had70, Had71, Had76a, HS78]. Glasser [Wei79]. Glassie [Wel70]. Glassmakers [Had78c]. Glassmaking [Had76c]. Glassworkers [Ste76a]. Glazes [Per70a]. Gleisberg [Str72]. Glenn [San72]. Glick [GD73, Str71b]. Glorious [Coh74]. Glowing [Dre76]. GM [Con72c]. Gnudi [Fer76b]. Go [Van71]. Goals [Per78a, Gol70b, Swe72]. Goddard [Emm72]. Godfrey [Gra72, Had76c]. gody [Lie74b, Lie76a, Lie78a]. Goist [Con71d, Fri78b]. Golcondas [Hud79]. Gold [Maj75, Fre74, Ros76a]. Goldberg [Win70]. Golden [Pos75b]. Goldenberg [Bak76]. Goldin [Coo77]. Goldman [Lie74e]. Goldstine [Tro75]. Goldwater [Mul74d]. Good [Kap77, San73c]. Good-bye [Kap77]. Goodstein [Hei77a]. Goodwin [Gra72]. Göran [Gie79]. Gordon [Blo77, Hig77, Koh72, She74]. Górewicz [CZ73]. Gorge [Hud75b]. Gothic [Han79a]. Gotlieb [Mar77c]. Gough [All79b]. Gouldner [McL79]. Goulet [Jam79b]. Government [Pur79a, Qui71, Man72]. govierno [Gli75a]. Gowans [Ger76]. Gowing [Dun76]. Grace [Buc75a, Cal70b, Hou78]. Graduate [Sha76a]. Grady [Con74i]. Graebner [McM77]. Graells [Gli73a]. Grafton [McM78]. Graham [Seg77c, Sha74, McC73]. Grahame [Mei71a]. Grammar [McL79]. Grand [All79b]. Grant [Con75b, Hok77, Mic70a]. Granville [Fel76b].

Graphic [Nes77]. Graphics [Le 72]. Graphics/Computer [Le 72]. Grass [Fit73a]. Grass-Roots [Fit73a]. Grattan [Gil73b]. Grattan-Guinness [Gil73b]. Gray [The74, Hou75]. Great [Abb72b, Buc78, Car72c, Car78d, Con77b, Fus74a, Har75b, Hil70c, Hol76, Lee78, Lie74b, Lie78a, Tha73b, Wat73, Pur79a, Hau76]. Greatness [Hol70a]. Greece [Sky70, Wer79]. Greek [Pac76, Pac78, Web74, Bur74a, Kah70, Dra70, Dra72, Nor77b]. Greeks [Dra76a]. Green [Dal72, Hun71, McM75, Fra74a, Goi71]. Green-Belt [Goi71]. Greenberg [Ham71, Koh70c]. Greene [Cra79a, Cra79a]. Greenhalgh [Web74]. Greenhill [Rey74, Ung78]. Greenslade [Had77c]. Greenwich [War76]. Gregori [Bar71b]. Gregory [Ros76b, Smi73b]. Greif [Con77c]. Greville [Maa78]. Greywater [Tar79]. Griessman [Hol79]. Griffin [Har78a, McM75]. Grigg [Ras75]. Grimwood [Sha79]. Grist [Gad71]. Gromeka [Lie78b]. Groups [Byr74, Del76]. Growth [Con71c, DeG71, Fra71, Fre71a, Gli79, Har79a, Hin79, Hog75, Jun70, Kob74, Mer76b, Mor72a, Ros79a, Sin79b, Sol76b, Str76, Dal70, De 74c, Kni77, Sel78, Use78, Wil73b]. Grubb [DeV75b]. Grube [Con74g]. Gruber [Hel71]. Guatemalan [Cla73]. Gudozhnik [Lie73a]. Guerickes [May71b]. Guerra [Rap75]. Guggenheim [Cla78]. Guide [Alt73a, Fra74b, Goo74a, Hil77, Hou74, Rop72, Rop76a, Hei77a, Mos75a]. Guilmartin [Tho76]. Guinness [Gil73b]. Guldbeck [Ber74]. Gun [Cla77]. Gunfounding [Smi76b]. Gunnar [Nut79c]. Gunnery [How74]. Gunns [Hei77a]. Gunpowder [Tho76]. Gunsmith [Wes76]. Guralnick [Mor77a]. Guss [Rop76d]. Gustafsson [Hor79]. Gustina [Con71a]. Gutman [Con73e]. Gutnov [Bri71a]. Guy [Buc75a, Hel77]. Gvishiani [Par79].

H [And74, Bal73a, Bal75b, Bed71, Bee70, Bis76b, Bla76a, Bry73b, Buc73, Buc75a, Car72b, Cli71, Coh74, Con72e, Con74a, Con74i, Con77f, Con79f, Cos71, Cot70, Cow75a, Cra79a, De 71, Dev76, Dib75, Dot75, Eat73, Ehr74, Emm72, Eze77, Flo72, Fre73, Fuk75a, Fus73, Gue77, Ham71, Har75a, Hec70, Hel71, Hig73, Hil75a, Hok79a, Hud75b, Hud79, Koh71a, Lak70, Lay71b, Le 75, Lie74d, Man72, Mid78, Nie70b, Per72, Pos76a, Rog75, Rol78, Rop76d, Ros75c, Sha72c, She78b, Smi72b, Smi73a, Sum75, tZS76, Tha73a, Tro75, Vac71, Wal73a, Wat72a, Wes76, Whe73, Wil76, Win70, dC70, dC72a]. H. [Car75a]. Haag [Con76d]. Habermas [Bun74]. Habitat [Bay78a]. Hacienda [Gli73b]. Hacker [Sha79]. Haden [Per72]. Hagan [Gla78]. Haggerty [Wei75]. Hagley [Bat74]. Hague [Ros79c]. Hahn [Bri72]. Half [CZ73, Tho75b]. Hall [Bry77a, Str78, Can74, Hac78, Kre74, Smi73a, Smi74b, Swe78a, Cha73a]. Halleux [Ste76e]. Hallion [Cla78]. Hamarneh [Ger74]. Hamilton [Sin70b]. Hammond [Con74i]. Hand [Mei72, Pen75]. Handbook [Whi73c]. Handloom [Lea70]. Hands [Swi78]. Hann [Bal75a]. Hanns [Han79a]. Hans [Bry71b, Con74d, May71b, Rob76d, Wei78b]. Hans-Joachim [Rob76d]. Hansen [Con74d]. Hap [Hal73]. Happ [Ros75c]. Hardin [Wei78a]. Hardrock [Jac75b]. Hardware [Rob75a]. Harless [Cal71b].

Harley [Gar72a]. Harness [Ber74, Gil79]. Harold [Bel78, Had79, Hig73, Koh76b, Mor71, Mul73b, Rea78, Sin70b]. Harpers [Eze78]. Harriett [War70]. Harrington [Spi78]. Harriot [Kre76]. Harrison [Ell77]. Harriss [Wat76b]. Harrisville [Bat71]. Harry [Ash71, Dev78, Emm76, Har71a, Rae73a]. Hartman [Dra78, Hac71]. Harvest [Fus71]. Harvey [Sko76, Sch78c]. Haskins [Mor70]. Hasn’t [Cra79b]. Hassan [Con74c]. Hasslof [Cha73b]. Haug [Aga78]. Hausen [May77]. Haute [Pic77]. Havens [Fuk75a]. Hawk [De 71]. Hawthorne [Dre75]. Haydn [Sch78a]. Hayter [Wik72]. Heal [Bru76]. Health [Gro72, Kra77b]. Heat [Bry73a, BR74, Hin71b, Mol79]. Hector [Adr70]. Heilbron [Car75a]. Heilbroner [Bra76a]. Heinrich [Ait79]. Heinz [Bou78, Lie76c]. Heisler [Nel79]. Held [Sha76b]. Helen [Bil79, Sch78a]. Helena [Tuc73]. Helfrich [Bed71]. Helfrich-Dörner [Bed71]. Hellie [Esp72]. Helmer [Kra76]. Helwan [Nee70]. Hempstone [Sum75]. Henderson [Hau76]. Hendricks [Abb72a]. Henning [Bec78]. Henrietta [Sal73]. Henry [Bic73, Bri74, Bri77b, Bul76, Con75d, Con78e, Con78c, Fri79a, Hea71, Kob74, Mil76, Mor72a, Par74, Rae71a, Wal71b, Wel70, Wel73a, Fit73a, Hin71a, Wik70, Woo72b]. Henry-Russell [Con78c]. Her [Buc70a, Cow75a, Hau76]. Heralds [Wil70b]. Herbert [Gol75, Le 72, Mic70b, San77, Sch71b, Sha72a, Shr73, Con79c]. Hereward [Vac72a]. Heritage [Bri77a, Bat77, Bra70b]. Herman [Bal74, Bir76, Dal70, Tro75, Woo77b]. Hermann [Hac73b, Str72]. Hermes [Woo79a]. Herndon [Koh71b]. Hero [Kre74]. héroe [Gli74]. Heron [Bil77]. Herrera [Gli75a]. herrerías [Gli75a]. Herrington [Ros79a]. Herselle [Con79d]. Herskowitz [Emm78]. Hertz [Ait79, Ait79]. Herzogtümern [Bec78]. Hessen [Hal72a]. Hetzler [DeG71, De 74c]. Heuvelmans [Rou75]. Hewlett [Rey75, San74]. Heyman [Gil73a]. Hiam [Sol76b]. Hiawatha [Ove71]. hidraúlicas [Gli75a]. Hieroglyphs [Cas76b]. hierro [Gli75a]. Higgs [Row76a, Wal74]. High [Cha71b, Fow77, Fra74a, Hig73, Hug76b, Kap74a, May75b]. High-flying [Cha71b]. High-Speed [Hig73, May75b]. High-Voltage [Hug76b]. Higham [Emm79, Emm74, Rop72, Rop76a]. Highland [Rol72]. Hildebrand [Con75b]. Hill [Kle77, Per72, Mul74c, dSP75]. Hilton [Bar71a, Mas73]. Hindle [Jen77a, McM76a]. Hinge [Dot75]. Hiram [She74]. Hired [Swi78]. hispana [Gli75b]. histoire [KF76]. Historia [Mos75b]. Historian [Smi70c]. Historians [Rür74, Slo70]. Historic [Bat70b, Bay76b, Con73e, Fra74b, Gar72a, Jac79, MCA74, Per78b, Smi77a, Wat79a, Woo77a, Gre72]. histórica [Gli70a, Gli75b]. Historical [Bed76, DeG77, Dra70, Emm79, Fra74b, Fuk77, Fus74a, Gre72, Koh71b, Lak70, Lei71, Lew73, Maj77, Mos79, Ros70, Smi70a, Smi70c, tZS76, Woo72a, Gor74, Tur76, Whe78]. Historiographic [De 74e, May76]. Historiography [Bur70, Mul74e]. History [Ada74, Alt70, Ano70f, Ano74g, Ano75n, Bal75a, Bal77, Bat70a, Bat71, Bay78b, Ber77, Bis70, Bis76b, Bor71a, Bra76c, Bra77a, Bra77d, Bra79a, Bri77a, Bri78b, Cah76, Cam74, Cha71b, Cla77, Con72a, Con73b, Con74d, Con77b, Con77a, Con78e, Con79a, Cra79c, Dan70a, De 74a, De 74b, De 74e, DeV75b, Dug79, Emm71, Fai71, Fen78, Fer73a, Fer74a, Fie77, For79b, Fri79b, Fus73, Ger76, Ger74, Gie78, Gil73a, Gol74a, Goo70c, Goo71, Goo72, Goo73, Goo74b, Goo74a, Goo75, Goo76, Goo77, Goo78, Goo79, Gra72, Hac71, Hac78, Hac79, Had77c, HS70, Hal74, HV78, Har78b, Hea72, Hec70, Hey74, Hil76, Hod79, Hok79b, Hop75, Hou74, Hud75a, Hug70, Hug79b, Hun75, Jac75b, Jen73a, Kop77, Lei70, Mas77a, Mid78, Mos75a, Mul74e, Mul74d, Nut79b, Nut79c, Ord70, Orł73]. History [Par71a, Par72a, Par78, Pau71, Pen77, Pen79, Per74, Pia71, Pon79, Por74, Ram70, Ras74, Ras76, Ris78, Ron78, Rop76d, Rop76b, Rou72, Rür74, San73b, San74, Sch74, Sha72a, Sha73, Sha76c, Sha78c, Sha79, Sha72c, Sma78, Smi75a, Smi76b, Ste76c, Swe78a, Tha73a, Tha76, Top72, Ung72, War76, Wes78, Whi77b, Wil71b, Win71, Woo72a, ZIL77, dC70, Gra75, Lie75c, Lie79a, Mos75b, Pia70, Con79a, Bil73, Buc71a, Clo70a, Coo77, Dru76, Eme78, Emm74, Fer73c, Had75b, Had76b, Had77b, Had78b, Jer78a, Jer79, Koh77, Lay77a, Lea76, Lea79, Mal78, Nor78, Pur71, Rav72, Rol73, Rop72, Rop76a, She78a, Smi72b, Tay75, Teb74, Wat72a, All79b, All79c, Aus76, Bry77a, Car71b, Cha70, Con74f, Con75d, Con78a, Fer71a, Gat73, Isa76, Lan73, May70, Ran72, Sal73, Sha72b, Smi73b]. History [Sta77a]. Historycznym [Mos79, Pia71]. Hitchcock [Con78c]. Hitt [Whi72b, Whi73a]. Hix [Pet75]. Hobsbawm [Sal77]. Hodas [May78a]. Hodges [Wal71b]. Hoelscher [De 71]. Hoffecker [Bay75]. Hoffmann [Con79a, tzS75]. Hogg [De 74d]. Hohenberg [Mok77]. Holbrook [Wik77]. Holcombe [Cow75b]. Holes [Goo70a]. Holland [Mok77]. Hollingworth [Jer70a]. Hollis [Sch78a]. Hollow [Jud74a]. Holt [Wik79]. Holtez [Hud76]. Holton [Cut74, Top79b]. Home [Cow76b]. Homeric [Web74]. Homesteader [Bal74]. homme [Row73]. Hommel [Siv72]. Homo [Row73]. Honor [Aga77, Str78]. Honour [Hau76]. Hoover [Dib72]. Hope [Bee79]. Hopkins [De 74b, Wat72a]. Hoppen [Tha72]. Horizons [Sal73]. Horizontal [Gad71, Mue74]. Hornung [Nes77]. Horologium [Kel70c]. Horse [Whi77b]. Horsecars [Pos76a]. Horsemen [Web74, Mei71b]. Horsford [Rez70]. Horwitz [Fri79b]. Hospital [Cat76]. Hot [Smi72a]. Hot-Rolled [Smi72a]. Houck [Nel79]. Hough [Fel76b]. House [Con79a, McK79, Bri78a, Pet75]. Household [Cow76b, Van78c]. Houses [Mos75a]. Housewares [Gol74a]. Housewife [Tuc73]. Housework [Van78c]. Housing [Gai75, Str73a]. Houston [Swe71]. Howard [Jac79, Mid78, Pau71, Rei71]. Howell [Pen79]. Howse [War76]. Høyrup [Ros79d]. Hudson [Har74a, Pos76b, Spi78, Alt73a, Bil73, Hit76, Wat72b]. Huggett [Fus76b]. Huggins [Le 75]. Hugh [Buc76a, Nor77a, Oka78, Pac74, Swe75, Dor79]. Hughes [Smi76d, Dan72, Sch77a]. Huguenot [Kel73]. Hulman [Pic77]. Human [Aga76, Bra76a, Fow77, Gha70b, Hea71, Hol79, Lew74, Ros75d, Sch71b, Sel77, Tar79, Win70, Zin79, Zup77, Per73]. Humanism [Fer71c]. Humanist [Kel70c]. Humanities [Caw77, FS78, Sto71, Win70]. Humanization [Nel79]. Humanizing [Gie76a]. Hume [Buc79, Har78b, Rae76]. Humelsine [Cot70]. Hunch [Fra73]. Hundert [Bry70]. Hundred [Bra77c, Hok79a, Lei70, Rea78, Sha78a, Wil70b]. Hunter [Jac77, Wil78]. Hunter-Gatherer [Wil78]. Huntington [Cha73a, Hin72]. Hutchings [Bla78a]. Hyde [Per78b, Sta78a]. Hydraulic [Lay79]. Hydraulics [Jac77]. Hydro [Bal75b]. Hydro-Electric [Bal75b]. Hydrology [Bis76a, Car71b]. Hydrostatic [Kel71b].

IAHS [Bis76a]. Ian [Aus74, Bay76a, Gar72b, Kul79, Mar76b, May75a, Seg79b, Wel73b]. iberoamericana [Gli75b]. IBM [Süs71]. Ibn [dSP75]. Icarus [Le 71]. Icohtec [Ano75g, Bal71]. Iconic [Le 75]. Idea [Hac74, Lub78, Ran72]. Ideal [Bri71a, Svi77, How72]. Idealism [Bur77]. Ideas [Bur77, Smi77a, Sta78b]. Ideologies [Lay76b]. Ideology [Cro76, Kap71, McL79, Ros72, McL79]. Idiom [Rei70]. Ignition [Sha73]. II [Bri79, Pos79, Ano73h, Ano73i, Ano73j, Ano73t, Hea71, Buc75a, McM76b]. III [tZS73b, Ano73k, Ano73l, Sha78c]. Illich [Kra77b]. Illinois [Koh72]. Illustrated [Sha78c, Siv72, Smi75a, Wat72a, Whi73c]. im [Geb71, Hal72c, Lie76c]. Image [Gor72, Lay71d]. Images [Con74i, Ste76b, Wat78b, Cha77]. Imagination [Top79b]. imeni [Lie75c]. Immediate [All79b]. Immigrants [Jer74, War73]. Imminent [Mee78]. Immortality [Maj75]. Impact [Bil75a, Bos71, Hol79, Lie77d, Mol78, Ove75, Rob76c, Swe78a, Web79]. Impacts [Por76]. Impart [Bla79a]. Imperialism [Noe77]. Implementing [Mon75]. Implements [Fen78, Fus71]. Implications [Lei73, Kel71d]. Import [Rol77b]. Improbable [McM76b]. Improved [Bee75, She77]. Improvement [Lie79b]. Incentives [Hil78]. Including [Koh71b, Bee75]. Incomparable [Bil75b]. Incompleat [Ihd77]. Incorporate [Aga70]. Independence [Dun76]. Index [Ano70c, Ano71e, Ano72f, Ano73m, Ano74h, Ano75h, Ano76f, Ano77g, Ano78f, Ano79d, Fer76a, Fer79b]. India [Cha72b, Gha70a, Gha70b, Rah77, Wei78b, Bag75, Gor71a, Bla76b, Bos71, Gha73b, Dal72]. Indian [Gha73a]. Indiana [Pic77]. Indicators [Fel78]. Indigenous [DeW78, Smi77a]. Indigo [Cok72, Sha71]. Individual [Aga70, Bla76a]. Individualist [Sin70a]. Indonesia [DeG75]. indústria [Gli73a]. Industrial [Alt73a, Bay75, Ben78, Buc71b, Buc74, Buc76b, Cam74, Car72e, Car79c, CZ73, Dau72, De 73b, Dub77, Emm76, Fra71, Gie77, Gol78, Hal76, Har75b, Hin79, Hor79, Hud76, Jer78a, Kur73, Lay79, Lea70, Lie74a, Mar72b, Mok77, Pei72, Per78b, Rae76, Ran76, Rol70b, San73c, Seg78, Smi76a, Sta78c, Ste75a, Str73a, Wat79a, Mul74b, Cow76b, Koh76a, Hor79, Bat77, Bru76, Con74g, Lew73, Lie74a, Mei70, Noe77, Rol70a, San73a]. Industrialisation [Fel76a]. Industrialism [Wik72]. Industrialist [May78a]. Industrialization [Clo70a, De 79a, Lie73b, Oka78, Vol78, De 71]. Industriegeschichte [Son78]. Industrielle [Fre71b]. Industries [Siv72, Ros73a]. Industry [Abb72a, Bal77, Bla75, Bru76, Cal71b, Cha77, Cok72, Con75b, Con75d, Edw72, Fel74, Fel76a, Fri75, Har78a, Hit76, Hol70a, How78, Jer70a, Jer70b, Kob77, Lea70, dLM76, Mat75, McC70b, Per72, Pur76b, Qui71, Rae79a, Rob76b, SA77, Sta78a, Ste76c, Vol78, Buc76a, Car71a, Con77b, Gie76c, Gol74a, Gor74, Har79a, Hud75a, Joh73, Lie77c, Lie78c, Pen75, Rae79b, Sim71]. Inflation [Mur79]. Influence [Gor72, Koh74, Rez70]. Influences [Rey79b, Sin79a]. Informal [Sha73]. Information [Fer73a, Hil77, Nev79, Pei72, Del76]. Ingenierus [Abr79]. Ingenieure [Lie76c]. ingenio [Gli75a]. Ingenious [dSP75, Sta78b]. Ingenuity [Bur77]. Inglis [Ste71]. initiating [Bal71]. Inkeles [Bla76a]. Inland [Goo70b]. Innis [Mul73b]. Innovation [Bar70b, Bat74, DeW78, Har79b, Hil78, Jer73, Lie77c, Mag76, Man73, Mei74, Nie70b, Smi73a, Str73a, Str76, Bar70a, Lay71c, Per73]. Innovationen [Rap77]. Innovations [Mas77b, Gil72a]. Inquiries [Nel79]. Inquiry [Boz77, Bra76a, Con79b]. Insect [Per78a]. Insect-Control [Per78a]. Installment [Smi74c]. Institute [Bel78, FS78, Hei77a, Pic77, Sha76a, Swe71, Tha76, Sew72]. Institution [All79a, Bri72, Sha76a, Sha76b, Tha73b, Lew77]. Institutional [Cah76]. Institutionally [Ros75d]. Institutions [Kap71, Hin71b, Jen73a]. Instruction [Day74]. Instrument [Don77, Nut78]. Instrument-Maker [Nut78]. Instrument-Makers [Don77]. Instruments [Bis70, Nut79c, Sch78a, WK71, And74, Bis71]. Insulin [Maj77]. Intellectual [San73c, SG75]. Intellectuals [Seg78]. Intelligence [Sok75]. Intelligentsia [Par79]. Interaction [Bla78a, Hug76b, Jev76, Koh76a, Lay70b, Pur74, Smi70a, Wei71, Fel76a]. Interborough [Con72a]. Interchangeable [How78]. Interdependence [Ros79b, Dev78]. Interest [Van71]. INTERFACE [FS78]. Interference [Jac70]. Interferences [Cha71b]. Interiors [Mar76a]. Interlude [Mar76a]. Internal [Pos75a, Bry77b]. Internal-Combustion [Pos75a]. International [Ano70f, Ano75n, Bra77d, Buc76b, Coo71, De 79b, Emm71, Gol77c, Hac78, Hal74, Jer76b, Kel74, Koh78, Pos79, Sha78a, Sha78b, Sko76, Smi76d]. internazionale [Kel70a]. Interplanetary [Swe73]. Interpretations [Con71d, Mul74b]. Interpretive [Bla74b]. Interrelations [McG76, WK71]. Introduction [Bri70, Cha73b, De 75a, Fow78, Jon78, Maq70, MM73, Pur79b, RM76, Sky73, Rob71, Sch78a]. Inventing [Fri77]. Invention [Cal70b, Hou78, Jac70, Jev76, Maj75, Maj77, May75a, Pos72, Tar74, Wik79, Fer70a, Tav74, Bis71, Fre71a, Mag76]. Inventions [Buc70a, Dib79, Fer78a, Jer70b, Kel70c, She77, Con71a, dC72a, She75b]. Inventor [Dan72, Jac71, Kel73, Rae73a, Sch77a]. Inventors [She77, Wis79]. Inventory [Per78b, Wat79a]. investigación [Gli75b]. Investigation [Bos71]. Investigations [Kap74c]. Investment [De 79a, Syl72]. Invisible [Jer74, Sto74]. Involvement [Mar77a]. Inyokern [All79b, Sha72b]. Ionescu [Sal70]. Iraj [Mat76]. Ireland [Gre79, McC70b]. Irene [Kel71d]. Irish [Jer78a]. Irma [Hog75]. Iron [Har75a, Laf72, Lie75b, Mei71b, SA77, Sch78c, Smi72a, Sta78a, Pia71, Hud79, Mur79, Sch78b, Wer70]. Iron-Age [Mei71b]. iron-carbon [Pia71]. Ironbridge [Hud75b]. ironwork [Tyl77]. Ironworks [Hil76, Hud76]. Irrigation [Fra74a, GD73, Str71b]. Irwin [Wal76b].

Isambard [Emm77]. Ishwaran [Cha72b]. ISIS [Bri78b, Jen73a, Bri78b, Jen73a]. Islamic [Con74b, Hil74]. Isles [Gre72]. Isolation [Bai74]. Israel [Sol76b, DeG75]. Issekutz [Geb72]. Issues [Mar77c]. istorii [Gra75, Lie75c, Lie79a]. Istoriia [Lie75c]. Italia [Rap75, Clo75]. Italian [Wat78a]. italiana [Kel70a]. italiani [Clo70b]. Italy [Bon71, Fle77, Mok77, Noe77]. Ithiel [Mol78]. IV [Ano73n]. Ivan [Kra77b]. Ives [Smi72c]. Ivo [BR74].

J [All79b, Bee79, Bla79a, Bla79b, Bri76a, Bru73, Bry71a, Cal70a, Car75a, Car79b, Car79c, Chu78a, Coh74, Con72b, Con74i, Con76b, Con78e, De 75a, Dev75a, Don77, Edw72, Ell76, Fer70a, Fer71a, Fer76c, Fer74b, Fus71, Fus74a, Gal75, Gar72a, Gil70, Gla78, Gli73b, Goi71, Gol75, Gol77c, Goo74a, Hac72, Had78d, Har78b, Hil75b, Hok79b, Kle77, Lie74a, Lie74d, Moe71, Mul72a, Mul74d, Nel79, Nor75, Nor77a, Pac78, PS79, Pos76b, Pos79, Rae73b, Rav72, Ray77, Rob71, Rog75, Ros79a, Ros79c, Row79, Sal77, Sch71b, Sin79b, Smi76b, Sol76a, Spi74, Ste75a, Wal73b, Wat72a, Wel73a, Whi73c, Whi76a, Whi77b, Wil70a, dC78, vS72]. Jac. [Mul72a]. Jack [Car72c, Con71d, Mei70, Per72, Vac70]. Jackson [Con71c, Smi76b, Con73d, The74]. Jacob [Eis75a, Emm79, Sin70b, Tar74]. Jacobs [Wri76]. Jacques [Aga78, Boz75, De 73a, Gil73a, Mor72b, Sol76a, FM79, Kel73]. Jahre [Bry70]. Jahrhundert [Fuk75b, Geb71, Hal72c]. Jahrhunderten [Bed71]. Jahrhunderts [Bec78, Hal72b, Rob76d]. Jairazbhoy [Con74b]. James [Ber70b, Bri78a, Bry71a, Con73c, Con74e, Eat73, Ehr73, Hea72, Hil77, Hol76, Kap77, Law71, Lev79, Nor78, San73c, Sch74, Seg78, Sha79, Smi76d, Ste75b, Tav72a, Wat79b, Whi74, Whi76a, Car76, Rob72, Cra75, Fer70a, Hei77b, Kle77, Pur71]. Jan [Fel76a, Sko75]. Jane [Had75a, Süs70, Wel71]. Janet [Ber74]. Jannis [Maj74]. Janota [Bry76a]. January [Bat74, Ano70g, Ano71g, Ano72h, Ano73r, Ano74j, Ano75j, Ano76h, Ano77i, Ano78i, Ano79g]. Japan [Fuk77, Fuk75a, Hac77, Siv75, Aus76, Uch75]. Japanese [Dru76, Bed71, Ino76, Kob74, Kob77, Oka78]. Japans [Fuk75b]. Jardim [Rae71a]. Jay [De 79a, Goi70]. Jazari [dSP75]. Jean [Boz75, Hud75a, Jen70]. Jean-Jacques [Boz75]. Jeannette [Wer78a]. Jeffcott [dC72a]. Jefferson [Cra78]. Jeffrey [Dru76, Teb74]. Jenkins [Cha77, dLM76]. Jentina [Coo74]. Jeon [Hen76a]. Jeremy [Hin71a]. Jerold [Mul74a]. Jerome [Hun75, Sta78d]. Jerry [Kap74a]. Jersey [Sal73]. Jervis [Rou71]. Jerzy [CZ73, Par72a, Smi76c]. Jesse [Cal74]. Jethro [Koh74]. Jevons [SG75, Sta77b]. Jim [Ove71]. Jimmy [Hal73]. Jo [Row72a]. Joachim [Rob76d]. Joan [Kel71a, Mar79a, Pal79, Ste76a, Wil71b]. Joel [Cai78, Ger79]. Johan [Alt70]. Johanna [Ait79]. Johannes [Dob78]. John [Bat71, Bri73, Bry77c, Buc79, Bud78, Cas73, Cla77, Coh74, Con70b, Con70a, Con71c, Con73d, Con73e, Con74a, Con77f, Con79f, Dra76c, Dre75, Dur79, Eys74, Fei77, Fer70b, Fer72, Fer74b, Fre76, Fus74b, Gai75, Gol76, Gro72, Hin71a, Hol70a, Kap78, Kra76, Kre76, Lam71b, Lei71, Lie73b, Lie73c, Min77a, Mon75, Nel79, Nev79, Nut78, Ols77, Pet75, Pos76a, Pos78, Pur78, Ran76, Ras74, Ras76, Rol77a, Rou71, Sch70, Seg77a, Sha73, Sha76c, Smi71b, Smi76d, Spi75a, Spi78, Str75, Tho76, Vac72b, Van71, Wei78a, Wer78b, Wik70, Wik73, Wil76, Wol78, Zin79, Lol78, Smi73a, Buc78, Han79b, Kel70b, Min79b, Mor74]. Johnson [Had74b, Bla71, Dev78, Mas77b, Nes77]. Johnston [Dal70]. Joiner [Bul76]. Jon [Ihd77, Vol78]. Jonathan [Kap74b, Wei73]. Jones [Mar77b, She75b]. Jorge [Tag72]. Jorpes [Mul72a]. José [Gli70a]. Joseph [All79c, Bak76, Bla73a, Bri74, Bri77b, DeG77, Fai71, Har76a, Hol70b, Law71, Lie77c, Maj75, Maj77, Pen77, PS71a, Sch78a, Sch77b, Sin72, Smi72c, Sta78b, tZS73b, Wat76b, The74, Gil73b, May75a]. Joule [Car76, Mol79]. Journal [Hey74, Hin71a, Jen72, Had70, Had71, Had73, Had74a, Had75a, Had76a, Had77a, Had78a]. Journals [Fri79a, Sch78c]. Jozeph [DeG75]. Jr [Bri78a, Mon72, Sin70a, Wer79, Whi75a, Str78]. Jr. [Bla79a, Bla76b, Bry77b, Bud78, Con74a, Con77f, Con79f, Cra75, Dre76, Ell77, Fit73b, Fra73, Gie78, Hab72, Hil70b, Jac79, Joh78, Kap72, Kni77, Kop77, Lay70a, Lay71a, Pos76a, Rav70, Rog77, Seg77c, Slo72, Smi75b, Smi77d, Tho76, Top72, Wes76, dC72a]. Juan [Gli75a]. Juanelo [GD76]. Judgment [Aga76]. Judith [Hei77a, Kel71d]. Judson [Eat75]. July [Bal71, Bel78, Kra77a, Ano70i, Ano71i, Ano72j, Ano73u, Ano74l, Ano75l, Ano76j, Ano77k, Ano78k, Ano79i]. June [Bal71, Bel78, Bon71, Bra77d]. Jupiter [Nie70b]. Jurgen [Bun74, Con74d]. Justice [Gli79]. Justus [Bee76].

K. [Par70]. Kahn [Woo77b, Con75b]. Kakar [Lay71c]. Kallab [Hel77]. Kanter [Gar78]. Karin [May77]. Kariya [Kob77]. Karl [Lie76c, Par70]. Karl-Heinz [Lie76c]. Kash [Bra77b, Fla75]. Kasper [Del76, Tri72]. Kasson [Seg77a]. Katherine [Cha70, Sch78c]. Kauffman [Wel73a]. Kaufman [Aga70]. Kay [De 73b]. Kaye [Gol77a]. Kazushi [Kob74]. Keating [Gol76]. Keefer [Bal73a]. Keith [Fra73, McM75, Ste76d]. Keive [Teb74]. Keller [Pen79]. Kelley [Dru76]. Kemeny [Cas73]. Kemper [Fei77]. Ken [Gla78]. Kennedy [Rol72]. Kenneth [Alt73a, Bil73, Bla77, Bla79b, Cha72a, Cra79a, Le 71, Rae79a, Smi72c, Sti77, Tan78, Wat72b, Wel71, Wel77]. Kenricks [Rob75a]. Kensington [Buc70b]. Kenya [De 78]. Kette [Fre73]. Kettering [Les79]. Key [Dre75]. Khare [Fer74b]. Khatri [Fer79b]. Khramoi [May70]. Kibbutz [Swa70]. Kidder [Con79a]. Kiddo [Kap77]. Kidney [Cra76a]. Kilbridge [Mon71]. Killers [Rou75]. Killian [Bri78a]. Kim [Lie79a]. King [Rea78, Fri78a]. Kingdom [CZ73, Emm77]. Kingsford [Hig79a]. Kingsford-Smith [Hig79a]. Kintner [Sko76]. Kip [Wat79b]. Kipp [Emm79]. Kirker [Mor71]. Kistiakowsky [Bri78a]. Kit [Ber77]. Klapper [Pos75b]. Klass [Wei79, Hac72]. Klein [Gat73]. Knaack [Bri79]. Knight [Hod79, Mon72, Ros79c]. Knitting [Jer70b]. Knodel [Fre76]. Know [Lay76a]. Know-how [Lay76a]. Knowledge [Chu79, Lay74, Ols77, dSP75, Sto74, Sta77b]. Knowles [Bis70, Bis71]. Koch [Whi72a]. Kohle [Lie75b]. Kohlmeyer [Fra76]. Kohlstedt [Ros76b]. Koivumaki [Kel71d]. Kolankiewicz [Byr74]. Kolin [BR74].

Kolomenskogo [Lie75c]. Kolomna [Lie75c]. Konstantinovich [Par71b]. Koparek [Mos79]. Kopp [Con71e]. Korbin [De 79a]. Korea [Hen76a]. Korneyev [Car78d]. Kornwolf [Eat73]. Kostelanetz [Lew70]. Kostof [Cra79c]. Krafft [Sch71a]. Kraftfahrt [Bry71b]. Krankheit [Geb71]. Krantz [Mok77]. Kranzberg [Buc73, Har76a]. Kravchenko [Lie74b]. Krinsky [Con79d]. Krishan [Hor79]. Krohn [Kap71]. Kron [Ros75c]. Kuhn [Top79a]. Kuhns [Mul74b]. Kuibyshev [Lie75c]. Kuibysheva [Lie75c]. Kultur [May78b]. Kulturelle [Hal72c]. Kumanev [Lie78a]. Kumar [Hor79]. Kunst [Har73b]. Kunstuhrmacher [Bed71]. Kupferseigerns [Osp79]. Kushlin [Lie77a]. Kusum [Ols74]. Kuuse [Fel76a]. Kuzin [Par70].

L [Bee73, Bra76a, Bra79b, Car75a, Car72d, Cha72a, Cli71, Con77e, Cra79a, Cro71, Dau72, Dib75, Fer73c, Fla75, Fra71, Fra73, Gli79, Had78d, Hil76, Kel70b, Lie76a, Lie77b, Mar72a, Mas77b, Mat75, McC70a, McN78, Min77a, Nor77b, Pau71, Ram70, Rob70, Rop76b, Sch76, Sch77b, Seg77c, Sha76a, Sta77a, Swe75, Tav72a, Teb74, Tof72, Tuc73, Web74, Whe73, Whi73b, Win78, Fre71b]. Labor [Car71a, Jac75b, Kra75, Sau72, Swi78, PS71a]. Laboratory [All79c, Dev75a, Ihd77]. Labour [Min77b]. Labyrinths [dC70]. Lace [You72]. Lace-Boot [You72]. Lach [Nee79]. Lacunae [MGM78]. Ladies [Cow75b]. Ladislao [Dib74, Ste76f]. Laird [Fra76]. Lake [All79b, Sha72b, Cai74]. Lakefront [Con79e]. Lakes [Hil70c]. Lambent [Gol76]. Lancashire [Par74, Pen75, Jer76a, Jer76b]. Land [Cro71, Fus76b, Mat74, Ras74, Ros74, Rot72, Sky73, Fra74a, Koh70c]. Land-Surveyors [Sky73]. Landels [dC78]. Landscape [Bed76, Rot72]. landwirtschaftlichen [Fuk75b]. Lane [Byr74]. Lang [Con71f]. Langdon [Car78a]. Laroche [Cha71c]. Larry [Bil75b, Sel78]. Lars [Jer76b]. Larson [Sal73]. Lasby [Hac73a]. Last [Con72a, Min77a]. Laszlo [Ros73b, Ros75b]. Late [Kre77]. Later [Pac74]. Latin [Han79b, PS71a]. Latrobe [Fri79a]. Lattin [War70]. Laudan [Sel78]. Law [Bar73a, Jac71, Mag76, Rob72, Fer70a]. Lawless [Fer78b]. Lawrence [Bed71, Blo77, Edw72, Goo74a, HS78, Hei70, Hei77a, Wal73b]. Laws [Cok72, McL75, McL76, Ven76]. Layton [Hab72, Smi75b, De 74e, Lev75]. Lazare [Car73]. Lazerson [DeV75b]. Leadership [Rae71a]. Leap [Sca72]. Leary [Smi77d]. Leavitt [Jer70a, Vac77a]. Lebedev [Lie77a]. Lectures [Buc71a, Fai71, Sta77b, Sto71]. Lee [Cow75b, Con73a, HS78, Hil70c]. Leene [Coo74]. leeren [May71b]. Left [Lew70]. Legacy [Cla78]. Legal [Jac75a]. Legend [Bry71c]. Legislation [Kev71]. Legitimacy [Bai76]. Leighton [Bla74a]. Leiss [Wag73]. Leisure [DeG75]. Leisure-Time [DeG75]. Leitbilder [Hal71]. Leiter [Bed71]. Lenore [Cal74]. Leo [Por76]. Leon [Bai74, Kel71a]. Leonard [Mul74d, De 78]. Leonardo [Dib79, Mar72c, Ste76f, Ano70d, Ano71f, Ano72g, Ano73o, Ano74i, Ano75i, Ano76g, Ano77h, Ano78g, Ano79e, Bon71, Ret72, Con74h]. Lerna [Wal73b]. Lerner [Sol76b]. Lesley [Con75d, Vin79]. Leslie [Cra77]. Lessons [Sol76b, Dru76]. Lester [Hun71, Vac72a]. let [Bra72, Lie75c]. Letters [Ait79, Bal74, Car75a, Hei70, Jer70a, Law71, Had74b, Rol74]. Leuba [Con72b]. Leupold [Fer71b]. Level [Lie78c]. Levels [Hil71]. Levere [Dib75]. Lévi [Maq70]. Lewicki [De 75b]. Lewis [Hol79, Mul72b, Whe78, Bri73, De 79b, Sch78b, Sin70a, War71, Goi73]. Lex [Mer76a]. Liberalizers [Pos76d]. Library [Cha71a, Sha76b, Goo70b, Wil70b]. Liebhafsky [Bir76]. Liebig [Bee76, Koh74]. Life [Bri71b, Car75a, Con71c, Cow79, Dib75, Fus74b, Min77a, Min77b, Mon75, Mor74, Mul72a, Pur70, Siv72, Str73b, Van78b, Wal76b, Fer74b, Gie76b, Hec73, Kul79, Lay70a, Wik73, Win78, Gre79]. Lifesaving [Per70b]. Lifshey [Gol74a]. lifting [Hil74]. Light [Hin71b, Top74, Whi77a, Whi78, vO78]. Lighthouses [Bla78b]. Lighting [Whi77a, vO78]. Lillian [Van78b]. Limits [Kni77]. Lincoln [Dor79, Ash71]. Linda [Fer79b, Kel71d]. Lindsay [Fra74b, Hud75a]. Lindstrom [Hai79]. Line [Car79b]. Linear [Cas76b]. Lingenfelter [Jac75b]. Linguistic [Mea71]. Lionel [Whi71c, dc72b, Buc75b]. Liquid [Sha73]. List [Fus71]. Literary [Nee79]. Literature [Gie77]. Little [Dre75]. Lived [dC79]. Liverpool [Ste70, Hil75b]. Living [Van78c]. Livingston [Top74]. Lloyd [Con72e, Con79a, Ray77, Seg79a]. Loaded [Car79a]. Loading [Bri70]. Lobel [Bay76b, Gre72]. Location [Fel74]. Lockheed [Hig73]. Locomotive [Con71b, Pos72, Pos75a, Spo72, Whi72a, Lie75c, Pos75a, Con77f, Whi73b]. Locomotives [Con74a, Whi72c, Whi71c, Whi73d]. Logistics [Rey79a]. Logsdon [Van71]. Lon [Han79a]. Lonaconing [Sch78c]. London [Con71c, dLM73, Rob75b, Bra75, Con71c, Con73e, Bri76b]. Long [Bud78, Meu77, Tuc78]. Long-Distance [Tuc78]. Long-Term [Meu77]. Longitude [Tho75a]. Longview [Bud78]. Look [Bat73, Bat77]. Looks [Kel70c]. Lopata [Tuc73]. López [Gli70a]. Lora [Gli70b]. Lord [Spo73, Gor71a]. Lore [Smi75a]. Loren [McC73]. Loria [Bul75, Clo75]. Lothar [Osp79]. Lotz [Wat78a]. Louis [Bry73b, Con79e, Gie76b, Mor78]. Louise [Com74]. Louisville [Gat73]. Love [Wat78b]. Lovell [Sha78b]. Lovett [Fer73a]. Lovins [Wil76]. Lower [JR78, Per78b]. Loyd [Top72]. Lubos [Win71]. Lucas [Rea78]. Lucie [HS78]. Lucie-Smith [HS78]. Lucille [Pau71]. Lucius [Smi70c]. Ludwig [Lie76c]. Luigi [Clo70b]. Luís [Gli75a]. Lumber [Ste76c]. Lumbermen [Bul76]. Lunar [Hac79, Swe78a]. Lundeberg [Fri76]. Lundström [Alt70]. Lure [dC74a]. Luxury [Whi77a]. Lyle [Bry77b]. Lynch [Hig79a]. Lynn [Rav70, Sma78, Str78, Whi75a, Dub77].

M [Adr70, Aga77, Art70, Aus74, Bay76b, Bla75, Bra77a, Bra79a, Bry76a, Bud78, Cam74, Cha72a, Chu78b, Con71f, Dal70, De 71, Ehr73, Ehr74, Fei77, Fit73a, Fri77, Gar72b, Gla74, Gli70a, Gor74, Gor79, Gre72, Had77c, Har74a, Hau76, Hig79a, Jen73b, Koh78, Kul79, Lev79, Lie77b, Lie79a, Mar76b, Mok77, Mon72, Mor77a, Mul74a, Par79, Pen77, Ron78, Ros73a, Row79, Sal73, Sch76, Sha78a, Sha79, Sky75, Sma78, Smi77d, Sta78d, Sti77, Van71, War71, Wat79a, Wel73b, Whi73c, Wil72, vS72]. M. [Eat73]. M.I.T [Rog75]. M.I.T. [Gre73]. Ma [Par72b]. Maass [Gli79, Coh74]. Mac [Pos76c]. Maccagni [Clo78, Kel70a]. MacDonald [Con77e]. Machina [Rav70]. Machinarum [Fer71b]. Machine [Bat70a, Bat73, Cal71b, Cla77, Gor71b, Hou77, Hou78, Mul72b, Seg77a, Smi73a, Woo72b, Rob71, Cal70b, Vac71, Bic73]. Machinefabriek [Whi72c]. Machinery [Jer77, Mos79, Sin72]. Machines [Ada74, Bat70b, Bee73, Kel73, Kel78, Maa78, She75a, Shr72, Sta78b, Jer70b, Lay77b]. MacKenzie [Kra75]. Mackie [For79a]. MacRae [Rog77]. Mad [dC72a]. Maddison [Hok79b]. Made [Bat70b]. Maestre [Cla73]. Magda [Bri78b, Jen73a, Zin79]. Magdeburger [May71b]. Magisteries [Maj75]. magnetic [Mos75b]. Magnetism [Sha72a]. magnetycznego [Mos75b]. Magnitude [Bil75a, Kop77]. Magnus [dC74b]. Magnusson [dC74b]. Mail [Hec70, Hec73]. Maillart [MCA74]. Main [Fri78b]. Maine [Fri79b, Rol78]. Mainstone [Con76b]. Maintenance [Sta78d]. Maizels [Jun70]. Majesty [Buc70a]. Major [Bla79b]. Majority [Min77a]. Maker [Bul76, Nut78]. Makers [Don77, Smi73a, Wel73a]. Making [Bed76, Car72d, Con76e, Dun76, Lie74a, Mon75, Nee79]. Makinson [Cra79a, Cra79a]. Malcolm [Smi73b]. Mammals [Isa76]. Man [Bed71, Cah76, De 73b, Hig79a, How70, Kel71a, Lew71, Lie75a, Mei72, Shr72, Kah70, Rie70, Tav72a, Aga70, Bic73, Car77a, Cas73, Fer70b, Gie77, San72, Str75]. Man-Machine [Bic73]. Man-science-technology [Lie75a]. Management [Lie74a, Lie77d, Ros76a, Die73]. Manager [Tha73b]. Managerial [Cas74, Fre74]. Managers [Spa70, Jam76]. Managing [Wei78a]. Manchester [Hil75b, Sha76a, Ste70]. Mandelson [Cra77]. Manfred [Mag76]. Manfredo [Hay79]. Manhattan [Pos76b, Con75f]. Mankind [Ran72]. Manners [Laf72]. Manor [Pen79]. Mans [Cah76]. Mansfield [Man73]. Mantua [Wat78b]. Manual [Cot72]. Manuals [Pac74]. Manuel [Gli70b]. Manufactory [Dew76]. Manufacture [Bee70, Bee75, Moe71]. Manufactures [Buc70a, Sin72]. Manufacturing [Siv77]. Manuscript [Wil71b]. Manuscripts [Hou74]. Map [Qui77]. Mappae [Dre75]. Mapping [Cot72]. Maps [Gre72]. Maquet [De 73a]. Máquinas [Gli75a]. Mar [Ord70]. Marcelle [Bri79]. March [Bat74, Kel70a, O’B71, Mei74, Nic71]. Marcson [Tru71]. Marcus [Bry71c]. Marcuse [San73c, Shr73]. Margaret [Bee76, De 73b, Dun76, Jen73b]. Margery [Lei73]. Mariano [She75a]. Marie [Aga78, Kre74]. Marietta [FS78]. Marija [Row72b]. marina [Rap75]. Mariner [Kre73]. Mario [Clo75, Bul75]. Maritime [Cha73b, Ehr73]. Marjorie [She78a]. Mark [MM74, Wil78]. Market [Jen75, Laf72, Sim71]. Markets [Hou77, Ste76c]. Marks [Jen76, Par70, Spr76]. marksistskogo [Lie75a]. Marqueses [Gli73b]. Marquis [Hel71]. Marsden [Dra70, Dra72]. Marshall [Con79a, Lie74e]. Marston [Con73c]. Martha [Fer76b]. Martin [Con77c, Ham71, Nut78, Rey79a, Rou75, She74, Ste76b, Fra71, Hig73, Kle77, Sto73]. Martinus [Dib75]. Marton [Ram70]. Martyr [Dib76]. Marum [Dib75]. Marvin [DeV75b, Gil70]. Marwah [San78]. Marx [Par70, dC74a]. Marxist [Lie75a].

Mary [De 74f, Eis76, Lew74, Fus71]. Mason [Con74i]. Mass [Jen75, Pos78]. Masses [Siv72]. Master [Con78e, Had74b, Top74, Rae73b]. Masters [Fri77]. Materiaiy [Mos78]. Material [Wel70, Kul79, Smi70c]. Materials [Ano73j, Ano73q, Cah76, SA77, ZIL77, Cah76, Jac73, Cah76]. Mathematical [Dra76c]. Mathematics [Kel73, Ros79d]. Mathes [Han79a]. Matter [Nel79]. Matthews [dC70]. Matthias [Whi72b]. Mattias [Lie75b]. Matupit [Cro71]. Mature [Bla74c, Hyd73]. Maturity [Har75a]. Maudslay [Woo72b]. Maureen [Smi70c]. Maurice [Buc70a, Cas76b, Mon71]. Maury [Gat73]. Max [Hen76b, Ord78]. Maximilian [Bed71]. Maxwell [Abb72a, Hei77b, Win70]. May [Kel74]. Maya [Wal76b, Con76c]. Maybeck [Cra79a]. Maynard [Wil70a]. Mayr [Har70, Lay77b, Woo72a]. Maze [Bur77]. Mazes [dC70]. M’Bengue [DeG75]. McClelland [Bud78]. McCloskey [Har75a, Hyd73]. McCormmach [Whe78, Tur76]. McCullough [Wat73]. McCutcheon [Jer78a]. McDonald [Fit73b]. McGillem [Woo79a]. McHale [Zin79, Nev79, Sch70, Zin79]. McIlvaine [Bic73]. McKay [Lie73b, Pos78]. McKee [Gar72a]. McKie [Law71]. McLaughlin [Woo79a]. McLaurin [Sau72]. McLuhan’s [McL75, Ven76]. McNally [Hig79a]. McNeil [May75a]. McNeill [Dot75]. McShane [Mer76b]. Mead [Hil70b]. Meakin [Gie77]. Meaning [Sal70]. Means [Swe72]. Measurement [HP72, Zup77]. Measures [De 74c, Lid76, Nut79b]. Measuring [Bed71]. Mechanic [Sha76a]. Mechanical [dSP75]. Mechanics [Kra76, Tha76]. Mechanik [Sch71a]. Mechanisation [Fel76a]. Mechanism [Dra76a]. Mechanisms [Woo72a]. Mechanization [Hil72]. Mechanized [Koh74]. Medal [Ano70d, Ano71f, Ano72g, Ano73o, Ano74i, Ano75i, Ano76g, Ano77h, Ano78g, Ano79e]. Media [McL75, McL76, Vac77b]. Medical [Blo77, Mat78, Kra77b]. Medicinal [Lei71]. Medicine [Mul79b, Nut77, Jen76, Ris78, Ger79]. Medieval [Bac75, Bal78, Bla74a, Dob78, Dre75, Kre73, Str71b, Whi75b, Mue74, She78a]. Mediterranean [Tho76, Kre73]. Medizin [Har71c]. Meeting [Ano70e, Ano70k, Ano78h, Ano79f, AGS^{+}79, BMP^{+}75, BPV^{+}78, Fei78, HS70, HRM^{+}76, LDS^{+}72, LLC^{+}74, MBB^{+}71, PPH^{+}77, RHTH73]. Meetings [Tha73b]. Meier [Par72a]. Mellon [Tar77]. Melton [Sau72]. Melvin [Buc73, Har76a, Koh71b, Smi76b]. Members [Ano74g, Bra77b]. Memoir [Gil73a]. Memoirs [Ait79, HP76, Whi71b, Kar76]. Memories [Gol77a]. Men [Ada74, Bee73, Pos76d, Spa70, Sta78b, Sha76c, Ste71]. Menard [Mor72a]. Mendelsohn [Chu79, Per73, Lay75]. Mendelssohn [Süs76a]. Menéndez [Gli75b]. Mensch [Rap77]. Mercer [Bul76]. Merchant [May78a]. Mercury [Mul74d]. Merovingian [Gil73c]. Merritt [Eze78, Lay71b]. Merton [Kap74c]. Merzbach [Bra79c]. Messenia [Fit73b]. Metal [Hud79]. Metalldrehens [Har73b]. Metallurgical [Nut79a, Smi77a]. metallurgischen [Osp79]. Metallurgy [Par78]. Metals [Jen76, Par72a, Ste75b]. Metalwork [Ste76d]. Metaphor [Lew74]. Metaphysical [Ano73n]. métaux [Ste76e]. Meteorological [Bis70, Bis71]. Meteorology [Mid78]. Meters [Bis76b]. Methodological [MGM78].

Methodologici [Kel70a]. Methodology [Vin79]. Methods [Ada74, Bed71, Bee70, Bee75, Qui71, Ras71, ZIL77, Gla74]. metodologici [Kel70a]. Metropolis [Con79e]. Metropolitan [Mon72]. Mexican [Ham71]. Mexico [PS79, DeW78]. Meyer [Hal72d, Sha72a]. Michael [Bed76, Buc70b, Buc79, Chu78a, Con74i, Gie76c, Hil71, Jer76a, Nic71, Nie70b, Seg79b, Sol76a, Tay75, Whi72a, Wri76, Bri70, Dat71]. Michelle [Cra78]. Michelson [Top72, Top74, Top74]. Michigan [Per78b, Cai74]. Michman [DeG75]. Mickey [Emm78]. Microeconomic [Por74]. Microelectronics [Bra79c]. Microscope [Ram70]. Microscopy [Nut79a]. Mid [Par72b, Pos72, Ste73]. Mid-19th-Century [Par72b, Pos72]. Mid-Victorian [Ste73]. Middle [Cow75b, De 75b, Kre77, Mas77a, Ste76d]. Middleton [Bis70, Bis71, Con75g]. Midland [Con71c]. Midlands [Car75b]. Midwest [Con72e, Swi78]. Mieczysław [Cas76a]. Mighty [Pos76b]. Migration [Lay71a]. Miguel [Cla73]. Miles [Kni77, Bee77, Hud75b]. Militär [Bec78]. Military [Bac79, Esp72, Gil73c, Gre73, Hac77, Hel73, Jon78, Kel70b, Kev75, Per78a, Rop72, Rop76a]. Miliutin [Lie76b]. Mill [Car72a, Pen79, San73b, Sau72]. Millburn [Nut78]. Millennium [Wal73b]. Miller [Rae71b, Top72, Jac75a, Rei71]. Millikan [Hei77a]. Milling [Bat73, Gad71, Smi73a, Pen79]. Millions [Ada74]. Mills [HP72, Min79a, Mue74, Pen79, Rol70b, Str71a, Tan78, Ste76c]. Milsom [Lie73c]. Milwaukee [Mer76b]. Mine [Abb72b]. Miner [Bry73b]. Mineral [Ste75b]. minería [Gli75b]. Miners [Jac75b, Mul79a]. Mines [All74, Bal75b, Hud79, Rol70b]. Mining [All74, Jac75b, Min79b, Pau71, Smi75a, Smi75a, You72, McM77]. Ministerio [Gli70b]. Minnesota [Fit73b]. Minority [Min77a]. Minutes [Tha73b]. mirovoi [Lie76a]. Mirrit [Mul74a]. Mirror [Buc71a, Lay71d, Wil73a, Gor72]. Mirror-Image [Lay71d]. Mirrors [Sim77]. Mirsky [Wer78a]. Mishan [Sin79b]. Mishmash [Woo77c]. Missile [Bri79]. Missing [Kel73]. Missouri [Qui78]. Misunderstanding [McL76]. Mitchell [Per70b]. Mitte [Rob76d]. Model [Con73g, For71]. Models [Bat70b, Wal76a]. Modern [Bed71, Bla76a, Cal70a, Con71d, Con77c, Cra79b, Fuk77, Fuk75a, Fus71, Hud71, IS77, Kel76, Man73, Mar79b, Mok77, Mor72b, Pal79, Rae71b, Rür74, San70, Sch74, Sch71b, She77, Siv75, Smi71a, Str78, tZS76, Vac70, Vac72c, Wal76b, Vac72b]. moderne [Row73, May77]. Modernization [Gla78, Hac77, Kui78, Mat75, DeG71]. Mody [Bed71]. Mogilev [Lie75a]. Mohawk [Har74a]. Moisture [CZ73]. Molinology [Kel74]. molinos [Gli75a, Gli78]. Momadou [DeG75]. Mönche [Hal72d]. Monika [Hal72c]. Monod [Mor72b]. Monsanto [Bee79]. Monte [Per74]. montes [Gli75a]. Montgomery [Mon75]. Monuments [Buc76b, GD72]. Moogk [McK79]. Moon [Hac79, Van71]. Moore [Cal71a, Lew76, Sto70]. Moores [Koh72]. Moral [Gre73]. Moran [Sch74]. Moravcsik [Sol76a]. Morehouse [Gha70b, Gha73b]. Morgan [Con72b, Rol70b, Con72b, Con76e, Dra78, Pur78, Süs70]. Morison [Lay76a]. Morley [Top72]. Morris [All79a, Hog75]. Morse [Fer72, Smi71b]. Morton [Wei79]. Moseley [Car75a]. Moses [May78a]. Moshav [Swa70]. Moshava [Swa70]. Moshe [Gor79]. Moslem [Con78b]. Moss [Buc79, Gar78]. Most [Bee75, Per74]. Mote [Bra79b]. Mother [Jev76, She75b]. Motion [Har78b]. Motivation [Uch75]. Motives [FM79]. Motor [Bry77c, Rae79a]. Motorcycle [Sta78d]. Mountain [Gol77a]. Mousson [Bal71]. Movement [Eat75, Jac75b, Lie75e, Seg77b, Eat73]. Moves [Fra71]. Moyen [Fre71b]. Moyen-Age [Fre71b]. Mozans [Rog75]. Mr. [Sha75]. Muhly [Ste75b]. Mujica [Con78a]. Mule [Hil75a]. Mulkay [Chu78a]. Muller [Gol75, Sch71b]. Mulloy [Cra78]. Multhauf [De 74e]. Multinational [Gol77c]. Mumford [Goi73, Mul72b]. München [May78b]. Munich [Hac73b]. Murphy [Cha70]. Muscovite [Hel73]. Muscovy [Esp72]. Museology [vR78]. Museorum [Koh70a]. Muses [Ger74, Sky70]. Museum [Bat74, Bis70, Bis76b, Buc76b, Eis76, Had70, Had71, Had76a, Hok79b, Whi73f, Woo72a, Ber74, Hil70c, Per70a, Smi74a, Whi70, Jer77, Kle76]. Museums [Bag75, May78b, Ger74, Mor77c, Smi77c]. Music [Tur75, Ash71]. Musical [Sch78a, WK71]. Musson [Buc71b, Har79a, Pur71]. Mutz [Har73b]. Myra [Gol77a]. Mysticism [Sch76]. Mystique [Hol79]. Myth [Lew71, Mul72b, Sha72c]. Mythology [Bar70a, Maq70].

N [Bed71, Goo70b, Hig79a, Hyd73, Koh76b, Lie75d, Lie76b, McK79, Par71b, Pau71, Slo72, Ste75a, Swe73, Whi72b, Wil71a]. N.H [Bat71]. NACA [Kop77]. nach [Osp79]. Nagatsune [Siv77]. Nair [Ols74]. nakanune [Lie78a]. Nakayama [Fuk77]. Names [Bil79]. Nancy [Mat78]. Narrative [All79b]. Nartov [Par71b]. NASA [Bil79, Kop77]. Nascent [Bur74b]. Nashville [Gat73]. Natalie [Kel76]. Natchez [Cok72]. Nathan [Bri74, Bri77b, Lay77c, Sin72, tZS78, Use78, Wil73b, Wil78]. Nation [Fuk75a, Sta78b]. National [Abb72a, Bis76b, Bra77b, Bra79a, Cah76, Con74f, Gol70b, IS77, Kar76, Kev75, Koh78, Lie75b, Low76, Ols76, Ros72, Sal70, Van71, Woo72a, Gli70b, San73b, Eis76, Mil76, Wil71b]. Nationalism [Fuk75a]. Nations [Hen76b]. NATO [Bel78]. Natural [Hei77b, Mor72b, Pen77, Pur74, Vac71, Hal76]. Nature [Bur70, Cah76, Con79b, Cra77, Hin71b, Hod79, Lev79, Pau71, Smi73a, Ste75b, Str75, Wag73]. Naturwissenschaft [Har71c]. Nau [Mil76]. Nauchno [Lie78b, Lie74c, Lie75a, Lie73a, Lie77b, Lie77a]. Nauchno-tekhnicheskai [Lie77a]. nauchno-tekhnicheskaia [Lie74c]. Nauchno-tekhnicheskii [Lie78b, Lie77b]. Nauchno-teknicheskii [Lie73a]. nauchno-teknicheskoi [Lie75a]. nauka [Lie75a]. nauki [Mos78]. Nautics [tZS73b]. Naval [All79b, Buc78, Sha72b, Gla78]. Navigable [McC70a]. Navigation [Per74, Tho75a]. Navy [Rey75, Sha72b, vD75]. nazionale [Rap75]. Nduchno [Lie75d]. Nduchno-tekhnicheskaia [Lie75d]. Neal [Rou71, Wat79b]. Near [Mas77a, San78]. Near-Nuclear [San78]. Necessity [Mor72b, She75b, Sti77]. Need [Fer71b]. Needham [Maj77, Fai71, Maj75, tZS73b]. Needles [Win70]. Needs [Cah76, Zin79, Wei78b]. Negreponte [Vac71]. Neil [Rol70b]. Nelkin [Zin79, Gre73, Str73a]. nell [Clo70b]. Nelles [Bal73a, Bal75b]. nell’evoluzione [Rap75]. Nelson [Gai75, Wil71b, Jam76]. Nemesis [Kra77b]. Neolithic [Smi77a]. Nernst [Süs76a]. Nerra [Bac79]. ness [Fer79a]. Nest [Cai74]. Net [Wei79]. Netherlands [Kel74]. neue [May71b]. Neumann [Tro75]. Neumeyer [Jac71]. Neurometrics [Lol78]. Newcomen [Had78d]. Newlon [Jac79]. Newman [Goi73]. News [Wik77]. Newton [Hal72a, Rot72]. Next [Woo77b]. Niblett [Caw77]. Nice [She78a]. Nicholas [Aga77, Vac71]. Nichoria [Wer79]. Nieburg [Cli71, Whe73]. Niederhuemer [Har71b, Har73a, Har74b, Har76b, Har77]. Nigeria [De 74a]. Nikolaus [Con77a]. Nilsson [Gha70a]. Nineteen [Bar73b]. Nineteen-sixties [Bar73b]. Nineteenth [Con78e, Con79c, Fer79b, Jer74, Jer76a, Per72, Rei71, Ste76a, Str76, Tho75b, War73, Wel71]. Nineteenth-Century [Jer74, Rei71, Ste76a, War73]. Nishida [Aus76]. Nitric [Bee70]. Nixon [Rol70a, Seg77c]. No [Wil71b, Tur75]. Noble [Tyl77, Cha78b]. Nockolds [Rea78]. Noel [Smi77a]. Non [Bal71, Par72a, Wil76]. Non-ferrous [Par72a]. Non-initiating [Bal71]. Non-nuclear [Wil76]. Nordhaus [Fre71a]. Noria [Gli77]. Norman [Bry77a, Car77a, Con73b, Hud71, Rot72, Smi77b]. North [Bag79, Con74d, Con75a, Ehr73, Hen76b, Hud75a, McC70b, Rey74, Bat79, Smi73a]. Northeast [Slo70]. Northern [Bag79, Koh71a, Moe71]. Northwest [Kle77]. Northwestern [O’B71]. Norton [DeV75b, Fra76, Rez70]. Notable [Hud79]. Note [Hig79a]. Notebaart [Maj74]. Notes [Ber77, Gor74, Smi70a, Spi78]. Notizie [Clo78]. November [Bri77b, Spi75b]. Nový [Win71]. Nowhere [Lay76a]. NSWirtschaft [Lie75b]. Nuclear [Bra77e, Del76, Hew76, Mil76, Rey75, San78, Spo73, Nie70a, San78, Wil76]. Number [Ano70h, Ano70g, Ano70i, Ano70j, Ano71h, Ano71g, Ano71i, Ano71j, Ano72i, Ano72h, Ano72j, Ano72k, Ano73s, Ano73t, Ano73r, Ano73u, Ano73v, Ano74k, Ano74j, Ano74l, Ano74m, Ano75k, Ano75j, Ano75l, Ano75m, Ano76i, Ano76h, Ano76j, Ano76k, Ano77j, Ano77i, Ano77k, Ano77l, Ano78j, Ano78i, Ano78k, Ano78l, Ano79h, Ano79g, Ano79i, Ano79j, Smi74b]. Nurnberg [Hal72c]. Nutrition [Koh74]. Nuttall [And74].

O [Arn70, Chu78a, Fer73b, Hal73, Hau76, Rol74, Sky73, Smi76c]. Oak [All79c]. Oakham [Hil70a]. Oberth [Hac73b]. obrazów [Mos75b]. O’Brien [Smi70c]. Observations [Mul74e]. Observatory [War76]. Observed [SG75]. Obsession [Har78b]. obshchestvo [Lie75d]. Obsolete [Ihd77, Sti77]. Occupation [Tuc73]. Oceans [Fla75]. Ocherk [Lie75c]. October [AGS^{+}79, Bri74, BPV^{+}78, Emm71, FS78, HS70, HRM^{+}76, Mer76a, Ord70, SB78, Slo70, Swe71, Ung72, Ano70j, Ano71j, Ano72k, Ano73v, Ano74m, Ano75m, Ano76k, Ano77l, Ano78l, Ano79j]. Odds [Had79]. Oettinger [Bar70a]. Oeynhausen [Car72b]. Office [Pos76d, Cos71]. Officina [Tyl77]. Offshore [Kap77]. Ogden [Pur70]. Ogilvie [Hig77]. Ohio [Sca72, Whi70, Wil71a]. Ohkawa [Kob74]. Oil [Buc75a, Fla75, Gie78, OO73, Sal73, Siv77, Wil77]. Oils [Kel73]. Oklahoma [Bra77b]. Old [GD73, Rou71]. Oldenbourg [Hac73b]. Oliver [Sum75]. Olof [Cha73b]. Olson [Sha72c]. Olufemi [De 74a]. Omnibuses [Pos76a]. One [Flo72, Lei70]. O’Neill [Tuc73, Fow77, Gre79]. Onkar [San78]. Ontario [Bal75b]. Open [Bra76c]. Opening [Kle77]. Operation [Bal78, Gli70b]. Operations [Hac74, Hil75b, Kel71b, Fla75]. Opinions [Dor76, Fru75]. Oppenheimer [San77]. Opper [Eis75a]. Opportunities [Cah76]. Opportunity [Fer71b, Bra79d]. Optics [Jen70]. Optimal [Sew75]. Opus [Gli70b, Hal72d]. opyt [Lie75a]. Oranges [Pat75]. Orbital [Hac74]. Orbiter [Hac79]. Orchestra [Wat77]. Ord [Har78b]. Ord-Hume [Har78b]. Order [Hug77, De 79b, Whe73]. Orders [Kop77]. Ordnance [All79b, Sha72b]. Ordway [Sha78c]. Ore [Laf72]. Organisation [All79a, Gli70b]. Organization [All74, Dav75, De 78, Gil73c, Wol78, Cas74]. Organizational [Tay75]. Organizations [Chu78b, MGM78, Pei72]. Organized [Aga70, Sau72]. Orientations [Col71]. Origin [Con78b, Row73, Wil78]. Original [Rol77b]. Originality [Kap74a]. Origine [Row73]. Origins [Arn70, Bed71, Bla73a, Cut74, Jam76, Jen75, Lay79, Lev75, McL79, Nor77a, Sha72b, Sha78b, Smi77a, Sto73, Whi73e, Bil79]. Orthographia [Kel70c]. Osborn [Goi71]. O’Shaughnessy [Gor71a]. osnovnye [Lie73a]. osterreichischen [Bry71b]. Ostrander [Gor71b]. Ostrowski [Sko75]. Oswald [Con74g]. Oswaldo [Mat78]. Otechestvennoi [Lie74b]. Otechestvennoivoiny [Lie78a]. Other [Bal73a, Con72b, Sta78d, Wel71]. Otis [Seg77c]. Otto [Hal72d, Har70, Lay77b, Woo72a, May71b]. Ottoman [Gra72]. Our [Fer71c, Koh70c, Lew70, Mul73a, Bra77b, Con72c, Lew74]. Outer [Fla75]. Outline [Con74b, Ger76]. Overpopulation [Wil78]. Overshot [Rey79b]. Overskill [Vac72c]. Own [Lew74]. Oxford [Hok79b]. Ozawa [Uch75].

P [All79c, Art70, Cla78, Con79e, Cra78, DeG77, Dra78, Fri79b, Fus73, Fus74a, Gie76a, Hac71, Hey74, Hil77, Jac73, Lie73b, Lie75c, Lie79a, May78a, McM76b, Moe71, Mor70, Nes77, Pos78, Rey74, Ros71a, Sch77a, Siv72, Spi78, Vin79, Web74, Woo79a]. Pacey [Bur77]. Pacific [Pur76a, Ste76c]. Packaged [Ada74]. Packard [Bla74b]. Paddle [Pos75a]. Page [Pos72, McM78]. Pain [Van78b]. Paine [Fer73b]. Painter [Jen73c]. Painting [Vac71]. Pakistan [Gha70b]. Palace [Con74h]. Palaeoeconomy [Row76a]. Palazzo [Wat78b]. Palm [Gie79]. Pamphlets [De 74d, Wil70b]. Pantano [Gli75a]. Pantheon [Con77e]. Paolo [Smi71a, Whi71a]. Paper [For71, Rob76b]. Paperclip [Hac73a]. Papermaking [Smi77b]. Papers [Bri74, Bri77b, De 74e, Emm72, Jac79, Sin70b, Swe75, Tar79, Wal74]. para [Gli70b, Gli75a]. Paracelsian [Mul79b]. Paradigms [Lev79]. Parallel [Bry73b]. Parallels [Ger76]. Paris [Bri72]. Park [Bru73, Con71d, Fri78b, San73b]. Parke [Dan72]. Parks [Dob75]. Parnell [Hig79a]. Parsegian [Bee73]. Parsons [Bic73, Kel71c]. Part [Ano73s, Ano73t, Jen73a, tZS73b]. Partially [Ros79d]. Participation [Gie76c, Zin79]. Partners [Law71]. Partridge [Bee75]. Parts [How78, Cam74]. Pascal [Tro75]. Passenger [Con79f, Con70a]. Passmore [Dur79, Str75]. Past [Pic77, Whi73f, MM74]. Pastor [Kel73, Gli75a]. Paszek [Goo74a]. Patent [Cha71b, Jac70, Pos76d, Smi74b]. Patents [McM78, Rob72, She75b, Tar74]. Pater [Whi72c]. Paternalism [Sau72]. Patricia [Car79b, Hit76, Hud75b]. Patrick [Wei75, Oka78]. Patriotic [Lie78a]. Patrons [Rei71]. Patt [Rap77]. Pattern [MM74, Wel70]. Patterns [Ino76, Lay75, Meu77]. Patton [Rey79a]. Patty [Row72a]. Pauer [Fuk75b]. Paul [Dug79, Fie77, Gor73, Had72, Lie75b, Mar77b, Mok77, Mor79, Pos76c, Rog78, Spr76, Smi75a]. Payen [Rob71]. Pays [Ell77]. Payson [Ano70a, Ano71a, Ano72a, Ano73a, Ano74a, Ano75a, Ano76a, Ano77a, Ano78a, Ano79a]. Pazyryk [Mei71b]. Peace [Bra79a]. Peacefully [Hou77]. Pearse [Hig77]. Peasant [De 78, DeW78, tZS73a]. Peasants [Kui78]. Peccei [Gol70a]. Peckham [Whi73c]. Pedretti [Con74h]. Pedro [Gli75a]. Peil [De 73b]. Pelto [Lie74d]. Pelz [Chu78b]. Pendray [Emm72]. Penguins [Sew75]. Penicillin [Par71a]. Peninsula [Per78b]. Pennsy [Pos76b]. Pennsylvania [Car79b, Lew76]. Pentagon [Mul72b]. People [Lev75, Rae75, Vol78, Wal73b, Com74, Mat74, Sok75, Con73e, Sew75]. Perceptions [Ehr74]. Percussion [Cha72a]. Perel’man [Swe72]. Perfect [Whi77a]. Performance [War70]. Perier [Rob71]. Period [McM77, Lie74b, Ihd77, Nes77]. Periodical [Car77b, De 74d]. Peripheral [JR78]. Perkins [tZS76]. Pernambuco [Mat75]. Perpetual [Har78b]. Perry [Fus74a]. Persians [dC79]. Persistent [Cha71c]. Personal [Lew74, Pos79, Whi75b, Pos79]. Personalities [Jen73a, Sca72]. Personality [Lay71c, Rae71a]. Personnel [Par79]. Perspective [Dur72, Mos79, Cos71, tZS76]. Perspectives [Rie70, Ros70, Wil72, Blo77, DeG77, Lay77a, Tha73a]. Pertti [Lie74d]. Peru [PS79]. pervoi [Lie76a]. Peter [Bay78a, Con71c, Cra79b, De 74d, Fei78, Fei77, Fer76a, Lei73, Mat75, McK79, Ste75a, Vac77b, Pen75]. Peterson [Bal73b, Con77b]. Petroleum [Joh73]. Petronio [Wat77]. Petsche [Hun75]. Petulla [Pen77]. Pevsner [Con77a]. Pfeiffer [Wer78b]. Pferde [Bry71b]. Pharmacy [Ger74, Lei70, Lei70]. Phase [Hug76a]. Philadelphia [Bal73b, Hai79, Rop76d, PPH^{+}77, Tha76]. Philanthropy [Gai75]. Philip [Boz77, Bri76a, De 75a, De 76, Fri76, Hac72, Kra74, Lam71a, Ros75d, Van78a]. Philippines [DeG75]. Philipsburg [Pen79]. Phillips [Sim71]. Philo [Dra76b]. Philosopher [Dur72, Tha76, Hei77b]. Philosophers [Lay77b]. Philosophical [Ano73g, Lan73, Tha72, Sch77b]. Philosophies [Tho78]. Philosophy [Byr78b, Car78c, Fei78, McC73, Mor72b, Mor79, Mul79b, Rou72, Sin77, Bal73a, Smi71a]. Photograph [Jen73c]. Photographer [War71]. Photographic [Cha77, Con72d, Gie78]. Photographica [Jen77b]. Photographs [Pos76a]. Photography [Jen75, Jen77b, Vac78]. Photojournalist [Jen73b]. Physical [Gla74, Sha76b, tZS73b, Tur76, Whe78]. Physicist [Car75a, Hei77b, Kle71, Nor75]. Physics [Kap74a, McK73, Nut79c, tZS73b, Wil71b, Dev76, McM78]. Piaskowski [Smi76c]. Pickett [Byr78a]. Pickup [Sta78d]. Picon [HS78]. Pictorial [Con79a, Pau71]. Pilain [De 74f]. Pilkington [Had78c]. Pilot [Gli70b, Hud76]. Piñero [Gli70a]. Ping [Smi77a]. Ping-Ti [Smi77a]. Pintsch [Whi77a]. Pioneer [Bez79, Con74a, Con77f, Ord78, Tuc78, War71, Rae73a, Wal71a]. Pioneers [Koh70c, Con79c, Lie73b]. Piper [Sha75]. Pipping [Nut79c]. Pirsig [Sta78d]. Pisa [Kel70a]. Pistoia [Mue74]. Pistols [Cha72a]. Pithole [Joh73]. Pittsburgh [AGS^{+}79]. Place [War76, Lie75b, Dre76]. Placemaking [Con76c]. Plains [Fra74a]. Plan [Bra77b, Con75f]. Planemakers [Wel71]. Planes [Wel77, Hol76]. Planned [Bud78, Seg77c]. Planning [Bay74a, Bay74b, Blo77, Bur74a, Con71e, Rae73b, Con75c, Fel78]. Plans [Gre72]. Plata [Ord70]. Platt [Cra76a, Fit71]. Pleasures [Wat76a]. Pleiger [Lie75b]. Pleigers [Lie75b]. Plenty [Gol77b]. Plictho [Adr70]. Ploger [Pri75]. Plommer [Pac74]. Plough [Fen78]. Ploughing [Fen78]. Plowboys [Swi78]. Plowden [Con75a]. Plowed [Koh70c]. Plowshares [Bra78]. Plum [Hal76, Tho75b]. Plummer [dLM73]. Plus [Jer78b]. Pneumatica [Dra76b]. Pneumatics [Kre74]. Pocket [Cha72a]. Poem [Tyl77]. Point [Dev76, Byr78a]. Points [Fer73c, Nor78]. Polak [Tof72]. Poland [Ros74]. Pole [Bri71b]. Policies [Jac71, Gol74b]. Policy [Blo77, Bra77b, DeG75, Dun76, Gol70b, IS77, Lei73, Mat78, Mor78, Ols76, Pal79, Ros75a, Wei78b, Woo79a, Cha71a, Dav75, Dra78, Gar78, Kap72]. Polish [Byr74, CZ73, Tyl77, OO73, Orł73]. Polites [Fri79a]. política [Gli70b]. Political [Ano73h, Ano73i, Ano73j, Car78a, Cow75a, Dal72, De 73b, Kni77, McM75, McM77, Ste76a, Cli71]. Politics [Bai74, Bal75b, Boz77, Cro71, Dob75, Eze77, Gre73, Kob77, Mar79a, McM78, Mil76, Nie70b, Sko76, Str73a, Whe73, Bou78, Boz75, Wat78b]. Polity [Kel71d]. Pollution [Bra75, Bur79, Lie74e, Bry76a, Sew75]. polozhenie [Lie76a]. polskiej [Mos78]. Polyphony [Lev79]. Pont [Bal71]. Pont-à-Mousson [Bal71]. Ponting [Jer79, Lea76, Lea79, Mal78]. Pool [Mol78]. Poole [Gie76c]. Poor [Con74c]. Pope [Cas76b]. Poppino [Mat74]. Popular [Tav72b]. Populism [Sal70]. por [Gli75a]. Porter [May75b]. Portfolio [Gar72b, Mar76b, Wel73b]. Portland [Con75d]. Portrait [Bay75]. Position [Kra75, Per74, Lie76a]. Possessed [Kah70]. Post [Bri79, Bru76, Hor79, Mul74b, Sta78d, Hor79, McM78]. Post-industrial [Hor79, Mul74b, Hor79]. Post-technological [Sta78d]. Post-War [Bru76]. Postwar [All79b, Kev75]. Potential [Pos79, Lie78b]. Pots [Gli77]. Potter [Hin72]. Pottery [Spr76, Wal71a]. pour [KF76]. Power [Aga76, Bal75b, Com74, Del76, Emm74, Emm79, Fri78a, Hil72, HP72, Hug76b, Mul74a, Pur75, Sta77b, Bra77e, Hal73, Mul72b, Wil77, Com74]. powered [Bla79b]. Practical [Bee75, Gil70, Hin71b, Koh71b, Meu71, Svi77]. Practice [De 78, Koh74, May75b, Pei72, vR78, Bun74, Jac71, Row77]. Practices [Per78a]. Prager [Con71a, Dra76b, Rol77a]. Prairie [Con72e]. Prairies [Bal74]. Prakash [Ino76]. Pratt [War70]. praxeologie [Sko75]. Pre [De 76, Pau71, Str78]. Pre-Colonial [De 76]. Pre-Columbian [Pau71].

Pre-Modern [Str78]. Precede [Owe73]. Precious [Jen76]. Precision [Smi74a]. Preclassical [Wal73b]. precurseur [Sko75]. Prefabrication [Con79c]. Preface [Dra76c]. Prehistoric [For79a]. Prehistory [Wer78b, Wil78, Mei71a, Wal74, Wal76a]. preimushchestva [Lie77a]. Preindustrial [Bra75, Mei72]. Preliminary [Cha70, Swe75]. Prescott [Car76]. Present [Fus71, Pic77, Whi73f, Pau71]. Preservation [Dob75, Hud71]. Press [Mar79b, Wal76b]. Presses [Sch74]. Price [Com74, For71, Pur70, Dra76a, Wil76]. Priced [Fus71]. Priestley [Hol70b]. primacy [Lie77a]. Primary [Ano73h, Ano73k, Bal75a]. Primer [Sch71b]. Primitive [Mei72, Siv72, Spi75b]. primo [Kel70a]. Princeton [Bri77b]. Principia [Hal72a]. Principle [Ret70]. Principles [How74]. Printed [Ada74, Bor71a]. Printing [Alt73b, Bra79c, Mar79b, Nes77, Qui77, Sch74]. Prints [Alt73b, Fer72]. Priorities [Cah76, Nic71, Com74]. Priority [Jac70]. Pritchard [Nut79a]. Private [How78]. Prize [Ano70b, Ano71d, Ano72e, Ano73f, Ano74f, Ano75f, Ano76e, Ano77f, Ano78e, Ano79c, Ano70a, Ano71a, Ano72a, Ano73a, Ano74a, Ano75a, Ano76a, Ano77a, Ano78a, Ano79a]. Próba [Cas76a]. Probable [Per74]. Problem [Lie76b, May76, Tho75a, Web76, Whi75a]. Problème [Ste76e]. problemi [Kel70a]. Problems [Col71, Gra75, Min79a, Par70, Pos79, Rol77c, Rür74, Sel78, Str75, Wes78, Lie79a]. problemy [Lie79a, Par70]. Procedures [Fel76b]. Proceedings [Bel78, Fei78, Hac78, Jac70, Ras71, Dra78]. Process [Mon72, Cli71, De 70]. Processing [Wil70a, Gor74]. Prodipto [Bos71]. Producer [Wil78]. Producers [Spr76]. Production [Buc75a, Chu79, Mul71, Mul76, Par71a]. Productive [Chu78b, Wei75]. Productivity [Koh78, Spa70]. Prof. [Hal72d]. Profession [Cra79c, Hab72]. Professional [Sch77a]. Professions [Cal71a, Sky70, Kra74]. Profile [Bon79]. Profit [Lie73b]. Prognostics [Tof72]. Program [Bra77b, Mul71, Sha72b, Hil78, Hac79, Str73a]. Programs [Fel76b, Eze77]. progreso [Gli74]. Progress [Buc70a, Hor79, Lie74e, Pos79, San73c, Lie73a, Lie77b, Pur70, Sel78]. Progressive [McM77]. Project [Bry73b, Ols76, Sha79, Smi76e, Ste70, Swe78a, Van71, Smi76e, Hud76, Hac73a]. Projects [Pal79]. Proliferation [San78, Nie70a]. Prometheus [Kah70]. Promise [Jam79b, McN78]. Promising [Pos79]. Promoter [Fra73]. Promoting [De 74c]. Promotion [Tho75b]. Propaedeumata [Han79b]. Propellants [Sha73]. Propeller [Vin79]. Prophecy [Hor79]. Prophets [Mul74b]. Prospect [Bra76a, Har78a]. Protection [CZ73, Pat75]. Protest [Sau72]. Provinces [Moe71]. Przekroju [Mos79]. PSA [Fei78]. Psychology [Dav70]. Pt [Sha78a, Maj75, Maj77]. Public [Bra77b, Cai74, Del77, Dra78, Gol70b, Hin71b, Kap72, Lei71, Mor78, Pos79, She74, Sta77a, Zin79]. Publications [Car72e]. Published [Buc70a]. Pueblo [Con72d]. Pugsley [Emm77]. Pulp [Bur79]. Pupin [Bri70]. Purbank [Kel71d]. Puritan [Jud74b]. Purpose [Lew74]. Pursell [Lay70a, Lay71a]. Pursuit [San73c]. Pyenson [Whe78]. Pyne [Whi70].

Quality [Fer74b, Gie76b]. Quantitative [Coo77, Lol78]. Quarrymen [Mul79a]. Quarter [Con72a]. Quarterly [Fli72]. Queen [Gol77a]. Queries [De 75c]. Quest [Wei79]. Question [Fus76b, GD76]. Questionable [Fru75, Dor76]. Questions [Bra70a, Dan70a]. Quicksilver [Mul74d]. Quigley [Fei77]. Quimby [Aus74, Gar72b, Kul79, Mar76b, Wel73b].

R [Aga78, And74, Bar71b, Bil75a, Bla74b, Bri78a, Buc79, Cha73a, Con74f, Dal72, De 74a, Dor70, Ehr74, Eys74, Fer70a, Fer76c, Fit73b, Fuk75a, Fus73, Gol78, Gol70b, Hac73b, Hac78, Han79a, Har78a, Hil78, Hok79b, Hud79, Hun71, Jac71, Kap74b, Kap78, Kel70b, Le 75, Mar72b, McC73, Mul73a, Mul74c, Nut78, Par78, Per70b, PS71a, dSP75, Pri75, Pur70, Qui71, Rae76, Rob75a, Rol70b, San73a, San72, Sch76, SG75, She77, Sko76, Smi74a, Spa70, Sta77b, Ste76c, Sto71, Swe72, Swe78a, Syl72, Whi77b, Woo72b]. R. [Bud78]. Rabinowitch [Fel77, Fel77]. Race [Smi76d, Eys74]. Radical [Lew70]. Radio [Chu78a, Wik77, Nor77a]. Radioactivity [Hei70]. Rae [Hol70a, Wik70, Wik73]. Rafael [Gli75a]. Rafique [Con74b]. Rahman [Bla78c, Wei78b]. Rail [Con71f]. Railroad [Bay78b, Bez79, Bry73b, Car71a, Con79f, Sha72c, She78b, Whi73d, Whi76b, Lie78a, Car79b, Con71c, Gat73, Gol77a, She78b]. Railroads [Bal73a, Con75g, Sto73, Rop76c]. Rails [Pos76b]. Railway [Con73e, Hil75b, Ste70, Whi73a, Whi73f, Con71c, Car72c, Car72d, Rol72]. Railways [Mas77b, Mer76b, Syl72, Con74f, Car72b]. Rain [Fra74a]. Raising [Cai72]. Raistrick [Lew73]. Rakowski [Hil77]. Ralf [Rah79]. Ralph [Low76, Wei79]. Ramelli [Gnu74, Hal70, Ret72]. Ramos [PS71a]. Randell [Cra79a]. Ranger [Swe78a]. Raphael [Del76, Min77b, Mul79a, Tri72]. Rapid [Con72a, Con74e, Pos79, Whi77b]. Rapp [Fit73b, Wer79]. Rasmussen [Koh77]. Ratcliffe [Hau76]. Rate [Lie79b]. Rationality [De 75c, Hop75]. Raum [May71b]. Raumschiffahrt [Hac73b]. Ravenswaay [Qui78]. Ravindra [Fer74b]. Raw [Maq70, SA77]. Rawski [tZS73a]. Raybould [Gal75]. Raymond [Bla78a, Coo71, Gli79, Lay71b, Mar72c, McL78, Nor75]. razvitogo [Lie77b, Lie79a]. Razzaz [dSP75]. Reaching [Bra77b, De 78]. Reactor [Mil76]. Read [Con74i]. Reader [Bat70b]. Reading [Ras71]. Readings [Had75b, Had76b, Had77b, Had78b, Lay70a]. Real [Con72a]. Reale [Geb71]. Realities [Con74i]. Reality [Lew71]. rear [Lie78a]. Rearview [Wil73a]. Reason [Aga76, McL70, Sch76]. Rebuilding [Sta78d]. Reconstructing [Fit73b]. Reconstruction [Sle78]. Record [Con72a, Siv72, Woo77a]. recording [Mos75b, Gar72a]. Records [Hit76, Wil70a]. Recycle [Tar79]. Red [Sha78c]. Rees [Rol70b]. Reese [Cha77]. Reexamined [How78]. Reference [Cat76]. References [Ros79d]. Reflections [Rür74, Whi75b]. Reform [Mar79a, Mer76b, McM77]. Refuse [Fei77]. Regelungen [Har70]. Regency [Gol71]. Region [Cra77, Hai79, Rol70b, Hol79]. Regional [Con75f, Fit73b]. Regulated [Ros73a]. Regulation [Cok72, Mar77a]. Reich [Lie75b, Lie76c, Whi71b, Kra78]. Reid [Gol70b]. Reign [Buc70a, Ger79].

Reingold [Bri74, Bri77b, Lay77c]. Reinhard [May77]. Reiser [Ger79]. Rejoice [Gli79]. Rekonstrukcji [Cas76a]. Related [Cra79a, Ros75d, Ros79d]. Relating [Wil70a]. Relation [Ros75a]. Relations [Car78d, Vac71]. Relationship [May76]. Relationships [Eis75a]. Relevance [Ros73b]. Religious [Ano73k, Ano73l]. Reluctant [Bak72]. Remains [Smi77a]. Remington [Cal74]. Reminiscences [Rou71]. Renaissance [Kre76, Mok77, Wat78a, Kel70c, Kel78, Kel71c, Kel71a]. Rendezvous [Hac74]. Renfrew [Wal73b, Wal76a]. Reply [Dan70b, Flo70, Sta78c]. Report [Cah76, Har74a, Rog75, Sin72, Wil71b, Wei78b]. Reported [Pos79]. Reports [Gli70b, Sin72]. Repositories [Wil71b]. Represented [Wil70b]. Reprint [Koh71b]. Republic [DeG75, Vol78]. Republican [Seg77a]. Research [Bay78a, Bra77d, Bra77b, Cah76, Chu78b, Gar78, Gil79, Gol70b, Gre73, Hac71, Kev75, Lay79, Mat78, Nic71, Per78a, Rei71, Spa70, Dug79, Fie77, Koh78, Mor79, Dug79, Fie77, Man73, Mor79]. Reshaping [Per78a]. Residence [Van78c]. Resource [Dob75, Gol77c, Koh78]. Resources [Bra79b, Ste75b, Win70, Low76, Pen77]. Response [Fri75, Ino76, Ray77]. Responsibility [Con70d, Hab72, Kra77a, Str75, Fer74b]. Reston [McG76]. Restoration [GD72]. Restorations [Jud74a]. Results [Wal73b]. Reti [Dib74, Ste76f]. Retrospect [Har78a]. Retrospective [Nes77, Tar77, Cai78]. Returned [Sha78a]. Review [Fus71, Gar78, HV78, Spi74, Abb72b, Abb72a, Abr79, Ada74, Adr70, Aga70, Aga76, Aga77, Aga78, Ait79, All74, All79a, All79b, All79c, Alt70, Alt73a, Alt73b, And74, Arn70, Art70, Ash71, Aus74, Aus76, Aus77, Bag79, Bai74, Bak72, Bak76, Bal73a, Bal73b, Bal74, Bal75b, Bal75a, Bal77, Bar73a, Bar70a, Bar70b, Bar71a, Bar71b, Bar73b, Bat79, Bat70a, Bat70b, Bat71, Bat77, Bay74a, Bay74b, Bay75, Bay76b, Bay76a, Bay78a, Bay78b, Bec78, Bed76, Bed71, Bee70, Bee73, Bee75, Bee76, Bee77, Bee79, Bel78, Ben78, Ber77, Ber70b, Ber70a, Ber74, Bez79, Bic73, Bil73, Bil75b, Bil75a, Bil77, Bil79, Bir76, Bis70, Bis71, Bis76a, Bis76b, Bla71, Bla72, Bla75, Bla76a, Bla78a, Bla79a, Bla73a, Bla77, Bla73b, Bla74a, Bla74b, Bla79b, Bla74c, Bla76b, Bla78c, Blo77, Bor71a]. review [Bor71b, Bos71, Bou78, Boz75, Boz77, Bra76a, Bra77a, Bra77b, Bra77c, Bra79b, Bra79a, Bra79c, Bra70b, Bra72, Bra79d, Bri78a, Bri71a, Bri72, Bri73, Bri76a, Bri76b, Bri79, Bri71b, Bri74, Bri77a, Bri77b, Bri78b, Bru73, Bru76, Bry70, Bry71a, Bry71b, Bry73b, BR74, Bry76a, Bry77c, Bry77a, Bry77b, Buc70a, Buc70b, Buc71a, Buc71b, Buc73, Buc74, Buc75a, Buc76a, Buc78, Buc79, Bud78, Bul76, Bun74, Bur77, Bur74b, Byr78a, Byr74, Cah76, Cai78, Cal70a, Cal71a, Cal74, Cal70b, Cal71b, Cam74, Can74, Car72a, Car73, Car75a, Car77a, Car71a, Car72c, Car72d, Car72b, Car75b, Car79b, Car71b, Car75c, Car78a, Car78b, Car78d, Car79c, Cas73, Cas74, Cas76a, Cas76b, Cat76, Caw77, Ces70, Cha72a, Cha73a, Cha78a, Cha77, Cha78b, Cha73b, Cha70, Cha71a, Cha72b]. review [Chu78a, Chu78b, Chu79, Cla78, Cla77, Cli71, Clo70a, Clo70b, Clo75, Clo78, Coh74, Col71, Com74, Con70b, Con70a, Con71a, Con71b, Con71c, Con71e, Con71f, Con71d, Con72a, Con72c, Con72d, Con72e, Con72b, Con73c, Con73f, Con73d, Con73a, Con73b, Con73e, CZ73, Con74c, Con74a, Con74d, Con74e, Con74f, Con74g, Con74h, Con74b, Con74i, Con75a, Con75b, Con75c, Con75d, Con75e, Con75f, Con75g, Con76a, Con76b, Con76c, Con76e, Con76d, Con77b, Con77c, Con77d, Con77a, Con77f, Con77e, Con78d, Con78a, Con78b, Con78e, Con78c, Con79f, Con79a, Con79c, Con79d, Con79e, Con79b, Coo77, Coo71, Coo74, Cos71, Cot70, Cot72, Cow75a, Cow75b, Cra75, Cra76a, Cra76b, Cra77, Cra78, Cra79c, Cra79a, Cra79b, Cro71, Cut74, Dal70, Dal72, Dan72, Dat71, Dau72, Dav75, Day74, De 70, De 71, DeG71, De 73a, De 73b]. review [De 74c, De 74d, De 74a, De 74b, DeG75, De 75a, De 75b, De 76, DeG76, DeG77, De 78, De 79b, De 79a, De 74f, DeV75b, Del76, Dev75a, Dev76, Dev78, Dew76, Dib75, Dib79, Die73, Dob78, Dob75, Don77, Dor70, Dot75, Dra70, Dra72, Dra76a, Dra76b, Dra76c, Dra78, Dre75, Dre76, Dru74, Dru76, Dub77, Dug79, Dun76, Dur79, Eat73, Eat75, Edw72, Ehr73, Ehr74, Eis75a, Eis76, Ell77, Ell76, Eme78, Emm72, Emm74, Emm76, Emm78, Emm79, Emm77, Esp72, Eys74, Eze77, Eze78, Fai71, Fei78, Fei77, Fel78, Fel77, Fel76a, Fel76b, Fen78, Fer70a, Fer71a, Fer73a, Fer73b, Fer73c, Fer76a, Fer79b, Fer70b, Fer71c, Fer76c, Fer78b, Fer72, Fer74b, Fie77, Fit73a, Fit71, Fit73b, Fla75, Fli72, Flo72, For79a, Fow77, Fow78, Fow79, Fra71, Fra73, Fra74a]. review [Fra74b, Fra76, Fre71a, Fre76, Fre71b, Fre73, Fre74, Fri79a, Fri76, Fri77, Fri78a, Fri78b, Fri79b, Fuk77, Fuk75a, Fuk75b, Fus72, Fus73, Fus74a, Fus74b, Fus76a, Fus76b, Gai75, Gal75, GD73, Gar72a, Gar72b, Gat73, Geb71, Geb72, Ger76, Ger74, Ger79, Gha70a, Gha70b, Gha73a, Gha73b, Gia72, Gie78, Gie76a, Gie76b, Gie76c, Gie77, Gie79, Gil70, Gil72a, Gil72b, Gil73a, Gil73b, Gil73c, Gla78, Gla74, Gli70a, Gli70b, Gli73a, Gli73b, Gli74, Gli75a, Gli75b, Gli78, Gli79, Goi70, Goi71, Goi73, Gol75, Gol74a, Gol76, Gol77a, Gol77b, Gol78, Gol70a, Gol70b, Gol74b, Gol77c, Goo70b, Goo74a, Gor74, Gor79, Gor71b, Gor72, Gor73, Gra72, Gra75, Gre72, Gre79, Gre73, Gro72, Gue73, Gue77, Hab72, Hac71, Hac72, Hac73a, Hac73b, Hac78, Hac79, Had70, Had71]. review [Had72, Had73, Had74a, Had74b, Had75a, Had75b, Had76c, Had76a, Had76b, Had77a, Had77b, Had77c, Had78c, Had78a, Had78b, Had78d, Had79, Hai79, Hal71, Hal72b, Hal72c, Hal72a, Hal72d, Hal76, Hal73, Ham78, HS78, Ham71, Han79a, Han79b, Har71a, Har74a, Har75a, Har75b, Har76a, Har78a, Har79a, Har70, Har71b, Har71c, Har73a, Har73b, Har74b, Har76b, Har77, Har78b, Hau76, Hay79, Hea75, Hea71, Hea72, Hec70, Hec73, Hei70, Hei77a, Hei77b, Hel71, Hel77, Hen76a, Hen76b, Her75, Hey74, Hig73, Hig77, Hig79a, Hig79b, Hil74, Hil78, Hil70a, Hil71, Hil75b, Hil75a, Hil76, Hil70b, Hil70c, Hil77, Hin72, Hin71b, Hin71a, Hin79, Hit76, Hod79, Hog75, Hok77, Hok79a, Hok79b, Hol70a, Hol76, Hol79, Hol70b, Hor79, Hou77, Hou78, How74, How72, Hud71, Hud75b, Hud75a, Hud76, Hud79]. review [Hug70, Hun71, Hun75, Hyd73, Ihd77, Ino76, Isa76, Jac71, Jac73, Jac75b, Jac75a, Jac77, Jac79, Jam76, Jam79a, Jam79b, Jen70, Jen72, Jen73a, Jen73b, Jen73c, Jen76, Jen77a, Jer70a, Jer70b, Jer74, Jer76a, Jer76b, Jer77, Jer78a, Jer79, Joh73, Joh78, Jud74b, Jun70, Kap71, Kap72, Kap74a, Kap74b, Kap74c, Kap77, Kap78, Kar76, KF76, Kel70a, Kel70b, Kel71c, Kel71a, Kel76, Kel77, Kel71d, Kle71, Kle77, Kle76, Kni77, Kob74, Kob77, Koh70a, Koh70b, Koh70c, Koh71a, Koh71b, Koh72, Koh74, Koh76b, Koh77, Koh78, Kop77, Kra74, Kra75, Kra76, Kra77b, Kre74, Kre76, Kui78, Kul79, Kur73, Laf72, Lak70, Lam71b, Lam71a, Lan73, Law71, Lay70a, Lay71b, Lay71a, Lay71c, Lay75, Lay76a, Lay77a, Lay77b, Lay77c, Le 71, Le 72, Le 75, Lea70, Lea76]. review [Lea79, Led77, Lei70, Lei71, Lei73, Lev75, Lev79, Lew70, Lew71, Lew73, Lew74, Lew76, Lid76, Lie73a, Lie73b, Lie73c, Lie73d, Lie74a, Lie74b, Lie74c, Lie74e, Lie74d, Lie74f, Lie75a, Lie75b, Lie75c, Lie75d, Lie76a, Lie76b, Lie76c, Lie77c, Lie77b, Lie77a, Lie77d, Lie78a, Lie78b, Lie78c, Lie79a, Lol78, Lon78, Low76, Lub78, Mag76, MM74, Maj74, Maj75, Maj77, Mal78, dLM73, dLM76, Man72, Man73, Maq70, Mar72a, Mar72b, Mar72c, Mar77a, Mar76a, Mar76b, Mar77b, Mar79a, Mar77c, Mar79b, Mas73, Mas77a, Mat76, Mat74, Mat75, Mat78, May78a, May70, May71b, May75a, May77, May78b, McC73, McC70b, McC70a, McK73, McK79, McL70, McL78, McL79, McM76a, McM78, McM75, McM76b, McM77, McN78, Mei70, Mei71a, Mei71b, Mei72]. review [Mer76a, Mer76b, Mic70b, Mic70a, Mid78, Mil76, Min78, Min79b, Min77a, Min77b, Moe71, Mok77, Mol78, Mon71, Mon72, Mon75, Mor72a, Mor71, Mor77a, Mor70, Mor72b, Mor74, Mor77c, Mor79, Mos75a, Mos75b, Mos78, Mos79, Mul73a, Mul73b, Mul74b, Mul74a, Mul72a, Mul72b, Mul74c, Mul74d, Mul78, Mul79b, Mul79a, Mur79, Nee79, Nee70, Nel79, Nes77, Nev79, Nic71, Nie70a, Nie70b, Nob71, Noe77, Nor75, Nor77a, Nor78, Nor77b, Nut78, Nut79b, Nut79c, Oka78, Ols74, Ols76, Ols77, Ord78, Osp79, Ove71, Pac74, Pac76, Pac78, Pal79, Par71a, Par74, Par72a, Par78, Par79, Par70, Par71b, Pau71, Pei72, Pen77, Pen75, Pen79, Per73, Per70a, Per72, Per70b, Per74, Per78b, Pet75, PS71a, PS71b, PS79, Pia70, Pia71, Por74, Por76, Pos75a, Pos75b, Pos76a]. review [Pos76b, Pos76c, Pos78, Pos79, dSP75, Pri75, Pur70, Pur71, Pur74, Pur78, Qui74, Qui78, Qui71, Qui77, Rae71a, Rae71b, Rae73a, Rae73b, Rae73c, Rae74, Rae76, Rae79a, Rae79b, Rah79, Ram70, Ran72, Ran76, Rap75, Rap77, Ras71, Ras74, Ras75, Ras76, Rav70, Rav72, Ray77, Rea78, Rei71, Rey74, Rey75, Rey79a, Rie70, Ris78, Rob70, Rob71, Rob75a, Rob75b, Rob76c, Rob76a, Rob76b, Rob76d, Rog75, Rog77, Rog78, Rol77a, Rol78, Rol70a, Rol70b, Rol72, Rol73, Rol74, Rop72, Rop76d, Rop76a, Rop76b, Rop76c, Ros79a, Ros73a, Ros75a, Ros74, Ros71a, Ros76a, Ros73b, Ros75c, Ros75b, Ros75d, Ros79c, Ros72, Ros76b, Ros79d, Rot72, Rou71, Rou72, Rou75, Row77, Row72b, Row72a, Row73, Row76a, Row76b, Row79, Sal70, Sal71, Sal73, Sal77, San73a, San72, San74, San77, San78, San73c]. review [Sau72, Sca72, Sch78a, Sch77a, Sch78b, Sch78c, Sch70, Sch74, Sch71a, Sch71b, Sch76, Sch77b, SG75, See78, Seg77a, Seg77b, Seg77c, Seg78, Seg79b, Seg79a, Sel77, Sel78, Sew72, Sew75, Sha76a, Sha72a, Sha74, Sha76b, Sha72b, Sha73, Sha75, Sha76c, Sha78a, Sha78b, Sha78c, Sha79, Sha72c, She78a, She74, She75b, She77, She78b, Shr73, Sim71, Sin70a, Sin70b, Sin72, Sin79b, Siv72, Siv75, Siv77, Sko75, Sko76, Sky70, Sky73, Sky75, Slo72, Sma78, Smi70c, Smi71b, Smi71a, Smi72c, Smi72b, Smi73b, Smi74a, Smi75a, Smi75b, Smi76d, Smi76b, Smi76a, Smi76c, Smi77d, Smi77a, Smi77c, Smi77b, Sok75, Sol76b, Sol76a, Son78, Spa70, SR73, Spi75a, Spi78, Spo73, Spr76, Sta77a, Sta78b, Sta78a, Sta77b, Sta78d, Ste71, Ste73, Ste76a, Ste76b, Ste76c, Ste75a]. review [Ste75b, Ste76d, Ste76e, Ste70, Sti77, Sto70, Sto74, Sto71, Sto73, Str73a, Str76, Str75, Str72, Str71b, Str78, Str73b, Sum75, tZS73a, tZS73b, tZS76, tZS78, Süs70, Süs71, Süs76a, Svi77, Swa70, Swe72, Swe73, Swe75, Swe78a, Swe78b, Swi78, Syl72, Tag72, Tan78, Tar74, Tar79, Tav72a, Tav74, Tay75, Teb74, Tha72, Tha73b, Tha73a, Tha76, Tho75a, Tho75b, Tho76, Tof72, Top72, Top74, Top79a, Top79b, Tri72, Tri77, Tro72, Tro75, Tru71, Tuc73, Tur75, Tur76, Tyl77, Uch75, Ung78, Use78, Vac70, Vac71, Vac72a, Vac72b, Vac72c, Vac74, Vac77a, Vac77b, Vac78, Van71, Van78a, Van78b, Vol78, Wag73, Wal71a, Wal71b, Wal73b, Wal73a, Wal74, Wal76a, Wal76b, War70, War73, War76, Wat72b, Wat72a, Wat73, Wat76a, Wat76b, Wat77, Wat78b, Wat78a, Wat79a, Wat79b]. review [Web76, Web74, Wei71, Wei73, Wei78a, Wei78b, Wei79, Wei75, Wei76, Wei72, Wel70, Wel71, Wel73a, Wel73b, Wel77, Wer70, Wer78b, Wer78a, Wer79, Wes76, Wes77, Whe78, Whe73, Whi70, Whi71c, Whi71b, Whi72b, Whi72c, Whi72a, Whi73a, Whi73c, Whi73b, Whi73d, Whi75a, Whi76a, Whi77b, Wik70, Wik72, Wik73, Wik77, Wil72, Wil70a, Wil71a, Wil73b, Wil77, Wil70b, Wil71b, Wil76, Wil78, Win78, Win70, Win71, Wis79, Wol78, Woo79a, Woo72a, Woo72b, Woo77a, Woo77b, Woo79b, Woo79c, Wri76, You72, Zin79, dC70, dC72a, dC74a, dC74b, dC76, dC78, dC79, dc72b, vD75, vS72, tzS75, Emm79, Bar73a]. Revised [Bac75]. Revisited [Bur75, Hel73, How70]. revoliutsii [Lie75a]. revoliutsiia [Lie74c, Lie75d, Lie77a]. Revolt [Hab72, Shr73]. Revolution [Bou78, Cam74, Con71e, Con73g, Cow76b, Dal72, De 74e, Dub77, Hal76, Hun71, Lie77d, Pur71, Rah79, Ran76, Rog78, Sto70, Swe78b, Web76, Lie74c, Lie75a, Lie75d, Lie77a, Fre71b, Buc71b, Buc74, Emm76, Hin79, Lea70, McM75, Sch76, Smi76a, Ste75a]. Revolutionary [Art70]. révolutions [Her75]. Revolvers [Cha72a]. Revolving [Hal70]. Rex [Car75b]. Reynold [Fit73a]. Reynolds [Hig79b]. Rhee [Cos71]. Rhinebeck [Slo70]. Rhodehamel [Smi74a]. Ribeiro [De 70]. Rica [Bos71]. Rice [Sch78a, Ols74]. ricerche [Rap75]. Rich [Koh72]. Richard [Bai74, Cla78, Con71f, Esp72, Fri79b, Gar72b, Gor72, Hig77, Jac75b, Jac75a, Jam79a, Koh72, Lew70, Min78, Mon72, Noe77, Pei72, PS71b, Rey75, San74, Swe75, Whi71b, vD75]. Richardson [Dob75, Rae79a, Vac72b]. Rickards [Buc70a]. Rickover [Rol78]. ricognizione [Kel70a]. Riddle [Hig77]. Rider [Fer71a]. Ridiculous [She75b]. Riding [Hit76, dLM76]. Riedel [Lie75b]. Riefstahl [Per70a]. Rifle [Smi74b]. Rifles [Rop76c, Cha72a]. Righini [Sch76]. Right [Cow76a, Fit76, Lew70]. Riot [Jon78]. Ripoll [Gli73a]. Rise [Bak72, Cha78b, Dau72, Pos78, Süs76a]. Rising [Van78c]. Risk [Hil78]. Risks [Ols76]. Ritterbush [Ros75d]. Ritti [Mar72b]. Ritual [Whe73]. Rivalry [Sha78a]. River [Hey74, Rou75, Sle78]. Rivera [Cla73]. Rivers [Goo70b]. Rivoira [Con78b]. Road [Con72b, Hal76, Rae73c, Wik73, Rea78].

Roads [Ber70a]. Robbins [Buc70b]. Robert [Bat79, Bil77, Bra76a, Can74, Con73e, Con75d, Cot72, Dra78, Eat75, Emm72, Fer73a, Fra71, Har74a, Hei77a, Hil70c, Hol70b, Hou77, Kap74c, Kle76, McM78, Mei72, Por74, Pri75, Seg79a, Sha74, Smi74a, Sta78d, Ste76e, Ste70, Tav72a, Wal73a, dC74a, Par72b, Whi74, Pur78]. Roberta [Siv75]. Roberts [Wel71, Bla77, Eis76, Smi72c, Wel71, Wel77]. Robin [Car78b, Emm74, Emm79, Rop72, Rop76a]. Robins [How74]. Robinson [Bor71a, Con76d, Law71, Pur71, Bar73b, Bay76a, Bor71a, Buc71b, SB72, Woo79b]. Robrecht [Whi73c]. Rockdale [Hin79]. Rockefeller [Con79d]. Rocket [Per70b, Sha72b, Sha73, Sha78c]. Rocketry [Emm71, Hac78, Hal74, Sha78c]. Rockets [Sha72b]. Roderick [Ber77]. Rodgers [Ben78, Süs71]. Rodney [Whi72b, Whi73a]. Roe [Eze78]. Roebuck [Sky70]. Roegen [Aga77]. Roemer [Sti77]. Roger [All74, Bri72, Dev76, Kap71, Mar72a, Min79b]. Rogers [Cal70b, Hou78, Kel70b, Tan78]. roku [Siv77]. Roland [McM76b, Woo79c]. Role [Bra77b, Bry73a, Cow75a, DeG75, Hel71, Lie74a, Mul78, Zup77]. Roles [Cal71a]. Rolf [Har71b, Har73a, Har74b, Har76b, Har77]. Rolled [Smi72a]. Roller [Tha73a]. Rollie [Mat74]. Rolt [Had78d, Buc75b, Car72d, Hil76, McC70a, Rob70]. Roman [Dra70, Dra72, Fus72, Fus76a, Moe71, Mos75b, Nor77b, Pac74, Pac76, Sky73, Whi75a, Smi70b, Hil74]. Rome [Sky70]. Römern [Har73b]. Romorantin [Con74h]. Ron [Gor79]. Ronald [Bar73b, Gil72b, Lie78c, Nut79b, Rae71b, Ros72]. Rondo [Clo70a]. Roosevelt [Seg77c]. Root [Use74]. Roots [Bla76b, Cra77, Fit73a, Fra76, Hal72a]. Rope [Pur75]. Roriczer [Han79a]. Rosabeth [Gar78]. Rose [Pic77]. Rosemarie [Con76d]. Rosemary [Die73, Gue73]. Rosenberg [Sin72, Use78, Wil73b]. Rosenbloom [Pei72]. Rosenblum [Cal71a]. Rosetti [Adr70]. Rosovsky [Kob74]. Rossi [Smi71a]. Rossii [Lie76a]. Rossini [Lid76]. Rossiter [Bee76]. Rossmann [Fre73]. Rothblatt [Rie70]. Rotten [Esp70]. Rottensteiner [dC76]. Rouse [Jac77]. Rowe [Hil78]. Rowland [Con76b, Pos75a]. Rowlands [Row79]. Rowsome [Dre76]. Roy [Bak72, Bos71, Caw77, Gie76a, Lol78]. Royal [All79a, Con74h, Dob78, Mos75a, Nut79c, Sha76b, Smi76b, Tha73b]. Roysdon [Fer79b]. Roździeński [Tyl77]. Rozwoj [Mos79]. rozwoju [Pia70, Pia71]. Rubin [Van78b]. Rudenko [Mei71b]. Rudisill [Gor72]. Rudolf [Fre73, Siv72]. Rufus [Kni77]. Ruina [Rog75]. Rules [Bal73b, Cal71a]. Rumford [Hin71b]. Run [Bar70a, Bar70a]. Rundell [Gie78]. Running [Hil70b]. Rupert [Bry77a]. Rural [Bos71, Con74c, DeW78, Gre79, Kui78, Min77a, Pur70, Vol78, Wik72, Vol78, Bra79d, Vol78]. Rürup [May77]. Russell [Bal75a, Buc78, Con78c, Lie74d, Tur76, Whe78]. Russia [Lie76a, May70, Bla73a]. Russian [Bla72, Buc71a, Lie73b, Car77b, Lie73c]. Ruth [Vac77a]. Rutherford [Hei70, Hei77a]. Rutherfurd [War71]. Ryn [Tar79]. Rynin [Swe73]. Rysselberghe [Tuc78].

S [Ber77, Bil77, Bra77a, Bra72, Bri79, Bry76a, Buc78, Buc79, Car78d, Dau72, Fer73c, Fer74b, Gil73c, Gla74, Goi73, Had76c, Had78d, Hug70, Ino76, Isa76, Jam79a, Jer79, Lea76, Lea79, Lie73a, Lie74b, Lie77c, Mal78, Mar72c, McN78, Nic71, Pei72, Pos75a, Pur78, Rei71, Row76a, Sha76a, Smi73b, Smi77c, Str78, Tav74, Top72, Top79a, Vac72c, Van78a, Wal74, Wer79]. S.J. [Tav72a]. S.S. [Lee78]. Saad [Nee70]. Safeguarding [Lei71]. Safety [McM77]. Safford [Svi77]. Saga [Hig79b]. Sail [Rey74, Woo79c]. Sailors [Cha73b, Sha72b]. Sakakida [tZS73a]. Salisbury [PS71b]. Sally [Ros76b]. Salomon [Boz75]. Salt [Mul76]. Saltpeter [Mul71]. Saltworkers [Mul79a]. Salvage [Kel71b]. Salvation [Seg78]. Sami [Ger74]. samoletostroeniya [Bra72]. Samuel [Cha72a, Fer76a, Hel71, Pur70, Wat76a, Min77b, Mul79a, Fri76]. San [Bry73b, Cra77, GD73, Wat77, LLC^{+}74]. Sanborn [Hin71b]. Sand [Bal78]. Sandberg [Jer76b]. Sande [Bat77]. Sanders [Low76]. Sanford [Win78]. Sang [Hen76a]. Sang-woon [Hen76a]. Sanitary [Cai74]. Sanitation [Cai72, Con79e]. Santa [Cat76]. sapiens [Row73]. sapiens/Origine [Row73]. Sarnoff [Lon78]. Satō [Smi77a]. Satellites [Mar77a]. Sauer [Arn70]. Savant [Car73]. Savelle [Hen76b]. Sawers [Rae71b]. Sawyier [Mar77b]. Scaglia [Con71a]. Scale [Vol78]. Scenario [Woo77b]. Scenic [Con79b]. Schauer [Eze77]. Scheele [Hec70]. Scheffer [Con71f]. scheuten [Bry71b]. Scheutz [Bra79c]. Schimank [May71b]. Schiøler [Hil74]. Schlaifer [Bil77]. Schlebecker [Ras74, Ras76]. Schmauderer [Har71c]. Schmookler [Tar74]. Schmuck [Bed71]. Schmuttermayer [Han79a]. Schob [Swi78]. Schofield [Hol70b]. Scholar [Maa78]. Scholarly [Nee79]. School [Con72e, Lev75, Wal73b]. Schooler [Kap72]. Schreiber [Hal72d]. Schrifttum [Osp79]. Schulz [San78]. Schumacher [Fer74b]. Schuss [Fre73]. Schwartz [Vac72c]. Schwarzmann [Coh74]. Schwedmfestungen [Bec78]. Schwob [Hal72c]. Science [Ano70f, Ano75n, Bag75, Bee76, Bla78a, Bla78c, Bra77a, Bra77b, Bra79a, Bri78b, Cah76, Car76, Cha78b, Cha71a, Dav75, Fai71, Fei78, Fel77, Fer71a, Fer73c, Fow78, Gha73a, Gli70b, Gru71, Hal76, Hod79, Hok79b, Hug76b, Jen73a, Jer77, Jev76, Kah70, Kap71, Kap74a, Kap74c, Koh76a, Koh74, Lan73, Law71, Lay71d, Lay76b, Lev75, Lid76, Maq70, Mat78, May76, McG76, Mee78, Mul79b, O’B71, Pat75, Rei70, Rei71, Rog75, Ros72, Ros76b, Ros79d, SG75, Sha76a, Sha76b, Smi70a, SR73, Str78, Süs76a, Tha73a, Tob76, Ung72, Vac70, Wei71, Wil70a, Wil70b, dC76, Gra75, Lie75a, Ste76e, Con78e, Sol76a, Wei78b, Boz77, Jen77a, Kap71, Kap74b, Mor74, Rog75, Rog77, Art70, Bla72, Boz75, Buc71b, Cal70a, DeG76, Dur79, Eis75a, For79a, Fuk77, Gha70b, Gha73b]. Science [Ham78, Hen76a, Kap72, Lay75, Lay77c, Lev75, Low76, Maj75, Maj77, McC73, Mor72a, Mor77a, Rav72, Ros75a, SG75, Siv75, Smi72b, Sol76a, tZS73b, tZS78, Tav72a, Vac77b, Wal73a, Wei73]. Science-Technology [May76]. Sciences [Bri72, Buc70a, Cah76, Car78d, Caw77, Chu79, Col71, Mic70b, Nut79c, Ols76, Sha76b, Sto71, Vac71, Win71, Cah76, Kar76, Bay76a, Tur76, Whe78]. Scientia [Mer76a]. Scientific [All79a, And74, Bis76a, Bra79c, Bra70b, Bri72, Buc70a, Car78d, Chu79, Cut74, Del76, Don77, Gil72a, Gol70b, Hol70b, Lay79, Lei73, Lie73a, Lie77b, Lie77d, Ols77, Par79, Pos76d, Pur76b, Rey79b, Ros76b, Row72a, Sch76, Sel78, Sto74, Tho75a, Top79a, Top79b, Vac72b, Lie74c, Lie75a, Lie75d, Lie77a, Lie78b, Rah79, Wil70a]. Scientific-technical [Lie73a, Lie77b, Lie74c, Lie75a, Lie75d, Lie77a]. Scientific-Technological [Rah79]. Scientist [Bon71, Bri78a, Cla73, Gol70b, Tha72, Hal73, Kre76]. Scientist-Engineer [Cla73]. Scientists [Bri78a, Hac73a, Kap72, Kev75, Kra77a, Sha72b, Chu78b]. scienza [Clo70b, Kel70a]. Scope [Cah76]. Scotland [Rob76a, Rob76b, Buc79]. Scott [Buc78, Eat73, Mar72a, Bak72, Ste76a]. Scottish [Con71c, Eme78, How72, Jer74, Sch77b, Don77]. Scrapbook [Con72c]. Scribbins [Ove71]. Scriptorum [Hal72d]. Scully [Cra75]. Sculpture [Vac70]. Sea [Tho76, Tho75a]. Seaborg [San72]. Seale [Con78c, Mar76a]. Seamanship [dc72b]. Search [Mor70, Smi77c]. Sebree [Pos76c]. Secoli [Kel70a]. secolo [Rap75]. Second [All79b, Nee70, Smi74c, Buc76b]. Secondary [Ano73i, Ano73l]. Secret [Buc75a, Fri76, Hac72]. Sector [How78]. Security [Kev75]. Seeds [Hun71]. Seeger [Nor75]. Segrè [Kle71]. Seigerhüttenprozesz [Osp79]. Seijo [Gli78]. Seiyû [Siv77]. Select [Fer71a]. Selected [Con71d, Fuk77, Tar74, Top79a, Wil70b]. Selection [Jac79, Wil71b]. Self [Car78b]. Self-Sufficiency [Car78b]. Sellers [Cha72a]. Semiautomatic [Cla72]. Seminar [Bur75, Rah77, Eis75b]. Semmler [Rae74]. Senegal [DeG75]. Senegambia [De 76]. Sentries [Hac72]. Separate [Fer70a]. Seper [Bry71b]. sept [Her75]. September [Buc76b, Hal72d, Hal74, Kel70a, Rah77, Sol76c]. Sergei [Mei71b]. Serials [Orł73]. Series [Pur76b]. Servants [Bag79]. Service [Bra79a, Lie78a, San73b, Hec70]. Services [Cal74]. ses [Sko75]. Sethi [Ino76]. Settlement [Koh70c]. Settlements [Qui78]. Seventeenth [Hen76b, Mok77, Mul79b, Nut79b, Tha72]. Sewage [Tar79]. Sewing [Cal70b, Hou77, Hou78]. Seyni [DeG75]. Shabad [Bra79b]. Shadel [Had75a]. Shahar [Sol76b]. Shall [Gli79]. Shand [Hun71]. Shape [Con73c]. Shapes [Smi72a]. Shaping [Bri76b, Fow78, Kap71]. Shares [Fen78]. Sharon [Nor77b]. Sharpe [Per70b]. Shay [Whi72a]. Shea [Sch76]. Sheffield [Aus77]. Shelby [Han79a, Kel70b]. Shelf [Bla77, Fla75]. Shepard [Mor77c]. Shepherd [Ehr73]. Sherfield [Spo73]. Sherry [Sha72c]. Sherwood [Buc75a]. Shield [San74]. Shigeru [Fuk77]. Shils [Gol70b]. Ship [Bag79, San70]. Shipbuilding [Buc79, Slo72, Bak76]. Shipping [Ehr73]. Ships [Cha73b, dc72b]. Shipyards [Cha73b]. Shirley [Kre76]. Shock [Bor71b, Fer78b]. Short [Hec70, Meu77]. Short- [Meu77]. Shortline [Gol77a]. SHOT [Ano70e, Ano78h, Ano79f]. Shoulders [Sha79]. Shover [Min77a]. Showalter [Rop76c]. Showman [Nut78]. Shumaker [Han79b]. Siberia [Mei71b]. Siberian [Bra79b]. Sicherheit [Bry70]. Sicherman [Sko76]. Sickle [Fus71]. Sides [Bra77e]. Sidney [Adr70, Bla74b, Gor74, Gor79, Ron78]. Sidorov [Lie76a]. siècle [Rob71, Day74]. Siegel [Spi74]. Siegfried [Bry70]. Sigfrid [Son78].

Signitzer [Mar77a]. Sigurdson [Vol78]. Silk [Bee75]. Silverman [Kel71d]. Silvio [Str73b]. Sim [Tar79]. Simeon [Smi73a]. Simla [Rah77]. Simmons [Car72c]. Simon [Tru71, Mic70b, Wei72]. Simple [Kra76, Smi71b, Bat70b]. Simplified [Cot72]. Sims [Swe78b]. since [Edw72, Fli72, Hig79a, Hod79, Lay77c, McM76b, Nes77]. Sinclair [Con71b, Tha76]. Singer [Hou77, Wei78b]. Single [Con74a, Har79b]. Single-Control [Har79b]. Sinking [Par72b]. Sir [Bri71b, Emm77, Hig79a, Sha76b, Sha78b, Gor71a, Gil72b, Wer78a]. Site [Wal73b, Wer79]. Sites [Fra74b, Per78b, Wat79a]. Sivin [tZS78]. Six [Bla76a, Nel72]. Six-wheeled [Nel72]. Sixteenth [Hit76, Mul79b, Tho76]. Sixth [Hac78]. sixties [Bar73b]. Sixty [Con75d]. Skeat [Rol74]. Skehan [Tav72a]. Skein [Bee75]. Skill [Bou78]. Skilled [Kra75]. Sky [Cra78]. Skyscraper [Con78a, Con76d]. Slate [Hud75a]. Slave [De 74d, De 76]. Slaven [Rob76a]. Slavery [Coo77, Cro76, Web76]. Sleepy [Jud74a]. Slims [Cow79]. Sloan [How72]. Sloane [Pau71, Pau71]. sluzhbe [Lie78a]. Smaldone [All79c, DeG77]. Small [Fri75, Vol78, Cha78a, Ros79a]. Small-Arms [Fri75]. Small-Scale [Vol78]. Smith [Con79a, Dre75, HS78, Par74, Sum75, Bla76a, Bry77a, Car77a, Cha72a, Con73b, Dib79, Eze78, Hig79a, Hud71, Le 71, May71a, Swe75, vD75]. Smithson [Con71c, Con71c]. Smithsonian [Bis76b, Sum75, Lew77]. Smokey [Hig79b]. Smyth [DeG75]. Smythe [Sal71]. Snow [Por76]. Snyder [Rop76d]. Sobel [Por74]. Soberman [Con71f]. sobre [Gli70a]. Social [Bay76a, Bil73, Chu79, Cla77, Col71, Con70d, Cos71, Cow75a, Cow76b, DeG71, De 79a, Eme78, Fer70b, Fer78b, Fer74b, Fri78a, Gla78, Hab72, Hal72a, Hog75, Ino76, Kni77, Kra77a, Lei73, Lie74d, Lie79a, Mas77a, Mol78, Mul74a, Oka78, Rah79, Ris78, Rog77, Row79, SG75, Siv75, Smi72b, Smi75b, Spo73, Teb74, Tho78, Tri72, Van78c, Gil72a, All79a, Byr74, Dav75, Kap74b, Mar77c]. socialism [Lie77b, Lie77a, Lie79a, Bla79a, Bla78c]. Socialist [Lie75b, Lie76b]. Societies [Fee79, Cam74, Lak70]. Society [Bra78, Byr78b, Eis75b, For79a, Fuk77, Lay70b, Lay71b, Le 71, Lei73, Mok77, Mos75a, Pal79, Pur76a, Seg77c, Smi72b, Str71b, Tha72, Web79, Wei72, Wer78b, Lie75d, Ano74g, Hor79, Jen72, Mul78, Aga70, Bla74c, Byr74, Cal70a, Dat71, Fus76b, Gie77, Hor79, Kra76, Mar72a, Mat76, Mul78, Ols77, Pac76, PS71b, Rae74, Ste76b, Tur75, Wei75, Wei76, Kel76]. Socio [Bla79a, Fel78]. Socio-Cultural [Bla79a]. Socio-Economic [Fel78]. Sociological [DeG77, Swe78b, Wil72]. Sociology [Chu79, Hor79, Kap74c, Kra74, Lie75e, Tav74]. Söderfors [Alt70]. sogenannte [May71b]. Sokoto [DeG77]. Sola [Mol78]. Solberg [Wil77]. Soldiers [Rop76c]. Soleri [Whi71a]. Solitude [Sha74]. Solla [Dra76a]. Solving [Tho75a]. Some [Ber77, Bra70a, Fer71c, Min79a, Mul74e, Ros70, Con75e]. Somerset [Hil71, Tan78]. Something [Esp70]. Son [SB72, Tho78]. Sonnenberg [Bry70]. Sons [Jen70]. Sontag [Vac78]. Soot [Pat75]. Sotsgorod [Lie76b]. sotsializma [Lie77b, Lie77a, Lie79a]. Sotsial’no [Lie79a].

Sotsial’no-ekonomicheskie [Lie79a]. Source [Jac73]. Sources [Ano73h, Ano73i, Ano73k, Ano73l, Bal75a, Bee70, Rop72, Rop76a, Wil71b, KF76, De 75b, Fer73a, Fuk77, Goo74a, Hil77]. South [Bed76, BM77, Coo77, McM76b, tZS73a, Cok72, Sha71, Cra75, Buc70b]. Southern [FS78, Hol79, Sau72, Bar73a]. Southwest [Con72d, Fra74b, Wer79]. Southworth [Dal70]. Sovetskaia [Lie74b]. sovetskogo [Bra72]. Soviet [Bla75, Bra72, Con71e, Eze77, Lie73d, Lie74a, Lie74b, Lie74e, Lie74f, Lie77c, Lie78c, Mag76, McC73, Ano73j, Bai76, Bla78a, Emm79]. Sovremennaia [Lie74c]. Space [Con73d, Cra78, Eze77, Hac74, Ord78, Sha78a, Sha78b, Swe71, Fow77, Hac72, Swe72, Vac72a, Fow79]. Space-Travel [Hac74]. Spaceflight [Hac73b, Swe78b]. Spaey [Sol76a]. Spagyrical [Maj75, Maj77]. Spain [Gli70b, GD72, Gli77]. Span [Wat72a]. Spanish [Cok72]. Spans [Con75a]. Spark [Nor77a]. Speaks [Hel77]. Speas [Bil75a]. Special [Sin72]. Spectroscopist [War71]. Speed [Hig73, May75b]. Speer [Whi71b]. Spence [You72]. Sperry [Dan72]. Spier [Cot72, Mei72]. Spillman [Had75a]. Spindle [Jer78a]. Spinning [Hil75a, Jer70b]. Spirit [Smi75a]. Spiro [Cra79c]. Spoils [Lie74e]. Sponsorship [Pos72]. Sporn [Lam71a]. Spree [Son78]. Sputnik [Bri78a]. Spy [Har75b]. SShA [Lie78b]. SSSR [Lie74b, Lie79a]. SST [Smi76d]. St [Bry73b]. St. [Mor78]. Stacy [She75b]. Stadt [Geb71]. staff [Bra77b]. Stafford [Cas74, Had77c]. Stage [Bez79]. Stages [Clo70a]. Stained [HS78]. Stake [Del77]. stali [Smi76c]. Stalinism [Bai76]. Stanback [Mon72]. Stand [Maj74]. Standard [Sal73, Smi74c]. Standards [Van78c, Zup77]. Stanley [DeG71, De 74c, Dre75, Ger79, Mor77a]. Star [War70]. Starlifter [Hig73]. Starr [Mul74a]. State [Con78c, Fri78b, Joh78, Mul74e, San73c, Wel71]. Statement [Ger76]. States [Cra78, Fel76a, Gar78, Goo74a, Jac71, Jac77, Jam76, Koh71a, Koh77, Rop76a, San74, Sin72, Sma78, Spr76, Sta77a, Cal74, Con75d, Ham78, Ros75a, Wel70]. Statics [Gil73a]. Station [All79b, Hac74, Koh76b, Sha72b]. Stationary [Gol78, Lay71a]. Stations [Kev71, Con73e]. statische [Sch71a]. Statistics [De 73b]. Status [Van78c, Wei78b]. Stavrianos [McN78]. Steam [Bag79, Con75g, Gol78, Had78d, Hea75, HP72, Lay71a, May75b, Pur71, She78b, Whi73e, Kap78, Rey74]. Steam-driven [HP72]. Steamboat [Hun75]. Steamers [Pos75a]. Stearns [Ste75a]. Steed [Mur79]. Steeds [Bat70a]. Steel [Aus76, Bla75, Bru76, Har75a, Sch78b, Smi76c, Aus77, Hud79, Wer70]. Steele [Buc75a]. Steelmaking [Par72b]. Steensberg [Fen78]. Stein [Wer78a]. Steinhoff [Sha78a, Sha78a]. Stellung [Lie75b]. Sten [Gha70a]. Stepan [Mat78]. Stepchild [Pos76c]. Stephen [De 79a, Kap74b]. Stephenson [Pos76a, Rol74]. Stevens [Con74f, Cha72a]. Stewart [Fow79, McK73, Die73, Jam79a]. Still [Hil70b, Slo72]. Stilwell [Fra73]. Stine [Emm76]. Stites [Mar72c]. Stöber [Fer74b]. Stockholm [Web79, Nut79c]. Stöcklein [Hal71]. Stone [Cra78, Wat72b, Fen78].

Stoneman [Rog78]. stopów [Pia70]. storia [Clo70b, Kel70a, Rap75, Clo75]. Storm [Whe73]. Story [Bed71, Bee79, Bru73, Bud78, Cas76b, Fra76, Gol74a, Gol77a, Joh73, Pos76b, Ove71]. Stourbridge [Per72]. Stover [Con71c, Hig79b, Hig79b]. Strategy [Con79e, Ros75b, Wil76]. Stratification [Gue77, Kap74b]. Strauss [Maq70]. Streamlined [Cra76b]. Street [Bru73, Fri78b, Mas77b, Mer76b]. Stress [MCA74]. Strong [Hok77, Row77, Bee70]. Strongholds [Bac79]. Structural [Con76b, Con78e, Ros71a]. Structure [Jer76a, Par79, Sim71, Kra75]. Structures [MCA74]. Struggle [Ran76, Sew72, Svi77]. Stuart [Bor71a, Led77]. Stubs [Pen75]. Students [Fer71a]. Studi [Clo70b, Kel70a, Clo78]. Studia [Mos78]. Studie [Str72]. Studies [Ano73n, Con71a, Fel76a, Pri75, Str78, Top79a, Tur76, Wal73b, Wei78b, Whe78, Had70, Had71, Had75a, Had76a, Top79b, Car79c, Had73, Had74a, Had77a, Had78a, Wat78a]. Study [Bed76, Bel78, Bra78, Bro79, Bur74a, Byr78b, Clo70a, Eis75a, Fer73c, Fow78, Gal75, Jer76b, Kap74a, Kul79, Lay71a, Lay71c, Lea70, dLM76, Per72, Per70b, Rae71a, Sim71, Tay75, Tha72, Vac71, Vin79, Whi75b, Win70, ZIL77, Ham71, Lan73, Swe78b]. Stuewer [Dev76]. Style [Con76d, Con77c]. Styles [vR78]. Subculture [Kur73]. Subiaco [Smi70b]. Subjects [Bri78b]. Sublimations [Mar72c]. Sublime [She75b]. Submarine [Fri76, Rol77b]. Submarining [Fer73b]. Subway [Con72a]. Subways [Bru73]. Success [Lon78]. successeurs [Sko75]. Suchodolski [Ros74]. Sudhir [Lay71c]. Südosten [Hal72c]. Sufficiency [Car78b]. Sugar [Gli73b, Mat75]. Suggestion [Gil79]. Suhling [Osp79]. sui [Kel70a]. Sulzer [Fre73]. Summary [Cah76]. Summer [Eis75b]. Sundials [Nor77b]. Sunken [dC74a]. Superbomb [San77]. Superior [Abr79]. Superrealism [HS78]. Superwarriors [Smi76d]. Suppe [Fei78]. Supplement [Tav74]. Supplementary [Cah76]. Supply [Bur74a, SA77, Bra77b]. Supplying [Rey79a]. Suppression [De 74d]. Surrealism [HS78]. Surrealists [HS78]. Surrey [Pen75]. Survey [Buc70a, Cah76, Fra74b, Lak70, Maj77, Nes77, Pei72, Qui78, Tav72b, Wel73a, Woo77a, San73b, Sha78a, Har74a, Lew73]. Surveying [Cot72]. Surveyors [Sky73]. Survive [Mul73a]. Susan [Vac78]. sushchnost [Lie73a]. Susskind [Dru74, Wis79]. Sussman [Cra77]. Sutherland [Ber74]. Sutton [Lie73d, Lie73d, Lie74f]. Suzannah [Nob71]. Suzanne [Kel77]. Svedenstierna [Har75b, Har75b]. Sweat [Har76a]. Sweden [Fel76a]. Swedish [Nut79c, Hud76]. Swenson [Top72]. Swords [Bra78]. Sylvester [Cai72]. Sylvia [Lub78]. Symposia [Hac78]. Symposium [Abr78, Bal71, Emm71, Hal74, Kel74, Nob71, O’B71, Ord70, Ras71, Koh70b]. Synergetics [Fer76c]. Synthetic [Maj77]. Syntony [Nor77a]. Syracuse [Bur74a, Sim77]. Syrett [Sin70b]. Syria [Fen78]. System [Bic73, Bra77a, Bra79a, Cai72, Fri75, GD73, Gor71a, Hug79a, Jam76, May71a, Rae73c, Sin72, Use74]. Systems [Bri79, Hug76b, Ras75, Ros73b, Ros75c, Tar79, WK71, Bee73, Bla71, Row79].

T [Bal73a, Ben78, Bil79, Bla73a, Car72d, Car79b, Cha73a, Con78b, Dib75, Fer73b, Fra74b, Fri79a, Fus71, Gal75, Hab72, Had78c, Had78d, Har75b, Hig79b, Hil76, Hil70b, Hun71, dLM76, May75b, McC70a, Mor79, Ras74, Ras76, Rob70, Rot72, San72, Smi75b]. Taccola [She75a]. Tactical [Rop76b]. Tactics [Bac79]. Tadeusz [De 75b]. Taft [Hog75]. Tafuri [Hay79]. Tager [Con71d]. Tallest [Wat76b]. Tallulah [Gol77a]. Tamayo [Gli70b]. Tamotsu [Smi77a]. Tang [Aga77]. Tank [Sha76c]. Tanks [Lie73c, Sha76c, Gor73]. Tarn [Gai75]. Tarpley [Joh78]. Tarr [Cai78]. Taskers [Hil76]. Tasteful [Mar76a]. Tatham [Koh71b]. Taviss [Kel71d]. Taylor [Lay71c, May78a, Min79b, Bal77]. Te [Wat78b]. Teach [Fer76b]. Teacher [De 74f, Kel73]. Teams [Gli70b]. Tear [Jon78]. Tech [BM77]. Technical [Bac75, Buc70a, Buc70b, Cal74, Dra72, FS78, Jer70a, Rae71b, Spa70, Svi77, Lie73a, Lie74c, Lie75a, Lie75d, Lie77b, Lie77a, Lie78b, Pur76a, CZ73, De 74f, Str76]. Technics [Car78a]. Technics-out-of-Control [Car78a]. Technik [Bec78, Hal71, May78b, Son78, Har71c, Fuk75b, Lie76c]. Technik- [Son78]. Technik-Wirtschaft-Gesellschaft [Fuk75b]. Technikgeschichte [Har73a, Tro72, Har71b, Har74b, Har76b, Har77, May77]. techniky [Win71]. Technique [Abr79, Fus71]. Techniques [Con74d, Han79a, Hin71b, Smi77a, Fre71b, KF76, Cot72, Dre75, Rop76b, Wer79, KF76]. technischen [Bry70, Har70]. technischer [Hal71]. technischgeschichtliche [Str72]. Technocracy [Seg77b]. Technocrat [Seg77b]. Technoeconomic [Car77b]. Technologia [Abr79]. Technological [Bay78b, Bla75, Car75c, Caw77, Con73g, De 74c, Del76, Fel74, Fre71a, GD72, Gil72a, Han73, Hug76a, Lam71b, Lie77d, Lie78c, Lie79b, Man72, Mat76, Mea71, Mei74, Per70b, Ros79b, SA77, Tri72, Uch75, Vin79, Wil70a, Win70, Sta78d, Bat74, Hil78, IS77, Car71a, DeG71, Gil70, Lew71, Qui71, Rah79, Rog78, Ros73a, Ros76a, Sta78a, Zin79]. Technologie [Osp79]. Technologies [Jam79a, Mei70, SB78, Wol78, Mei72, Wei78b]. technologische [Rap77, Rob76d]. Technologist [Bon71]. Technology [Bag75, Bai76, Bal71, Bel78, Ben75, Bla78a, Bla79a, Blo77, Bra76c, Bra78, Bro79, Bry77a, Buc71b, Bur70, Byr78b, Cah76, Cai78, Car76, Car78a, Car78c, Cha77, Cha78b, Cha71a, Cha71b, Con71a, Coo71, Cow76b, Cow79, Dan70a, DeG76, De 74e, DeW78, Dra78, Dur72, Eis75b, Eze78, Fee79, Fel77, Fer71a, Fer73c, Fer74a, Fer78a, Fer79a, Fer70b, FS78, Fla75, For71, For79b, Gha73a, Gla78, Gol77c, Goo70c, Goo71, Goo72, Goo73, Goo74b, Goo75, Goo76, Goo77, Goo78, Goo79, Gru71, Hac77, Hal76, HV78, Ham78, Hel71, Hen76a, Hew76, Hol79, Hop75, Hud71, Hug76b, Hug79b, IS77, Jac75a, Jam79b, Jen75, JR78, Jer73, Jer76b, Jev76, Jon78, Kah70, Koh76a, Kou70, Lay70b, Lay70a, Lay71d, Lay74, Lay79, Lie73d, Lie74f, Lie75e, Lie76c, Low76, Maj75]. Technology [Maj77, May71a, May76, McG76, McL78, McL79, Mee78, Meu71, Mil76, MGM78, Mor78, Mor77b, Mor79, Mul74e, Nob71, Orł73, Par70, Per78a, Pic77, Pur79a, Rah77, Rei70, Rey79b, Rog75, Rop76c, Ros70, Ros71b, Ros79c, Rür74, RH73, San73b, Sch71b, Seg77a, Sew72, Sin79a, Siv75, Smi70a, Smi71a, Smi72c, Sol76c, Spi75b, Str73a, Str78, tZS73b, tZS78, Tar77, Tav72b, Tho76, Tho78, Tob76, Ung72, Vac70, Vac72c, Van78c, Web79, Wei73, Whi75b, Woo79a, Gra75, Lie75a, Win71, Ano74g, Fer74b, Fle77, Hei77a, How70, Jam79a, Kel71d, Ron78, Aga78, Bay74a, Bay74b, Bis70, Bis76b, Bur77, Con71f, De 75a, Dru74, Fai71, Ger79, Gue73, Hel71, Hug70, Kel71d, Lay71a, Lay76a, Lid76, McM76a, Mul74b, Sha76a, Smi76b, Tha73a, Wei71, Wil70a, Woo72a, Ano70c, Ano71e, Ano72f, Ano73m]. Technology [Ano74h, Ano75h, Ano76f, Ano77g, Ano78f, Ano79d, Bar70b, Bar71b, Bel78, Bla79a, Bou78, Buc73, Byr78a, Dev78, Fel76b, Fer71c, Fer78b, Fer74b, Jac73, Lam71a, Lie74d, Mer76b, Mic70a, Mon75, Mul74a, Pei72, Ros75d, Ros79c, Sel77, Sim71, Sko76, Smi75b, Tri72, Use78, Wal71b, Wil73b]. tecnica [Clo70b, Rap75, Gli70a]. Teece [Gol77c]. tekhnicheskai [Lie77a]. tekhnicheskaia [Lie74c, Lie75d]. tekhnicheskii [Lie78b, Lie77b]. tekhnika [Lie75a]. tekhniki [Gra75, Par70]. Tekhnologicheskii [Car77b]. teknicheskii [Lie73a]. teknicheskoi [Lie75a]. Telecommunications [Woo79a]. Telegraph [Gor71a, Teb74]. Telegraphy [Sha76b]. Telephone [Hou75, Mol78, Bra77c]. Telephony [Tuc78]. Television [Gol77b, McL78]. Tell [Kap77]. Teller [San77]. Temples [Con78c, Ger74]. tendencies [Lie73a]. tendentsii [Lie73a]. Tension [Top79a]. teplovozostroitel’nogo [Lie75c]. Term [Meu77]. Terminals [Con70a]. Termini [Con71c]. Termodynamika [Pia70]. Terms [Car79a, Ihd77]. Terre [Pic77]. Terry [Bla77, Smi76d]. Terutomo [Uch75]. Test [All79b, Sha72b]. Testing [Pur76b]. Tests [Vin79]. Tew [Hil70a]. Texas [Fra74a, Gie78, SB78, Swe71, Fra71, GD73]. Textile [Car72a, Fel74, HP72, Hit76, Jer70a, Jer70b, Jer73, Jer77, dLM76, Pon79, San73b, Coo74, Jer79, Lea76, Lea79, Mal78, Moe71]. Textiles [Bor71a, Bor71a]. Thackray [Lay75, Mor74]. Theaters [Kel78]. Theatre [Maa78]. Theatrum [Fer71b]. Their [Bal78, Car72e, Dob78, Koh70c, Smi70a, Spr76, Wel73a, dC70]. Thematic [Cut74]. Theme [Car78a]. Themes [Hug79b]. Themselves [Koh70c]. Theodore [Bat77, Bra79b, Jen73b, Tha72]. Theorems [Kel71b]. Theoretical [Fre71a, Kap74c, Tob76]. Theory [Car79a, De 78, May75b, Row77, Sel78, Ros73b, Ros75c, Bun74]. Thermodynamics [Bry73a, Dau72, Pia70, BR74]. Think [Gor73, Süs71]. Thinking [Fer76c]. Third [Bar73b, Emm76, Hac78, Kel74, Kra78, Lie75b, Lie76c, Wal73b, Whi71b]. Thirsk [Pal79]. Thirteenth [Ano70e, Ano70f]. Thirties [Con77c, Wik77]. Thirty [Bry73b]. Thirty-fifth [Bry73b]. Thomas [Art70, Bar71b, Bee70, Buc75b, Cra78, Dan72, Fuk75a, GD73, Had78d, Hil75b, Jer70a, Koh78, Mon72, Ros79c, Sch77a, Ste76c, Str71b, Top79a, Kre76, Sch77a]. Thomson [Rob76b]. Thor [Nie70b]. Thorkild [Hil74]. Thorold [Gia72]. Those [Bil75b]. Thought [CZ73, Hac74, Lew70, May71a, Car78a, Cut74, Vac72b]. Thousand [Fer73b]. Thrall [Mul74a]. Threat [Caw77]. Three [Fer73b, Ran76, Bis76a, Ols74]. Threshold [Row76b]. Thrive [Ras76]. Thus [Siv72]. Thy [Har76a]. Ti [Smi77a]. Tibi [Gli75a]. Tide [Min79a]. Tideman [vS72]. Tidings [San73c]. tierras [Gli78]. Timber [Jud74b, Sha72c, Whi72a, Bal77, Fra76]. Time [Aga77, Bed71, DeG75, Ols76, Bed71]. Times [Gre72, Pau71, Fer71c, Lew70, Sch74]. Timothy [Gre79]. Tin [Ste75b]. Tirols [Tro72]. tissage [Fre71b]. Titan [Whi72a]. Titans [Sha79]. Tite [Gla74]. Tobacco [Koh71b, Koh71b]. Tobey [Ros72]. Tod [Seg79b]. Today [Jev76, Wei73]. Todd [Lan73]. Toffler [Bor71b]. Together [Bul76]. Toil [Koh72, Siv72]. Toilet [Tar79]. Toilets [Tar79]. Tokaty [Rou72]. Tolai [Cro71]. Told [Koh70c]. Tom [Sca72, Mas77b]. Tombs [Mei71b]. Tomol [Spi78]. Tomorrow [Fow78, Wes77]. Tomorrows [Lak70]. Tongues [Gol70b]. Tony [Cra78]. Tool [Cal71b, Pen75, Wel71]. Tools [Bat70a, Bul76]. Topics [Ros79d]. Torrey [Wes77]. Tortelli [Kel70c]. Tour [Har75b]. Tower [Wat76b]. Town [Fri78b, Fri79b, Con71e]. Towns [Gre72, Bay76b, Ran76]. Toy [Spo72]. Toys [Fer79c]. tr. [dSP75, Whi71b]. tra [Rap75]. Tracing [Mea71]. Track [Wik79]. Track-Type [Wik79]. Traction [Hea75, Mas77b]. Tractor [Wik79]. Trade [Bag79, De 74d, Ehr73, Pur75, Ste75b, Tho75b, Coo71, De 76, Jer76b]. Tradition [Bla73b, Gre79, Hal71, Lay75, Top79a, Mul74a, Cra77, Wal76b]. Traditional [PS71b, PS79]. Traditions [Sin79a, Str75]. Tragedians [Kah70]. Training [De 74f]. Trains [Con75e, Con70a, Con75e]. traitement [KF76]. Tramways [Whi77b, Pos75b, Pos78]. Transatlantic [vD75]. Transcontinental [Bry73b]. Transfer [Bel78, Ben75, Bro79, Fee79, Hel71, Meu71, Pei72, Ros70, RH73, Sol76c, Hel71, Gol77c, Jam79b, Pos76b]. Transferring [JR78]. Transformation [PS71b, Whi75a]. Transformations [Mar79b]. Transformed [Chu78a, Sta78b]. Transforming [Min77a]. Transit [Con71f, Con72a, Con74e, Pos78, Pos79, Pos79, Whi77b, Pos76c]. Transition [Bri76a, Lei73, Ros79a]. Transitions [Mok77]. Translation [Hac73b, Mar72c]. Transmission [Hug76b, Pur75]. Transport [HS70, JR78, Lie78a, Eis76]. Transportation [Bar71a, Pos79, Whi70, Bla74a, Hil77]. Travel [Bil73, Hac74, Har75b, Ord78]. Traveler [Kre77]. Travis [Spi78]. trazione [Clo75]. Treasure [dC74a]. Treatise [Bee75]. Treatises [Dra72]. Treatment [Fre71a]. Trend [Kob74]. Trends [Hil72, IS77]. Treue [Geb71]. Triebwerke [Str72]. Triveni [Bla78c]. Trivulzianus [Mar72c]. Trolley [Pos76c]. Trolleys [Pos78]. Tropical [Bar71b]. Trotsky [Bai74]. Troubled [Wik72]. Truck [Sta78d, Hig73]. Truman [Dob75]. Tsander [Bra70b]. Tube [Gol77b]. Tubes [Pos76b]. Tudor [Kel70b]. Tull [Koh74]. Tuning [Har79b]. Tunnard [Con79b]. Tunnels [Pos76b]. Turbine [Lay79]. Turbojet [Con73g]. Turin [Kel70a]. Turner [Dib75, Hok79b, Kur73]. Turning [Bon79, Byr78a, Dev76, Fer79c, Fer73c, Nor78]. Turnpike [Rae73c]. Turriano [GD76]. Tushman [Tay75]. TVA [Con76e]. Twelfth [Bla74b]. Twelfth-Century [Bla74b]. Twentieth [Con78e, Hit76, Kob74, Le 71, Seg79a]. Twenty [Hig79a, Buc76a, Wis79]. Twenty-Dollar [Hig79a]. Twenty-five [Wis79]. Twins [Lay71d]. Two [Bra77e, Whi75a, Wil70b, Cam74, Sha78a, Hok79a, Wei76]. tyla [Lie78a]. Tylecote [Par78]. Type [Wik79]. Types [Con77a]. Typography [Har71a]. Tyranny [Bed71]. Tyrkiel [Pia70, Pia71].

U.S [Bri79, Cal71b, Pri75, Rop76d, vD75]. U.S.A [Lie78b]. überwinden [Rap77]. UCLA [Eis75b]. UFO [Wri76]. Uhr [Bed71]. UK [Rog78]. Ulla [Alt73b]. Unanswered [Bra70a]. Unbuilt [Cra78]. Uncertain [Jam79b]. Underdevelopment [Jam79a]. Underground [Fra74a]. Understanding [Fow78, Tri72, Mor70, Dru74]. Underview [Nel72]. Underwater [Bur74b, Woo79c]. UNESCO [Bis76a, DeG75, Col71, Dav75, DeG76, Fel78, Ham78, Ras71, Tur75]. Unfouling [Cai74]. Unification [Rop76c]. Union [Lie74a, Lie74e, McC73]. United [Cal74, Con75d, Cra78, Fel76a, Gar78, Jac71, Jac77, Jam76, Koh71a, Koh77, Rop76a, Ros75a, San74, Sin72, Sma78, Spr76, Sta77a, Wel70, Goo74a]. Universal [Ger76, Kel71a]. University [Bra77b, Dra78, Gre73, Koh72, Mor78, O’B71, Sha76a, Tar77, Wil70a, Bra76c, Kra77a, Ung72, Hok79b]. Unknown [Ste76f]. Unorthodox [Smi76e]. Unprepared [Dat71]. Upper [Pen79]. Urban [Bay74a, Bay74b, Bri76a, Con70c, Con71d, Con75c, Fri78a, Gai75, Lub78, Pos78, Sew72, Tar79, Mul73a, Con71f, Con75f, Coo77, Goi70, Mon71, Seg79a]. Urbanism [Kob77]. Urbanistic [Bay78b]. Urbanization [Tag72, Con76c]. Ursprung [Maj74]. USA [Con78c]. Usable [Smi77c]. Use [Bul76, Fel78, Ras74, Sha72c, Ber77]. Uselding [Dug79, Fie77]. Uses [Bed71, Del76, Bee70, Gol75]. Usher [Ano70a, Ano71a, Ano72a, Ano73a, Ano74a, Ano75a, Ano76a, Ano77a, Ano78a, Ano79a]. USSR [Lie74b, Lie79a, Par79, Lie75e, Car79c, Car78d]. Uta [Bra79c]. Ute [Hal72c]. Utilities [Com74]. Utilization [Lew76, Jac75a]. Utopia [Hay79, Seg79b]. Utopian [Sti77]. Utopias [Seg79a].

V [Bak72, Bal73a, Bal75b, Bee73, Had72, Hal73, Hud79, Lie75c, Lie77a, Lie77d, Lie78b, May70, Mon72, Sha74, She75b, Wer70, Wik72, Lie74b, Lie76a, Lie78a, Lie79a, Ano73q, Ano73p]. Valencia [Str71b]. Valeriana [Hel77]. Valerie [Mor77c]. Valier [Ord78]. Valle [Gli75b, Gli73b]. Valley [Hey74]. Value [Jam79b]. Values [Bay78a, Dur72, Eis75b, Hol79, Lay75, McG76, Seg77a, Win70, Kel71d, Con79b, Ros75d, Sch71b]. Vanderbilt [Fri77]. Vanished [Joh73]. Vanishing [Pur70, Qui78]. vapeur [Rob71]. Varagnac [Her75]. Variable [Han73]. Variations [Ehr74]. Vaucanson [FM79]. ved [Win71]. Vehicle [Bry76a]. Vehicles [Ber70a]. Velikoi [Lie74b, Lie78a]. Velocipede [Whi76b]. Venice [Kel71b, Dot75]. Vera [Gli75a]. Veränderungen [Fuk75b]. Verbruggens [Smi76b]. Verden [Bec78]. Verheyen [Wat78b]. Verlag [Hac73b]. Vernacular [Can74]. Vernon [Coo71]. Versuche [May71b]. Versus [Meu77, Pos76d]. Vickers [Had79]. Victor [Bra79b, Car75c, Fel77, Lew71]. Victoria [Had77c]. Victorian [Buc78, Con74i, Cow75b, Hud71, Rob70, Ste73]. video [Mos75b]. viento [Gli78]. Vietnam [Pri75]. View [Con79b, For79b, Har71a]. Views [Fel77]. Viking [dC74b]. Village [Hin79, Wal76b, Min77b]. Villages [Cha72b]. Villareal [Gli75a]. Vinci [Ano70d, Ano71f, Ano72g, Ano73o, Ano74i, Ano75i, Ano76g, Ano77h, Ano78g, Ano79e, Con74h, Mar72c, Bon71, Dib79]. Violence [Cli71]. Viravens [Gli75a]. Virginia [Dew76, Fri79a, Koh76b, McG76, Cow79, Wes76]. Vision [Con71d, Shr72, Whi71a]. Vital [Low76]. Vitruvius [Pac74]. Vizcaya [Gli75a]. Vlack [Con76a]. Vocabulary [Bac75]. Vocational [De 74f]. Vocationalism [DeV75b]. Vogel [Har74a]. voiny [Lie74b, Lie76a]. Vol [All79b, Bra79a, Bri79, Bri74, Bri77b, Bri78b, BR74, Cah76, Cam74, Dun76, Har73a, Hin71b, Kel70a, Maj75, Maj77, Mer76a, Nee79, Rea78, San74, Sha72b, tZS73b, Tur75, Wal73b, Wer79]. Volta [Wes77]. Voltage [Hug76b]. Volume [Ano70c, Ano70h, Ano70g, Ano70i, Ano70j, Ano71e, Ano71h, Ano71g, Ano71i, Ano71j, Ano72f, Ano72i, Ano72h, Ano72j, Ano72k, Ano73m, Ano73s, Ano73t, Ano73r, Ano73u, Ano73v, Ano74h, Ano74k, Ano74j, Ano74l, Ano74m, Ano75h, Ano75k, Ano75j, Ano75l, Ano75m, Ano76f, Ano76i, Ano76h, Ano76j, Ano76k, Ano77g, Ano77j, Ano77i, Ano77k, Ano77l, Ano78f, Ano78j, Ano78i, Ano78k, Ano78l, Ano79d, Ano79h, Ano79g, Ano79i, Ano79j, Bry77a, Str78]. Voprosy [Gra75]. vorindustriellen [Fuk75b]. Vorkommen [Maj74]. Vucinich [Bla72]. Vunamami [PS71b].

W [Bal77, Bar71a, Bat70a, Bay74a, Bay74b, Bay78b, Bee76, Bez79, Bis70, Bis71, Bru76, Cal71a, Caw77, Ces70, Cha77, Con72c, Con74g, Con75d, DeG75, DeV75b, Dra70, Dra72, Eis76, Emm79, Fel76b, Fer73a, Fer74b, For79a, Fra76, Goi70, Gol70b, Had77c, Har78b, Hau76, Hei77b, Hin72, Hol79, Hud79, Jer70a, Jer78a, Kop77, Kre76, Lak70, Lay70a, Lay71a, Le 72, Le 75, Mar72a, Mas73, McC70b, McL79, Min78, Mor72a, Nel79, Par72a, Pei72, Por76, Ran72, Rey74, Rol74, Ros75a, San73c, Sch78b, Sha72a, Sky73, Tav72a, Wei71, Wel71, Wer70, Wol78, dC70, Mos79, Pia70, Pia71]. W. [Vin79]. Waddington [Vac71]. Wagar [Ran72, San73c]. Waggons [Ber70b]. Wagoner [Cal71b]. Wajdowicz [Mos75b]. Walenty [Tyl77]. Wales [Cow75b, Hud75a, Ber70b, Fus73, Rol70b]. Walker [Con74e]. Walkowitz [Ste75a]. Wallace [Con72c, Hin79]. Wallach [Ste76a]. Wallenstein [Rey79a]. Wallis [Sin72]. Walter [Bla76b, Cra76a, Emm78, Fre71b, Gie78, Jer70b, Nie70a, Row76b]. Walther [Süs76a]. Walton [Ehr73]. Wansey [Hin71a]. War [All79b, Bra79a, Bri79, Bru76, Buc75a, How78, Kev71, Lie74b, Lie76a, Lie78a, McM76b, Rey79a, Hac73a]. Ward [Gha70b, Gha73b, Gli73b, Hig79a, Pos76c, War73]. Warehouse [MCA74]. Warfare [Tho76, Web74, Woo79c, Sha76c, DeG77]. Warhol [Jen73c]. Warner [Ces70, Wol78, Jen70]. Warnow [Wil71b]. Warren [Ran72, San73c, Sto70]. Wartime [Per78a]. Was [Rol78]. Washington [HRM^{+}76, Ano70k, BPV^{+}78, RHTH73]. Waste [Tar79]. Watch [Par74].

Watches [Hok79a, Hok79b]. Water [Bee70, Bur79, Bur74a, Cai74, Hil74, Tar79, Car77a, Bis76b, Fra74b]. Water-lifting [Hil74]. Watercraft [Spi78]. Watering [Cai72]. Waterloo [Hil76]. Waterways [McC70a]. Waterwheel [Gad71, Rey79b]. Watkins [Car72a, Gol78]. Watson [Row72a]. Watsons [Süs71]. Watt [Fer70a, Law71, Pur71, Rob72, Rob71]. Watts [Dau72]. Wayne [Han79b, Hec73, Koh77]. Ways [Hac73b]. Wealth [Pau71]. Weapons [All79b, Gol70b, Hac77, Nie70b, Sha72b, Sto70]. Weavers [Lea70, dLM73]. Weaving [Jer70b]. Webb [Pos75a]. Weber [Kui78]. Webmaschine [Fre73]. Webster [Ber77, Noe77]. Wedderburn [Gue73]. Wedlake [Fus71]. Weeks [Sha76c]. Weg [Son78]. Wege [Hac73b]. Wegiel [Pia71]. Wehmer [Hal72b]. Weights [Nut79b]. Weiher [Son78]. Weintraub [Swa70]. Weirs [GD76]. Weisbecker [Por76]. Welfare [Aga77, Fre71a, Led77]. Wells [Bil79]. Wentz [Nie70a]. Werk [Hal72d]. Werkbund [Mar79a]. Werner [Hal76, Tho75b]. Wernher [Sha78c, Tri77]. West [Bac75, Bee75, Bla73b, Buc79, Car75b, De 74b, Emm71, Fai71, Fra71, Hit76, Jac75b, dLM76, McL70, Rob76a, Sol76b, Süs70, Uch75, You72, Hac77, Ret70, Smi76a, Sta77b, Str78, De 75b]. Western [Car72c, Fer73c, Har76a, Kap78, Mil76, Rav70, Str75, Lie73d, Lie73d, Lie74f]. Westwards [dC74b]. Wet [Gor74]. Wet-processing [Gor74]. Whaling [Kap78]. Wheel [Min78, Jer78a]. wheeled [Nel72]. Wheeler [Seg79b]. Wheeling [Sum75]. Wheels [Str71a, Hil74, Sum75]. Wheelwrights [Rob75b]. Whereby [Ras76]. Which [Hey74]. Whirlwind [Smi76e]. White [All79c, Bri78a, Con74a, Con77f, Con79f, Jen73b, Pos76a, Rav70, Seg77a, Str78, Edw72, Fla75, Fus72, Fus76a]. Whiteman [Abb72a]. Whitney [Bat73, Woo77a, Sta78c]. Whitrow [Bri78b, Jen73a]. Whittick [Con75c]. Whitworth [Sin72]. Who [Ell77]. Whose [Ada74, Siv72]. Wiener [Whi71c]. Wiezenbaum [Aga76]. Wik [Fit73a]. Wilbert [Cal71a]. Wilcox [Wil71a]. Wild [Moe71]. Wilfred [Wei76]. Wilhelm [Geb71, Mag76]. Wilkinson [Smi77b]. Willard [Woo79b]. Willem [Bal74]. William [Bee75, Bri71b, Buc73, Byr78a, Con72d, Con75g, Con77e, Con78c, Cow75a, Dot75, Eze77, Fer71c, Fit73b, Flo72, Fre71a, Gor71a, Gue77, Hea75, Hel71, Hil78, Jen76, Joh73, Kel71c, Kel71d, Lan73, Lie74a, Mar76a, McM77, Mul74b, Rae73c, Rob76c, Ros73a, Sal71, San72, Sch76, Sch77b, Seg77b, Sew75, She78b, Sko76, Slo72, Smi77d, Süs71, Swe78b, Tuc73, Wag73, Whi70, Woo79a, Bir76, Had74b, Koh71b]. Williams [Bed76, Hil71, McL78]. Williamsburg [Cot70]. Williamson [Dru76]. Wilmington [Bay75]. Wilson [Bri73, She78b]. Wiltshire [Tan78]. Wind [Wes77]. Wind-Catchers [Wes77]. Windmills [Wes77, Kel77]. Windmühlen [Maj74]. Wings [Smi77d]. Winner [Car78a]. Winston [Whi71b, Whi71b]. Winterthur [Mei74, Gar72b, Mar76b, Wel73b]. Wire [Pur75]. Wireless [Sha76b]. Wirtschaft [Fuk75b]. wirtschaftlicher [Fuk75b]. Wishing [Sko76]. Wislicki [Mos79]. Wissenschaft [Har71c]. within [Ehr74]. without [Mat75, Seg78, Str71a]. Wittlin [Smi77c]. WMO [Bis76a]. Wolek [Pei72]. Wolff [Con74i]. Wolfgang [Smi75a, Wat78a].

Woman [Cow75a, Rog75]. Women [Cow75b, Cow79, Rog75, Ros79d, Tuc73]. Wonder [Nee79]. Wood [Bur79, Con74d, Qui74]. Woodbury [Sta78c]. Woodcroft [Kre74]. Wooden [Bri77a, Jud74a, McM76a, Wel77]. Woodruff [Rob76c]. Woods [Rop76b]. Woodward [Buc75a, Buc75a, Qui77]. Wool [Hea72, Hit76, dLM76]. Woolen [Bee75]. Woollen [Tan78]. woon [Hen76a]. Word [Ada74]. Work [Aga70, Bal73b, Ben78, Car76, Cow75b, Ehr74, Fri79b, Gie77, Har76a, Lie75e, Mol79, Siv72, Sky70, Kel71d, Bir76, Dib75, Gie79, Mul72a, Nel79, Pac78, Tuc73, Aga78, Gar78, Seg78, Wei72]. Worked [dC79, Cra79b]. Worker [De 73b, tzS75, Gie76c, Gue73]. Workers [Jam76, Sau72, Ste75a]. Working [Bal75a, Bat70b, Cow75b, Gie76b, Hou77, Van78b, Ste76b]. Working-Class [Van78b, Ste76b]. Workplace [Gie76a]. Works [Ano73g, Emm77, Hin71b, Lie75c, Orł73, Sta77a, Alt70]. Workshop [McG76, Ung72, Rog75, Buc79]. World [All79b, Bar73b, Bat79, Bee73, Bri76a, Bri79, Buc75a, Con70b, Con79b, Dre75, GD73, Hau76, Hel77, Hen76b, Hou77, Laf72, Lie76a, MN73, McM76b, Nor75, Rae79b, Ras75, Süs76a, Whi75a, Con74i, Kev71, Cha78a, Fow78, Fus76a, Har76a, Rob76c, She77, Wal71b, Woo77b, dC78, dc72b, Fit71]. Worlds [Van78b]. Worn [Sta78d]. Worn-out [Sta78d]. Worshipful [Rob75b]. Wrenn [Cra78]. Wright [Con72e, Con79a, Seg79a, Bed71, Per74]. Writings [Sti77]. Wykres [Pia71]. Wyndham [Bee77].

XIV [Kel70a]. XIVth [Ano75n]. XIX [Day74]. XIXth [CZ73]. XV [Rap75]. XVI [Kel70a]. XVII [Rap75]. XVIII [Rob71].

Yakovlev [Bra72]. Yankee [May75b, Hok77]. Yankees [Sta78b]. Yearbook [Chu79, Mor77c]. Years [Bra77a, Bra72, Con73d, Con74e, Con75d, HS70, Hok79a, Kel73, Lei70, Nes77, Rea78, Sha78a, Woo77b, Lie74b, Lie75c, Lie76a, Lie78a, All79b, Bra77c, Bri74, Bri77b, Süs70]. Yesterday [Wes77, Hud79, Lak70]. Yoder [Hey74]. York [Con72a, Con76a, Cra75, LDS^{+}72, Slo70, Wel71, San77, Con76d]. Young [Com74, Fus74b, Koh76b]. Yucatan [Wal76b]. Yves [Wei72].

Z [Tuc73, Mos78]. za [Lie75c]. Zagorsky [Par71b]. Zaki [Nee70]. Zandi [Mat76]. zapisu [Mos75b]. zavoda [Lie75c]. Zealand [DeG75]. Zeitmesser [Bed71]. Zeitschrift [May78b]. Zemon [Kel76]. Zen [Sta78d]. Zheleznodorozhnyi [Lie78a]. zhurnal [Car77b]. Ziman [Ols77]. Zimmerman [Hal72d]. Zouche [Can74]. zum [Hal72d, Rob76d]. Zupko [Nut79b]. zur [Bry70, Hac73b, Son78, Tro72, Har70]. zwischen [Hal72c, Rob76d]. Zygmunta [Cas76a].

References

Abbott:1972:CAH

[Abb72a] Collamer M. Abbott. Copper for America: The Hendricks Family and a National Industry, 1755–1939 by Maxwell Whiteman (review). Technology and Culture, 13(3):503–505, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893939/pdf.

Abbott:1972:CAF

[Abb72b] Collamer M. Abbott. The Cliff: America’s First Great Copper Mine by Donald Chaput (review). Technology and Culture, 13(4):649–650, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893394/pdf.

Abrams:1978:GAS

[Abr78] John W. Abrams. General assembly and symposium, Edinburgh, August 1977. Technology and Culture, 19(3):600–601, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892161/pdf.

Abrams:1979:TBA

[Abr79] John W. Abrams. Technologia Bruxellensis by L’Association des Ingenierus Diplomes de L’Enseignement Superior Technique de L’Etat de Bruxelles (review). Technology and Culture, 20(2):397, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890518/pdf.

Adams:1974:BMH

[Ada74] J. Michael Adams. Books for the Millions: a History of the Men Whose Methods and Machines Packaged the Printed Word by Frank E. Comparato (review). Technology and Culture, 15(3):503–505, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892755/pdf.

Adrosko:1970:PGR

[Adr70] Rita J. Adrosko. The Plictho of Gioanventura Rosetti ed. by Sidney M. Edelstein, Hector C. Borghetty (review). Technology and Culture, 11(4):623–625, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894053/pdf.

Agassi:1970:MII

[Aga70] Judith B. Agassi. Man Incorporate: The Individual and His Work in an Organized Society by Carl B. Kaufman (review). Technology and Culture, 11(4):649–652, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894069/pdf.

Agassi:1976:CPH

[Aga76] Joseph Agassi. Computer Power and Human Reason: From Judgment to Calculation by Joseph Wiezenbaum (review). Technology and Culture, 17(4):813–816, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891859/pdf.

Agassi:1977:EWT

[Aga77] Joseph Agassi. Evolution, Welfare and Time in Economics: Essays in Honor of Nicholas Georgescu-Roegen ed. by Anthony M. Tang, et al. (review). Technology and Culture, 18(3):577–578, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891983/pdf.

Agassi:1978:WTE

[Aga78] Judith Buber Agassi. Work and Technology ed. by Marie R. Haug, Jacques Dofny (review). Technology and Culture, 19(3):585–589, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892156/pdf.

Austin:1979:PMO

[AGS^{+}79] James B. Austin, Martha Goodway, G. Terry Sharrer, Joel A. Tarr, A. Hunter Dupree, Douglas Gomery, Brooke Hindle, John F. Kasson, Patrick M. Malone, Michael Brewster Folsom, J. Z. Fullmer, and Hugo A. Meier. The Pittsburgh Meeting, October 18–21, 1978. Technology and Culture, 20(3):590–613, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890529/pdf.

Aitken:1979:HHM

[Ait79] Hugh G. J. Aitken. Heinrich Hertz: Memoirs, Letters, Diaries ed. by Johanna Hertz (review). Technology and Culture, 20(2):388–389, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890514/pdf.

Allan:1974:CME

[All74] D. G. C. Allan. Cornish Mining: Essays on the Organization of Cornish Mines and the Cornish Mining Economy ed. by Roger Burt (review). Technology and Culture, 15(4):633–634, October 1974. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892709/pdf.

Allan:1979:SCS

[All79a] D. G. C. Allan. Social Change and Scientific Organisation: The Royal Institution, 1799–1844 by Morris Berman (review). Technology and Culture, 20(1):211–212, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890468/pdf.

Allison:1979:HNW

[All79b] David K. Allison. History of the Naval Weapons Center, China Lake, California. Vol. 2: The Grand Experiment at Inyokern: Narrative of the Naval Ordnance Test Station during the Second World War and the Immediate Postwar Years by J. D. Gerrard-Gough, Albert B. Christman (review). Technology and Culture, 20(3):674– 675, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890560/pdf.

Allison:1979:HWO

[All79c] David K. Allison. History of the White Oak Laboratory, 1945– 1975 by Joseph P. Smaldone (review). Technology and Culture, 20 (4):824–825, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890591/pdf.

Althin:1970:HSA

[Alt70] Torsten Althin. The History of the Söderfors Anchor-Works by Johan Lundström (review). Technology and Culture, 11(4):625–627, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894054/pdf.

Althin:1973:GIA

[Alt73a] Torsten Althin. A Guide to the Industrial Archaeology of Europe by Kenneth Hudson (review). Technology and Culture, 14(2 (part I)):352–354, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893293/pdf.

Althin:1973:PPC

[Alt73b] Torsten Althin. Prints and Printing in Color by Dr. Ulla Ehrensvärd (review). Technology and Culture, 14(3):415–417, July 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893211/pdf.

Anderson:1974:ESI

[And74] R. G. W. Anderson. Early Scientific Instruments comp. by R. H. Nuttall (review). Technology and Culture, 15(4):673–674, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892729/pdf.

Anonymous:1970:APU

[Ano70a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 11(2):201–202, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894149/pdf.

Anonymous:1970:DP

[Ano70b] Anonymous. The Dexter prize. Technology and Culture, 11(2):203– 204, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894150/pdf.

Anonymous:1970:IVT

[Ano70c] Anonymous. Index to volume 11 (1970) of Technology and Culture. Technology and Culture, 11(4):673–686, October 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894078/pdf.

Anonymous:1970:LVM

[Ano70d] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 11(2):205–213, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894151/pdf.

Anonymous:1970:TAM

[Ano70e] Anonymous. Thirteenth annual meeting of SHOT. Technology and Culture, 11(4):669–671, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894077/pdf.

Anonymous:1970:TIC

[Ano70f] Anonymous. Thirteenth international Congress of the history of science. Technology and Culture, 11(3):490–492, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894141/pdf.

Anonymous:1970:VNJa

[Ano70g] Anonymous. Volume 11, number 1, January 1970. Technology and Culture, 11(1):??, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1970:VNA

[Ano70h] Anonymous. Volume 11, number 2, April 1970. Technology and Culture, 11(2):??, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1970:VNJb

[Ano70i] Anonymous. Volume 11, number 3, July 1970. Technology and Culture, 11(3):??, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1970:VNO

[Ano70j] Anonymous. Volume 11, number 4, October 1970. Technology and Culture, 11(4):??, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1970:WMD

[Ano70k] Anonymous. The Washington Meeting, December 26–30,1969. Technology and Culture, 11(2):214–236, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/894152/pdf.

Anonymous:1971:APU

[Ano71a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 12(2):231–232, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893958/pdf.

Anonymous:1971:Aa

[Ano71b] Anonymous. Announcements. Technology and Culture, 12(3):539– 541, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893623/pdf.

Anonymous:1971:Ab

[Ano71c] Anonymous. Announcements. Technology and Culture, 12(4): 688–695, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893590/pdf.

Anonymous:1971:DP

[Ano71d] Anonymous. The Dexter prize. Technology and Culture, 12(2):233– 234, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893959/pdf.

Anonymous:1971:IVT

[Ano71e] Anonymous. Index to volume 12 (1971) of Technology and Culture. Technology and Culture, 12(4):697–711, October 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893591/pdf.

Anonymous:1971:LVM

[Ano71f] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 12(2):235–237, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893960/pdf.

Anonymous:1971:VNJa

[Ano71g] Anonymous. Volume 12, number 1, January 1971. Technology and Culture, 12(1):??, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1971:VNA

[Ano71h] Anonymous. Volume 12, number 2, April 1971. Technology and Culture, 12(2):??, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1971:VNJb

[Ano71i] Anonymous. Volume 12, number 3, July 1971. Technology and Culture, 12(3):??, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1971:VNO

[Ano71j] Anonymous. Volume 12, number 4, October 1971. Technology and Culture, 12(4):??, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1972:APU

[Ano72a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 13(3):430–431, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893926/pdf.

Anonymous:1972:Aa

[Ano72b] Anonymous. Announcements. Technology and Culture, 13(2):346– 351, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893495/pdf.

Anonymous:1972:Ab

[Ano72c] Anonymous. Announcements. Technology and Culture, 13(3):527– 529, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893952/pdf.

Anonymous:1972:Ac

[Ano72d] Anonymous. Announcements. Technology and Culture, 13(4): 693–697, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893418/pdf.

Anonymous:1972:DP

[Ano72e] Anonymous. The Dexter prize. Technology and Culture, 13(3):432– 433, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893927/pdf.

Anonymous:1972:IVT

[Ano72f] Anonymous. Index to volume 13 (1972) of Technology and Culture. Technology and Culture, 13(4):699–711, October 1972. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893419/pdf.

Anonymous:1972:LVM

[Ano72g] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 13(3):434–439, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893928/pdf.

Anonymous:1972:VNJa

[Ano72h] Anonymous. Volume 13, number 1, January 1972. Technology and Culture, 13(1):??, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1972:VNA

[Ano72i] Anonymous. Volume 13, number 2, April 1972. Technology and Culture, 13(2):??, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1972:VNJb

[Ano72j] Anonymous. Volume 13, number 3, July 1972. Technology and Culture, 13(3):??, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1972:VNO

[Ano72k] Anonymous. Volume 13, number 4, October 1972. Technology and Culture, 13(4):??, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1973:APU

[Ano73a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 14(3):418–420, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893212/pdf.

Anonymous:1973:Aa (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893375/pdf.

Anonymous:1973:Ab

[Ano73c] Anonymous. Announcements. Technology and Culture, 14(2 (part I)):361–365, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893297/pdf.

Anonymous:1973:Ac

[Ano73d] Anonymous. Announcements. Technology and Culture, 14(3):533– 534, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893245/pdf.

Anonymous:1973:Ad

[Ano73e] Anonymous. Announcements. Technology and Culture, 14(4): 685–688, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892911/pdf.

Anonymous:1973:DP

[Ano73f] Anonymous. The Dexter prize. Technology and Culture, 14(3):421– 422, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893213/pdf.

Anonymous:1973:CPW

[Ano73g] Anonymous. I. Comprehensive philosophical works. Technology and Culture, 14(2 (part II)):S1–S20, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893246/pdf.

Anonymous:1973:IEPa

[Ano73h] Anonymous. II. Ethical and political critiques: A. Primary sources. Technology and Culture, 14(2 (part II)):S20–S47, April 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893247/pdf.

Anonymous:1973:IEPb

[Ano73i] Anonymous. II. Ethical and political critiques: B. Secondary sources. Technology and Culture, 14(2 (part II)):S47–S105, April 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893248/pdf.

Anonymous:1973:IEPc

[Ano73j] Anonymous. II. Ethical and political critiques: C. Appendix: Soviet and East European materials. Technology and Culture, 14(2 (part II)):S105–S123, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893249/pdf.

Anonymous:1973:IRCa

[Ano73k] Anonymous. III. Religious critiques: A. Primary sources. Technology and Culture, 14(2 (part II)):S123–S130, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893250/pdf.

Anonymous:1973:IRCb

[Ano73l] Anonymous. III. Religious critiques: B. Secondary sources. Technology and Culture, 14(2 (part II)):S130–S141, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893251/pdf.

Anonymous:1973:IVT

[Ano73m] Anonymous. Index to volume 14 (1973) of Technology and Culture. Technology and Culture, 14(4):689–701, October 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892912/pdf.

Anonymous:1973:IME

[Ano73n] Anonymous. IV. Metaphysical and epistemological studies. Technology and Culture, 14(2 (part II)):S141–S184, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893252/pdf.

Anonymous:1973:LVM

[Ano73o] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 14(3):423–428, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893214/pdf.

Anonymous:1973:VAC

[Ano73p] Anonymous. V. Appendix: A. Classical documents. Technology and Culture, 14(2 (part II)):S184–S188, April 1973. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893253/pdf.

Anonymous:1973:VAB

[Ano73q] Anonymous. V. Appendix: B. Background materials. Technology and Culture, 14(2 (part II)):S188–S205, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893254/pdf.

Anonymous:1973:VNJa

[Ano73r] Anonymous. Volume 14, number 1, January 1973. Technology and Culture, 14(1):??, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1973:VNAa

[Ano73s] Anonymous. Volume 14, number 2, April 1973, part I. Technology and Culture, 14(2 (part I)):??, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1973:VNAb

[Ano73t] Anonymous. Volume 14, number 2, April 1973, part II. Technology and Culture, 14(2 (part II)):??, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1973:VNJb

[Ano73u] Anonymous. Volume 14, number 3, July 1973. Technology and Culture, 14(3):??, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1973:VNO

[Ano73v] Anonymous. Volume 14, number 4, October 1973. Technology and Culture, 14(4):??, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1974:APU

[Ano74a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 15(3):443–444, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892740/pdf.

Anonymous:1974:Aa (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892865/pdf.

Anonymous:1974:Ab

[Ano74c] Anonymous. Announcements. Technology and Culture, 15(2):370– 372, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892819/pdf.

Anonymous:1974:Ac

[Ano74d] Anonymous. Announcements. Technology and Culture, 15(3):541– 542, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892775/pdf.

Anonymous:1974:Ad

[Ano74e] Anonymous. Announcements. Technology and Culture, 15(4): 678–681, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892732/pdf.

Anonymous:1974:DP

[Ano74f] Anonymous. The Dexter prize. Technology and Culture, 15(3):445– 446, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892741/pdf.

Anonymous:1974:DMS

[Ano74g] Anonymous. Directory of members: The Society for the History of Technology. Technology and Culture, 15(1):1A–27A, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892866/pdf.

Anonymous:1974:IVT

[Ano74h] Anonymous. Index to volume 15 (1974) of Technology and Culture. Technology and Culture, 15(4):683–697, October 1974. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892733/pdf.

Anonymous:1974:LVM

[Ano74i] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 15(3):447–452, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892742/pdf.

Anonymous:1974:VNJa

[Ano74j] Anonymous. Volume 15, number 1, January 1974. Technology and Culture, 15(1):??, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1974:VNA

[Ano74k] Anonymous. Volume 15, number 2, April 1974. Technology and Culture, 15(2):??, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1974:VNJb

[Ano74l] Anonymous. Volume 15, number 3, July 1974. Technology and Culture, 15(3):??, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1974:VNO

[Ano74m] Anonymous. Volume 15, number 4, October 1974. Technology and Culture, 15(4):??, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1975:APU

[Ano75a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 16(3):419–420, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891564/pdf.

Anonymous:1975:Aa

[Ano75b] Anonymous. Announcements. Technology and Culture, 16(1): 129–131, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891527/pdf.

Anonymous:1975:Ab

[Ano75c] Anonymous. Announcements. Technology and Culture, 16(2):335, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891559/pdf.

Anonymous:1975:Ac

[Ano75d] Anonymous. Announcements. Technology and Culture, 16(3):516– 518, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891594/pdf.

Anonymous:1975:Ad

[Ano75e] Anonymous. Announcements. Technology and Culture, 16(4): 711–718, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891647/pdf.

Anonymous:1975:DP

[Ano75f] Anonymous. The Dexter prize. Technology and Culture, 16(3):421– 423, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891565/pdf.

Anonymous:1975:I

[Ano75g] Anonymous. Icohtec. Technology and Culture, 16(2):192–194, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891533/pdf.

Anonymous:1975:IVT

[Ano75h] Anonymous. Index to volume 16 (1975) of Technology and Culture. Technology and Culture, 16(4):719–732, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891648/pdf.

Anonymous:1975:LVM

[Ano75i] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 16(3):424–428, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891566/pdf.

Anonymous:1975:VNJa

[Ano75j] Anonymous. Volume 16, number 1, January 1975. Technology and Culture, 16(1):??, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1975:VNA

[Ano75k] Anonymous. Volume 16, number 2, April 1975. Technology and Culture, 16(2):??, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1975:VNJb

[Ano75l] Anonymous. Volume 16, number 3, July 1975. Technology and Culture, 16(3):??, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1975:VNO

[Ano75m] Anonymous. Volume 16, number 4, October 1975. Technology and Culture, 16(4):??, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1975:XIC

[Ano75n] Anonymous. XIVth international Congress of the history of science. Technology and Culture, 16(2):191–192, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891532/pdf.

Anonymous:1976:APU

[Ano76a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 17(3):489–490, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891756/pdf.

Anonymous:1976:Aa

[Ano76b] Anonymous. Announcements. Technology and Culture, 17(1): 197–199, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891701/pdf.

Anonymous:1976:Ab

[Ano76c] Anonymous. Announcements. Technology and Culture, 17(2):418– 421, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891749/pdf.

Anonymous:1976:Ac

[Ano76d] Anonymous. Announcements. Technology and Culture, 17(3):616– 619, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891812/pdf.

Anonymous:1976:DP

[Ano76e] Anonymous. The Dexter prize. Technology and Culture, 17(3):491– 492, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891757/pdf.

Anonymous:1976:IVT

[Ano76f] Anonymous. Index to volume 17 (1976) of Technology and Culture. Technology and Culture, 17(4):817–834, October 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891860/pdf.

Anonymous:1976:LVM

[Ano76g] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 17(3):493–498, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891758/pdf.

Anonymous:1976:VNJa

[Ano76h] Anonymous. Volume 17, number 1, January 1976. Technology and Culture, 17(1):??, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1976:VNA

[Ano76i] Anonymous. Volume 17, number 2, April 1976. Technology and Culture, 17(2):??, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1976:VNJb

[Ano76j] Anonymous. Volume 17, number 3, July 1976. Technology and Culture, 17(3):??, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1976:VNO

[Ano76k] Anonymous. Volume 17, number 4, October 1976. Technology and Culture, 17(4):??, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1977:APU

[Ano77a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 18(3):467–468, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891940/pdf.

Anonymous:1977:Aa

[Ano77b] Anonymous. Announcements. Technology and Culture, 18(1): 154–156, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891907/pdf.

Anonymous:1977:Ab

[Ano77c] Anonymous. Announcements. Technology and Culture, 18(2):393– 394, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891936/pdf.

Anonymous:1977:Ac

[Ano77d] Anonymous. Announcements. Technology and Culture, 18(3):591– 592, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891989/pdf.

Anonymous:1977:Ad

[Ano77e] Anonymous. Announcements. Technology and Culture, 18(4): 752–754, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892033/pdf.

Anonymous:1977:DP

[Ano77f] Anonymous. The Dexter prize. Technology and Culture, 18(3):469– 470, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891941/pdf.

Anonymous:1977:IVT

[Ano77g] Anonymous. Index to volume 18 (1977) of Technology and Culture. Technology and Culture, 18(4):755–772, October 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892034/pdf.

Anonymous:1977:LVM

[Ano77h] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 18(3):471–478, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891942/pdf.

Anonymous:1977:VNJa

[Ano77i] Anonymous. Volume 18, number 1, January 1977. Technology and Culture, 18(1):??, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1977:VNA

[Ano77j] Anonymous. Volume 18, number 2, April 1977. Technology and Culture, 18(2):??, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1977:VNJb

[Ano77k] Anonymous. Volume 18, number 3, July 1977. Technology and Culture, 18(3):??, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1977:VNO

[Ano77l] Anonymous. Volume 18, number 4, October 1977. Technology and Culture, 18(4):??, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1978:APU

[Ano78a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 19(3):461–462, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892113/pdf.

Anonymous:1978:Aa

[Ano78b] Anonymous. Announcements. Technology and Culture, 19(1): 148–149, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892066/pdf.

Anonymous:1978:Ab

[Ano78c] Anonymous. Announcements. Technology and Culture, 19(2):357– 359, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892107/pdf.

Anonymous:1978:Ac

[Ano78d] Anonymous. Announcements. Technology and Culture, 19(4): 807–808, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892218/pdf.

Anonymous:1978:DP

[Ano78e] Anonymous. The Dexter prize. Technology and Culture, 19(3):463– 464, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892114/pdf.

Anonymous:1978:IVT

[Ano78f] Anonymous. Index to volume 19 (1978) of Technology and Culture. Technology and Culture, 19(4):809–827, October 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892219/pdf.

Anonymous:1978:LVM

[Ano78g] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 19(3):465–471, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892115/pdf.

Anonymous:1978:SAM

[Ano78h] Anonymous. SHOT annual meeting, 1978. Technology and Culture, 19(3):612–614, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892167/pdf.

Anonymous:1978:VNJa

[Ano78i] Anonymous. Volume 19, number 1, January 1978. Technology and Culture, 19(1):??, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1978:VNA

[Ano78j] Anonymous. Volume 19, number 2, April 1978. Technology and Culture, 19(2):??, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1978:VNJb

[Ano78k] Anonymous. Volume 19, number 3, July 1978. Technology and Culture, 19(3):??, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1978:VNO

[Ano78l] Anonymous. Volume 19, number 4, October 1978. Technology and Culture, 19(4):??, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1979:APU

[Ano79a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 20(3):578–579, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890526/pdf.

Anonymous:1979:A

[Ano79b] Anonymous. Announcements. Technology and Culture, 20(1):256, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890983/pdf.

Anonymous:1979:DP

[Ano79c] Anonymous. The Dexter prize. Technology and Culture, 20(3):580– 582, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890527/pdf.

Anonymous:1979:IVT

[Ano79d] Anonymous. Index to volume 20 (1979) of Technology and Culture. Technology and Culture, 20(4):861–877, October 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890601/pdf.

Anonymous:1979:LVM

[Ano79e] Anonymous. The Leonardo Da Vinci Medal. Technology and Culture, 20(3):583–589, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890528/pdf.

Anonymous:1979:SAM

[Ano79f] Anonymous. SHOT annual meeting, 1979. Technology and Culture, 20(3):693–695, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890984/pdf.

Anonymous:1979:VNJa

[Ano79g] Anonymous. Volume 20, number 1, January 1979. Technology and Culture, 20(1):??, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1979:VNA

[Ano79h] Anonymous. Volume 20, number 2, April 1979. Technology and Culture, 20(2):??, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1979:VNJb

[Ano79i] Anonymous. Volume 20, number 3, July 1979. Technology and Culture, 20(3):??, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1979:VNO

[Ano79j] Anonymous. Volume 20, number 4, October 1979. Technology and Culture, 20(4):??, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Arnott:1970:AOD

[Arn70] Margaret L. Arnott. Agricultural Origins and Dispersals: The Domestication of Animals and Foodstuffs by Carl O. Sauer (review). Technology and Culture, 11(1):94–95, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894194/pdf.

Artz:1970:SFR

[Art70] Frederick B. Artz. Science in France in the Revolutionary Era, Described by Thomas Bugge ed. by M. P. Crosland (review). Technology and Culture, 11(3):434, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894108/pdf.

Ashley:1971:CME

[Ash71] Patricia Ashley. The Computer and Music ed. by Harry B. Lincoln (review). Technology and Culture, 12(2):384–386, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893992/pdf.

Austin:1974:CAE

[Aus74] John C. Austin. Ceramics in America ed. by Ian M. G. Quimby (review). Technology and Culture, 15(3):493–494, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892749/pdf.

Austin:1976:HSJ

[Aus76] James B. Austin. History of Steel in Japan ed. by Eiho Nishida (review). Technology and Culture, 17(3):559, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891782/pdf.

Austin:1977:SSK

[Aus77] James B. Austin. Sheffield Steel by K. C. Barraclough (review). Technology and Culture, 18(2):244–247, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891919/pdf.

Bachrach:1975:EMF

[Bac75] Bernard S. Bachrach. Early medieval fortifications in the “West” of France: a revised technical vocabulary. Technology and Culture, 16(4):531–569, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891596/pdf.

Bachrach:1979:FMT

[Bac79] Bernard S. Bachrach. Fortifications and military tactics: Fulk Nerra’s strongholds circa 1000. Technology and Culture, 20(3):531– 549, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890523/pdf.

Bagchi:1975:STM

[Bag75] Samar K. Bagchi. Science and technology museums in India. Technology and Culture, 16(4):607–614, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891599/pdf.

Bagwell:1979:SNA

[Bag79] Philip S. Bagwell. Servants of the North: Adventures on the Coastal Trade with the Northern Steam Ship Company by Cliff Furniss (review). Technology and Culture, 20(4):804–805, October 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890583/pdf.

Bailes:1974:LTP

[Bai74] Kendall E. Bailes. Leon Trotsky and the Politics of Economic Isolation by Richard B. Day (review). Technology and Culture, 15 (4):652–655, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892720/pdf.

Bailes:1976:TLS

[Bai76] K. E. Bailes. Technology and legitimacy: Soviet aviation and Stalinism in the 1930s. Technology and Culture, 17(1):55–81, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891651/pdf.

Baker:1972:RFR

[Bak72] Gladys L. Baker. The Reluctant Farmer: The Rise of Agricultural Extension to 1914 by Roy V. Scott (review). Technology and Culture, 13(4):642–644, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893391/pdf.

Baker:1976:SCA

[Bak76] William A. Baker. Shipbuilding in Colonial America by Joseph A. Goldenberg (review). Technology and Culture, 17(4):750–751, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891826/pdf.

Ballon:1971:ATN

[Bal71] Robert J. Ballon. The acquisition of technology by non-initiating countries: ICOHTEC Symposium, Pont-à-Mousson, June 29–July 4, 1970. Technology and Culture, 12(2):217–227, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893956/pdf.

Ball:1973:PRO

[Bal73a] Norman R. Ball. Philosophy of Railroads and Other Essays by T. C. Keefer ed. by H. V. Nelles (review). Technology and Culture, 14 (4):632–635, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892882/pdf.

Ball:1973:RWC

[Bal73b] Norman R. Ball. The Rules of Work of the Carpenters’ Company of the City and County of Philadelphia, 1786 ed. by Charles E. Peterson (review). Technology and Culture, 14(2 (part I)):300–301, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893270/pdf.

Ball:1974:DHP

[Bal74] Norman R. Ball. A Dutch Homesteader on the Prairies: The Letters of Willem De Gelder by Herman Ganzevoort (review). Technology and Culture, 15(3):499–501, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892752/pdf.

Ball:1975:PSC

[Bal75a] Norman R. Ball. Primary Sources in Canadian Working Class History, 1860–1930 by Russell G. Hann, et al. (review). Technology and Culture, 16(3):511–514, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891591/pdf.

Ball:1975:PDF

[Bal75b] Norman R. Ball. The Politics of Development: Forests, Mines, and Hydro-Electric Power in Ontario, 1849–1941 by H. V. Nelles (review). Technology and Culture, 16(4):667–670, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891625/pdf.

Ball:1977:THF

[Bal77] Norman R. Ball. Timber: History of the Forest Industry in British Columbia by Geoffrey W. Taylor (review). Technology and Culture, 18(3):573–574, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891980/pdf.

Balmer:1978:OSC

[Bal78] R. T. Balmer. The operation of sand clocks and their medieval development. Technology and Culture, 19(4):615–632, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892168/pdf.

Barden:1970:RCR

[Bar70a] John Barden. Run, Computer, Run: The Mythology of Educational Innovation by Anthony G. Oettinger (review). Technology and Culture, 11(2):337–338, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894181/pdf.

Barden:1970:TIE

[Bar70b] John Barden. Technology and Innovation in Education by the Aerospace Education Foundation (review). Technology and Culture, 11(1):129–130, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894209/pdf.

Barloon:1971:TAD

[Bar71a] Marvin J. Barloon. The Transportation Act of 1958: a Decade of Experience by George W. Hilton (review). Technology and Culture, 12(1):115–120, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894020/pdf.

Barrows:1971:TED

[Bar71b] Leland Barrows. Technology and the Economic Development of the Tropical African Frontier by Thomas R. De Gregori (review). Technology and Culture, 12(2):363–366, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893982/pdf.

Baram:1973:SCL

[Bar73a] Michael S. Baram. Southern California Law Review (review). Technology and Culture, 14(2 (part I)):343–345, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893289/pdf.

Barrows:1973:DTW

[Bar73b] Leland Barrows. Developing the Third World: The Experience of the Nineteen-sixties ed. by Ronald Robinson (review). Technology and Culture, 14(2 (part I)):347–351, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893291/pdf.

Battison:1970:HMT

[Bat70a] Edwin A. Battison. A History of Machine Tools, 1700–1910 by W. Steeds (review). Technology and Culture, 11(2):305–307, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894163/pdf.

Battison:1970:SWM

[Bat70b] Edwin A. Battison. Simple Working Models of Historic Machines (Easily Made by the Reader) by Aubrey F. Burstall (review). Technology and Culture, 11(3):478–480, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/894133/pdf.

Battison:1971:FUE

[Bat71] Edwin A. Battison. Factory under the Elms: a History of Harrisville, N.H., 1774–1969 by John Borden Armstrong (review). Technology and Culture, 12(4):644–646, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893568/pdf.

Battison:1973:NLW

[Bat73] Edwin A. Battison. A new look at the “Whitney” milling machine. Technology and Culture, 14(4):592–598, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892870/pdf.

Battison:1974:TID

[Bat74] Edwin A. Battison. “Technological Innovation and the Decorative Arts”: Hagley Museum, March 29, 1973–January 1, 1974. Technology and Culture, 15(2):242–245, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892794/pdf.

Battison:1977:IAN

[Bat77] Edwin A. Battison. Industrial Archeology — a New Look at the American Heritage by Theodore Anton Sande (review). Technology and Culture, 18(3):571–572, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891979/pdf.

Bates:1979:WCR

[Bat79] Seth Pardee Bates. The World That Could Be by Robert C. North (review). Technology and Culture, 20(4):822–824, October 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890590/pdf.

Bayor:1974:CBPa

[Bay74a] Ronald H. Bayor. Chicago, 1910–1929: Building, Planning, and Urban Technology by Carl W. Condit (review). Technology and Culture, 15(2):342–343, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892806/pdf.

Bayor:1974:CBPb

[Bay74b] Ronald H. Bayor. Chicago, 1930–70: Building, Planning, and Urban Technology by Carl W. Condit (review). Technology and Culture, 15(4):648–649, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892718/pdf.

Bayor:1975:WDP

[Bay75] Ronald H. Bayor. Wilmington, Delaware: Portrait of an Industrial City, 1830–1910 by Carol E. Hoffecker (review). Technology and Culture, 16(4):639–640, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891610/pdf.

Bayor:1976:CSS

[Bay76a] Ronald H. Bayor. Computers and the Social Sciences by Alan Brier, Ian Robinson (review). Technology and Culture, 17(3):580– 581, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891793/pdf.

Bayor:1976:AHT

[Bay76b] Ronald H. Bayor. The Atlas of Historic Towns ed. by M. D. Lobel (review). Technology and Culture, 17(4):787–788, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891844/pdf.

Bayor:1978:AVH

[Bay78a] Ronald H. Bayor. American Values and Habitat: a Research Agenda by Peter Bearse, et al. (review). Technology and Culture, 19(2):224– 226, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892083/pdf.

Bayor:1978:RCT

[Bay78b] Ronald H. Bayor. The Railroad and the City: a Technological and Urbanistic History of Cincinnati by Carl W. Condit (review). Technology and Culture, 19(4):747–748, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892188/pdf.

Becker:1978:MTS

[Bec78] Peter W. Becker. Militär und Technik: Schwedmfestungen des 17. Jahrhunderts in den Herzogtümern Bremen und Verden by Henning Eichberg (review). Technology and Culture, 19(3):554–556, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892143/pdf.

Bedini:1971:KAA

[Bed71] Silvio A. Bedini. Kunstuhrmacher in Alt-Augsburg by Maximilian Bobinger, and: Die Uhr, Zeitmesser und Schmuck in fünf Jahrhunderten by Alfred Leiter, Alma Helfrich-Dörner, and: Time and Clocks: a Description of Ancient and Modern Methods of Measuring Time by H. H. Cunynghame, and: Japanese Clocks by N. H. N. Mody, and: Clockwork Man: The Story of Time, Its Origins, Uses and Tyranny by Lawrence Wright (review). Technology and Culture, 12(3):490–493, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893601/pdf.

Bederman:1976:MSA

[Bed76] Sanford H. Bederman. The Making of the South Australian Landscape: a Study in the Historical Geography of Australia by Michael Williams (review). Technology and Culture, 17(2):408–410, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891743/
pdf.

Beer:1970:SWN

[Bee70] John J. Beer. Strong Water, Nitric Acid: Sources, Methods of Manufacture and Uses by Thomas H. Chilton (review). Technology and Culture, 11(3):458, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894121/pdf.

Beer:1973:CWM

[Bee73] Stafford Beer. This Cybernetic World of Men, Machines and Earth Systems by V. L. Parsegian (review). Technology and Culture, 14(2 (part I)):331–333, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893285/pdf.

Beer:1975:PTD

[Bee75] John J. Beer. A Practical Treatise on Dying of Woolen, Cotton, and Skein Silk with the Manufacture of Broadcloth and Cassimere including the Most Improved Methods in the West of England by William Partridge (review). Technology and Culture, 16(1):90–93, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891507/pdf.

Beer:1976:EAS

[Bee76] John J. Beer. The Emergence of Agricultural Science: Justus Liebig and the Americans, 1840–1880 by Margaret W. Rossiter (review). Technology and Culture, 17(4):755–758, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891829/pdf.

Beer:1977:ACC

[Bee77] John J. Beer. American Chemists and Chemical Engineers ed. by Wyndham D. Miles (review). Technology and Culture, 18(3):557– 558, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891971/pdf.

Beer:1979:FHS

[Bee79] John J. Beer. Faith, Hope and $5,000: The Story of Monsanto by Dan J. Forrestal (review). Technology and Culture, 20(1):236–237, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890481/pdf.

Belfield:1978:TTP

[Bel78] Robert Belfield. Technology Transfer: Proceedings of the NATO Advanced Study Institute on Technology Transfer (Evry, France, June 24–July 6, 1973) ed. by Harold F. Davidson, et al (review). Technology and Culture, 19(1):138–140, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892062/pdf.

Benson:1975:EAT

[Ben75] Barbara E. Benson. The engineer as an agent in technology transfer. Technology and Culture, 16(1):67–69, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891498/pdf.

Benson:1978:WEI

[Ben78] Ronald M. Benson. The Work Ethic in Industrial America, 1850– 1920 by Daniel T. Rodgers (review). Technology and Culture, 19 (4):756–758, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892193/pdf.

Berkebile:1970:RVA

[Ber70a] Don H. Berkebile. Roads and Vehicles by Anthony Bird (review). Technology and Culture, 11(3):439–441, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894111/pdf.

Berkebile:1970:FWE

[Ber70b] Don H. Berkebile. The Farm Waggons of England and Wales by James Arnold (review). Technology and Culture, 11(4):638–639, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894062/pdf.

Berkebile:1974:CHM

[Ber74] Don H. Berkebile. The Carriage and Harness Museum by Janet Sutherland Guldbeck (review). Technology and Culture, 15(3):537– 538, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892771/pdf.

Berggren:1977:AKA

[Ber77] J. L. Berggren. Astrolabe Kit and The Astrolabe: Some Notes on Its History, Construction and Use by Roderick S. Webster (review). Technology and Culture, 18(1):151–153, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891906/pdf.

Bezilla:1979:PSR

[Bez79] Michael Bezilla. The Pioneer Stage of Railroad Electrification by Carl W. Condit (review). Technology and Culture, 20(2):374–375, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890506/pdf.

Bickner:1973:MMS

[Bic73] Robert E. Bickner. Man-Machine System Experiments by Henry McIlvaine Parsons (review). Technology and Culture, 14(3):521– 522, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893240/pdf.

Bilstein:1973:ATS

[Bil73] Roger E. Bilstein. Air Travel: a Social History by Kenneth Hudson (review). Technology and Culture, 14(4):651–653, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892893/pdf.

Bilstein:1975:MEI

[Bil75a] Roger E. Bilstein. The Magnitude and Economic Impact of General Aviation, 1968–1980 by R. Dixon Speas Associates (review). Technology and Culture, 16(1):107–109, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891516/pdf.

Bilstein:1975:IBL

[Bil75b] Roger E. Bilstein. Those Incomparable Bonanzas by Larry A. Ball (review). Technology and Culture, 16(1):109–110, January 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891517/pdf.

Bilstein:1977:DAE

[Bil77] Roger E. Bilstein. Development of Aircraft Engines and Fuels by Robert Schlaifer, S. D. Heron (review). Technology and Culture, 18 (1):117–118, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891888/pdf.

Bilstein:1979:ONN

[Bil79] Roger E. Bilstein. Origins of NASA Names ed. by Helen T. Wells, et al. (review). Technology and Culture, 20(1):238, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890482/pdf.

Birr:1976:WDC

[Bir76] Kendall Birr. William David Coolidge: a Centenarian and His Work by Herman A. Liebhafsky (review). Technology and Culture, 17 (1):153–154, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891679/pdf.

Biswas:1970:CMI

[Bis70] Asit K. Biswas. Catalog of Meteorological Instruments in the Museum of History and Technology by W. E. Knowles Middleton (review). Technology and Culture, 11(3):488–489, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894140/pdf.

Biswas:1971:IMI

[Bis71] Asit K. Biswas. Invention of the Meteorological Instruments by W. E. Knowles Middleton (review). Technology and Culture, 12(2):340– 342, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893969/pdf.

Biswas:1976:TCS

[Bis76a] Asit K. Biswas. Three Centuries of Scientific Hydrology by UNESCO–WMO–IAHS (review). Technology and Culture, 17(1): 121–122, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891660/pdf.

Biswas:1976:WCM

[Bis76b] Asit K. Biswas. Water Current Meters in the Smithsonian Collections of the National Museum of History and Technology by Arthur H. Frazier (review). Technology and Culture, 17(2):415–417, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891748/pdf.

Black:1971:CCS

[Bla71] C. E. Black. Change in Communist Systems ed. by Chalmers Johnson (review). Technology and Culture, 12(3):521–523, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893616/pdf.

Black:1972:SRC

[Bla72] Cyril E. Black. Science in Russian Culture, 1861–1917 by Alexander Vucinich (review). Technology and Culture, 13(2):328–330, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893484/pdf.

Blackwell:1973:OCR

[Bla73a] William L. Blackwell. The Origins of Capitalism in Russia by Joseph T. Fuhrmann (review). Technology and Culture, 14(2 (part I)):315– 317, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893280/pdf.

Blaine:1973:CTW

[Bla73b] Bradford B. Blaine. The Classical Tradition in West European Farming by G. E. Fussell (review). Technology and Culture, 14(3):493– 495, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893225/pdf.

Blaine:1974:TCEa

[Bla74a] Bradford B. Blaine. Transportation and Communication in Early Medieval Europe, AD 500–1100 by Albert C. Leighton (review). Technology and Culture, 15(3):481–484, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892744/pdf.

Blaine:1974:TCEb

[Bla74b] Bradford B. Blaine. Twelfth-Century Europe: an Interpretive Essay by Sidney R. Packard (review). Technology and Culture, 15 (4):631–632, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892708/pdf.

Blanpied:1974:MSD

[Bla74c] William A. Blanpied. The Mature Society by Dennis Gabor (review). Technology and Culture, 15(2):352–354, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892812/pdf.

Black:1975:DCS

[Bla75] C. E. Black. Development of China’s Steel Industry and Soviet Technological Aid by M. Gardner Clark (review). Technology and Culture, 16(4):670–671, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891626/pdf.

Black:1976:BMI

[Bla76a] Cyril E. Black. Becoming Modern: Individual Change in Six Developing Countries by Alex Inkeles, David H. Smith (review). Technology and Culture, 17(3):568–569, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891788/pdf.

Blanpied:1976:RAI

[Bla76b] William A. Blanpied. The Roots of Ancient India by Walter Fairservis, Jr. (review). Technology and Culture, 17(4):785–787, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891843/pdf.

Blackwell:1977:ETC

[Bla77] D. J. Blackwell. Eli Terry and the Connecticut Shelf Clock by Kenneth D. Roberts (review). Technology and Culture, 18(4):695–697, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892003/pdf.

Black:1978:SST

[Bla78a] Cyril E. Black. Soviet Science, Technology, Design: Interaction and Convergence by Raymond Hutchings (review). Technology and Culture, 19(2):236–238, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892089/pdf.

Blanchard:1978:FL

[Bla78b] Leslie Blanchard. Focus on lighthouses. Technology and Culture, 19(4):804–806, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892217/pdf.

Blanpied:1978:TSD

[Bla78c] William A. Blanpied. Triveni: Science, Democracy and Socialism by A. Rahman (review). Technology and Culture, 19(3):539–541, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892134/pdf.

Black:1979:TCC

[Bla79a] Cyril E. Black. Technology and Communist Culture: The Socio-Cultural Impart of Technology under Socialism ed. by Frederic J. Fleron, Jr. (review). Technology and Culture, 20(1):252–255, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890490/pdf.

Blaine:1979:APE

[Bla79b] Bradford B. Blaine. Animal-powered Engines by J. Kenneth Major (review). Technology and Culture, 20(2):359–360, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890498/pdf.

Bloom:1977:DMT

[Blo77] Walter L. Bloom. The Diffusion of Medical Technology: Policy and Planning Perspectives ed. by Gerald Gordon, G. Lawrence Fisher (review). Technology and Culture, 18(4):727–728, October 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892021/pdf.

Brittain:1977:ENS

[BM77] James E. Brittain and Robert C. McMath, Jr. Engineers and the New South Creed: The formation and early development of Georgia Tech. Technology and Culture, 18(2):175–201, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891909/pdf.

Brittain:1975:CMD

[BMP^{+}75] James E. Brittain, Robert E. McGinn, Robert C. Post, Thomas Parke Hughes, Kenneth J. Hagan, Melvin Kranzberg, and Arthur L. Norberg. The Chicago Meeting, December 27–29, 1974. Technology and Culture, 16(3):429–453, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891567/pdf.

Bonelli:1971:LST

[Bon71] Maria Luisa Righini Bonelli. Leonardo as scientist and as technologist: Florence-Vinci, Italy, June 1969. Technology and Culture, 12(1):52–55, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894000/pdf.

Bond:1979:PT

[Bon79] Kener E. Bond, Jr. Profile turning. Technology and Culture, 20(3): 690, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890568/pdf.

Borghetty:1971:HDT

[Bor71a] Hector C. Borghetty. A History of Dyed Textiles by Stuart Robinson, and: A History of Printed Textiles by Stuart Robinson (review). Technology and Culture, 12(1):90–92, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894006/pdf.

Bornstein:1971:FSA

[Bor71b] Marc Bornstein. Future Shock by Alvin Toffler (review). Technology and Culture, 12(3):532–536, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893621/pdf.

Bostian:1971:ICR

[Bos71] Lloyd R. Bostian. The Impact of Communication on Rural Development: an Investigation in Costa Rica and India by Prodipto Roy, et al. (review). Technology and Culture, 12(2):368–369, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893984/pdf.

Boulding:1978:TCS

[Bou78] Kenneth E. Boulding. Technology and Civility: The Skill Revolution in Politics by Heinz Eulau (review). Technology and Culture, 19(2):253–254, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892098/pdf.

Bozeman:1975:SPJ

[Boz75] Barry Bozeman. Science and Politics by Jean-Jacques Salomon (review). Technology and Culture, 16(3):506–508, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891588/pdf.

Bozeman:1977:ABB

[Boz77] Barry Bozeman. America’s Brain Bank: an Inquiry into the Politics of Science by Philip Boffey (review). Technology and Culture, 18 (1):131–133, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891896/pdf.

Bryant:1978:WMO

[BPV^{+}78] Lynwood Bryant, Theodore A. Penn, Robert M. Vogel, Donald E. Drake II, Reese V. Jenkins, Carl F. Barnes, Jr., Kenneth J. Hagen, Arthur L. Norberg, and Bernard S. Finn. The Washington Meeting, October 19–22, 1977. Technology and Culture, 19(3):472– 490, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892116/pdf.

Bryant:1974:TAV

[BR74] Lynwood Bryant and A. R. Rogowski. Thermodynamics Atlas. Vol. 2: The Evolution of the Heat Engine by Ivo Kolin (review). Technology and Culture, 15(4):641–644, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892714/pdf.

Brainerd:1970:SUQ

[Bra70a] John G. Brainerd. Some unanswered questions. Technology and Culture, 11(4):601–603, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894047/pdf.

Bramwell:1970:SHF

[Bra70b] A. R. S. Bramwell. From a Scientific Heritage by F. A. Tsander (review). Technology and Culture, 11(2):318–319, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894169/pdf.

Bramwell:1972:FYS

[Bra72] A. R. S. Bramwell. Fifty Years of Soviet Aircraft Construction [50 let. sovetskogo samoletostroeniya] by A. S. Yakovlev (review). Technology and Culture, 13(1):111–112, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893458/pdf.

Brake:1975:APF

[Bra75] William H. Te Brake. Air pollution and fuel crises in preindustrial London, 1250–1650. Technology and Culture, 16(3):337–359, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891560/pdf.

Brainerd:1976:IHP

[Bra76a] John G. Brainerd. An Inquiry into the Human Prospect by Robert L. Heilbroner (review). Technology and Culture, 17(1):187–188, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891696/pdf.

Brainerd:1976:GE

[Bra76b] John G. Brainerd. Genesis of the Eniac. Technology and Culture, 17(3):482–488, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891755/pdf.

Brainerd:1976:OUH

[Bra76c] John G. Brainerd. The Open University and the history of technology. Technology and Culture, 17(2):235–244, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891713/pdf.

Brainerd:1977:HES

[Bra77a] John G. Brainerd. A History of Engineering and Science in the Bell System: The Early Years (1875–1925) ed. by M. S. Fagen (review). Technology and Culture, 18(1):113–116, January 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891886/pdf.

Brainerd:1977:NPE

[Bra77b] John G. Brainerd. A National Plan for Energy Research, Development and Demonstration: Creating Energy Choices for the Future by the staff of the Energy Research, Development Administration, and: Our Energy Future: The Role of Research, Development and Demonstration in Reaching a National Consensus on Energy Supply by Don E. Kash, members of the Science and Public Policy Program at the University of Oklahoma (review). Technology and Culture, 18(4):720–723, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892018/pdf.

Brainerd:1977:TFH

[Bra77c] John G. Brainerd. Telephone: The First Hundred Years by John Brooks (review). Technology and Culture, 18(4):706, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892008/pdf.

Brainerd:1977:IRC

[Bra77d] John G. Brainerd. International research conference on the history of computing, Los Alamos, June 10–15, 1976. Technology and Culture, 18(2):218–221, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891911/pdf.

Brainerd:1977:NPT

[Bra77e] John G. Brainerd. Nuclear power: Two sides. Technology and Culture, 18(3):503–504, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891945/pdf.

Bray:1978:SPS

[Bra78] Francesca Bray. Swords into plowshares: a study of agricultural technology and society in Early China. Technology and Culture, 19(1):1–31, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892038/pdf.

Brainerd:1979:HES

[Bra79a] John G. Brainerd. A History of Engineering and Science in the Bell System. Vol. 2: National Service in War and Peace (1925–1975) ed. by M. D. Fagen (review). Technology and Culture, 20(4):817–822, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890589/pdf.

Brainerd:1979:GSR

[Bra79b] John G. Brainerd. Gateway to Siberian Resources: The BAM by Theodore Shabad, Victor L. Mote (review). Technology and Culture, 20(1):250–252, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890489/pdf.

Brainerd:1979:MSA

[Bra79c] John G. Brainerd. Microelectronics by Scientific American, and: George Scheutz and the First Printing Calculator by Uta C. Merzbach (review). Technology and Culture, 20(2):390–392, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890515/
pdf.

Bray:1979:RAC

[Bra79d] Francesca Bray. Rural Asia: Challenge and Opportunity by the Asian Development Bank (review). Technology and Culture, 20(2):386– 388, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890513/pdf.

Brittain:1970:ILC

[Bri70] James E. Brittain. The introduction of the loading coil: George A. Campbell and Michael I. Pupin. Technology and Culture, 11(1):36– 57, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894189/pdf.

Briggs:1971:ICC

[Bri71a] Asa Briggs. The Ideal Communist City by A. Gutnov, et al. (review). Technology and Culture, 12(4):676–677, October 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893584/pdf.

Brittain:1971:LSW

[Bri71b] James Brittain. The Life of Sir William Fairbairn, Bart ed. by William Pole (review). Technology and Culture, 12(3):504–506, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893606/pdf.

Briggs:1972:ASI

[Bri72] J. Morton Briggs, Jr. The Anatomy of a Scientific Institution: The Paris Academy of Sciences, 1666–1803 by Roger Hahn (review). Technology and Culture, 13(2):312–314, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893473/pdf.

Briggs:1973:CCC

[Bri73] Asa Briggs. The City in Communist China ed. by John Wilson Lewis (review). Technology and Culture, 14(1):95–96, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893366/pdf.

Brittain:1974:PJH

[Bri74] James E. Brittain. The Papers of Joseph Henry. Vol. 1: December 1797–October 1832: The Albany Years ed. by Nathan Reingold (review). Technology and Culture, 15(1):103–105, January 1974. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892839/pdf.

Briggs:1976:CTA

[Bri76a] Asa Briggs. Cities in Transition: From the Ancient World to Urban America ed. by Frank J. Coppa, Philip C. Dolce (review). Technology and Culture, 17(3):548–549, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891775/pdf.

Briggs:1976:LSC

[Bri76b] Asa Briggs. London, 800–1216: The Shaping of a City by Christopher Brooke (review). Technology and Culture, 17(1):123–125, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891662/pdf.

Brittain:1977:AWB

[Bri77a] James E. Brittain. American Wooden Bridges ed. by the Committee on History and Heritage of American Civil Engineering (review). Technology and Culture, 18(1):100–101, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891877/pdf.

Brittain:1977:PJH

[Bri77b] James H. Brittain. The Papers of Joseph Henry. Vol. 2: November 1832–December 1833: The Princeton Years ed. by Nathan Reingold (review). Technology and Culture, 18(1):101–102, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891878/pdf.

Brickman:1978:SSE

[Bri78a] Ronald Brickman. Sputnik, Scientists, and Eisenhower by James R. Killian, Jr., and: A Scientist at the White House by George B. Kistiakowsky (review). Technology and Culture, 19(4):779–781, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892206/pdf.

Brittain:1978:ICB

[Bri78b] James E. Brittain. ISIS Cumulative Bibliography: a Bibliography of the History of Science Formed from ISIS Critical Bibliographies 1–90, 1913–1965. Vol. 3: Subjects ed. by Magda Whitrow (review). Technology and Culture, 19(2):266–267, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892105/pdf.

Bright:1979:EUA

[Bri79] Charles D. Bright. Encyclopedia of U.S. Air Force Aircraft and Missile Systems. Vol. 1: Post World War II Fighters, 1945–1973 by Marcelle S. Knaack (review). Technology and Culture, 20(4):826–827, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890593/pdf.

Brown:1979:EFS

[Bro79] Shannon R. Brown. The ewo filature: a study in the transfer of technology to China in the 19th century. Technology and Culture, 20(3):550–568, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890524/pdf.

Bruce:1973:CPS

[Bru73] Robert V. Bruce. Change at Park Street Under: The Story of Bostons Subways by Brian J. Cudahy (review). Technology and Culture, 14(4):642–643, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892887/pdf.

Brunner:1976:IBS

[Bru76] Elizabeth Brunner. Industrial Britain: The Steel Industry in Post-War Britain by David W. Heal (review). Technology and Culture, 17(1):165–167, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891686/pdf.

Bryant:1970:HJS

[Bry70] Lynwood Bryant. Hundert Jahre Sicherheit: Beiträge zur technischen und administrativen Entwicklung des Dampfkesselwesens in Deutschland 1810 bis 1910 by Gerhard Siegfried Sonnenberg (review). Technology and Culture, 11(3):451–454, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894119/pdf.

Bryant:1971:AAA

[Bry71a] Lynwood Bryant. America Adopts the Automobile, 1895–1910 by James J. Flink (review). Technology and Culture, 12(4):661–664, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893577/pdf.

Bryant:1971:DPS

[Bry71b] Lynwood Bryant. Damals als die Pferde scheuten: Die Geschichte der osterreichischen Kraftfahrt by Hans Seper (review). Technology and Culture, 12(4):659–661, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893576/pdf.

Bryant:1971:EML

[Bry71c] Lynwood Bryant. The end of the Marcus legend. Technology and Culture, 12(4):618–620, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893559/pdf.

Bryant:1973:RTE

[Bry73a] Lynwood Bryant. The role of thermodynamics in the evolution of heat engines. Technology and Culture, 14(2 (part I)):152–165, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893258/pdf.

Bryant:1973:SLS

[Bry73b] Keith L. Bryant, Jr. The St. Louis–San Francisco Transcontinental Railroad: The Thirty-fifth Parallel Project, 1853–1890 by H. Craig Miner (review). Technology and Culture, 14(4):639–640, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892885/pdf.

Bryant:1976:VEF

[Bry76a] Lynwood Bryant. Vehicle Engines: Fuel Consumption and Air Pollution ed. by M. S. Janota (review). Technology and Culture, 17 (1):175–177, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891691/pdf.

Bryant:1976:DDE

[Bry76b] Lynwood Bryant. The development of the diesel engine. Technology and Culture, 17(3):432–446, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891751/pdf.

Bryant:1977:HTF

[Bry77a] Lynwood Bryant. History of Technology: First Annual Volume, 1976 ed. by A. Rupert Hall, Norman Smith (review). Technology and Culture, 18(4):738–742, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892027/pdf.

Bryant:1977:IFC

[Bry77b] Lynwood Bryant. Internal Fire by C. Lyle Cummins, Jr. (review). Technology and Culture, 18(4):703–705, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892007.

Bryant:1977:BBM

[Bry77c] Lynwood Bryant. The Bosch Book of the Motor Car: Its Evolution and Engineering Development by John Day (review). Technology and Culture, 18(1):110–112, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891884/pdf.

Buchanan:1970:NIC

[Buc70a] R. A. Buchanan. New Inventions: a Comprehensive Survey of Scientific and Technical Progress in the Arts, Sciences, and Manufactures as Published during the Reign of Her Majesty ed. by Maurice Rickards (review). Technology and Culture, 11(3):443–444, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894113/pdf.

Buchanan:1970:SKR

[Buc70b] R. A. Buchanan. South Kensington to Robbins by Michael Argles, and: The Development of Technical Education in France, 1500–1850 by Frederick B. Artz (review). Technology and Culture, 11(2):308– 311, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894165/pdf.

Buchanan:1971:ERM

[Buc71a] H. Ray Buchanan. Europe in the Russian Mirror: Four Lectures in Economic History by Alexander Gerschenkron (review). Technology and Culture, 12(2):351–354, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893975/pdf.

Buchanan:1971:STI

[Buc71b] R. A. Buchanan. Science and Technology in the Industrial Revolution by A. E. Musson and Eric Robinson (review). Technology and Culture, 12(1):95–96, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894009/pdf.

Buchanan:1973:TCA

[Buc73] R. A. Buchanan. Technology and Culture: an Anthology ed. by Melvin Kranzberg, William H. Davenport (review). Technology and Culture, 14(3):474–477, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893217/pdf.

Buchanan:1974:AIR

[Buc74] R. A. Buchanan. The Archaeology of the Industrial Revolution by Brian Bracegirdle, et al. (review). Technology and Culture, 15 (4):669–671, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892727/pdf.

Buchanan:1975:SSF

[Buc75a] R. A. Buchanan. The Secret of Sherwood Forest: Oil Production in England during World War II by Guy H. Woodward, Grace Steele Woodward (review). Technology and Culture, 16(2):306–307, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891544/pdf.

Buchanan:1975:LTC

[Buc75b] R. A. Buchanan. Lionel Thomas Caswall Rolt (1910–1974). Technology and Culture, 16(3):454–457, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891568/pdf.

Buchanan:1976:TCB

[Buc76a] R. A. Buchanan. Twenty Centuries of British Industry by Hugh Bodey (review). Technology and Culture, 17(3):598–599, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891802/pdf.

Buchanan:1976:SIC

[Buc76b] R. A. Buchanan. Second International Congress on the Conservation of Industrial Monuments: Bergbau-Museum, Bochum, Germany, September 3–9, 1975. Technology and Culture, 17(2):245– 248, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891714/pdf.

Buchanan:1978:JSR

[Buc78] R. A. Buchanan. John Scott Russell: a Great Victorian Engineer and Naval Architect by George S. Emmerson (review). Technology and Culture, 19(4):767–769, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892199/pdf.

Buchanan:1979:WBE

[Buc79] R. A. Buchanan. Workshop of the British Empire: Engineering and Shipbuilding in the West of Scotland by Michael S. Moss, John R. Hume (review). Technology and Culture, 20(2):363–365, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890501/pdf.

Buder:1978:RLP

[Bud78] Stanley Buder. R. A. Long’s Planned City: The Story of Longview by John M. McClelland, Jr. (review). Technology and Culture, 19(3):560–562, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892147/pdf.

Bulferetti:1975:ML

[Bul75] Luigi Bulferetti. Mario Loria (1892–1971). Technology and Culture, 16(3):458–459, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891569/pdf.

Bullough:1976:ACT

[Bul76] Vern L. Bullough. Ancient Carpenters’ Tools: Together with Lumbermen’s, Joiners’ and Cabinet Makers’ Tools in Use in the Eighteenth Century by Henry C. Mercer (review). Technology and Culture, 17 (4):754–755, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891828/pdf.

Bunge:1974:TPJ

[Bun74] Mario Bunge. Theory and Practice by Jurgen Habermas (review). Technology and Culture, 15(3):539, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892773/pdf.

Burke:1970:CCN

[Bur70] John G. Burke. Comment: The complex nature of explanations in the historiography of technology. Technology and Culture, 11(1):22– 26, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894186/pdf.

Burns:1974:AGW

[Bur74a] Alfred Burns. Ancient Greek water supply and city planning: a study of Syracuse and acragas. Technology and Culture, 15(3):389– 412, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892736/pdf.

Burstyn:1974:UAN

[Bur74b] Harold L. Burstyn. Underwater Archaeology: a Nascent Discipline (review). Technology and Culture, 15(1):151–153, January 1974. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892861/pdf.

Burke:1975:SR

[Bur75] John G. Burke. The seminar revisited: 1974. Technology and Culture, 16(2):189–190, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891531/pdf.

Burke:1977:MII

[Bur77] John G. Burke. The Maze of Ingenuity: Ideas and Idealism in the Development of Technology by Arnold Pacey (review). Technology and Culture, 18(1):86–88, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891869/pdf.

Burke:1979:WPW

[Bur79] John G. Burke. Wood pulp, water pollution, and advertising. Technology and Culture, 20(1):175–195, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890458/pdf.

Byrnes:1974:SGP

[Byr74] Robert F. Byrnes. Social Groups in Polish Society ed. by David Lane, George Kolankiewicz (review). Technology and Culture, 15(2):349– 351, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892810/pdf.

Byrne:1978:TTP

[Byr78a] Edmund Byrne. Technology at the Turning Point ed. by William B. Pickett (review). Technology and Culture, 19(4):795–796, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892213/pdf.

Byrne:1978:SSP

[Byr78b] Edmund Byrne. Society for the study of philosophy and technology: Chicago, April 1977. Technology and Culture, 19(1):100–103, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892043/pdf.

Cahn:1976:MMN

[Cah76] R. W. Cahn. Materials and Man’s Needs: Supplementary Report of the Committee on the Survey of Materials Science and Engineering (COSMAT). Vol. 1: The History, Scope, and Nature of Materials Science and Engineering. Vol. 2: The Needs, Priorities, and Opportunities for Materials Research. Vol. 3: The Institutional Framework for Materials Science and Engineering. Vol. 4: Aspects of Materials Technology Abroad by National Academy of Sciences, and: Materials and Mans Needs: Summary Report of COSMAT by National Academy of Sciences (review). Technology and Culture, 17(3):570– 575, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891789/pdf.

Cain:1972:RWC

[Cai72] Louis P. Cain. Raising and watering a city: Ellis Sylvester Chesbrough and Chicago’s first sanitation system. Technology and Culture, 13(3):353–372, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893921/pdf.

Cain:1974:UPN

[Cai74] Louis P. Cain. Unfouling the public’s nest: Chicago’s sanitary diversion of Lake Michigan water. Technology and Culture, 15(4):594–613, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892703/pdf.

Cain:1978:RTA

[Cai78] Louis P. Cain. Retrospective Technology Assessment — 1976 ed. by Joel A. Tarr (review). Technology and Culture, 19(4):725–728, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892176/pdf.

Caldwell:1970:SMS

[Cal70a] Lynton K. Caldwell. Science in Modern Society by J. G. Crowther (review). Technology and Culture, 11(2):335–337, April 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894180/pdf.

Calvert:1970:ISM

[Cal70b] Monte A. Calvert. The Invention of the Sewing Machine by Grace Rogers Cooper (review). Technology and Culture, 11(3):449–451, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894118/pdf.

Calhoun:1971:PRR

[Cal71a] Daniel H. Calhoun. The Professions: Roles and Rules by Wilbert E. Moore, Gerald W. Rosenblum (review). Technology and Culture, 12(2):394–395, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893995/pdf.

Calvert:1971:UMT

[Cal71b] Monte A. Calvert. The U.S. Machine Tool Industry from 1900– 1950 by Harless D. Wagoner (review). Technology and Culture, 12 (1):123–125, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894022/pdf.

Callahan:1974:TSC

[Cal74] Leslie G. Callahan. The Technical Services — the Corps of Engineers: Construction in the United States by Lenore Fine, Jesse A. Remington (review). Technology and Culture, 15(1):114–116, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892845/pdf.

Cameron:1974:FEH

[Cam74] Rondo Cameron. The Fontana Economic History of Europe, Vol. 3: The Industrial Revolution, Vol. 4 (in two parts): The Emergence of Industrial Societies ed. by Carlo M. Cipolla (review). Technology and Culture, 15(3):485–487, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892746/pdf.

Candee:1974:BVA

[Can74] Richard Candee. A Bibliography on Vernacular Architecture ed. by Robert de Zouche Hall (review). Technology and Culture, 15 (4):671–672, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892728/pdf.

Carlson:1971:TCL

[Car71a] Robert E. Carlson. Technological Change and Labor in the Railroad Industry by Fred Cottrell (review). Technology and Culture, 12(1):113–115, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894019/pdf.

Carozzi:1971:HHA

[Car71b] Albert V. Carozzi. History of Hydrology by Asit K. Biswas (review). Technology and Culture, 12(3):493–495, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893602/pdf.

Cardwell:1972:TME

[Car72a] D. S. L. Cardwell. The Textile Mill Engine by George Watkins (review). Technology and Culture, 13(1):76–77, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893440/pdf.

Carlson:1972:REC

[Car72b] Robert E. Carlson. Railways in England, 1826 and 1827 by C. von Oeynhausen, H. von Dechen (review). Technology and Culture, 13(4):635, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893386/pdf.

Carlson:1972:BGW

[Car72c] Robert E. Carlson. The Birth of the Great Western Railway ed. by Jack Simmons (review). Technology and Culture, 13(3):500– 501, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893935/pdf.

Carlson:1972:MRC

[Car72d] Robert E. Carlson. The Making of a Railway by L. T. C. Rolt (review). Technology and Culture, 13(2):318, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893478/pdf.

Carpenter:1972:EIE

[Car72e] Kenneth E. Carpenter. European industrial exhibitions before 1851 and their publications. Technology and Culture, 13(3):465–486, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893930/pdf.

Cardwell:1973:LCS

[Car73] D. S. L. Cardwell. Lazare Carnot, Savant by C. C. Gillispie (review). Technology and Culture, 14(4):624–626, October 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892878/pdf.

Cardwell:1975:HGJ

[Car75a] D. S. L. Cardwell. H. G. J. Moseley: The Life and Letters of an English Physicist, 1887–1915 by J. L. Heilbron (review). Technology and Culture, 16(4):658–660, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891620/pdf.

Carlson:1975:WMR

[Car75b] Robert E. Carlson. The West Midlands by Rex Christiansen (review). Technology and Culture, 16(1):97–98, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891510/pdf.

Carpenter:1975:FTC

[Car75c] Stanley R. Carpenter. The Future of Technological Civilization by Victor Ferkiss (review). Technology and Culture, 16(4):696–699, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891639/pdf.

Cardwell:1976:STW

[Car76] Donald Cardwell. Science and technology: The work of James prescott Joule. Technology and Culture, 17(4):674–687, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891818/pdf.

Cardwell:1977:MWN

[Car77a] Donald S. L. Cardwell. Man and Water by Norman Smith (review). Technology and Culture, 18(4):692–693, October 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892000/pdf.

Carver:1977:TZE

[Car77b] J. Scott Carver. “Tekhnologicheskii zhurnal”: an early Russian technoeconomic periodical. Technology and Culture, 18(4):622–643, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891991/pdf.

Carpenter:1978:ATT

[Car78a] Stanley R. Carpenter. Autonomous Technology: Technics-out-of-Control as a Theme in Political Thought by Langdon Winner (review). Technology and Culture, 19(1):142–145, January 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892064/pdf.

Carpenter:1978:BSS

[Car78b] Stanley R. Carpenter. Building for Self-Sufficiency by Robin Clarke (review). Technology and Culture, 19(2):262–263, April 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892102/pdf.

Carpenter:1978:DPT

[Car78c] Stanley R. Carpenter. Developments in the philosophy of technology in America. Technology and Culture, 19(1):93–99, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892042/pdf.

Carver:1978:UAS

[Car78d] Scott Carver. USSR Academy of Sciences: Scientific Relations with Great Britain ed. by S. C. Korneyev (review). Technology and Culture, 19(4):769–771, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892200/pdf.

Cardwell:1979:TLT

[Car79a] Donald Cardwell. Theory loaded terms. Technology and Culture, 20(3):688–689, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890567/pdf.

Carlson:1979:ELA

[Car79b] Robert E. Carlson. End of the Line: Alexander J. Cassatt and the Pennsylvania Railroad by Patricia T. Davis (review). Technology and Culture, 20(1):228–230, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890477/pdf.

Carver:1979:SIG

[Car79c] J. Scott Carver. Studies in Industrial Geography: The USSR by J. C. Dewdney (review). Technology and Culture, 20(2):384–386, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890512/pdf.

Casey:1973:MCJ

[Cas73] William L. Casey, Jr. Man and the Computer by John G. Kemeny (review). Technology and Culture, 14(4):665–667, October 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892901/pdf.

Casey:1974:BFM

[Cas74] William L. Casey, Jr. Brain of the Firm: The Managerial Cybernetics of Organization by Stafford Beer (review). Technology and Culture, 15(1):129–131, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892852/pdf.

Casson:1976:GZA

[Cas76a] Lionel Casson. Galeona Zygmunta Augusta: Próba Rekonstrukcji by Mieczysław Boczar (review). Technology and Culture, 17(4):790–791, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891846/pdf.

Casson:1976:SAD

[Cas76b] Lionel Casson. The Story of Archaeological Decipherment, from Egyptian Hieroglyphs to Linear B by Maurice Pope (review). Technology and Culture, 17(3):530–531, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891767/pdf.

Catapano:1976:ECC

[Cat76] Fred T. Catapano. An Essay on Church Clocks, with Reference to the Clock from the Hospital Santa Cruz by Gideon Berman (review). Technology and Culture, 17(1):135–136, January 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891668/pdf.

Caws:1977:SHT

[Caw77] Peter Caws. The Sciences, the Humanities, and the Technological Threat ed. by W. Roy Niblett (review). Technology and Culture, 18(3):515–518, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891950/pdf.

Cesari:1970:ECE

[Ces70] Gene S. Cesari. The Environment of Change ed. by Aaron W. Warner, et al. (review). Technology and Culture, 11(4):652–653, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894070/pdf.

Chapin:1970:HAA

[Cha70] Seymour L. Chapin. History of Aeronautics and Astronautics, a Preliminary Bibliography by Katherine Murphy Dickson (review). Technology and Culture, 11(3):485–487, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/894138/pdf.

Chapin:1971:AAC

[Cha71a] Seymour L. Chapin. Astronautics and Aeronautics, 1968: a Chronology on Science, Technology, and Policy by Science and Technology Division, Library of Congress (review). Technology and Culture, 12(2):361–362, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893980/pdf.

Chapin:1971:PIH

[Cha71b] Seymour L. Chapin. Patent interferences and the history of technology: a high-flying example. Technology and Culture, 12(3):414– 446, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893593/pdf.

Chapin:1971:PGA

[Cha71c] Seymour L. Chapin. The persistent ghost: The artisan laroche. Technology and Culture, 12(1):69–74, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/894003/pdf.

Chamberlain:1972:SPP

[Cha72a] W. H. J. Chamberlain. Stevens Pistols and Pocket Rifles by Kenneth L. Cope, and: American Percussion Revolvers by Frank M. Sellers, Samuel E. Smith (review). Technology and Culture, 13(2):321–324, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893481/pdf.

Chatterjee:1972:CCI

[Cha72b] Pranab Chatterjee. Change and Continuity in India’s Villages by K. Ishwaran (review). Technology and Culture, 13(1):95–96, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893450/pdf.

Chamberlain:1973:HBR

[Cha73a] W. H. J. Chamberlain. Hall’s Breechloaders by R. T. Huntington (review). Technology and Culture, 14(1):86–87, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893360/pdf.

Chapelle:1973:SSS

[Cha73b] Howard I. Chapelle. Ships and Shipyards, Sailors and Fishermen: Introduction to Maritime Ethnology ed. by Olof Hasslof, et al. (review). Technology and Culture, 14(2 (part I)):296–300, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893269/pdf.

Chandler:1977:IET

[Cha77] Alfred D. Chandler, Jr. Images and Enterprises: Technology and the American Photographic Industry, 1839–1925 by Reese W. Jenkins (review). Technology and Culture, 18(1):107–110, January 1977.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891883/pdf.

Chamberlain:1978:SAW

[Cha78a] William H. J. Chamberlain. Small Arms of the World ed. by Edward C. Ezell (review). Technology and Culture, 19(4):758–760, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892194/pdf.

Chandler:1978:ADS

[Cha78b] Alfred D. Chandler, Jr. America by Design: Science, Technology, and the Rise of Corporate Capitalism by David F. Noble (review). Technology and Culture, 19(3):569–572, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892151/pdf.

Chubin:1978:ATE

[Chu78a] Daryl E. Chubin. Astronomy Transformed: The Emergence of Radio Astronomy in Britain by David O. Edge, Michael J. Mulkay (review). Technology and Culture, 19(3):580–583, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892154/pdf.

Chubin:1978:SOP

[Chu78b] Daryl E. Chubin. Scientists in Organizations: Productive Climates for Research and Development by Donald C. Pelz, Frank M. Andrews (review). Technology and Culture, 19(4):786–788, October 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892209/pdf.

Chubin:1979:SPS

[Chu79] Daryl E. Chubin. The Social Production of Scientific Knowledge: Sociology of the Sciences Yearbook, 1977 ed. by Everett Mendelsohn, et al. (review). Technology and Culture, 20(1):221–223, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890473/pdf.

Clark:1972:DSF

[Cla72] Charles H. Clark. The development of the semiautomatic freightcar coupler, 1863–1893. Technology and Culture, 13(2):170–208, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893464/pdf.

Claxton:1973:MRM

[Cla73] Robert H. Claxton. Miguel rivera maestre: Guatemalan scientistengineer. Technology and Culture, 14(3):384–403, July 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893209/pdf.

Clarke:1977:SHM

[Cla77] Jeffrey J. Clarke. The Social History of the Machine Gun by John Ellis (review). Technology and Culture, 18(2):262–263, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891929/pdf.

Clark:1978:LFG

[Cla78] Paul Clark. Legacy of Flight: The Guggenheim Contribution to American Aviation by Richard P. Hallion (review). Technology and Culture, 19(3):572–575, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892152/pdf.

Cliffe:1971:PVB

[Cli71] Frank B. Cliffe, Jr. Political Violence: The Behavioral Process by H. L. Nieburg (review). Technology and Culture, 12(1):155–158, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894038/pdf.

Clough:1970:BES

[Clo70a] Shepard B. Clough. Banking in the Early Stages of Industrialization: a Study in Comparative Economic History by Rondo Cameron (review). Technology and Culture, 11(2):307–308, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894164/pdf.

Clough:1970:SIS

[Clo70b] Shepard B. Clough. Gli Studi italiani di storia della scienza e della tecnica nell’ età contemporanea by Luigi Bulferetti (review). Technology and Culture, 11(3):483–484, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894136/pdf.

Clough:1975:SDT

[Clo75] Shepard B. Clough. Storia della trazione electrica ferroviaria in Italia by Mario Loria (review). Technology and Culture, 16(4):646, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891613/pdf.

Clough:1978:SNE

[Clo78] Shepard B. Clough. Studi e Notizie ed. by Carlo Maccagni (review). Technology and Culture, 19(2):214, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892076/pdf.

Cohen:1974:GEC

[Coh74] Marilyn S. Cohen. The Glorious Enterprise: The Centennial Exhibition of 1876 and H. J. Schwarzmann, Architect-in-Chief by John Maass (review). Technology and Culture, 15(2):338–339, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892803/pdf.

Coker:1972:SRN

[Cok72] William S. Coker. Spanish regulation of the Natchez indigo industry, 1793–1794: The South’s first antipollution laws? Technology and Culture, 13(1):55–58, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893431/pdf.

Cole:1971:SSP

[Col71] Jonathan R. Cole. The Social Sciences: Problems and Orientations by UNESCO (review). Technology and Culture, 12(1):153–155, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894037/pdf.

Comp:1974:PPL

[Com74] T. Allan Comp. Power over People by Louise B. Young, and: The Price of Power: Electric Utilities and the Environment by the Council on Economic Priorities (review). Technology and Culture, 15 (1):127–129, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892851/pdf.

Condit:1970:PTT

[Con70a] Carl W. Condit. Passenger Terminals and Trains by John A. Droege (review). Technology and Culture, 11(3):462–464, July 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894124/pdf.

Condit:1970:NWD

[Con70b] Carl W. Condit. The New World Dutch Barn by John Fitchen (review). Technology and Culture, 11(1):106–107, January 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894199/pdf.

Condit:1970:EUF

[Con70c] Carl W. Condit. The evolution of urban form. Technology and Culture, 11(3):428–433, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894107/pdf.

Conference:1970:SRE

[Con70d] Santa Clara Conference. The social responsibility of engineers. Technology and Culture, 11(2):241, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894155/pdf.

Condit:1971:BSH

[Con71a] Carl W. Condit. Brunelleschi: Studies of His Technology and Inventions by Frank D. Prager, Gustina Scaglia (review). Technology and Culture, 12(3):498–501, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893604/pdf.

Condit:1971:DLE

[Con71b] Carl W. Condit. Development of the Locomotive Engine by Angus Sinclair (review). Technology and Culture, 12(4):653–656, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893574/pdf.

Condit:1971:LDA

[Con71c] Carl W. Condit. The Life and Decline of the American Railroad by John F. Stover, and: London’s Termini by Alan A. Jackson, and: The Euston Arch and the Growth of the London, Midland and Scottish Railway by Alison Smithson and Peter Smithson (review). Technology and Culture, 12(1):109–113, January 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894018/pdf.

Condit:1971:UVS

[Con71d] Carl W. Condit. The Urban Vision: Selected Interpretations of the Modern American City ed. by Jack Tager, Park Dixon Goist (review). Technology and Culture, 12(3):518–519, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893613/pdf.

Condit:1971:TRS

[Con71e] Carl W. Condit. Town and Revolution: Soviet Architecture and City Planning, 1917–1935 by Anatole Kopp (review). Technology and Culture, 12(3):520–521, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893615/pdf.

Condit:1971:URT

[Con71f] Carl W. Condit. Urban Rail Transit: Its Economics and Technology by A. Scheffer Lang, Richard M. Soberman (review). Technology and Culture, 12(3):519–520, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893614/pdf.

Condit:1972:HRE

[Con72a] Carl W. Condit. A History of Real Estate, Building, and Architecture in New York City during the Last Quarter Century (1898) comp. by the Real Estate Record Association, and: The New York Subway: Its Construction and Equipment by the Interborough Rapid Transit Company (review). Technology and Culture, 13(3):516–517, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893945/pdf.

Condit:1972:RCA

[Con72b] Carl W. Condit. A Road to Creativity: Arthur E. Morgan, Engineer, Educator, Administrator by Clarence J. Leuba, and: Dams and Other Disasters: a Century of the Army Corps of Engineers by Arthur E. Morgan (review). Technology and Culture, 13(4):645–649, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893393/pdf.

Condit:1972:OGS

[Con72c] Carl. W. Condit. Our GM Scrapbook by Wallace W. Abbey, et al. (review). Technology and Culture, 13(4):637–639, October 1972. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893388/pdf.

Condit:1972:PAS

[Con72d] Carl W. Condit. Pueblo Architecture of the Southwest: a Photographic Essay by William Current (review). Technology and Culture, 13(1):83–85, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893445/pdf.

Condit:1972:PSF

[Con72e] Carl W. Condit. The Prairie School: Frank Lloyd Wright and His Midwest Contemporaries by H. Allen Brooks (review). Technology and Culture, 13(4):651–653, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893396/pdf.

Condit:1973:BDA

[Con73a] Carl W. Condit. A Biographical Dictionary of American Civil Engineers by Antoinette Lee (review). Technology and Culture, 14 (1):108–109, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893371/pdf.

Condit:1973:HDN

[Con73b] Carl W. Condit. A History of Dams by Norman Smith (review). Technology and Culture, 14(4):621–622, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892876/pdf.

Condit:1973:ABE

[Con73c] Carl W. Condit. American Building: The Environmental Forces That Shape It by James Marston Fitch (review). Technology and Culture, 14(3):509–511, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893234/pdf.

Condit:1973:ASC

[Con73d] Carl W. Condit. American Space: The Centennial Years, 1865–1876 by John Brinckerhoff Jackson (review). Technology and Culture, 14 (4):643–644, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892888/pdf.

Condit:1973:PBE

[Con73e] Carl W. Condit. People and Buildings ed. by Robert Gutman, and: London’s Historic Railway Stations by John Betjeman (review). Technology and Culture, 14(2 (part I)):303–304, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893272/pdf.

Condit:1973:ACB

[Con73f] Carl W. Condit. The Architect and the Computer by Boyd Auger (review). Technology and Culture, 14(4):670–672, October 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892904/pdf.

Constant:1973:MTC

[Con73g] Edward W. Constant II. A model for technological change applied to the turbojet revolution. Technology and Culture, 14(4):553–572, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892868/pdf.

Condit:1974:ASL

[Con74a] Carl W. Condit. American Single Locomotives and the “Pioneer” by John H. White, Jr. (review). Technology and Culture, 15(2):337– 338, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892802/pdf.

Condit:1974:OIA

[Con74b] Carl W. Condit. An Outline of Islamic Architecture by Rafique Ali Jairazbhoy (review). Technology and Culture, 15(4):636–637, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892711/pdf.

Condit:1974:APE

[Con74c] Carl W. Condit. Architecture for the Poor: an Experiment in Rural Egypt by Hassan Fathy (review). Technology and Culture, 15(2):343– 345, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892807/pdf.

Condit:1974:AWH

[Con74d] Carl W. Condit. Architecture in Wood: a History of Wood Building and its Techniques in Europe and North America ed. by Hans Jürgen Hansen (review). Technology and Culture, 15(4):637–639, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892712/pdf.

Condit:1974:FYR

[Con74e] Carl W. Condit. Fifty Years of Rapid Transit, 1864–1917 by James Blaine Walker (review). Technology and Culture, 15(1):109–110, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892842/pdf.

Condit:1974:HCN

[Con74f] Carl W. Condit. History of the Canadian National Railways by G. R. Stevens (review). Technology and Culture, 15(1):110–112, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892843/pdf.

Condit:1974:IBF

[Con74g] Carl W. Condit. Industrial Buildings and Factories by Oswald W. Grube (review). Technology and Culture, 15(1):112–113, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892844/pdf.

Condit:1974:LVR

[Con74h] Carl W. Condit. Leonardo da Vinci: The Royal Palace al Romorantin by Carlo Pedretti (review). Technology and Culture, 15 (4):634–635, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892710/pdf.

Condit:1974:VCI

[Con74i] Carl W. Condit. The Victorian City: Images and Realities ed. by H. J. Dyos, Michael Wolff, and: The City in the Ancient World by Mason Hammond, and: Close-Up: How to Read the American City by Grady Clay (review). Technology and Culture, 15(3):489– 493, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892748/pdf.

Condit:1975:BSN

[Con75a] Carl W. Condit. Bridges: The Spans of North America by David Plowden (review). Technology and Culture, 16(2):310–313, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891546/pdf.

Condit:1975:DIA

[Con75b] Carl W. Condit. Designing for Industry: The Architecture of Albert Kahn by Grant Hildebrand (review). Technology and Culture, 16(2):313–315, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891547/pdf.

Condit:1975:EUP

[Con75c] Carl W. Condit. Encyclopedia of Urban Planning ed. by Arnold Whittick (review). Technology and Culture, 16(2):315–316, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891548/
pdf.

Condit:1975:HPC

[Con75d] Carl W. Condit. History of the Portland Cement Industry in the United States by Robert W. Lesley, and: Sixty Years a Builder: The Autobiography of Henry Ericsson by Henry Ericsson (review). Technology and Culture, 16(3):500–501, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891585/pdf.

Condit:1975:SCT

[Con75e] Carl W. Condit. Some Classic Trains, More Classic Trains by Arthur D. Dubin (review). Technology and Culture, 16(4):641–643, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891611/pdf.

Condit:1975:UDM

[Con75f] Carl W. Condit. Urban Design: Manhattan by Regional Plan Association (review). Technology and Culture, 16(1):105–106, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891514/pdf.

Condit:1975:WSR

[Con75g] Carl W. Condit. When the Steam Railroads Electrified by William D. Middleton (review). Technology and Culture, 16(4):646–649, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891614/pdf.

Condit:1976:ADA

[Con76a] Carl W. Condit. Art Deco Architecture in New York, 1920–1940 by Don Vlack (review). Technology and Culture, 17(1):161–162, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891683/pdf.

Condit:1976:DSF

[Con76b] Carl W. Condit. Developments in Structural Form by Rowland J. Mainstone (review). Technology and Culture, 17(2):396–399, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891737/pdf.

Condit:1976:MCP

[Con76c] Carl W. Condit. Maya Cities: Placemaking and Urbanization by George F. Andrews (review). Technology and Culture, 17(3):546– 548, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891774/pdf.

Condit:1976:SSA

[Con76d] Carl W. Condit. Skyscraper Style: Art Deco in New York by Cervin Robinson, Rosemarie Haag Bletter (review). Technology and Culture, 17(2):399–400, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891738/pdf.

Condit:1976:MTA

[Con76e] Carl W. Condit. The Making of TVA by Arthur E. Morgan (review). Technology and Culture, 17(1):162–163, January 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891684/pdf.

Condit:1977:HBT

[Con77a] Carl W. Condit. A History of Building Types by Nikolaus Pevsner (review). Technology and Culture, 18(2):239–241, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891916/pdf.

Condit:1977:BEA

[Con77b] Carl W. Condit. Building Early America: Contributions toward a History of a Great Industry ed. by Charles E. Peterson (review). Technology and Culture, 18(3):535–538, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891960/pdf.

Condit:1977:DMT

[Con77c] Carl W. Condit. Depression Modern: The Thirties Style in America by Martin Greif (review). Technology and Culture, 18(3):567– 569, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891977/pdf.

Condit:1977:EGR

[Con77d] Carl W. Condit. Electrification by GE (review). Technology and Culture, 18(3):554–557, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891970/pdf.

Condit:1977:PWM

[Con77e] Carl W. Condit. The Pantheon by William L. MacDonald (review). Technology and Culture, 18(2):241–243, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891917/pdf.

Condit:1977:PCF

[Con77f] Carl W. Condit. The Pioneer: Chicago’s First Locomotive by John H. White, Jr. (review). Technology and Culture, 18(1):102, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891879/pdf.

Condit:1978:HSF

[Con78a] Carl W. Condit. History of the Skyscraper by Francisco Mujica (review). Technology and Culture, 19(3):533–534, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892131/pdf.

Condit:1978:MAO

[Con78b] Carl W. Condit. Moslem Architecture: Its Origin and Development by Giovanni T. Rivoira (review). Technology and Culture, 19 (1):112–114, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892048/pdf.

Condit:1978:TDS

[Con78c] Carl W. Condit. Temples of Democracy: The State Capitols of the USA by Henry-Russell Hitchcock, William Seale (review). Technology and Culture, 19(1):126–127, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892056/pdf.

Condit:1978:AEB

[Con78d] Carl W. Condit. The Architecture of the Ecole des Beaux-Arts ed. by Arthur Drexler (review). Technology and Culture, 19(4):738–740, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892182/pdf.

Condit:1978:MBH

[Con78e] Carl W. Condit. The Master Builders: a History of Structural and Environmental Design from Ancient Egypt to the Nineteenth Century by Henry J. Cowan, and: Science and Building: Structural and Environmental Design in the Nineteenth and Twentieth Centuries by Henry J. Cowan (review). Technology and Culture, 19(3):525– 528, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892127/pdf.

Condit:1979:PHA

[Con79a] Carl W. Condit. A Pictorial History of Architecture in America by G. E. Kidder Smith, Marshall B. Davidson, and: Frank Lloyd Wright’s Fallingwater: The House and Its History by Donald Hoffmann (review). Technology and Culture, 20(3):645–649, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890547/pdf.

Condit:1979:WVI

[Con79b] Carl W. Condit. A World with a View: an Inquiry into the Nature of Scenic Values by Christopher Tunnard (review). Technology and Culture, 20(1):204–205, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890464/pdf.

Condit:1979:PPB

[Con79c] Carl W. Condit. Pioneers of Prefabrication: The British Contribution in the Nineteenth Century by Gilbert Herbert (review). Technology and Culture, 20(1):206–207, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890465/pdf.

Condit:1979:RCC

[Con79d] Carl W. Condit. Rockefeller Center by Carol Herselle Krinsky (review). Technology and Culture, 20(1):207–210, January 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890466/pdf.

Condit:1979:SSL

[Con79e] Carl W. Condit. Sanitation Strategy for a Lakefront Metropolis: The Case of Chicago by Louis P. Cain (review). Technology and Culture, 20(4):813–815, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890587/pdf.

Condit:1979:ARP

[Con79f] Carl W. Condit. The American Railroad Passenger Car by John H. White, Jr. (review). Technology and Culture, 20(2):370–374, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890505/pdf.

Cooper:1971:TFI

[Coo71] Richard N. Cooper. The Technology Factor in International Trade ed. by Raymond Vernon (review). Technology and Culture, 12 (3):523–525, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893617/pdf.

Cooper:1974:TCJ

[Coo74] Grace W. Cooper. Textile Conservation by Jentina E. Leene (review). Technology and Culture, 15(2):363–365, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892816/pdf.

Cook:1977:USA

[Coo77] Charles Orson Cook. Urban Slavery in the American South, 1820– 1860: a Quantitative History by Claudia Dale Goldin (review). Technology and Culture, 18(3):543–544, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891964/pdf.

Costello:1971:OAS

[Cos71] D. F. Costello. Office Automation in Social Perspective by H. A. Rhee (review). Technology and Culture, 12(1):131–134, January 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894027/pdf.

Cotter:1970:CWC

[Cot70] John L. Cotter. Colonial Williamsburg by Carlisle H. Humelsine, et al. (review). Technology and Culture, 11(3):417–427, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894106/pdf.

Cotter:1972:SMM

[Cot72] John L. Cotter. Surveying and Mapping: a Manual of Simplified Techniques by Robert F. G. Spier (review). Technology and Culture, 13(1):114, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893461/pdf.

Cowan:1975:AWH

[Cow75a] Ruth Schwartz Cowan. The American Woman: Her Changing Social, Economic, and Political Role, 1920–1970 by William H. Chafe (review). Technology and Culture, 16(2):322–324, April 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891553/pdf.

Cowan:1975:VLW

[Cow75b] Ruth Schwartz Cowan. Victorian Ladies at Work: Middle Class Working Women in England and Wales, 1850–1914 by Lee Holcombe (review). Technology and Culture, 16(3):493–496, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891582/pdf.

Cowan:1976:R

[Cow76a] Ruth Cowan. Right on — and on! Technology and Culture, 17(3): 527, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891765/pdf.

Cowan:1976:IRH

[Cow76b] Ruth Schwartz Cowan. The “Industrial revolution” in the home: Household technology and social change in the 20th century. Technology and Culture, 17(1):1–23, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891649/pdf.

Cowan:1979:VDV

[Cow79] Ruth Schwartz Cowan. From Virginia Dare to Virginia Slims: Women and technology in American life. Technology and Culture, 20(1):51–63, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890452/pdf.

Craig:1975:JDA

[Cra75] Robert M. Craig. James Dakin, Architect, His Career in New York and the South by Arthur Scully, Jr. (review). Technology and Culture, 16(2):308–310, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891545/pdf.

Craig:1976:ACE

[Cra76a] Robert M. Craig. The Architecture of Choice: Eclecticism in America, 1880–1930 by Walter C. Kidney, and: Americas Gilded Age, Its Architecture and Decoration by Frederick Platt (review). Technology and Culture, 17(4):776–779, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891839/pdf.

Craig:1976:SDD

[Cra76b] Robert M. Craig. The Streamlined Decade by Donald Bush (review). Technology and Culture, 17(4):779–781, October 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891840/pdf.

Craig:1977:BNR

[Cra77] Robert M. Craig. Building with Nature: Roots of the San Francisco Bay Region Tradition by Leslie Mandelson Freudenheim, Elizabeth Sussman (review). Technology and Culture, 18(1):122–124, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891891/pdf.

Craig:1978:AFA

[Cra78] Robert M. Craig. America’s Forgotten Architecture by Tony P. Wrenn, Elizabeth D. Mulloy, and: Unbuilt America: Forgotten Architecture in the United States from Thomas Jefferson to the Space Age by Alison Sky, Michelle Stone (review). Technology and Culture, 19(4):733–738, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892181/pdf.

Craig:1979:BMA

[Cra79a] Robert M. Craig. Bernard Maybeck, Artisan, Architect, Artist by Kenneth H. Cardwell, and: Greene and Greene: Architecture as a Fine Art by Randell L. Makinson, and: Greene and Greene: Furniture and Related Designs by Randell L. Makinson (review). Technology and Culture, 20(4):807–810, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890585/pdf.

Craig:1979:FFF

[Cra79b] Robert M. Craig. Form Follows Fiasco: Why Modern Architecture Hasn’t Worked by Peter Blake (review). Technology and Culture, 20(2):394–396, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890517/pdf.

Craig:1979:ACH

[Cra79c] Robert M. Craig. The Architect: Chapters in the History of the Profession ed. by Spiro Kostof (review). Technology and Culture, 20(3):650–651, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890548/pdf.

Crocombe:1971:MLP

[Cro71] Ron Crocombe. Matupit: Land, Politics, and Change among the Tolai of New Britain by A. L. Epstein (review). Technology and Culture, 12(1):151–153, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894036/pdf.

Crowe:1976:SIC

[Cro76] Charles Crowe. Slavery, ideology, and “Cliometrics”. Technology and Culture, 17(2):271–285, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891719/pdf.

Cutting:1974:TOS

[Cut74] Gary Cutting. Thematic Origins of Scientific Thought by Gerald Holton (review). Technology and Culture, 15(4):657–659, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892722/pdf.

Condit:1973:TTP

[CZ73] Carl W. Condit and Joseph Zbilut. Technical Thought and Protection against Moisture in Industrial Building in the Polish Kingdom of the First Half of the XIXth Century by Jerzy Górewicz (review). Technology and Culture, 14(3):505–506, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893232/pdf.

Dalrymple:1970:ADE

[Dal70] Dana G. Dalrymple. Agricultural Development and Economic Growth ed. by Herman M. Southworth, Bruce F. Johnston (review). Technology and Culture, 11(1):111–113, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894203/pdf.

Dalrymple:1972:IGR

[Dal72] Dana G. Dalrymple. India’s Green Revolution: Economic Gains and Political Costs by Francine R. Frankel (review). Technology and Culture, 13(2):330–331, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893485/pdf.

Daniels:1970:BQH

[Dan70a] George H. Daniels. The big questions in the history of American technology. Technology and Culture, 11(1):1–21, January 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894185/pdf.

Daniels:1970:RDA

[Dan70b] George H. Daniels. The reply: Differences and agreements. Technology and Culture, 11(1):32–35, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/894188/pdf.

Daniels:1972:ESI

[Dan72] George H. Daniels. Elmer Sperry: Inventor and Engineer by Thomas Parke Hughes (review). Technology and Culture, 13(2):326–328, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893483/pdf.

Dator:1971:USD

[Dat71] James Allen Dator. The Unprepared Society by Donald Michael (review). Technology and Culture, 12(1):141–143, January 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894031/pdf.

Daub:1972:WCR

[Dau72] Edward E. Daub. From Watts to Clausius: The Rise of Thermodynamics in the Early Industrial Age by Donald S. L. Cardwell (review). Technology and Culture, 13(2):315, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893475/pdf.

Davis:1970:BAP

[Dav70] Robert C. Davis. The Brass Age of psychology. Technology and Culture, 11(4):604–612, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894048/pdf.

Davis:1975:SSO

[Dav75] Robert C. Davis. Social Science Organization and Policy by UN-ESCO (review). Technology and Culture, 16(2):321–322, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891552/pdf.

Day:1974:IDE

[Day74] C. R. Day. Instruction et développement économique au XIXème siècle ed. by Carlo Cipolla (review). Technology and Culture, 15 (3):518–520, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892764/pdf.

deCamp:1970:MLG

[dC70] L. Sprague de Camp. Mazes and Labyrinths: a General Account of Their History and Developments by W. H. Matthews (review). Technology and Culture, 11(3):464–465, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/894125/pdf.

deCamp:1972:AMI

[dC72a] L. Sprague de Camp. Absolutely Mad Inventions by A. E. Brown, H. A. Jeffcott, Jr. (review). Technology and Culture, 13(3):525, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893950/pdf.

decamp:1972:SSA

[dc72b] L. Sprague de camp. Ships and Seamanship in the Ancient World by Lionel Casson (review). Technology and Culture, 13(1):67, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893434/pdf.

deCamp:1974:LST

[dC74a] L. Sprague de Camp. The Lure of Sunken Treasure by Robert F. Marx (review). Technology and Culture, 15(3):538–539, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892772/pdf.

deCamp:1974:VEW

[dC74b] L. Sprague de Camp. Viking Expansion Westwards by Magnus Magnusson (review). Technology and Culture, 15(3):484–485, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892745/pdf.

deCamp:1976:SFB

[dC76] L. Sprague de Camp. The Science Fiction Book by Franz Rottensteiner (review). Technology and Culture, 17(4):781–782, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891841/pdf.

deCamp:1978:EAW

[dC78] L. Sprague de Camp. Engineering in the Ancient World by J. G. Landels (review). Technology and Culture, 19(4):728, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892177/pdf.

deCamp:1979:APH

[dC79] L. Sprague de Camp. The Ancient Persians (How They Lived and Worked) by Brian Dicks (review). Technology and Culture, 20(4):789, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890576/pdf.

DeGregori:1970:CPD

[De 70] Thomas R. De Gregori. The Civilizational Process by Darcy Ribeiro (review). Technology and Culture, 11(4):667–668, October 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894076/pdf.

DeGregori:1971:IDE

[De 71] Thomas R. De Gregori. Industrialization and Development ed. by H. E. Hoelscher, M. C. Hawk (review). Technology and Culture, 12(2):362–363, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893981/pdf.

DeGregori:1973:CBA

[De 73a] Thomas R. De Gregori. Civilizations of Black Africa by Jacques Maquet (review). Technology and Culture, 14(1):96–97, January 1973.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893367/pdf.

DeGregori:1973:GFW

[De 73b] Thomas R. De Gregori. The Ghanaian Factory Worker: Industrial Man in Africa by Margaret Peil, and: The Political Economy of Colonialism in Ghana: a Collection of Documents and Statistics, 1900–1960 ed. by G. B. Kay (review). Technology and Culture, 14(3):517–518, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893238/pdf.

DeGregori:1974:EHN

[De 74a] Thomas R. De Gregori. An Economic History of Nigeria, 1860– 1960 by R. Olufemi Ekundare (review). Technology and Culture, 15(1):150, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892860/pdf.

DeGregori:1974:EHW

[De 74b] Thomas R. De Gregori. An Economic History of West Africa by A. G. Hopkins (review). Technology and Culture, 15(3):514–516, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892761/pdf.

DeGregori:1974:AMP

[De 74c] Thomas R. De Gregori. Applied Measures for Promoting Technological Growth by Stanley A. Hetzler (review). Technology and Culture, 15(3):517–518, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892763/pdf.

DeGregori:1974:AST

[De 74d] Thomas R. De Gregori. The African Slave Trade and Its Suppression: a Classified and Annotated Bibliography of Books, Pamphlets, and Periodical Articles by Peter C. Hogg (review). Technology and Culture, 15(3):516–517, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892762/pdf.

DeSollaPrice:1974:HRH

[De 74e] Derek De Solla Price. On the historiographic revolution in the history of technology: Commentary on the papers by Multhauf, Ferguson, and Layton. Technology and Culture, 15(1):42–48, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/

892823/pdf. See [Mul74e, Fer74a, Lay74].

DeVore:1974:TVT

[De 74f] Paul De Vore. Technical and Vocational Teacher Education and Training by Mary Ann Calkins Pilain (review). Technology and Culture, 15(3):520–522, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892765/pdf.

DeGregori:1975:IBA

[De 75a] Thomas R. De Gregori. An Introduction to Benin Art and Technology by Philip J. C. Dark (review). Technology and Culture, 16 (4):701–702, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891642/pdf.

DeGregori:1975:WAF

[De 75b] Thomas R. De Gregori. West African Food in the Middle Ages: According to Arabic Sources by Tadeusz Lewicki (review). Technology and Culture, 16(4):627–628, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891605/pdf.

DeGregori:1975:QER

[De 75c] Thomas R. De Gregori. Queries on economic rationality. Technology and Culture, 16(3):460–461, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891570/pdf.

DeGregori:1976:ECP

[De 76] Thomas R. De Gregori. Economic Change in Pre-Colonial Africa: Senegambia in the Era of the Slave Trade by Philip D. Curtin (review). Technology and Culture, 17(2):406–408, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891742/pdf.

DeGregori:1978:RPF

[De 78] Thomas R. De Gregori. Reaching the Peasant Farmer: Organization Theory and Practice in Kenya by David K. Leonard (review). Technology and Culture, 19(3):547–548, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892138/pdf.

DeGregori:1979:FDI

[De 79a] Thomas R. De Gregori. Foreign Direct Investment, Industrialization, and Social Change by Stephen Jay Korbin (review). Technology and Culture, 20(3):662–663, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890553/pdf.

DeGregori:1979:EIE

[De 79b] Thomas R. De Gregori. The Evolution of the International Economic Order by Arthur Lewis (review). Technology and Culture, 20(1):225, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890475/pdf.

DeGregori:1971:TGS

[DeG71] Thomas R. DeGregori. Technological Growth and Social Change: Achieving Modernization by Stanley A. Hetzler (review). Technology and Culture, 12(1):145–146, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894033/pdf.

DeGregori:1975:CPP

[DeG75] Thomas R. DeGregori. Cultural Policy in the Philippines by UNESCO, and: Cultural Policy in Senegal by Momadou Seyni M’Bengue, and: Cultural Policy in Indonesia by UNESCO, and: Cultural Policy in Israel by Jozeph Michman, and: Cultural Policy in the Federal Republic of Germany by UNESCO, and: The Role of Culture in Leisure-Time in New Zealand by Bernard W. Smyth (review). Technology and Culture, 16(2):320–321, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891551/pdf.

DeGregori:1976:STA

[DeG76] Thomas R. DeGregori. Science and Technology in African Development by UNESCO (review). Technology and Culture, 17(1):182–183, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891694/pdf.

DeGregori:1977:WSC

[DeG77] Thomas R. DeGregori. Warfare in the Sokoto Caliphate: Historical and Sociological Perspectives by Joseph P. Smaldone (review). Technology and Culture, 18(4):730–731, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892023/pdf.

DelSesto:1976:CGN

[Del76] Steven L. Del Sesto. Citizen Groups and the Nuclear Power Controversy: Uses of Scientific and Technological Information by Steven Ebbin, Raphael Kasper (review). Technology and Culture, 17(2):401– 404, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891740/pdf.

DelSesto:1977:PS

[Del77] Steven L. Del Sesto. The public’s stake. Technology and Culture, 18(3):505–507, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891946/pdf.

Devons:1975:CLJ

[Dev75a] Samuel Devons. The Cavendish Laboratory, 1874–1974 by J. G. Crowther (review). Technology and Culture, 16(4):655–658, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891619/pdf.

DeVore:1975:AEV

[DeV75b] Paul W. DeVore. American Education and Vocationalism: a Documentary History, 1870–1970 ed. by Marvin Lazerson and W. Norton Grubb (review). Technology and Culture, 16(1):113–114, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891519/pdf.

Devons:1976:CET

[Dev76] Samuel Devons. The Compton Effect: Turning Point in Physics by Roger H. Stuewer (review). Technology and Culture, 17(1):154–157, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891680/pdf.

Devor:1978:TEI

[Dev78] David S. Devor. Technology and Economic Interdependence by Harry G. Johnson (review). Technology and Culture, 19(2):241–242, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892091/pdf.

Dew:1976:VMA

[Dew76] Charles B. Dew. The Virginia Manufactory of Arms by Giles Cromwell (review). Technology and Culture, 17(1):138–140, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891670/pdf.

DeWalt:1978:ATR

[DeW78] Billie R. DeWalt. Appropriate technology in rural Mexico: Antecedents and consequences of an indigenous peasant innovation. Technology and Culture, 19(1):32–52, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892039/pdf.

Dibner:1972:HA

[Dib72] Bern Dibner. Hoover and Agricola. Technology and Culture, 13 (3):417–425, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893924/pdf.

Dibner:1974:LR

[Dib74] Bern Dibner. Ladislao Reti (1901–1973). Technology and Culture, 15(3):440–442, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892739/pdf.

Dibner:1975:MVM

[Dib75] Bern Dibner. Martinus van Marum: Life and Work by G. L. Turner, T. H. Levere (review). Technology and Culture, 16(2):295– 297, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891539/pdf.

Dibner:1976:FME

[Dib76] Bern Dibner. First martyr of electricity. Technology and Culture, 17(4):743–745, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891823/pdf.

Dibner:1979:ILV

[Dib79] Bern Dibner. The Inventions of Leonardo da Vinci by Charles Gibbs-Smith (review). Technology and Culture, 20(3):625–626, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890534/pdf.

Diebold:1973:HCA

[Die73] John Diebold. How Computers Affect Management by Rosemary Stewart (review). Technology and Culture, 14(3):518–520, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893239/pdf.

Mann:1973:LWC

[dLM73] J. de L. Mann. The London Weavers Company, 1600 to 1971 by Alfred Plummer (review). Technology and Culture, 14(4):620–621, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892875/pdf.

Mann:1976:WRW

[dLM76] J. de L. Mann. The West Riding Wool Textile Industry, 1770– 1835: a Study in Fixed Capital Formation by D. T. Jenkins (review). Technology and Culture, 17(4):797–798, October 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891850/pdf.

Dobney:1975:DPP

[Dob75] Fredrick J. Dobney. Dams, Parks, and Politics: Resource Development and Preservation in the Truman–Eisenhower Era by Elmo Richardson (review). Technology and Culture, 16(1):101–103, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891512/pdf.

Dobbs:1978:AMA

[Dob78] Betty Jo Dobbs. Alchemy: The Medieval Alchemists and Their Royal Art by Johannes Fabricius (review). Technology and Culture, 19(3):516–518, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892123/pdf.

Donovan:1977:SSI

[Don77] Arthur Donovan. Scottish Scientific Instrument-Makers, 1600–1900 by D. J. Bryden (review). Technology and Culture, 18(4):694–695, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892002/pdf.

Dorwart:1970:FFA

[Dor70] Reinhold A. Dorwart. Fire Fighting Appliances by K. R. Gilbert (review). Technology and Culture, 11(3):478, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894132/pdf.

Dorn:1976:QOD

[Dor76] Harold Dorn. “Questionable opinions” defended. Technology and Culture, 17(3):523–525, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891763/pdf.

Dorn:1979:HLC

[Dor79] Harold Dorn. Hugh Lincoln Cooper and the first Détente. Technology and Culture, 20(2):322–347, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890494/pdf.

Dotson:1975:VHE

[Dot75] John E. Dotson. Venice: The Hinge of Europe, 1081–1797 by William H. McNeill (review). Technology and Culture, 16(1):83–85, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891503/pdf.

Drachmann:1970:GRA

[Dra70] A. G. Drachmann. Greek and Roman Artillery: Historical Development by E. W. Marsden (review). Technology and Culture, 11 (4):620–623, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894052/pdf.

Drachmann:1972:GRA

[Dra72] A. G. Drachmann. Greek and Roman Artillery: Technical Treatises by E. W. Marsden (review). Technology and Culture, 13(3):487– 494, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893931/pdf.

Drachmann:1976:GGA

[Dra76a] A. G. Drachmann. Gears from the Greeks: The Antikythera Mechanism — a Calendar Computer from ca. 80 B.C. by Derek de Solla Price (review). Technology and Culture, 17(1):112–116, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL http://www.jstor.org/stable/3103259;   https:
//muse.jhu.edu/pub/1/article/891657/pdf.

Drachmann:1976:PBP

[Dra76b] A. G. Drachmann. Philo of Byzantium: Pneumatica by Frank David Prager (review). Technology and Culture, 17(1):116–119, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891658/pdf.

Drake:1976:MPL

[Dra76c] Stillman Drake. The Mathematical Preface to … Euclid (1570) by John Dee (review). Technology and Culture, 17(1):128–129, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891665/pdf.

Drake:1978:PCU

[Dra78] Donald E. Drake II. Proceedings, Conference on University Education for Technology and Public Policy ed. by Eric B. Hartman, Robert P. Morgan (review). Technology and Culture, 19(3):594– 596, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892159/pdf.

Dresbeck:1975:MCL

[Dre75] LeRoy Dresbeck. Mappae Clavicula: a Little Key to the World of Medieval Techniques ed. by Cyril Stanley Smith, John G. Hawthorne (review). Technology and Culture, 16(2):287–289, April 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891535/pdf.

Dresbeck:1976:BGP

[Dre76] LeRoy J. Dresbeck. The Bright and Glowing Place by Frank Rowsome, Jr. (review). Technology and Culture, 17(3):575–577, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891790/pdf.

Drucker:1974:UTC

[Dru74] Peter F. Drucker. Understanding Technology by Charles Susskind (review). Technology and Culture, 15(1):80–82, January 1974. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892828/pdf.

Drucker:1976:LJD

[Dru76] Peter F. Drucker. Lessons from Japanese Development: an Analytical Economic History by Allen C. Kelley, Jeffrey G. Williamson (review). Technology and Culture, 17(3):559–561, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891783/pdf.

Price:1975:BKI

[dSP75] Derek de Solla Price. The Book of Knowledge of Ingenious Mechanical Devices by Ibn al-Razzaz al-Jazari tr. by Donald R. Hill (review). Technology and Culture, 16(1):81–83, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891502/pdf.

Dublin:1977:CCI

[Dub77] Thomas Dublin. Class and Community: The Industrial Revolution in Lynn by Alan Dawley (review). Technology and Culture, 18(3):545–547, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891965/pdf.

Duggan:1979:REH

[Dug79] Edward P. Duggan. Research in Economic History: an Annual Compilation of Research ed. by Paul Uselding (review). Technology and Culture, 20(1):223–224, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890474.

Duncan:1976:IDB

[Dun76] Francis Duncan. Independence and Deterrence, Britain and Atomic Energy, 1945–1952. Vol. 1: Policy Making. Vol. 2: Policy Execution by Margaret Gowing (review). Technology and Culture, 17 (1):167–169, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891687/pdf.

Durbin:1972:TVP

[Dur72] Paul T. Durbin. Technology and values: a philosopher’s perspective. Technology and Culture, 13(4):556–576, October 1972. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893377/pdf.

Durbin:1979:SCJ

[Dur79] Paul T. Durbin. Science and Its Critics by John Passmore (review). Technology and Culture, 20(4):829–835, October 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890595/pdf.

Eaton:1973:MHB

[Eat73] Leonard K. Eaton. M. H. Baillie Scott and the Arts and Crafts Movement by James D. Kornwolf (review). Technology and Culture, 14(2 (part I)):301–303, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893271/pdf.

Eaton:1975:ACM

[Eat75] Leonard K. Eaton. The Arts and Crafts Movement in America ed. by Robert Judson Clark (review). Technology and Culture, 16(2):316– 318, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891549/pdf.

Edwards:1972:AIS

[Edw72] Charles E. Edwards. The Automobile Industry since 1945 by Lawrence J. White (review). Technology and Culture, 13(3):510– 511, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893942/pdf.

Ehrlich:1973:SMT

[Ehr73] Richard L. Ehrlich. Shipping, Maritime Trade, and the Economic Development of Colonial North America by James F. Shepherd, Gary M. Walton (review). Technology and Culture, 14(4):626–627, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892879/pdf.

Ehrlich:1974:PWV

[Ehr74] Richard L. Ehrlich. Perceptions of Work: Variations within a Factory by H. Beynon, R. M. Blackburn (review). Technology and Culture, 15(3):527–529, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892768/pdf.

Eisenman:1975:SAS

[Eis75a] H. J. Eisenman. Science and, the Arts: a Study in Relationships from 1600–1900 by Jacob Opper (review). Technology and Culture, 16(1):124–126, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891525/pdf.

Eisenman:1975:TSV

[Eis75b] H. J. Eisenman. Technology, society, and values in 20th-Century America: The UCLA 1973 Summer Seminar. Technology and Culture, 16(2):182–188, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891530/pdf.

Eisenman:1976:NMT

[Eis76] H. J. Eisenman. National Museum of Transport by Mary W. Roberts (review). Technology and Culture, 17(1):93–103, January 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891653/pdf.

Ellsworth:1976:ADE

[Ell76] Lucius F. Ellsworth. The Americans: The Democratic Experience by Daniel J. Boorstin (review). Technology and Culture, 17(2):385– 387, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891731/pdf.

Elliott:1977:WPC

[Ell77] Ward Elliott. Who Pays for Clean Air: The Cost and Benefit Distribution of Federal Automobile Emission Controls by David Harrison, Jr. (review). Technology and Culture, 18(1):135–138, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891898/pdf.

Emerson:1978:SES

[Eme78] Roger L. Emerson. The Scottish Enlightenment: a Social History by Anand Chitnis (review). Technology and Culture, 19(3):524– 525, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892126/pdf.

Emme:1971:IHR

[Emm71] Eugene M. Emme. International history of rocketry and astronautics symposium: Constance, West Germany, October 1970. Technology and Culture, 12(3):477–486, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893599/pdf.

Emme:1972:PRH

[Emm72] Eugene M. Emme. The Papers of Robert H. Goddard, 1898–1945 ed. by Esther C. Goddard, G. Edward Pendray (review). Technology and Culture, 13(4):674–677, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893408/pdf.

Emme:1974:APC

[Emm74] Eugene M. Emme. Air Power: a Concise History by Robin Higham (review). Technology and Culture, 15(2):345–346, April 1974. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892808/pdf.

Emme:1976:TIR

[Emm76] Eugene M. Emme. The Third Industrial Revolution by G. Harry Stine (review). Technology and Culture, 17(4):809–811, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891857/pdf.

Emmerson:1977:WIK

[Emm77] G. S. Emmerson. The Works of Isambard Kingdom Brunel ed. by Sir Alfred Pugsley (review). Technology and Culture, 18(3):529– 532, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891957/pdf.

Emme:1978:AAB

[Emm78] Eugene M. Emme. The All-American Boys by Walter Cunningham, Mickey Herskowitz (review). Technology and Culture, 19(4):782–784, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892207/pdf.

Emme:1979:SAA

[Emm79] Eugene M. Emme. Soviet Aviation and Air Power: a Historical Review ed. by Robin Higham, Jacob W. Kipp (review). Technology and Culture, 20(4):828–829, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890594/pdf.

Espenschied:1970:SRD

[Esp70] Lloyd Espenschied. “Something rotten in Denmark” … in Boston. Technology and Culture, 11(4):596–597, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894045/pdf.

Esper:1972:EMC

[Esp72] Thomas Esper. Enserfment and Military Change in Muscovy by Richard Hellie (review). Technology and Culture, 13(2):310–312, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893471/pdf.

Eysenck:1974:RJR

[Eys74] H. J. Eysenck. Race by John R. Baker (review). Technology and Culture, 15(4):666–669, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892726/pdf.

Ezell:1977:PSC

[Eze77] Edward C. Ezell. The Politics of Space: a Comparison of the Soviet and American Space Programs by William H. Schauer (review). Technology and Culture, 18(3):559–561, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891973/pdf.

Ezell:1978:HFA

[Eze78] Edward C. Ezell. Harpers Ferry Armory and the New Technology: The Challenge of Change by Merritt Roe Smith (review). Technology and Culture, 19(1):119–121, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892052/pdf.

Fairbank:1971:CCC

[Fai71] John K. Fairbank. Clerks and Craftsmen in China and the West: Lectures and Addresses on the History of Science and Technology by Joseph Needham (review). Technology and Culture, 12(2):328– 331, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893963/pdf.

Feenberg:1979:TTC

[Fee79] Andrew Feenberg. Technology transfer and cultural change in Communist societies. Technology and Culture, 20(2):348–354, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890495/pdf.

Feinman:1977:ERC

[Fei77] Stephen Feinman. The Economics of Refuse Collection by Peter Kemper, John M. Quigley (review). Technology and Culture, 18(2):270–271, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891933/pdf.

Feibleman:1978:PPB

[Fei78] James K. Feibleman. PSA 1976. Proceedings of the 1976 Biennial Meeting of the Philosophy of Science Association ed. by Frederick Suppe, Peter D. Asquith (review). Technology and Culture, 19(4):793–794, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892212/pdf.

Feller:1974:DLT

[Fel74] Irwin Feller. The diffusion and location of technological change in the American cotton-textile industry, 1890–1970. Technology and Culture, 15(4):569–593, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892702/pdf.

Feller:1976:IBA

[Fel76a] Irwin Feller. Interaction between Agriculture and Industry: Case Studies of Farm Mechanisation and Industrialisation in Sweden and the United States, 1830–1930 by Jan Kuuse (review). Technology and Culture, 17(4):761–763, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891831/pdf.

Feller:1976:TDF

[Fel76b] Irwin Feller. Technology Diffusion: Federal Programs and Procedures by Granville W. Hough (review). Technology and Culture, 17(3):578– 580, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891792/pdf.

Fell:1977:VST

[Fel77] James E. Fell, Jr. Views of Science, Technology and Development ed. by Eugene Rabinowitch, Victor Rabinowitch (review). Technology and Culture, 18(3):580–581, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891985/pdf.

Felix:1978:USE

[Fel78] David Felix. The Use of Socio-Economic Indicators in Development Planning by UNESCO (review). Technology and Culture, 19(2):242– 244, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892092/pdf.

Fenton:1978:SSP

[Fen78] Alexander Fenton. Stone Shares of Ploughing Implements from the Bronze Age of Syria: a Contribution to the Early History of the Ard-Plough by Axel Steensberg (review). Technology and Culture, 19(3):514–515, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892121/pdf.

Ferguson:1970:JWS

[Fer70a] Eugene S. Ferguson. James Watt and the Separate Condenser: an Account of the Invention by R. J. Law (review). Technology and Culture, 11(4):627–628, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894055/pdf.

Ferkiss:1970:MCT

[Fer70b] Victor C. Ferkiss. Man and the Computer: Technology as an Agent of Social Change by John Diebold (review). Technology and Culture, 11(2):338–340, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894182/pdf.

Ferguson:1971:HST

[Fer71a] Eugene S. Ferguson. History of Science and Technology: a Select Bibliography for Students by K. J. Rider (review). Technology and Culture, 12(4):687, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893589/pdf.

Ferguson:1971:LTM

[Fer71b] Eugene S. Ferguson. Leupold’s theatrum machinarum: a need and an opportunity. Technology and Culture, 12(1):64–68, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894002/pdf.

Ferkiss:1971:THS

[Fer71c] Victor C. Ferkiss. Technology and Humanism: Some Exploratory Essays for Our Times by William G. Carleton (review). Technology and Culture, 12(3):536–538, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893622/pdf.

Fern:1972:PAE

[Fer72] Alan Fern. Prints in and of America to 1850 ed. by John D. Morse (review). Technology and Culture, 13(2):319–321, April 1972. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893480/pdf.

Ferguson:1973:AEB

[Fer73a] Eugene S. Ferguson. American Economic and Business History Information Sources by Robert W. Lovett (review). Technology and Culture, 14(1):113–115, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893374/pdf.

Ferguson:1973:STT

[Fer73b] Eugene S. Ferguson. Submarining: Three Thousand Books and Articles by T. O. Paine (review). Technology and Culture, 14(1):112–113, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893373/pdf.

Ferguson:1973:TPW

[Fer73c] Eugene S. Ferguson. Turning Points in Western Technology: a Study of Technology, Science and History by D. S. L. Cardwell (review). Technology and Culture, 14(3):477–479, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893218/pdf.

Ferguson:1974:TDH

[Fer74a] Eugene S. Ferguson. Toward a discipline of the history of technology. Technology and Culture, 15(1):13–30, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892821/pdf. See com-

ment [De 74e].

Ferwerda:1974:TAQ

[Fer74b] Vernon L. Ferwerda. Technology Assessment and Quality of Life ed. by Gerhard J. Stöber, Dieter Schumacher, and: Technology Assessment by John W. Dickey, et al., and: Environmental Quality and Social Responsibility ed. by Ravindra S. Khare, et al. (review). Technology and Culture, 15(3):529–534, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892769/pdf.

Ferguson:1976:BIB

[Fer76a] Eugene S. Ferguson. A Biographical Index of British Engineers in the 19th Century comp. by Samuel Peter Bell (review). Technology and Culture, 17(3):611–612, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891809/pdf.

Ferguson:1976:MTG

[Fer76b] Eugene S. Ferguson. Martha Teach Gnudi (1908–76). Technology and Culture, 17(3):521–522, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891762/pdf.

Ferkiss:1976:SEG

[Fer76c] Victor Ferkiss. Synergetics: Explorations in the Geometry of Thinking by R. Buckminster Fuller, E. J. Applewhite (review). Technology and Culture, 17(1):104–108, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891654/pdf.

Ferguson:1978:EIA

[Fer78a] Eugene S. Ferguson. Elegant inventions: The artistic component of technology. Technology and Culture, 19(3):450–460, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892112/pdf.

Ferkiss:1978:TSS

[Fer78b] Victor Ferkiss. Technology and Social Shock by Edward B. Lawless (review). Technology and Culture, 19(4):796–797, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892214/pdf.

Ferguson:1979:ANA

[Fer79a] Eugene S. Ferguson. The American-ness of American technology. Technology and Culture, 20(1):3–24, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890450/pdf.

Ferguson:1979:AEN

[Fer79b] Eugene S. Ferguson. American Engineers of the Nineteenth Century: a Biographical Index by Christine Roysdon, Linda A. Khatri (review). Technology and Culture, 20(3):681–682, July 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890564/pdf.

Ferguson:1979:TT

[Fer79c] Eugene S. Ferguson. Toys and turning. Technology and Culture, 20(3):691, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890569/pdf.

Field:1977:REH

[Fie77] Alexander J. Field. Research in Economic History: an Annual Compilation of Research ed. by Paul Uselding (review). Technology and Culture, 18(4):744–746, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892029/pdf.

Fitting:1971:WAE

[Fit71] James E. Fitting. World Archaeology ed. by Colin Platt (review). Technology and Culture, 12(1):159–161, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894040/pdf.

Fite:1973:HFG

[Fit73a] Gilbert C. Fite. Henry Ford and Grass-Roots America by Reynold M. Wik (review). Technology and Culture, 14(4):645–647, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892890/pdf.

Fitting:1973:MME

[Fit73b] James E. Fitting. The Minnesota Messenia Expedition: Reconstructing a Bronze Age Regional Environment ed. by William A. McDonald, George R. Rapp, Jr. (review). Technology and Culture, 14(3):487–488, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893222/pdf.

Fitzsimons:1976:R

[Fit76] Mayvis Fitzsimons. Right on! Technology and Culture, 17(3): 526, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891764/pdf.

Flajser:1975:EUO

[Fla75] Steven H. Flajser. Energy under the Oceans: a Technology Assessment of Outer Continental Shelf Oil and Gas Operations by D. E. Kash and I. L. White (review). Technology and Culture, 16 (1):122–124, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891524/pdf.

Fleron:1977:TCC

[Fle77] Frederic J. Fleron, Jr. Technology and Communist Culture: Bellagio, Italy, August 22–28, 1975. Technology and Culture, 18(4):659–665, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891995/pdf.

Flink:1972:ACS

[Fli72] James J. Flink. The American Car since 1775 by the Editors of Automobile Quarterly (review). Technology and Culture, 13(2):325– 326, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893482/pdf.

Florman:1970:R

[Flo70] Samuel C. Florman. Reply. Technology and Culture, 11(2):240, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894154/pdf.

Florman:1972:OCW

[Flo72] Samuel C. Florman. The One Culture by William H. Davenport (review). Technology and Culture, 13(2):343–344, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893493/pdf.

Fryer:1979:MJV

[FM79] David M. Fryer and John C. Marshall. The motives of Jacques de Vaucanson. Technology and Culture, 20(2):257–269, April 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890491/pdf.

Fores:1971:PTP

[For71] Michael Fores. Price, technology, and the paper model. Technology and Culture, 12(4):621–627, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893560/pdf.

Ford:1979:SSP

[For79a] J. Glyn Ford. Science and Society in Prehistoric Britain by Euan W. Mackie (review). Technology and Culture, 20(2):356–359, April 1979.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890497/pdf.

Fores:1979:HTA

[For79b] Michael Fores. The history of technology: an alternative view. Technology and Culture, 20(4):853–860, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890600/pdf.

Fowles:1977:HFH

[Fow77] Jib Fowles. The High Frontier: Human Colonies in Space by Gerard K. O’Neill (review). Technology and Culture, 18(4):718–720, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892017/pdf.

Fowles:1978:SFI

[Fow78] Jib Fowles. The Study of the Future: an Introduction to the Art and Science of Understanding and Shaping Tomorrow’s World by Edward Cornish (review). Technology and Culture, 19(3):592–593, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892158/pdf.

Fowles:1979:SCE

[Fow79] Jib Fowles. Space Colonies ed. by Stewart Brand (review). Technology and Culture, 20(1):240–241, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890484/pdf.

Frantz:1971:CMW

[Fra71] Joe B. Frantz. The City Moves West: Economic and Industrial Growth in Central West Texas by Robert L. Martin (review). Technology and Culture, 12(1):128–129, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894025/pdf.

Frantz:1973:ASP

[Fra73] Joe B. Frantz. Arthur E. Stilwell, Promoter with a Hunch by Keith L. Bryant, Jr. (review). Technology and Culture, 14(1):88–89, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893361/pdf.

Frantz:1974:LUR

[Fra74a] Joe B. Frantz. Land of the Underground Rain: Irrigation on the Texas High Plains, 1910–1970 by Donald E. Green (review). Technology and Culture, 15(2):339–340, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892804/pdf.

Frantz:1974:WSH

[Fra74b] Joe B. Frantz. Water for the Southwest: Historical Survey and Guide to Historic Sites by T. Lindsay Baker, et al. (review). Technology and Culture, 15(3):501–502, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892753/pdf.

Frantz:1976:TRL

[Fra76] Joe B. Frantz. Timber Roots: The Laird, Norton Story, 1855–1905 by Fred W. Kohlmeyer (review). Technology and Culture, 17(1):158, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891681/pdf.

Freeman:1971:IGW

[Fre71a] C. Freeman. Invention, Growth, and Welfare: a Theoretical Treatment of Technological Change by William D. Nordhaus (review). Technology and Culture, 12(2):375–376, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893988/pdf.

Freudenberger:1971:ETD

[Fre71b] Herman Freudenberger. L’ evolution des techniques du filage et du tissage du Moyen-Age à la Révolution Industrielle by Walter Endrei (review). Technology and Culture, 12(1):92–93, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894007/pdf.

Freudenberger:1973:SKG

[Fre73] Herman Freudenberger. Schuss und Kette: Geschichte der “Sulzer”- Webmaschine by Rudolf H. Rossmann (review). Technology and Culture, 14(2 (part I)):312–313, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/901455/pdf.

Freudenberger:1974:EME

[Fre74] Herman Freudenberger. Explorations in Managerial Economics by Bela Gold (review). Technology and Culture, 15(2):351–352, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892811/pdf.

Frejka:1976:DFG

[Fre76] Tomas Frejka. The Decline of Fertility in Germany, 1871–1939 by John E. Knodel (review). Technology and Culture, 17(2):389–390, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891733/pdf.

Fries:1975:BRA

[Fri75] Russell I. Fries. British response to the American system: The case of the small-arms industry after 1850. Technology and Culture, 16(3):377–403, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891562/pdf.

Fries:1976:SCS

[Fri76] Russell I. Fries. Samuel Colt’s Submarine Battery — the Secret and the Enigma by Philip K. Lundeberg (review). Technology and Culture, 17(1):148–149, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891674/pdf.

Fries:1977:IHM

[Fri77] Russell I. Fries. Inventing: How the Masters Did It by Byron M. Vanderbilt (review). Technology and Culture, 18(1):105–107, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891882/pdf.

Fries:1978:CUD

[Fri78a] Sylvia Doughty Fries. Colonial Urban Development: Culture, Social Power and Environment by Anthony D. King (review). Technology and Culture, 19(2):217–219, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892079/pdf.

Fries:1978:MSS

[Fri78b] Sylvia Doughty Fries. From Main Street to State Street: Town, City, and Community in America by Park Dixon Goist (review). Technology and Culture, 19(4):750–752, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892190/pdf.

Frick:1979:VJB

[Fri79a] George F. Frick. The Virginia Journals of Benjamin Henry Latrobe, 1795–1798 ed. by Edward T. Carter, Angelina Polites (review). Technology and Culture, 20(1):226–228, January 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890476/pdf.

Fries:1979:ATH

[Fri79b] Sylvia Doughty Fries. Anthropology toward History: Culture and Work in a 19th-Century Maine Town by Richard P. Horwitz (review). Technology and Culture, 20(2):367–369, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890503/pdf.

Fruchtbaum:1975:QO

[Fru75] Harold Fruchtbaum. Questionable opinions. Technology and Culture, 16(4):615, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891600/pdf.

Fischer:1978:IHT

[FS78] Robert Fischer and Amos St. Germain. INTERFACE ’77 — Humanities and Technology: Southern Technical Institute, Marietta, Georgia, October 20–22, 1977. Technology and Culture, 19(4):703–708, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892172/pdf.

Fukushima:1975:FNM

[Fuk75a] Y. Fukushima. Farm and Nation in Modern Japan: Agrarian Nationalism, 1870–1940 by Thomas R. H. Havens (review). Technology and Culture, 16(3):504–505, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891587/pdf.

Fukushima:1975:TWG

[Fuk75b] Y. Fukushima. Technik-Wirtschaft-Gesellschaft: Der Einfluss wirtschaftlicher und gesellschaftlicher Veränderungen auf die Entwicklung der landwirtschaftlichen Geräte in der vorindustriellen Epoche Japans ab dem 17. Jahrhundert by Eric Pauer (review). Technology and Culture, 16(4):628–630, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891606/pdf.

Fukashima:1977:SSM

[Fuk77] Y. Fukashima. Science and Society in Modern Japan: Selected Historical Sources ed. by Shigeru Nakayama, et al. (review). Technology and Culture, 18(3):574–576, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891981/pdf.

Fussell:1971:MWP

[Fus71] G. E. Fussell. Mary Wedlake’s Priced List of Modern Farm Implements, and: Sickle to Combine: a Review of Harvest Technique from 1800 to the Present Day by E. J. T. Collins (review). Technology and Culture, 12(2):350–351, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893974/pdf.

Fussell:1972:BRA

[Fus72] G. E. Fussell. A Bibliography of Roman Agriculture by K. D. White (review). Technology and Culture, 13(1):113, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893459/pdf.

Fussell:1973:AHE

[Fus73] G. E. Fussell. The Agrarian History of England and Wales ed. by H. P. R. Finberg (review). Technology and Culture, 14(3):491– 493, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893224/pdf.

Fussell:1974:BFG

[Fus74a] G. E. Fussell. British Farming in the Great Depression, 1870–1914: an Historical Geography by P. J. Perry (review). Technology and Culture, 15(4):645–646, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892716/pdf.

Fussell:1974:LAY

[Fus74b] G. E. Fussell. The Life of Arthur Young, 1741–1820 by John G. Gazley (review). Technology and Culture, 15(3):487–489, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892747/pdf.

Fussell:1976:FER

[Fus76a] G. E. Fussell. Farm Equipment of the Roman World by K. D. White (review). Technology and Culture, 17(3):541–543, July 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891771/pdf.

Fussell:1976:LQE

[Fus76b] G. E. Fussell. The Land Question and European Society by Frank E. Huggett (review). Technology and Culture, 17(2):370, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891723/pdf.

Gade:1971:GMH

[Gad71] Daniel W. Gade. Grist milling with the horizontal waterwheel in the central Andes. Technology and Culture, 12(1):43–51, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893999/pdf.

Gaines:1975:FPC

[Gai75] David I. Gaines. Five Per Cent Philanthropy: an Account of Housing in Urban Areas between 1840 and 1914 by John Nelson Tarn (review). Technology and Culture, 16(3):496–498, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891583/pdf.

Gale:1975:EEB

[Gal75] W. K. V. Gale. The Economic Emergence of the Black Country: a Study of the Dudley Estate by T. J. Raybould (review). Technology and Culture, 16(3):484–486, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891577/pdf.

Garvan:1972:RHB

[Gar72a] Anthony N. B. Garvan. Recording Historic Buildings comp. by Harley J. McKee (review). Technology and Culture, 13(1):113–114, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893460/pdf.

Garvan:1972:WPE

[Gar72b] Anthony N. B. Garvan. Winterthur Portfolio 6 ed. by Richard K. Doud, Ian M. G. Quimby (review). Technology and Culture, 13(1):85–87, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893446/pdf.

Garside:1978:WFU

[Gar78] Patricia Garside. Work and Family in the United States: a Critical Review and Agenda for Research and Policy by Rosabeth Moss Kanter (review). Technology and Culture, 19(1):129–131, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892058/pdf.

Gates:1973:HLN

[Gat73] Paul. W. Gates. History of the Louisville & Nashville Railroad by Maury Klein (review). Technology and Culture, 14(4):637–638, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892884/pdf.

Garcia-Diego:1972:RTM

[GD72] José A. Garcia-Diego. Restoration of technological monuments in Spain. Technology and Culture, 13(3):426–429, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893925/pdf.

Garcia-Diego:1973:OWB

[GD73] Jose A. Garcia-Diego. The Old World Background of the Irrigation System of San Antonio, Texas by Thomas F. Glick (review). Technology and Culture, 14(4):622–624, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892877/pdf.

Garcia-Diego:1976:CWC

[GD76] J. A. Garcia-Diego. The chapter on weirs in the Codex of Juanelo Turriano: a question of authorship. Technology and Culture, 17 (2):217–234, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891712/pdf.

Gebhard:1971:RSK

[Geb71] Bruno Gebhard. Die Reale Stadt und die Krankheit im 17. Jahrhundert by Wilhelm Treue (review). Technology and Culture, 12(1): 127–128, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894024/pdf.

Gebhard:1972:GAB

[Geb72] Bruno Gebhard. Geschichte der Arzneimittelforschung by Béla Issekutz (review). Technology and Culture, 13(1):109–111, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893457/pdf.

Gerstner:1974:TMH

[Ger74] Patsy A. Gerstner. Temples of the Muses and a History of Pharmacy Museums by Sami K. Hamarneh (review). Technology and Culture, 15(1):156–157, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892864/pdf.

Germain:1976:PUH

[Ger76] Paul Germain. On Parallels in Universal History Discoverable in Arts and Artifacts: an Outline Statement by Alan Gowans (review). Technology and Culture, 17(1):191–193, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891699/pdf.

Gerstner:1979:MRT

[Ger79] Patsy A. Gerstner. Medicine and the Reign of Technology by Stanley Joel Reiser (review). Technology and Culture, 20(1):214–216, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890470/pdf.

Ghaswala:1970:EAI

[Gha70a] S. K. Ghaswala. European Architecture in India, 1750–1850 by Sten Nilsson (review). Technology and Culture, 11(4):642–643, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894065/pdf.

Ghaswala:1970:SHC

[Gha70b] S. K. Ghaswala. Science and the Human Condition in India and Pakistan ed. by Ward Morehouse (review). Technology and Culture, 11(2):333–334, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894179/pdf.

Ghaswala:1973:IST

[Gha73a] S. K. Ghaswala. Indian Science and Technology in the Eighteenth Century by Dharampal (review). Technology and Culture, 14(2 (part I)):293–294, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893267/pdf.

Ghaswala:1973:SIW

[Gha73b] S. K. Ghaswala. Science in India by Ward Morehouse (review). Technology and Culture, 14(2 (part I)):351–352, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893292/pdf.

Giannetti:1972:DCT

[Gia72] Louis D. Giannetti. A Discovery of Cinema by Thorold Dickinson (review). Technology and Culture, 13(2):340–343, April 1972. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893492/pdf.

Gies:1976:HWE

[Gie76a] Joseph Gies. Humanizing the Workplace ed. by Roy P. Fairfield (review). Technology and Culture, 17(3):583–585, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891795/pdf.

Gies:1976:QWL

[Gie76b] Joseph Gies. The Quality of Working Life ed. by Louis E. Davis, Albert B. Cherns (review). Technology and Culture, 17(3):581–583, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891794/pdf.

Gies:1976:WPI

[Gie76c] Joseph Gies. Worker’s Participation in Industry by Michael Poole (review). Technology and Culture, 17(1):172–173, January 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891689/pdf.

Gies:1977:MWL

[Gie77] Joseph Gies. Man and Work: Literature and Culture in Industrial Society by David Meakin (review). Technology and Culture, 18(3):563–565, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891975/pdf.

Giebelhaus:1978:ETO

[Gie78] August W. Giebelhaus. Early Texas Oil: a Photographic History, 1866–1936 by Walter Rundell, Jr. (review). Technology and Culture, 19(1):125–126, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892055/pdf.

Gies:1979:FWG

[Gie79] Joseph Gies. The Flight from Work by Göran Palm (review). Technology and Culture, 20(3):668–670, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890557/pdf.

Gilfillan:1970:TFP

[Gil70] S. C. Gilfillan. Technological Forecasting: a Practical Approach by Marvin J. Cetron (review). Technology and Culture, 11(4):659–661, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894073/pdf.

Gilfillan:1972:IST

[Gil72a] S. Colum Gilfillan. Innovations: Scientific, Technological, and Social by Dennis Gabor (review). Technology and Culture, 13(1):107–109, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893456/pdf.

Gillmor:1972:SEA

[Gil72b] C. Stewart Gillmor. Sir Edward Appleton by Ronald Clark (review). Technology and Culture, 13(4):670–672, October 1972. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893405/pdf.

Gillmor:1973:CMS

[Gil73a] C. Stewart Gillmor. Coulomb’s Memoir on Statics: an Essay in the History of Civil Engineering by Jacques Heyman (review). Technology and Culture, 14(3):500–501, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893229/pdf.

Gillmor:1973:JFG

[Gil73b] C. Stewart Gillmor. Joseph Fourier, 1768–1830 by I. Grattan-Guinness (review). Technology and Culture, 14(3):501–503, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/893230/
pdf.

Gillmor:1973:MMO

[Gil73c] Carroll Gillmor. Merovingian Military Organization, 481–751 by Bernard S. Bachrach (review). Technology and Culture, 14(3):497– 499, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893227/pdf.

Gilfillan:1979:ABH

[Gil79] S. Colum Gilfillan. The ancient breast harness: a suggestion for research. Technology and Culture, 20(2):355, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890496/pdf.

Glass:1974:MPE

[Gla74] Stephen L. Glass. Methods of Physical Examination in Archaeology by M. S. Tite (review). Technology and Culture, 15(1):153–155, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892862/pdf.

Glasow:1978:NTS

[Gla78] Richard D. Glasow. Naval Technology and Social Modernization in the 19th Century by Ken J. Hagan, et al. (review). Technology and Culture, 19(3):556–557, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892144/pdf.

Glick:1970:BHC

[Gli70a] Thomas F. Glick. Bibliografía histórica sobre la ciencia y la técnica en España ed. by José M. López Piñero, et al. (review). Technology and Culture, 11(1):131, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894211/pdf.

Glick:1970:CCM

[Gli70b] Thomas F. Glick. Un clima para la ciencia by Manuel Lora Tamayo, and: El Opus Dei en España by Daniel Artigues, and: National Reports of the Pilot Teams, Science and Development, Spain by Organisation for Economic Co-operation and Development, and: La educación en España: Bases para una política educativa by Ministerio de Educación y Ciencia (review). Technology and Culture, 11 (1):113–118, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894204/pdf.

Glick:1973:IDC

[Gli73a] Thomas F. Glick. La indústria dels claus a Ripoll. Contribució a l’estudi de la farga catalana by Eudald Graells (review). Technology and Culture, 14(3):499–500, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893228/pdf.

Glick:1973:SHM

[Gli73b] Thomas F. Glick. The Sugar Hacienda of the Marqueses del Valle by Ward J. Barrett (review). Technology and Culture, 14(2 (part I)):292–293, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893266/pdf.

Glick:1974:HEN

[Gli74] Thomas F. Glick. Un héroe español del progreso: Agustín de Betancourt by Alekséi Bogoliúbov (review). Technology and Culture, 15(1):97–98, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892836/pdf.

Glick:1975:ICJ

[Gli75a] Thomas F. Glick. El “ingenio” creado por Juan de Herrera para cortar hierro by Luís Cervera Vera, and: Máquinas hidraúlicas de molinos y herrerías y govierno de los árboles y montes de Vizcaya by Pedro Bernardo Villareal de Berriz, and: El Pantano de Tibi by Rafael Viravens Pastor (review). Technology and Culture, 16(2):293– 294, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891538/pdf.

Glick:1975:MHI

[Gli75b] Thomas F. Glick. La minería hispana e iberoamericana. Contribución a su investigación histórica. Estudios, fuentes, bibliografia ed. by Antonio del Valle Menéndez (review). Technology and Culture, 16(2):291–293, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891537/pdf.

Glick:1977:NPS

[Gli77] Thomas F. Glick. Noria pots in Spain. Technology and Culture, 18(4):644–650, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891992/pdf.

Glick:1978:MVT

[Gli78] Thomas F. Glick. Molinos de viento en tierras de Alicante by Francisco G. Seijo Alonso (review). Technology and Culture, 19(3):516, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892122/pdf.

Glick:1979:LDS

[Gli79] Thomas F. Glick. … And the Desert Shall Rejoice: Conflict, Growth, and Justice in Arid Environments by Arthur Maass, Raymond L. Anderson (review). Technology and Culture, 20(3):652–655, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890549/pdf.

Gnudi:1974:ARA

[Gnu74] Martha Teach Gnudi. Agostino ramelli and ambroise bachot. Technology and Culture, 15(4):614–625, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892704/pdf.

Goist:1970:UDJ

[Goi70] Park Dixon Goist. Urban Dynamics by Jay W. Forrester (review). Technology and Culture, 11(1):109–111, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894202/pdf.

Goist:1971:GBC

[Goi71] Park Dixon Goist. Green-Belt Cities by Frederic J. Osborn (review). Technology and Culture, 12(1):130–131, January 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894026/pdf.

Goist:1973:LMB

[Goi73] Park Dixon Goist. Lewis Mumford: a Bibliography, 1914–1970 ed. by Elmer S. Newman (review). Technology and Culture, 14(2 (part I)):359–360, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893296/pdf.

Goldstein:1970:CAA

[Gol70a] Walter Goldstein. The Chasm Ahead by Aurelio Peccei (review). Technology and Culture, 11(3):471–473, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894128/pdf.

Goldstein:1970:TCW

[Gol70b] Walter Goldstein. Tongues of Conscience: Weapons Research and the Scientists’ Dilemma by R. W. Reid, and: Criteria for Scientific Development: Public Policy and National Goals ed. by Edward Shils (review). Technology and Culture, 11(4):643–645, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894066/pdf.

Goldfarb:1971:RGB

[Gol71] Stephen J. Goldfarb. A regency gas burner. Technology and Culture, 12(3):471–476, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893598/pdf.

Goldfarb:1974:HSH

[Gol74a] Stephen Goldfarb. The Housewares Story: a History of the American Housewares Industry by Earl Lifshey (review). Technology and Culture, 15(3):502–503, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892754/pdf.

Goldstein:1974:CDE

[Gol74b] Walter Goldstein. Cultural Development: Experience and Policies by Augustin Girard (review). Technology and Culture, 15(1):142–144, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892857/pdf.

Goldberg:1975:UFH

[Gol75] Maxwell H. Goldberg. Uses of the Future by Herbert J. Muller (review). Technology and Culture, 16(4):693–696, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891638/pdf.

Goldfarb:1976:LFJ

[Gol76] Stephen Goldfarb. The Lambent Flame by John D. Keating (review). Technology and Culture, 17(1):149–150, January 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891675/pdf.

Goldfarb:1977:MMS

[Gol77a] Stephen Goldfarb. Memories of a Mountain Shortline: The Story of the Tallulah Falls Railroad ed. by Kaye Carver, Myra Queen (review). Technology and Culture, 18(3):553–554, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891969/pdf.

Goldfarb:1977:TPE

[Gol77b] Stephen Goldfarb. Tube of Plenty: The Evolution of American Television by Erik Barnouw (review). Technology and Culture, 18 (1):116–117, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891887/pdf.

Goldstein:1977:MCR

[Gol77c] Walter Goldstein. The Multinational Corporation and the Resource Cost of International Technology Transfer by David J. Teece (review). Technology and Culture, 18(4):735–736, October 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892025/pdf.

Goldfarb:1978:IAS

[Gol78] Stephen J. Goldfarb. The Industrial Archaeology of the Stationary Steam Engine by R. A. Buchanan, George Watkins (review). Technology and Culture, 19(1):116–117, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892050/pdf.

Goodman:1970:DHB

[Goo70a] W. L. Goodman. Drilling holes with the bow-drill. Technology and Culture, 11(1):69, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894191/pdf.

Goodwin:1970:CIR

[Goo70b] Jack Goodwin. Catalog of the Inland Rivers Library comp. by Clyde N. Bowden (review). Technology and Culture, 11(1):130–131, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894210/pdf.

Goodwin:1970:CBH

[Goo70c] Jack Goodwin. Current bibliography in the history of technology (1968). Technology and Culture, 11(2):242–288, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894156/pdf.

Goodwin:1971:CBH

[Goo71] Jack Goodwin. Current bibliography in the history of technology (1969). Technology and Culture, 12(2):269–327, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893962/pdf.

Goodwin:1972:CBH

[Goo72] Jack Goodwin. Current bibliography in the history of technology (1970). Technology and Culture, 13(2):226–294, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893466/pdf.

Goodwin:1973:CBH

[Goo73] Jack Goodwin. Current bibliography in the history of technology (1971). Technology and Culture, 14(2 (part I)):175–286, April 1973.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901456/pdf.

Goodwin:1974:USA

[Goo74a] Jack Goodwin. United States Air Force History: a Guide to Documentary Sources comp. by Lawrence J. Paszek (review). Technology and Culture, 15(4):675, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892730/pdf.

Goodwin:1974:CBH

[Goo74b] Jack Goodwin. Current bibliography in the history of technology (1972). Technology and Culture, 15(2):246–324, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892795/pdf.

Goodwin:1975:CBH

[Goo75] Jack Goodwin. Current bibliography in the history of technology (1973). Technology and Culture, 16(2):195–286, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891534/pdf.

Goodwin:1976:CBH

[Goo76] Jack Goodwin. Current bibliography in the history of technology (1974). Technology and Culture, 17(2):286–364, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891720/pdf.

Goodwin:1977:CBH

[Goo77] Jack Goodwin. Current bibliography in the history of technology (1975). Technology and Culture, 18(2):272–390, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891934/pdf.

Goodwin:1978:CBH

[Goo78] Jack Goodwin. Current bibliography in the history of technology (1976). Technology and Culture, 19(2):268–356, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892106/pdf.

Goodwin:1979:CBH

[Goo79] Jack Goodwin. Current bibliography in the history of technology (1977). Technology and Culture, 20(2):403–514, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890521/pdf.

Gorman:1971:SWO

[Gor71a] Mel Gorman. Sir William O’Shaughnessy, Lord Dalhousie, and the establishment of the telegraph system in India. Technology and Culture, 12(4):581–601, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893557/pdf.

Gorr:1971:ACF

[Gor71b] Louis F. Gorr. American Civilization in the First Machine Age, 1890–1940 by Gilman Ostrander (review). Technology and Culture, 12(3):508–511, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893608/pdf.

Gorr:1972:MII

[Gor72] Louis F. Gorr. Mirror Image: The Influence of the Daguerreotype on American Culture by Richard Rudisill (review). Technology and Culture, 13(4):631–635, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893385/pdf.

Gorr:1973:TTP

[Gor73] Louis F. Gorr. Think Tanks by Paul Dickson (review). Technology and Culture, 14(3):522–525, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893241/pdf.

Gorman:1974:HNW

[Gor74] Mel Gorman. Historical Notes on the Wet-processing Industry by Sidney M. Edelstein (review). Technology and Culture, 15(2):328– 329, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892798/pdf.

Gorman:1979:SME

[Gor79] Mel Gorman. The Sidney M. Edelstein Collection: From Alchemy to Atoms by Moshe Ron (review). Technology and Culture, 20(3):682– 685, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890565/pdf.

Grabar:1972:HOA

[Gra72] Oleg Grabar. A History of Ottoman Architecture by Godfrey Goodwin (review). Technology and Culture, 13(1):67–69, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893435/pdf.

Graham:1975:VIE

[Gra75] Loren Graham. Voprosy istorii estestvoznnniia i tekhniki [Problems of the history of science and technology] (review). Technology and Culture, 16(2):328–329, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891555/pdf.

Green:1972:HTM

[Gre72] Peter Green. Historic Towns: Maps and Plans of Towns and Cities in the British Isles with Historical Commentaries, from Earliest Times to circa 1800 ed. by M. D. Lobel (review). Technology and Culture, 13(2):335–336, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893489/pdf.

Greenberg:1973:UMR

[Gre73] Daniel Greenberg. The University and Military Research: Moral Politics at M.I.T. by Dorothy Nelkin (review). Technology and Culture, 14(1):99–100, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893369/pdf.

Green:1979:LTR

[Gre79] E. R. R. Green. Life and Tradition in Rural Ireland by Timothy O’Neill (review). Technology and Culture, 20(3):638–639, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890542/pdf.

Gross:1972:AHE

[Gro72] Walter E. Gross. The American Health Empire by Barbara, John Ehrenreich (review). Technology and Culture, 13(2):333–335, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893488/pdf.

Gruender:1971:DST

[Gru71] C. David Gruender. On distinguishing science and technology. Technology and Culture, 12(3):456–463, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893595/pdf.

Guest:1973:WAT

[Gue73] Robert H. Guest. Workers’ Attitudes and Technology by Dorothy Wedderburn, Rosemary Crompton (review). Technology and Culture, 14(3):515–517, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893237/pdf.

Guest:1977:BCS

[Gue77] Robert H. Guest. Blue Collar Stratification: Autoworkers in Four Countries by William H. Form (review). Technology and Culture, 18(1):124–125, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891892/pdf.

Haber:1972:RES

[Hab72] Samuel Haber. The Revolt of the Engineers: Social Responsibility and the American Engineering Profession by Edwin T. Layton, Jr. (review). Technology and Culture, 13(1):100–104, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893453/pdf.

Hacker:1971:ARH

[Hac71] Barton C. Hacker. Adventures in Research: a History of Ames Research Center, 1940–1965 by Edwin P. Hartman (review). Technology and Culture, 12(4):669–671, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893581/pdf.

Hacker:1972:SSS

[Hac72] Barton C. Hacker. Secret Sentries in Space by Philip J. Klass (review). Technology and Culture, 13(4):678–680, October 1972. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893410/pdf.

Hacker:1973:PPG

[Hac73a] Barton C. Hacker. Project Paperclip: German Scientists and the Cold War by Clarence G. Lasby (review). Technology and Culture, 14(4):659–660, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892898/pdf.

Hacker:1973:WST

[Hac73b] Barton C. Hacker. Ways to Spaceflight: Translation of “Wege zur Raumschiffahrt,” R. Oldenbourg Verlag, Munich–Berlin, 1929 by Hermann Oberth (review). Technology and Culture, 14(4):657–659, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892897/pdf.

Hacker:1974:IRS

[Hac74] Barton C. Hacker. The idea of rendezvous: From space station to orbital operations in space-travel thought, 1895–1951. Technology and Culture, 15(3):373–388, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892735/pdf.

Hacker:1977:WWM

[Hac77] Barton C. Hacker. The weapons of the West: Military technology and modernization in 19th-Century China and Japan. Technology and Culture, 18(1):43–55, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891863/pdf.

Hacker:1978:EHR

[Hac78] Barton C. Hacker. Essays on the History of Rocketry and Astronautics: Proceedings of the Third through the Sixth History Symposia of the International Academy of Astronautics ed. by R. Cargill Hall (review). Technology and Culture, 19(4):775–777, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892204/pdf.

Hacker:1979:DMH

[Hac79] Barton C. Hacker. Destination Moon: a History of the Lunar Orbiter Program by Bruce K. Byers (review). Technology and Culture, 20(2):381–382, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890510/pdf.

Haden:1970:JGS

[Had70] H. J. Haden. Journal of Glass Studies by Corning Museum of Glass (review). Technology and Culture, 11(4):664–667, October 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894075/pdf.

Haden:1971:JGS

[Had71] H. J. Haden. Journal of Glass Studies by Corning Museum of Glass (review). Technology and Culture, 12(4):681–684, October 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893587/pdf.

Haden:1972:GFC

[Had72] H. Jack Haden. The Glass of Frederick Carder by Paul V. Gardner (review). Technology and Culture, 13(4):656–659, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893398/pdf.

Haden:1973:JGS

[Had73] H. J. Haden. Journal of Glass Studies (review). Technology and Culture, 14(1):110–112, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893372/pdf.

Haden:1974:JGS

[Had74a] H. J. Haden. Journal of Glass Studies (review). Technology and Culture, 15(2):367–369, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892818/pdf.

Haden:1974:WCM

[Had74b] H. J. Haden. William Caslon: Master of Letters by Johnson Ball (review). Technology and Culture, 15(2):329–332, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892799/pdf.

Haden:1975:JGS

[Had75a] H. J. Haden. Journal of Glass Studies ed. by Jane Shadel Spillman (review). Technology and Culture, 16(4):619–621, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891603/pdf.

Haden:1975:RGH

[Had75b] H. J. Haden. Readings in Glass History ed. by Anita Engle (review). Technology and Culture, 16(3):479–481, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891574/pdf.

Haden:1976:JGS

[Had76a] H. J. Haden. Journal of Glass Studies by Corning Museum of Glass (review). Technology and Culture, 17(3):601–602, July 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891804/pdf.

Haden:1976:RGH

[Had76b] H. J. Haden. Readings in Glass History ed. by Anita Engle (review). Technology and Culture, 17(2):412–413, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891745/pdf.

Haden:1976:DEG

[Had76c] H. J. Haden. The Development of English Glassmaking, 1560– 1640 by Eleanor S. Godfrey (review). Technology and Culture, 17 (4):788–790, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891845/pdf.

Haden:1977:JGS

[Had77a] H. J. Haden. Journal of Glass Studies (review). Technology and Culture, 18(4):746–749, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892030/pdf.

Haden:1977:RGH

[Had77b] H. J. Haden. Readings in Glass History ed. by Anita Engle (review). Technology and Culture, 18(3):520–522, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891952/pdf.

Haden:1977:VHC

[Had77c] H. J. Haden. The Victoria History of the County of Stafford ed. by M. W. Greenslade (review). Technology and Culture, 18(3):526– 528, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891955/pdf.

Haden:1978:JGS

[Had78a] H. J. Haden. Journal of Glass Studies (review). Technology and Culture, 19(4):728–730, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892178/pdf.

Haden:1978:RGH

[Had78b] H. J. Haden. Readings in Glass History ed. by Anita Engle (review). Technology and Culture, 19(3):548–550, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892139/pdf.

Haden:1978:GPC

[Had78c] H. J. Haden. The Glassmakers: Pilkington, 1826–1976 by T. C. Barker (review). Technology and Culture, 19(4):765–767, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892198/pdf.

Haden:1978:SET

[Had78d] H. J. Haden. The Steam Engine of Thomas Newcomen by L. T. C. Rolt, J. S. Allen (review). Technology and Culture, 19(3):550– 551, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892140/pdf.

Haden:1979:VAO

[Had79] H. J. Haden. Vickers: Against the Odds, 1956–77 by Harold Evans (review). Technology and Culture, 20(3):672–674, July 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890559/pdf.

Haines:1979:EDP

[Hai79] Michael R. Haines. Economic Development in the Philadelphia Region, 1810–1850 by Diane Lindstrom (review). Technology and Culture, 20(3):642–644, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890545/pdf.

Hall:1970:RBA

[Hal70] Bert S. Hall. A revolving bookcase by agostino ramelli. Technology and Culture, 11(3):389–400, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894102/pdf.

Hall:1971:LTB

[Hal71] Bert S. Hall. Leitbilder der Technik; Biblische Tradition und technischer Fortschritt by Ansgar Stöcklein (review). Technology and Culture, 12(3):495–498, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893603/pdf.

Hall:1972:SER

[Hal72a] A. Rupert Hall. The Social and Economic Roots of Newton’s “Principia” by B. Hessen (review). Technology and Culture, 13(4):629–630, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893383/pdf.

Hall:1972:DBF

[Hal72b] Bert S. Hall. Deutsche Buchdrucker des fünfzehnten Jahrhunderts ed. by Carl Wehmer (review). Technology and Culture, 13(4):627–629, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893382/pdf.

Hall:1972:KBZ

[Hal72c] Bert S. Hall. Kulturelle Beziehungen zwischen Nurnberg und den Deutschen im Südosten im 14. bis 16. Jahrhundert by Ute Monika Schwob (review). Technology and Culture, 13(2):309–310, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893470/pdf.

Hall:1972:SOS

[Hal72d] Bert S. Hall. Scriptorum Opus: Schreiber–Mönche am Werk ed. by Fridolin Dressier, and: Prof. Dr. Otto Meyer zum 65. Geburtstag am 21. September 1971 ed. by Gerd Zimmerman (review). Technology and Culture, 13(3):494–495, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893932/pdf.

Hallion:1973:HAA

[Hal73] Richard P. Hallion. Hap Arnold: Architect of American Air Power by Flint O. DuPre, and: Jimmy Doolittle: Daredevil Aviator and Scientist by Carroll V. Clines (review). Technology and Culture, 14 (4):653–655, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892894/pdf.

Hall:1974:FIS

[Hal74] R. Cargill Hall. Fifth international symposium on the history of rocketry and astronautics: Brussels, September 1971. Technology and Culture, 15(1):70–75, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892826/pdf.

Hall:1976:NST

[Hal76] A. Rupert Hall. Natural Science and Technology on the Road to the “Industrial Revolution” by Werner Plum (review). Technology and Culture, 17(1):127–128, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891664/pdf.

Hamill:1971:BDM

[Ham71] Hugh M. Hamill, Jr. Bureaucracy and Development: a Mexican Case Study by Martin H. Greenberg (review). Technology and Culture, 12(2):366–368, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893983/pdf.

Hamarneh:1978:STD

[Ham78] Sami K. Hamarneh. Science and Technology in the Development of the Arab States by UNESCO (review). Technology and Culture, 19 (4):800–803, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892216/pdf.

Hanieski:1973:AEV

[Han73] John F. Hanieski. The airplane as an economic variable: Aspects of technological change in aeronautics, 1903–1955. Technology and Culture, 14(4):535–552, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892867/pdf.

Hansen:1979:GDT

[Han79a] Bert Hansen. Gothic Design Techniques: The Fifteenth-Century Design Booklets of Mathes Roriczer and Hanns Schmuttermayer by Lon R. Shelby (review). Technology and Culture, 20(3):627–630, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890535/pdf.

Hansen:1979:JDA

[Han79b] Bert Hansen. John Dee on Astronomy: Propaedeumata Aphoristica (1558 and 1568), Latin and English ed. by Wayne Shumaker (review). Technology and Culture, 20(4):790–792, October 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890577/pdf.

Hartenberg:1970:FTR

[Har70] R. S. Hartenberg. Zur Frühgeschichte der technischen Regelungen by Otto Mayr (review). Technology and Culture, 11(2):321–324, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894173/pdf.

Harris:1971:VET

[Har71a] Elizabeth Harris. A View of Early Typography by Harry Carter (review). Technology and Culture, 12(2):342–344, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893970/pdf.

Hartenberg:1971:BTE

[Har71b] R. S. Hartenberg. Blätter für Technikgeschichte ed. by Rolf Niederhuemer (review). Technology and Culture, 12(4):684–686, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893588/pdf.

Hartenberg:1971:BWB

[Har71c] Richard S. Hartenberg. Buch und Wissenschaft: Beispiele aus der Geschichte der Medizin, Naturwissenschaft und Technik ed. by Eberhard Schmauderer (review). Technology and Culture, 12(2):395–397, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893996/pdf.

Hartenberg:1973:BTV

[Har73a] Richard S. Hartenberg. Blätter für Technikgeschichte. Vol. 32/33 (1970/71) ed. by Rolf Niederhuemer (review). Technology and Culture, 14(2 (part I)):356–358, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893295/pdf.

Hartenberg:1973:KMR

[Har73b] Richard S. Hartenberg. Die Kunst des Metalldrehens bei den Römern by Alfred Mutz (review). Technology and Culture, 14(3):488–491, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893223/pdf.

Harris:1974:RMH

[Har74a] John R. Harris. A Report of the Mohawk–Hudson Area Survey ed. by Robert M. Vogel (review). Technology and Culture, 15(3):534– 536, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892770/pdf.

Hartenberg:1974:BTE

[Har74b] Richard S. Hartenberg. Blätter für Technikgeschichte ed. by Rolf Niederhuemer (review). Technology and Culture, 15(4):675–677, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892731/pdf.

Harris:1975:EME

[Har75a] J. R. Harris. Economic Maturity and Entrepreneurial Decline: British Iron and Steel, 1870–1913 by Donald H. McCloskey (review). Technology and Culture, 16(1):94–97, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891509/pdf.

Harris:1975:STG

[Har75b] J. R. Harris. Svedenstierna’s Tour of Great Britain, 1802–3: The Travel Diary of an Industrial Spy by Eric T. Svedenstierna (review). Technology and Culture, 16(2):297–299, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891540/pdf.

Harris:1976:STB

[Har76a] J. R. Harris. By the Sweat of Thy Brow: Work in the Western World by Melvin Kranzberg, Joseph Gies (review). Technology and Culture, 17(1):170–172, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891688/pdf.

Hartenberg:1976:BTE

[Har76b] R. S. Hartenberg. Blätter für Technikgeschichte ed. by Rolf Niederhuemer (review). Technology and Culture, 17(1):193–196, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891700/pdf.

Hartenberg:1977:BTE

[Har77] Richard S. Hartenberg. Blätter für Technikgeschichte ed. by Rolf Niederhuemer (review). Technology and Culture, 18(4):742–744, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892028/pdf.

Harris:1978:BCI

[Har78a] J. R. Harris. The British Coalmining Industry: Retrospect and Prospect by A. R. Griffin (review). Technology and Culture, 19 (2):228–230, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892085/pdf.

Hartenberg:1978:PMH

[Har78b] Richard S. Hartenberg. Perpetual Motion: The History of an Obsession by Arthur W. J. G. Ord-Hume (review). Technology and Culture, 19(3):518–522, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892124/pdf.

Harris:1979:GBI

[Har79a] John R. Harris. The Growth of British Industry by A. E. Musson (review). Technology and Culture, 20(4):797–798, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890579/pdf.

Harrison:1979:SCT

[Har79b] Arthur P. Harrison, Jr. Single-control tuning: an analysis of an innovation. Technology and Culture, 20(2):296–321, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890493/pdf.

Hausman:1976:GBH

[Hau76] William J. Hausman. Great Britain and Her World, 1750–1914: Essays in Honour of W. O. Henderson ed. by Barrie M. Ratcliffe (review). Technology and Culture, 17(4):795–797, October 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891849/pdf.

Hayden:1979:AUD

[Hay79] Dolores Hayden. Architecture and Utopia: Design and Capitalist Development by Manfredo Tafuri (review). Technology and Culture, 20(1):203–204, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890463/pdf.

Heald:1971:HEB

[Hea71] Morrell Heald. The Human Environment and Business by Henry Ford II (review). Technology and Culture, 12(2):386–388, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893993/pdf.

Heaton:1972:HWW

[Hea72] Herbert Heaton. The History of Wool and Wool Combing by James Burnley (review). Technology and Culture, 13(3):495–498, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893933/pdf.

Heal:1975:EAS

[Hea75] Anthony S. Heal. English and American Steam Carriages and Traction Engines by William Fletcher (review). Technology and Culture, 16(2):304–306, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891543/pdf.

Hecht:1970:SHM

[Hec70] Arthur Hecht. A Short History of the Mail Service by Carl H. Scheele (review). Technology and Culture, 11(4):640–641, October 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894063/pdf.

Hecht:1973:AME

[Hec73] Arthur Hecht. The American Mail: Enlarger of the Common Life by Wayne E. Fuller (review). Technology and Culture, 14(4):628–630, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892880/pdf.

Heilbron:1970:RBL

[Hei70] J. L. Heilbron. Rutherford and Boltwood: Letters on Radioactivity ed. by Lawrence Badash (review). Technology and Culture, 11(3):460– 462, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894123/pdf.

Heilbron:1977:RCC

[Hei77a] J. L. Heilbron. Rutherford Correspondence Catalog comp. by Lawrence Badash, and: Guide to the Robert Andrews Millikan Collection at the California Institute of Technology comp. by Albert F. Gunns, Judith Goodstein (review). Technology and Culture, 18 (4):749–750, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892031/pdf.

Heimann:1977:JCM

[Hei77b] P. M. Heimann. James Clerk Maxwell: Physicist and Natural Philosopher by C. W. F. Everitt (review). Technology and Culture, 18(2):263–264, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891930/pdf.

Heller:1971:FTT

[Hel71] Terry Sovel Heller. Factors in the Transfer of Technology ed. by William H. Gruber, Donald C. Marquis, and: The Role of the Federal Agencies in Technology Transfer by Samuel I. Doctors (review). Technology and Culture, 12(2):370–372, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893985/pdf.

Hellie:1973:MMC

[Hel73] Richard Hellie. Muscovite military change revisited. Technology and Culture, 14(2 (part I)):170–172, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893260/pdf.

Hellinger:1977:BDD

[Hel77] Douglas Hellinger. Beyond Dependency: The Developing World Speaks Out ed. by Guy F. Erb, Valeriana Kallab (review). Technology and Culture, 18(3):582–584, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891986/pdf.

Henderson:1976:STK

[Hen76a] Gregory Henderson. Science and Technology in Korea by Sang-woon Jeon (review). Technology and Culture, 17(2):404–406, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891741/pdf.

Henretta:1976:NAW

[Hen76b] James A. Henretta. The North Atlantic World in the Seventeenth Century by K. G. Davies, and: Empires to Nations: Expansion in America, 1713–1824 by Max Savelle (review). Technology and Culture, 17(1):130–132, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891666/pdf.

Herlea:1975:CES

[Her75] Alexandre Herlea. La conquête des énergies: Les sept révolutions énergétiques by André Varagnac (review). Technology and Culture, 16(1):79–80, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891501/pdf.

Hewlett:1976:BDN

[Hew76] Richard G. Hewlett. Beginnings of development in nuclear technology. Technology and Culture, 17(3):465–478, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891753/pdf.

Heydinger:1974:DCJ

[Hey74] Earl J. Heydinger. Delaware Canal Journal: a Definitive History of the Canal and the River Valley through Which It Flows by C. P. “Bill” Yoder (review). Technology and Culture, 15(1):106–108, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892841/pdf.

Higham:1973:SCL

[Hig73] Robin Higham. Starlifter: The C-141, Lockheed’s High-Speed Flying Truck by Harold H. Martin (review). Technology and Culture, 14(4):655, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892895/pdf.

Higham:1977:RRP

[Hig77] Robin Higham. The Riddle of Richard Pearse by Gordon Ogilvie (review). Technology and Culture, 18(1):119–120, January 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891889/pdf.

Higham:1979:AAF

[Hig79a] Robin Higham. Australian Air Force since 1911 by N. M. Parnell, C. A. Lynch, and: The Man on the Twenty-Dollar Note: Sir Charles Kingsford-Smith by Ward McNally (review). Technology and Culture, 20(2):379–381, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890509/pdf.

Higham:1979:SSS

[Hig79b] Robin Higham. The Saga of Smokey Stover by E. T. Stover, Clark G. Reynolds (review). Technology and Culture, 20(4):825–826, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890592/pdf.

Hills:1970:OCD

[Hil70a] R. L. Hills. The Oakham Canal by D. Tew (review). Technology and Culture, 11(2):304–305, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894162/pdf.

Hilton:1970:BSR

[Hil70b] George W. Hilton. Busted and Still Running by Edgar T. Mead, Jr. (review). Technology and Culture, 11(1):107–108, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894200/pdf.

Hilton:1970:DGL

[Hil70c] George W. Hilton. The Dossin Great Lakes Museum by Robert E. Lee (review). Technology and Culture, 11(2):289–297, April 1970.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894157/pdf.

Hills:1971:DSL

[Hil71] R. L. Hills. The Draining of the Somerset Levels by Michael Williams (review). Technology and Culture, 12(2):345–347, April 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893972/pdf.

Hilliard:1972:DPR

[Hil72] Sam B. Hilliard. The dynamics of power: Recent trends in mechanization on the American farm. Technology and Culture, 13(1):1–24, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893428/pdf.

Hill:1974:RIW

[Hil74] Donald R. Hill. Roman and Islamic Water-lifting Wheels by Thorkild Schiøler (review). Technology and Culture, 15(4):628–630, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892706/pdf.

Hills:1975:SMH

[Hil75a] Richard Hills. The Spinning Mule by H. Catling (review). Technology and Culture, 16(1):88–90, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891506/pdf.

Hills:1975:LMR

[Hil75b] Richard L. Hills. Liverpool and Manchester Railway Operations, 1831–1845 by Thomas J. Donaghy (review). Technology and Culture, 16(3):490–492, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891581/pdf.

Hills:1976:WIH

[Hil76] R. L. Hills. Waterloo Ironworks: a History of Taskers of Andover, 1809–1968 by L. T. C. Rolt (review). Technology and Culture, 17(2):378–380, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891728/pdf.

Hilton:1977:TEG

[Hil77] George W. Hilton. Transportation Economics: a Guide to Information Sources by James P. Rakowski (review). Technology and Culture, 18(4):750–751, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892032/pdf.

Hill:1978:ARW

[Hil78] Lewis E. Hill. An Anatomy of Risk by William D. Rowe, and: Technological Innovation: The Experimental R & D Incentives Program by Donald E. Cunningham, et al. (review). Technology and Culture, 19(4):788–790, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892210/pdf.

Hindle:1971:HWH

[Hin71a] Brooke Hindle. Henry Wansey and His American Journal, 1794 ed. by David John Jeremy (review). Technology and Culture, 12 (4):642–644, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893567/pdf.

Hindle:1971:CWC

[Hin71b] Brooke Hindle. The Collected Works of Count Rumford. Vol. 1, The Nature of Heat. Vol. 2, Practical Application of Heat. Vol. 3, Devices and Techniques. Vol. 4, Light and Armament. Vol. 5, Public Institutions ed. by Sanborn C. Brown (review). Technology and Culture, 12(4):640–642, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893566/pdf.

Hinckley:1972:CPH

[Hin72] Ted C. Hinckley. Collis Potter Huntington by Cerinda W. Evans (review). Technology and Culture, 13(4):639–641, October 1972. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893389/pdf.

Hindle:1979:RGA

[Hin79] Brooke Hindle. Rockdale: The Growth of an American Village in the Early Industrial Revolution by Anthony F. C. Wallace (review). Technology and Culture, 20(3):640–642, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890544/pdf.

Hitz:1976:WRW

[Hit76] Elizabeth Hitz. The West Riding Wool Textile Industry: a Catalogue of Business Records from the Sixteenth to the Twentieth Century by Patricia Hudson (review). Technology and Culture, 17(3):610– 611, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891808/pdf.

Hodge:1979:NSH

[Hod79] M. J. S. Hodge. The Nature of Science: The History of Science since 1600 by David Knight (review). Technology and Culture, 20(3):630– 631, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890537/pdf.

Hoggatt:1975:EGS

[Hog75] Austin Hoggatt. Economic Growth and Social Equity in Developing Countries by Irma Adelman, Cynthia Taft Morris (review). Technology and Culture, 16(4):678–681, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891630/pdf.

Hoke:1977:YDD

[Hok77] Donald Hoke. Yankee Dreamers and Doers by Ellsworth Strong Grant (review). Technology and Culture, 18(2):253–254, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891923/pdf.

Hoke:1979:THY

[Hok79a] Donald Hoke. Two Hundred Years of American Clocks and Watches by Chris H. Bailey (review). Technology and Culture, 20(4):805–807, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890584/pdf.

Hoke:1979:UOM

[Hok79b] Donald Hoke. University of Oxford, Museum of the History of Science: Catalogue 2 — Watches by F. R. Maddison, A. J. Turner (review). Technology and Culture, 20(2):360–361, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890499/pdf.

Holley:1970:CGA

[Hol70a] I. B. Holley, Jr. Climb to Greatness: The American Aircraft Industry, 1920–1960 by John B. Rae (review). Technology and Culture, 11(2):313–316, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894167/pdf.

Holmes:1970:SAJ

[Hol70b] Frederic L. Holmes. A Scientific Autobiography of Joseph Priestley (1733–1804) ed. by Robert E. Schofield (review). Technology and Culture, 11(3):454–457, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894120/pdf.

Holley:1976:GPJ

[Hol76] I. B. Holley. The Great Planes by James Gilbert (review). Technology and Culture, 17(4):769–771, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891836/pdf.

Holley:1979:SMI

[Hol79] I. B. Holley. The Southern Mystique: The Impact Of Technology on Human Values in a Changing Region ed. by W. David Lewis, B. Eugene Griessman (review). Technology and Culture, 20(1):242–244, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890485/pdf.

Hopkins:1975:RTA

[Hop75] A. G. Hopkins. Rationality and technology in African history. Technology and Culture, 16(3):462–465, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891571/pdf.

Horowitz:1979:PPS

[Hor79] Irving Louis Horowitz. Prophecy and Progress: The Sociology of Industrial and Post-industrial Society by Krishan Kumar, and: Postindustrial Society ed. by Bo Gustafsson (review). Technology and Culture, 20(4):848–852, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890599/pdf.

Hounshell:1974:GME

[Hou74] David A. Hounshell. A guide to manuscripts in electrical history. Technology and Culture, 15(4):626–627, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892705/pdf.

Hounshell:1975:EGT

[Hou75] David A. Hounshell. Elisha Gray and the telephone: On the disadvantages of being an expert. Technology and Culture, 16(2):133– 161, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891528/pdf.

Hounshell:1977:PWC

[Hou77] David A. Hounshell. Peacefully Working to Conquer the World: The Singer Sewing Machine in Foreign Markets, 1854–1920 by Robert Bruce Davies (review). Technology and Culture, 18(4):700–703, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892006.

Hounshell:1978:SMI

[Hou78] David A. Hounshell. The Sewing Machine: Its Invention and Development by Grace Rogers Cooper (review). Technology and Culture, 19(3):559–560, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892146/pdf.

Howland:1970:TMR

[How70] W. E. Howland. “Technology for man” revisited. Technology and Culture, 11(2):237–239, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894153/pdf.

Howe:1972:SEA

[How72] Daniel Walker Howe. The Scottish Enlightenment and the American College Ideal by Douglas Sloan (review). Technology and Culture, 13(1):74–75, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893439/pdf.

Howard:1974:NPG

[How74] Robert A. Howard. New Principles of Gunnery (1742) by Benjamin Robins (review). Technology and Culture, 15(1):95–97, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892835/pdf.

Howard:1978:IPR

[How78] Robert A. Howard. Interchangeable parts reexamined: The private sector of the American arms industry on the eve of the civil war. Technology and Culture, 19(4):633–649, October 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892169/pdf.

Hills:1972:MPE

[HP72] R. L. Hills and A. J. Pacey. The measurement of power in early steam-driven textile mills. Technology and Culture, 13(1):25–43, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893429/pdf.

Harris:1976:MDD

[HP76] J. R. Harris and C. Pris. The memoirs of Delaunay deslandes. Technology and Culture, 17(2):201–216, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891711/pdf.

Hindle:1976:BMW

[HRM^{+}76] Brooke Hindle, John B. Rae, Robert P. Multhauf, Melvin Kranzberg, Merritt Roe Smith, Carroll Pursell, and Arthur L. Norberg. The Bicentennial Meeting, Washington, October 16–19, 1975. Technology and Culture, 17(3):499–514, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891759/pdf.

Hall:1970:FYA

[HS70] R. Cargill Hall and James J. Sloan. Fifty years of air transport history: AIAA Meeting, Anaheim, California October 1969. Technology and Culture, 11(3):401–404, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894103/pdf.

Hamel-Schwulst:1978:SSG

[HS78] Mary Hamel-Schwulst. Surrealists and Surrealism, 1919–1939 by Gaëtan Picon, and: Art Now: From Abstract Expressionism to Superrealism by Edward Lucie-Smith, and: The Appreciation of Stained Glass by Lawrence Lee (review). Technology and Culture, 19(3):536–539, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892133/pdf.

Hudson:1971:VTP

[Hud71] Kenneth Hudson. Victorian Technology and Its Preservation in Modern Britain by Norman A. F. Smith (review). Technology and Culture, 12(1):101–102, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894013/pdf.

Hudson:1975:HNW

[Hud75a] Kenneth Hudson. A History of the North Wales Slate Industry by Jean Lindsay (review). Technology and Culture, 16(3):486–488, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891578/pdf.

Hudson:1975:DIG

[Hud75b] Kenneth Hudson. The Darbys and the Ironbridge Gorge by Brian Bracegirdle, Patricia H. Miles (review). Technology and Culture, 16(3):483–484, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891576/pdf.

Hudson:1976:SIA

[Hud76] Kenneth Hudson. Swedish Industrial Archaeology: Engelsberg Ironworks: a Pilot Project ed. by Bengt Holtez, et al. (review). Technology and Culture, 17(3):599–600, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891803/pdf.

Hudson:1979:ISW

[Hud79] Kenneth Hudson. Iron and Steel by W. K. V. Gale, and: Yesterday’s Golcondas: Notable British Metal Mines by R. H. Bird (review). Technology and Culture, 20(4):800–802, October 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890581/pdf.

Hughes:1970:BHT

[Hug70] Thomas P. Hughes. Bibliography of the History of Technology by Eugene S. Ferguson (review). Technology and Culture, 11(2):298– 299, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894158/pdf.

Hughes:1976:DPT

[Hug76a] Thomas P. Hughes. The development phase of technological change. Technology and Culture, 17(3):423–431, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891750/pdf.

Hughes:1976:STI

[Hug76b] Thomas Parke Hughes. The science–technology interaction: The case of high-voltage power transmission systems. Technology and Culture, 17(4):646–662, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891816/pdf.

Hughes:1977:CDO

[Hug77] Thomas P. Hughes. Conceptual definitions: Order out of chaos. Technology and Culture, 18(3):512–514, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891949/pdf.

Hughes:1979:EAS

[Hug79a] Thomas P. Hughes. The electrification of America: The system builders. Technology and Culture, 20(1):124–161, January 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890456/pdf.

Hughes:1979:ETH

[Hug79b] Thomas P. Hughes. Emerging themes in the history of technology. Technology and Culture, 20(4):697–711, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890571/pdf.

Hunsberger:1971:ADA

[Hun71] Warren S. Hunsberger. Agricultural Development in Asia ed. by R. T. Shand, and: Seeds of Change. The Green Revolution and Development in the 1970’s by Lester R. Brown (review). Technology and Culture, 12(1):146–150, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894034/pdf.

Hunter:1975:SBH

[Hun75] Louis C. Hunter. The Steamboat “Bertrand”: History, Excavation, and Architecture by Jerome E. Petsche (review). Technology and Culture, 16(4):705–707, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891645/pdf.

Halley:1978:TFH

[HV78] Richard B. Halley and Harold G. Vatter. Technology and the future as history: a critical review of futurism. Technology and Culture, 19(1):53–82, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892040/pdf.

Hyde:1973:EME

[Hyd73] Charles K. Hyde. Essays on a Mature Economy: Britain after 1840 ed. by Donald N. McCloskey (review). Technology and Culture, 14(2 (part I)):313–315, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893279/pdf.

Ihde:1977:ICB

[Ihd77] Aaron J. Ihde. The Incompleat Chymist: Being an Essay on the Eighteenth Century Chemist in His Laboratory, with a Dictionary of Obsolete Chemical Terms of the Period by Jon Eklund (review). Technology and Culture, 18(1):95–96, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891874/pdf.

Inohara:1976:JBS

[Ino76] Hideo Inohara. Japanese Business and Social Conflict: a Comparative Analysis of Response Patterns with American Business by S. Prakash Sethi (review). Technology and Culture, 17(1):173–175, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891690/pdf.

Inouye:1977:TTN

[IS77] Arlene Inouye and Charles Süsskind. “Technological trends and national policy,” 1937: The first modern technology assessment. Technology and Culture, 18(4):593–621, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891990/pdf.

Isaac:1976:HDM

[Isa76] Erich Isaac. History of Domestic Mammals in Central and Eastern Europe by S. Bökönyi (review). Technology and Culture, 17 (1):108–110, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891655/pdf.

Jackson:1970:PIP

[Jac70] Joseph Gray Jackson. Patent interference proceedings and priority of invention. Technology and Culture, 11(4):598–600, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894046/pdf.

Jackson:1971:EIU

[Jac71] Joseph Gray Jackson. The Employed Inventor in the United States: R & D Policies, Law, and Practice by Frederik Neumeyer (review). Technology and Culture, 12(4):673–675, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893583/pdf.

Jackson:1973:TCA

[Jac73] Joseph Gray Jackson. Technology and Copyright: Annotated Bibliography and Source Materials ed. by George P. Bush (review). Technology and Culture, 14(4):684, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892910/pdf.

Jackson:1975:LAT

[Jac75a] Joseph Gray Jackson. Legal Aspects of Technology Utilization by Richard I. Miller (review). Technology and Culture, 16(4):702–703, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891643/pdf.

Jackson:1975:HMH

[Jac75b] W. Turrentine Jackson. The Hardrock Miners: a History of the Mining Labor Movement in the American West, 1863–1893 by Richard E. Lingenfelter (review). Technology and Culture, 16(4):637–639, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891609/pdf.

Jackson:1977:HUS

[Jac77] Donald C. Jackson. Hydraulics in the United States, 1776–1976 by Hunter Rouse (review). Technology and Culture, 18(4):693–694, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892001/pdf.

Jackson:1979:SHA

[Jac79] Donald C. Jackson. A Selection of Historic American Papers on Concrete, 1876–1926 ed. by Howard Newlon, Jr. (review). Technology and Culture, 20(4):811–813, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890586/pdf.

James:1976:MWO

[Jam76] Dilmus James. Managers and Workers: Origins of the New Factory System of the United States, 1880–1920 by Daniel Nelson (review). Technology and Culture, 17(4):765–767, October 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891834/pdf.

James:1979:ATD

[Jam79a] Dilmus D. James. Appropriate Technologies for Developing Countries by Richard S. Eckaus, and: Technology and Underdevelopment by Frances Stewart (review). Technology and Culture, 20(3):659– 661, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890552/pdf.

James:1979:UPV

[Jam79b] Dilmus D. James. The Uncertain Promise: Value Conflicts in Technology Transfer by Denis Goulet (review). Technology and Culture, 20(3):657–659, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890551/pdf.

Jenkins:1970:ACS

[Jen70] Reese V. Jenkins. Alvan Clark & Sons, Artists in Optics by Deborah Jean Warner (review). Technology and Culture, 11(3):447– 449, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894116/pdf.

Jenkins:1972:NDJ

[Jen72] Reese V. Jenkins. The New Daguerreian Journal by Daguerreian Society (review). Technology and Culture, 13(3):525–526, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893951/pdf.

Jenkins:1973:ICB

[Jen73a] Reese V. Jenkins. ISIS Cumulative Bibliography: a Bibliography of the History of Science Formed from ISIS Critical Bibliographies 1– 90, 1913–1965. Part I: Personalities. Part 2: Personalities and Institutions ed. by Magda Whitrow (review). Technology and Culture, 14(2 (part I)):354–356, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893294/pdf.

Jenkins:1973:MBW

[Jen73b] Reese V. Jenkins. Margaret Bourke-White, Photojournalist by Theodore M. Brown (review). Technology and Culture, 14(4): 669–670, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892903/pdf.

Jenkins:1973:PPD

[Jen73c] Reese V. Jenkins. The Painter and the Photograph from Delacroix to Warhol by Van Deren Coke (review). Technology and Culture, 14(3):485–486, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893221/pdf.

Jenkins:1975:TMG

[Jen75] Reese V. Jenkins. Technology and the market: George Eastman and the origins of mass amateur photography. Technology and Culture, 16(1):1–19, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891495/pdf.

Jenkins:1976:PMM

[Jen76] Glen Jenkins. The Precious Metals of Medicine by Geoffrey Marks, William K. Beatty (review). Technology and Culture, 17(1):119–120, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891659/pdf.

Jenkins:1977:EAS

[Jen77a] Reese V. Jenkins. Early American Science ed. by Brooke Hindle (review). Technology and Culture, 18(3):534–535, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891959/pdf.

Jenkins:1977:PAP

[Jen77b] Reese V. Jenkins. Photographica: The artifacts of photography. Technology and Culture, 18(2):231–236, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891913/pdf.

Jeremy:1970:HLT

[Jer70a] David J. Jeremy. The Hollingworth Letters: Technical Change in the Textile Industry, 1826–1831 ed. by Thomas W. Leavitt (review). Technology and Culture, 11(4):637–638, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894061/pdf.

Jeremy:1970:TIA

[Jer70b] David J. Jeremy. The Textile Industry: an Account of the Early Inventions of Spinning, Weaving, and Knitting Machines by Walter English (review). Technology and Culture, 11(4):636–637, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894060/pdf.

Jeremy:1973:IAT

[Jer73] David J. Jeremy. Innovation in American textile technology during the early 19th century. Technology and Culture, 14(1):40–76, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893353/pdf.

Jeremy:1974:IIA

[Jer74] David J. Jeremy. Invisible Immigrants: The Adaptation of English and Scottish Immigrants in Nineteenth-Century America by Charlotte Erickson (review). Technology and Culture, 15(3):497– 498, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892751/pdf.

Jeremy:1976:FSN

[Jer76a] D. J. Jeremy. Family Structure in Nineteenth Century Lancashire by Michael Anderson (review). Technology and Culture, 17(3):549– 551, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891776/pdf.

Jeremy:1976:LDS

[Jer76b] David J. Jeremy. Lancashire in Decline: a Study in Entrepreneurship, Technology, and International Trade by Lars G. Sandberg (review). Technology and Culture, 17(1):158–160, January 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891682/pdf.

Jeremy:1977:TMC

[Jer77] David J. Jeremy. The Textile Machinery Collection, Science Museum (review). Technology and Culture, 18(1):70–79, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891867/pdf.

Jeremy:1978:WSA

[Jer78a] D. J. Jeremy. Wheel and Spindle: Aspects of Irish Industrial History by W. Alan McCutcheon (review). Technology and Culture, 19(3):551–552, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892141/pdf.

Jeremy:1978:PCC

[Jer78b] David J. Jeremy. La plus ça change . . .. Technology and Culture, 19(3):608, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892165/pdf.

Jeremy:1979:THE

[Jer79] David J. Jeremy. Textile History ed. by K. G. Ponting, S. D. Chapman (review). Technology and Culture, 20(2):361–363, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890500/pdf.

Jevons:1976:IST

[Jev76] F. R. Jevons. The interaction of science and technology today, or, is science the mother of invention? Technology and Culture, 17 (4):729–742, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891822/pdf.

Johnson:1973:PVC

[Joh73] Arthur M. Johnson. Pithole, the Vanished City: a Story of the Early Days of the Petroleum Industry by William Culp Darrah (review). Technology and Culture, 14(2 (part I)):311–312, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893277/pdf.

Johnson:1978:NSE

[Joh78] Arthur M. Johnson. The New State of the Economy by Fred C. Allvine, Fred A. Tarpley, Jr. (review). Technology and Culture, 19 (4):743–744, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892185/pdf.

Jones:1978:MCT

[Jon78] Daniel P. Jones. From military to civilian technology: The introduction of tear gas for civil riot control. Technology and Culture, 19(2):151–168, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892068/pdf.

Jensen:1978:TTP

[JR78] John H. Jensen and Gerhard Rosegger. Transferring technology to a peripheral economy: The case of lower Danube transport development, 1856–1928. Technology and Culture, 19(4):675–702, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892171/pdf.

Judd:1974:AWA

[Jud74a] Jacob Judd. America’s Wooden Age: Sleepy Hollow restorations, April 1973. Technology and Culture, 15(1):64–69, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892825/pdf.

Judd:1974:TEP

[Jud74b] Jacob Judd. The Timber Economy of Puritan New England by Charles F. Carroll (review). Technology and Culture, 15(3):494– 496, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892750/pdf.

Junker:1970:EEG

[Jun70] Louis Junker. Exports and Economie Growth of Developing Countries by A. Maizels (review). Technology and Culture, 11(2):330– 333, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894178/pdf.

Kahn:1970:EAP

[Kah70] Arthur D. Kahn. “Every art possessed by Man comes from Prometheus”: The Greek tragedians and science and technology. Technology and Culture, 11(2):133–162, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894146/pdf.

Kaplan:1971:SSI

[Kap71] Norman Kaplan. The Shaping of Science: Institutions, Ideology, and Careers in Science by Roger G. Krohn (review). Technology and Culture, 12(4):671–673, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893582/pdf.

Kaplan:1972:SSP

[Kap72] Norman Kaplan. Science, Scientists, and Public Policy by Dean Schooler, Jr. (review). Technology and Culture, 13(2):332–333, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893487/pdf.

Kaplan:1974:OCS

[Kap74a] Norman Kaplan. Originality and Competition in Science: a Study of the British High Energy Physics Community by Jerry Gaston (review). Technology and Culture, 15(4):664–666, October 1974. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892725/pdf.

Kaplan:1974:SSS

[Kap74b] Norman Kaplan. Social Stratification in Science by Jonathan R. Cole, Stephen Cole (review). Technology and Culture, 15(4):662–664, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892724/pdf.

Kaplan:1974:SST

[Kap74c] Norman Kaplan. The Sociology of Science: Theoretical and Empirical Investigations by Robert K. Merton (review). Technology and Culture, 15(4):659–662, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892723/pdf.

Kapstein:1977:TIG

[Kap77] Ethan B. Kapstein. “Tell It Good-bye, Kiddo”: The Decline of the New England Offshore Fishery by David Boeri, James Gibson (review). Technology and Culture, 18(4):716–717, October 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892015/pdf.

Kapstein:1978:SWW

[Kap78] Ethan B. Kapstein. Steam Whaling in the Western Arctic by John R. Bockstoce (review). Technology and Culture, 19(1):133–135, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892060/pdf.

Kargon:1976:BMN

[Kar76] R. Kargon. Biographical Memoirs by the National Academy of Sciences (review). Technology and Culture, 17(2):395–396, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891736/pdf.

Keller:1970:APC

[Kel70a] A. Keller. Atti del primo convegno internazionale di ricognizione delle fonti per la storia della scienza italiana. I. Secoli XIV–XVI. (Vol. 1) Domus Galilaeana, Pisa, September 14–16, 1966. Atti del convegno sui problemi metodologici di storia della scienza. (Vol. 2) Centro di Studi Methodologici, Turin, March 29–31, 1967 ed. by Carlo Maccagni (review). Technology and Culture, 11(1):97–101, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894196/pdf.

Keller:1970:JRT

[Kel70b] A. Keller. John Rogers, Tudor Military Engineer by L. R. Shelby (review). Technology and Culture, 11(1):102–104, January 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894197/pdf.

Keller:1970:RHL

[Kel70c] Alex Keller. A Renaissance humanist looks at “New” inventions: The article “Horologium” in Giovanni Tortelli’s De Orthographia. Technology and Culture, 11(3):345–365, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/894100/pdf.

Keller:1971:LBA

[Kel71a] A. G. Keller. Leon Battista Alberti: Universal Man of the Renaissance by Joan Gadol (review). Technology and Culture, 12 (4):632–635, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893564/pdf.

Keller:1971:AHT

[Kel71b] Alex Keller. Archimedean hydrostatic theorems and salvage operations in 16th-Century Venice. Technology and Culture, 12(4): 602–617, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893558/pdf.

Keller:1971:EER

[Kel71c] Alex Keller. Engineers and Engineering in the Renaissance by William Barclay Parsons (review). Technology and Culture, 12 (2):337–340, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893968/pdf.

Kelly:1971:IBT

[Kel71d] Patrick Kelly. Implications of Biomedical Technology ed. by Irene Taviss and Judith Koivumaki, and: Technology and Work ed. by Irene Taviss and William Gerber, and: Technology and Values ed. by Irene Taviss and Linda Silverman, and: Technology and the Polity ed. by Irene Taviss and Judith Purbank (review). Technology and Culture, 12(1):158–159, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894039/pdf.

Keller:1973:MYJ

[Kel73] Alex Keller. The missing years of Jacques Besson, inventor of machines, teacher of mathematics, distiller of oils, and Huguenot pastor. Technology and Culture, 14(1):28–39, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893352/pdf.

Keller:1974:TIS

[Kel74] Alex Keller. The third international symposium on molinology: The Netherlands, May 1973. Technology and Culture, 15(2):226–227, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892792/pdf.

Keller:1976:SCE

[Kel76] Alex Keller. Society and Culture in Early Modern France by Natalie Zemon Davis (review). Technology and Culture, 17(2):368–370, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891722/pdf.

Keller:1977:WSB

[Kel77] Alex Keller. Windmills by Suzanne Beedell (review). Technology and Culture, 18(1):92–94, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891872/pdf.

Keller:1978:RTM

[Kel78] Alex Keller. Renaissance theaters of machines. Technology and Culture, 19(3):495–508, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892118/pdf.

Kevles:1971:FLE

[Kev71] Daniel J. Kevles. Federal legislation for engineering experiment stations: The episode of World War I. Technology and Culture, 12(2):182–189, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893954/pdf.

Kevles:1975:SMC

[Kev75] Daniel J. Kevles. Scientists, the military, and the control of postwar defense research: The case of the research board for national security, 1944–46. Technology and Culture, 16(1):20–47, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891496/pdf.

Kearns-Faure:1976:EPT

[KF76] Françoise Kearns-Faure. Etudes pour un traitement automatique des sources en histoire des techniques by the Centre de documentation d’histoire des techniques (review). Technology and Culture, 17(2):410–411, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891744/pdf.

Klein:1971:EFP

[Kle71] Martin J. Klein. Enrico Fermi, Physicist by Emilio Segrè (review). Technology and Culture, 12(3):507–508, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL http://www.jstor.org/stable/3103011;   https:
//muse.jhu.edu/pub/1/article/893607/pdf.

Kley:1976:MCC

[Kle76] Ronald J. Kley. Museum Cataloging in the Computer Age by Robert G. Chenhall (review). Technology and Culture, 17(3):614– 615, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891811/pdf.

Klein:1977:JJH

[Kle77] Maury Klein. James J. Hill and the Opening of the Northwest by Albro Martin (review). Technology and Culture, 18(4):698–699, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892005/pdf.

Knight:1977:AAD

[Kni77] Thomas J. Knight. Awakening from the American Dream: The Social and Political Limits to Growth by Rufus E. Miles, Jr. (review). Technology and Culture, 18(1):140–142, January 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891900/pdf.

Kobayashi:1974:JEG

[Kob74] Tatsuya Kobayashi. Japanese Economic Growth: Trend Acceleration in the Twentieth Century by Kazushi Ohkawa, Henry Rosovsky (review). Technology and Culture, 15(2):347–349, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892809/pdf.

Kobayashi:1977:JUI

[Kob77] Tatsuya Kobayashi. Japanese Urbanism: Industry and Politics in Kariya, 1872–1972 by Gary D. Allinson (review). Technology and Culture, 18(2):257–259, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891926/pdf.

Kohlmeyer:1970:AMA

[Koh70a] Fred W. Kohlmeyer. Acta Museorum Agriculturae (review). Technology and Culture, 11(3):477–478, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894131/pdf.

Kohlmeyer:1970:ECA

[Koh70b] Fred W. Kohlmeyer. Eighteenth-Century Agriculture: a Symposium (review). Technology and Culture, 11(2):301–302, April 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894160/pdf.

Kohlmeyer:1970:LOF

[Koh70c] Fred W. Kohlmeyer. Land That Our Fathers Plowed: The Settlement of Our Country as Told by the Pioneers Themselves and Their Contemporaries ed. by David G. Greenberg (review). Technology and Culture, 11(2):303–304, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894161/pdf.

Kohlmeyer:1971:CAN

[Koh71a] Fred W. Kohlmeyer. Change in Agriculture: The Northern United States, 1820–1870 by Clarence H. Danhof (review). Technology and Culture, 12(1):105–107, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894016/pdf.

Kohlmeyer:1971:WTC

[Koh71b] Fred W. Kohlmeyer. William Tatham and the Culture of Tobacco: Including a Facsimile Reprint of an Historical and Practical Essay on the Culture and Commerce of Tobacco by G. Melvin Herndon (review). Technology and Culture, 12(1):104–105, January 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894015/pdf.

Kohlmeyer:1972:FRT

[Koh72] Fred W. Kohlmeyer. Fields of Rich Toil: The Development of the University of Illinois College of Agriculture by Richard Gordon Moores (review). Technology and Culture, 13(4):641–642, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893390/pdf.

Kohlmeyer:1974:CNS

[Koh74] Fred W. Kohlmeyer. Crop Nutrition: Science and Practice before Liebig by G. E. Fussell, and: Jethro Tull: His Influence on Mechanized Agriculture by G. E. Fussell (review). Technology and Culture, 15(1):98–100, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892837/pdf.

Kohler:1976:IST

[Koh76a] Robert E. Kohler. The interaction of science and technology in the Industrial Age. Technology and Culture, 17(4):621–623, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891813/pdf.

Kohlmeyer:1976:VAE

[Koh76b] Fred W. Kohlmeyer. The Virginia Agricultural Experiment Station, 1886–1966 by Harold N. Young (review). Technology and Culture, 17(3):545–546, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891773/pdf.

Kohlmeyer:1977:AUS

[Koh77] Fred W. Kohlmeyer. Agriculture in the United States: a Documentary History ed. by Wayne D. Rasmussen (review). Technology and Culture, 18(1):96–98, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891875/pdf.

Kohlmeyer:1978:RAP

[Koh78] Fred W. Kohlmeyer. Resource Allocation and Productivity in National and International Agricultural Research ed. by Thomas M. Arndt, et al. (review). Technology and Culture, 19(1):135–138, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892061/pdf.

Koppes:1977:OMH

[Kop77] Clayton R. Koppes. Orders of Magnitude: a History of NACA and NASA, 1915–1976 by Frank W. Anderson, Jr. (review). Technology and Culture, 18(3):558–559, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891972/pdf.

Kouwenhoven:1970:AET

[Kou70] John A. Kouwenhoven. Architecture as environmental technology. Technology and Culture, 11(1):85–93, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894193/pdf.

Kranzberg:1973:F

[Kra73] Melvin Kranzberg. Foreword. Technology and Culture, 14(2 (part II)):v–vi, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893255/pdf.

Krause:1974:SPP

[Kra74] Elliott A. Krause. The Sociology of the Professions by Philip Elliott (review). Technology and Culture, 15(3):524–526, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892767/pdf.

Krause:1975:ALP

[Kra75] Elliott A. Krause. The Aristocracy of Labor: The Position of Skilled Craftsmen in the American Class Structure by Gavin MacKenzie (review). Technology and Culture, 16(1):114–117, January 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891520/pdf.

Krause:1976:DSM

[Kra76] Elliott A. Krause. The Deadly Simple Mechanics of Society by John Helmer (review). Technology and Culture, 17(3):586–589, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891797/pdf.

Krajewski:1977:SSR

[Kra77a] Kathleen M. Krajewski. Scientists and social responsibility: University of Delaware, July 1975. Technology and Culture, 18(1):56–61, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891864/pdf.

Krause:1977:MNE

[Kra77b] Elliott A. Krause. Medical Nemesis: The Expropriation of Health by Ivan Illich (review). Technology and Culture, 18(4):725–727, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892020/pdf.

Krammer:1978:FTR

[Kra78] Arnold Krammer. Fueling the third Reich. Technology and Culture, 19(3):394–422, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892110/pdf.

Kreutz:1973:MCM

[Kre73] Barbara M. Kreutz. Mediterranean contributions to the medieval mariner’s compass. Technology and Culture, 14(3):367–383, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893208/pdf.

Kren:1974:PHA

[Kre74] Claudia Kren. The “Pneumatics” of Hero of Alexandria: a Facsimile of the 1851 Woodcroft Edition by Marie Boas Hall (review). Technology and Culture, 15(2):326–328, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892797/pdf.

Kren:1976:THR

[Kre76] Claudia Kren. Thomas Harriot: Renaissance Scientist ed. by John W. Shirley (review). Technology and Culture, 17(4):791–793, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891847/pdf.

Kren:1977:TDL

[Kre77] Claudia Kren. The traveler’s dial in the Late Middle Ages: The chilinder. Technology and Culture, 18(3):419–435, July 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891938/pdf.

Kuisel:1978:PFM

[Kui78] Richard F. Kuisel. Peasants into Frenchmen: The Modernization of Rural France, 1870–1914 by Eugen Weber (review). Technology and Culture, 19(2):232–234, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892087/pdf.

Kullikoff:1979:MCS

[Kul79] Allan Kullikoff. Material Culture and the Study of American Life ed. by Ian M. G. Quimby (review). Technology and Culture, 20(2):366– 367, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890502/pdf.

Kurzweil:1973:EIS

[Kur73] Edith Kurzweil. Exploring the Industrial Subculture by Barry A. Turner (review). Technology and Culture, 14(4):672–674, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892905/pdf.

Lafleur:1972:CWM

[Laf72] P. Lafleur. The Changing World Market for Iron Ore, 1950–1980 by Gerald Manners (review). Technology and Culture, 13(4):650–651, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893395/pdf.

Lakoff:1970:YTH

[Lak70] Sanford A. Lakoff. Yesterday’s Tomorrows: a Historical Survey of Future Societies by W. H. G. Armytage (review). Technology and Culture, 11(1):124–127, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894207/pdf.

Lamberton:1971:TEE

[Lam71a] D. M. Lamberton. Technology, Engineering, and Economics by Philip Sporn (review). Technology and Culture, 12(1):143–144, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894032/pdf.

Lamberton:1971:BDT

[Lam71b] Donald M. Lamberton. Business Decisions and Technological Change by John Diebold (review). Technology and Culture, 12 (3):528–529, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893619/pdf.

Langan:1973:HAS

[Lan73] Thomas Langan. History as Applied Science: a Philosophical Study by William Todd (review). Technology and Culture, 14(3):525–527, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893242/pdf.

Law:1971:PSL

[Law71] R. J. Law. Partners in Science: Letters of James Watt and Joseph Black ed. by Eric Robinson and Douglas McKie (review). Technology and Culture, 12(1):96–98, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894010/pdf.

Layton:1970:RTA

[Lay70a] Edwin Layton. Readings in Technology and American Life ed. by Carroll W. Pursell, Jr. (review). Technology and Culture, 11 (1):104–105, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894198/pdf.

Layton:1970:CIT

[Lay70b] Edwin Layton. Comment: The interaction of technology and society. Technology and Culture, 11(1):27–31, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894187/pdf.

Layton:1971:ESS

[Lay71a] Edwin Layton. Early Stationary Steam Engines in America: a Study in the Migration of a Technology by Carroll W. Pursell, Jr. (review). Technology and Culture, 12(1):99–101, January 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894012/pdf.

Layton:1971:EAS

[Lay71b] Edwin Layton. Engineering in American Society, 1850–1875 by Raymond H. Merritt (review). Technology and Culture, 12(1):120–123, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894021/pdf.

Layton:1971:FTS

[Lay71c] Edwin Layton. Frederick Taylor: a Study in Personality and Innovation by Sudhir Kakar (review). Technology and Culture, 12 (4):666–668, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893579/pdf.

Layton:1971:MIT

[Lay71d] Edwin Layton. Mirror-image twins: The communities of science and technology in 19th-Century America. Technology and Culture, 12 (4):562–580, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893556/pdf.

Layton:1974:TK

[Lay74] Edwin T. Layton, Jr. Technology as knowledge. Technology and Culture, 15(1):31–41, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892822/pdf. See comment [De 74e].

Layton:1975:SVP

[Lay75] Edwin Layton. Science and Values, Patterns of Tradition and Change ed. by Arnold Thackray, Everett Mendelsohn (review). Technology and Culture, 16(4):651–653, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891617/pdf.

Layton:1976:KHN

[Lay76a] Edwin Layton. From Know-how to Nowhere: The Development of American Technology by Elting E. Morison (review). Technology and Culture, 17(4):746–748, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891824/pdf.

Layton:1976:AIS

[Lay76b] Edwin T. Layton, Jr. American ideologies of science and engineering. Technology and Culture, 17(4):688–701, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891819/pdf.

Layton:1977:PAH

[Lay77a] Edwin T. Layton, Jr. Perspectives in American History ed. by Donald Fleming, Bernard Bailyn (review). Technology and Culture, 18(3):532–534, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891958/pdf.

Layton:1977:PME

[Lay77b] Edwin T. Layton, Jr. Philosophers and Machines ed. by Otto Mayr (review). Technology and Culture, 18(1):89–91, January 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891870/pdf.

Layton:1977:SAS

[Lay77c] Edwin T. Layton, Jr. Science in America since 1820 ed. by Nathan Reingold (review). Technology and Culture, 18(1):98–100, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891876/pdf.

Layton:1979:STH

[Lay79] Edwin T. Layton, Jr. Scientific technology, 1845–1900: The hydraulic turbine and the origins of American industrial research. Technology and Culture, 20(1):64–89, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890453/pdf.

Layton:1972:NYM

[LDS^{+}72] Edwin T. Layton, Robert C. Davis, Bruce Sinclair, Nathan Rosenberg, and Joseph Gray Jackson. The New York Meeting, December 28–30, 1971. Technology and Culture, 13(3):440–464, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893929/pdf.

LeCorbeiller:1971:DIA

[Le 71] Jean Le Corbeiller. The Dream of Icarus: Art and Society in the Twentieth Century by Kenneth Coutts-Smith (review). Technology and Culture, 12(2):378–380, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893990/pdf.

LeCorbeiller:1972:CGC

[Le 72] Jean Le Corbeiller. Computer Graphics/Computer Art by Herbert W. Franke (review). Technology and Culture, 13(4):660–662, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893400/pdf.

LeCorbeiller:1975:ICA

[Le 75] Jean Le Corbeiller. Iconic Communication: an Annotated Bibliography by W. H. Huggins and Doris R. Entwisle (review). Technology and Culture, 16(1):126–128, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891526/pdf.

Leavitt:1970:HWS

[Lea70] Thomas W. Leavitt. The Handloom Weavers: a Study in the English Cotton Industry during the Industrial Revolution by Duncan Bythell (review). Technology and Culture, 11(3):441–443, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894112/pdf.

Leavitt:1976:THE

[Lea76] Thomas W. Leavitt. Textile History ed. by K. G. Ponting, S. D. Chapman (review). Technology and Culture, 17(3):603–604, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891805/pdf.

Leavitt:1979:THE

[Lea79] Thomas M. Leavitt. Textile History ed. by K. G. Ponting, S. D. Chapman (review). Technology and Culture, 20(4):799–800, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890580/pdf.

Lederer:1977:AWC

[Led77] Norman Lederer. American Welfare Capitalism, 1880–1940 by Stuart D. Brandes (review). Technology and Culture, 18(4):706–708, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892009.

Leek:1978:GB

[Lee78] Michael E. Leek. The S.S. Great Britain. Technology and Culture, 19(2):199–203, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892071/pdf.

Leicester:1970:OHY

[Lei70] Henry M. Leicester. One Hundred Years of Pharmacy in Canada by Canadian Academy of the History of Pharmacy (review). Technology and Culture, 11(2):321, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894172/pdf.

Leicester:1971:SPH

[Lei71] Henry M. Leicester. Safeguarding the Public: Historical Aspects of Medicinal Drug Control ed. by John B. Blake (review). Technology and Culture, 12(3):515–516, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893611/pdf.

Leitenberg:1973:EST

[Lei73] Milton Leitenberg. Environment and Society in Transition: Scientific Developments, Social Consequences, Policy Implications ed. by Peter Albertson, Margery Barnett (review). Technology and Culture, 14(2 (part I)):345–347, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893290/pdf.

Leslie:1979:CFK

[Les79] Stuart W. Leslie. Charles F. Kettering and the copper-cooled engine. Technology and Culture, 20(4):752–776, October 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890573/pdf.

Levine:1975:SPO

[Lev75] Arnold S. Levine. Science for the People: The Origins of the School Science Curriculum in England by David Layton (review). Technology and Culture, 16(1):119–120, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891522/pdf.

Levinson:1979:CPE

[Lev79] Paul Levinson. Culture as Polyphony: an Essay on the Nature of Paradigms by James M. Curtis (review). Technology and Culture, 20(4):835–837, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890596/pdf.

Lewis:1970:BLR

[Lew70] W. David Lewis. Beyond Left and Right: Radical Thought for Our Times ed. by Richard Kostelanetz (review). Technology and Culture, 11(1):120–124, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894206/pdf.

Lewis:1971:TMM

[Lew71] W. David Lewis. Technological Man: The Myth and the Reality by Victor C. Ferkiss (review). Technology and Culture, 12(2):388– 394, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893994/pdf.

Lewis:1973:IAH

[Lew73] M. J. T. Lewis. Industrial Archaeology: an Historical Survey by Arthur Raistrick (review). Technology and Culture, 14(4):616–618, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892873/pdf.

Lewis:1974:OOM

[Lew74] W. David Lewis. Our Own Metaphor: a Personal Account of a Conference on the Effects of Conscious Purpose on Human Adaptation by Mary Catherine Bateson (review). Technology and Culture, 15 (1):146–150, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892859/pdf.

Lewis:1976:CAE

[Lew76] W. David Lewis. Coal Age Empire: Pennsylvania Coal and Its Utilization to 1860 by Frederick Moore Binder (review). Technology and Culture, 17(1):140–143, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891671/pdf.

Lewis:1977:SIE

[Lew77] W. David Lewis. The Smithsonian Institution’s “1876” Exhibit. Technology and Culture, 18(4):670–684, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891997/pdf.

Lide:1976:FMC

[Lid76] David R. Lide, Jr. Fundamental Measures and Constants for Science and Technology by Frederick D. Rossini (review). Technology and Culture, 17(1):122–123, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891661/pdf.

Lieberstein:1973:NTP

[Lie73a] S. Lieberstein. Nauchno-teknicheskii progress: sushchnost’ osnovnye tendentsii [Scientific-technical progress: essence of basic tendencies] by G. S. Gudozhnik (review). Technology and Culture, 14(2 (part I)):323–325, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893282/pdf.

Lieberstein:1973:PPF

[Lie73b] Samuel Lieberstein. Pioneers for Profit: Foreign Entrepreneurship and Russian Industrialization 1885–1913 by John P. McKay (review). Technology and Culture, 14(4):647–649, October 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892891/pdf.

Lieberstein:1973:RTJ

[Lie73c] Samuel Lieberstein. Russian Tanks, 1900–1970 by John Milsom (review). Technology and Culture, 14(4):649–651, October 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892892/pdf.

Lieberstein:1973:WTS

[Lie73d] Samuel Lieberstein. Western Technology and Soviet Economic Development, 1917 to 1930 by Antony C. Sutton, and: Western Technology and Soviet Economic Development, 1930 to 1945 by Antony C. Sutton (review). Technology and Culture, 14(1):92–93, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893364/pdf.

Lieberstein:1974:IMS

[Lie74a] Samuel Lieberstein. Industrial Management in the Soviet Union: The Role of the CPSU in Industrial Decision-Making, 1917–1970 by William J. Conyngham (review). Technology and Culture, 15 (1):121–122, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892848/pdf.

Lieberstein:1974:SEV

[Lie74b] Samuel Lieberstein. Sovetskaia ekonomika v period Velikoi Otechestvennoi voiny 1941–1945 gg. [The Soviet economy in the period of the Great Fatherland War, 1941–1945], and: Ekonomika SSSR v gody Velikoi Otechestvennoi voiny (1941–1945 gg.) [The economy of the USSR in the years of the Great Fatherland War (1941–1945)] by G. S. Kravchenko (review). Technology and Culture, 15(1):122–125, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892849/pdf.

Lieberstein:1974:SNT

[Lie74c] Samuel Lieberstein. Sovremennaia nauchno-tekhnicheskaia revoliutsiia [The contemporary scientific-technical revolution] (review). Technology and Culture, 15(2):354–356, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892813/pdf.

Lieberstein:1974:TSC

[Lie74d] Samuel Lieberstein. Technology and Social Change ed. by H. Russell Bernard, Pertti J. Pelto (review). Technology and Culture, 15 (1):144–146, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892858/pdf.

Lieberstein:1974:SPE

[Lie74e] Samuel Lieberstein. The Spoils of Progress: Environmental Pollution in the Soviet Union by Marshall I. Goldman (review). Technology and Culture, 15(1):125–127, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892850/pdf.

Lieberstein:1974:WTS

[Lie74f] Samuel Lieberstein. Western Technology and Soviet Economic Development, 1945 to 1965 by Antony C. Sutton (review). Technology and Culture, 15(3):508–510, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892758/pdf.

Lieberstein:1975:CNT

[Lie75a] Samuel Lieberstein. Chelovek-nauka-tekhnika (opyt marksistskogo analisa nauchno-teknicheskoi revoliutsii) [Man-science-technology: toward a Marxist analysis of the scientific-technical revolution] ed. by A. I. Mogilev (review). Technology and Culture, 16(4):691–693, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891637/pdf.

Lieberstein:1975:EKD

[Lie75b] Samuel Lieberstein. Eisen und Kohle für das Dritte Reich: Paul Pleigers Stellung in der NSWirtschaft [Iron and coal for the Third Reich: Paul Pleiger’s place in the National Socialist economy] by Mattias Riedel (review). Technology and Culture, 16(4):649–650, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891615/pdf.

Lieberstein:1975:IKZ

[Lie75c] Samuel Lieberstein. Istoriia Kolomenskogo zavoda. Ocherk istorii Kolomenskogo teplovozostroitel’nogo zavoda imeni V. V. Kuibysheva za 110 let (1863–1973) [The history of the Kolomna Works. An account of the history of the Kolomna V. V. Kuibyshev diesel locomotive works covering 110 years (1863–1973)] by G. P. Efremtsev (review). Technology and Culture, 16(4):643–645, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891612/pdf.

Lieberstein:1975:NTR

[Lie75d] Samuel Lieberstein. Nduchno-tekhnicheskaia revoliutsiia i obshchestvo [The scientific-technical revolution and society] ed. by N. I. Driakhlov, et al. (review). Technology and Culture, 16(2):329–331, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891556/pdf.

Lieberstein:1975:TWS

[Lie75e] Samuel Lieberstein. Technology, work, and sociology in the USSR: The NOT movement. Technology and Culture, 16(1):48–66, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891497/pdf.

Lieberstein:1976:EPR

[Lie76a] Samuel Lieberstein. Ekonomicheskoe polozhenie Rossii v gody pervoi mirovoi voiny [The economic position of Russia in the years of the First World War] by A. L. Sidorov (review). Technology and Culture, 17(2):390–392, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891734/pdf.

Lieberstein:1976:SPB

[Lie76b] Samuel Lieberstein. Sotsgorod: The Problem of Building Socialist Cities by N. A. Miliutin (review). Technology and Culture, 17 (4):808–809, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891856/pdf.

Lieberstein:1976:TII

[Lie76c] Samuel Lieberstein. Technik und Ingenieure im Dritten Reich [Technology and engineers in the Third Reich] by Karl-Heinz Ludwig (review). Technology and Culture, 17(2):392–395, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891735/pdf.

Lieberstein:1977:NTR

[Lie77a] Samuel Lieberstein. Nauchno-tekhnicheskai revoliutsiia i preimushchestva sotsializma [The scientific-technical revolution and the primacy of socialism] ed. by V. G. Lebedev, V. I. Kushlin (review). Technology and Culture, 18(3):576–577, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891982/pdf.

Lieberstein:1977:NTP

[Lie77b] Samuel Lieberstein. Nauchno-tekhnicheskii progress i ekonomika razvitogo sotsializma [Scientific-technical progress and the economics of developed socialism] by L. M. Gatovskii (review). Technology and Culture, 18(2):261–262, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891928/pdf.

Lieberstein:1977:IDS

[Lie77c] Samuel Lieberstein. The Innovation Decision in Soviet Industry by Joseph S. Berliner (review). Technology and Culture, 18(4):733–734, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892024/pdf.

Lieberstein:1977:STR

[Lie77d] Samuel Lieberstein. The Scientific and Technological Revolution — Its Impact on Management and Education by V. G. Afanasyev (review). Technology and Culture, 18(2):259–260, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891927/pdf.

Lieberstein:1978:SFT

[Lie78a] Samuel Lieberstein. Na sluzhbe fronta i tyla: Zheleznodorozhnyi transport nakanune i v gody Velikoi Otechestvennoivoiny 1938–1945 [At the service of the front and rear: railroad transport on the eve and during the years of the Great Patriotic War, 1938–1945] by Georgii A. Kumanev (review). Technology and Culture, 19(2):238– 241, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892090/pdf.

Lieberstein:1978:SNT

[Lie78b] Samuel Lieberstein. SShA: Nauchno-tekhnicheskii potential [U.S.A.: scientific and technical potential] by V. I. Gromeka (review). Technology and Culture, 19(4):771–772, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892201/pdf.

Lieberstein:1978:TLS

[Lie78c] Samuel Lieberstein. The Technological Level of Soviet Industry ed. by Ronald Amann, et al. (review). Technology and Culture, 19 (4):772–774, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892202/pdf.

Lieberstein:1979:SEP

[Lie79a] Samuel Lieberstein. Sotsial’no-ekonomicheskie problemy istorii razvitogo sotsializma v SSSR [Social and economic problems of the history of developed socialism in the USSR] ed. by M. P. Kim (review). Technology and Culture, 20(1):248–250, January 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890488/pdf.

Lienhard:1979:RTI

[Lie79b] John H. Lienhard. The rate of technological improvement before and after the 1830s. Technology and Culture, 20(3):515–530, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890522/pdf.

Lewis:1974:SFM

[LLC^{+}74] W. David Lewis, Edwin T. Layton, Rondo Cameron, Melvin Kranzberg, Reynold M. Wik, Bradford B. Blaine, Theodore Ropp, Richard N. Sinkin, Melvin Kranzberg, and Thomas P. Hughes. The San Francisco Meeting, December 27–30, 1973. Technology and Culture, 15(3):453–480, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892743/pdf.

Lolas:1978:NCA

[Lol78] Fernando Lolas. Neurometrics: Clinical Applications of Quantitative Electrophysiology by E. Roy John (review). Technology and Culture, 19(3):583–585, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892155/pdf.

Long:1978:SAS

[Lon78] Stewart L. Long. Sarnoff: an American Success by Carl Dreher (review). Technology and Culture, 19(3):567–569, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892150/pdf.

Lowrance:1976:STV

[Low76] William W. Lowrance. Science and Technology: Vital National Resources ed. by Ralph Sanders (review). Technology and Culture, 17(3):577–578, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891791/pdf.

Lubove:1978:UIC

[Lub78] Roy Lubove. The Urban Idea in Colonial America by Sylvia Doughty Fries (review). Technology and Culture, 19(4):748–750, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892189/pdf.

Maass:1978:GBT

[Maa78] Eleanor A. Maass. Greville Bathe’s “Theatre of Machines”: The evolution of a scholar and his collection. Technology and Culture, 19(4):713–723, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892174/pdf.

Maggs:1976:IIU

[Mag76] Peter B. Maggs. Invention and Innovation under Soviet Law: a Comparative Analysis by Manfred Wilhelm Balz (review). Technology and Culture, 17(3):561–562, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891784/pdf.

Major:1974:WSF

[Maj74] J. Kenneth Major. Windmühlen: Der Stand der Forschung über das Vorkommen und den Ursprung by Jannis C. Notebaart (review). Technology and Culture, 15(4):630–631, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892707/pdf.

Major:1975:SCC

[Maj75] John S. Major. Science and Civilisation in China. Vol. 5: Chemistry and Chemical Technology; pt. 2: Spagyrical Discovery and Invention: Magisteries of Gold, and Immortality by Joseph Needham (review). Technology and Culture, 16(4):621–627, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891604/pdf.

Major:1977:SCC

[Maj77] J. Kenneth Major. Science and Civilisation in China. Vol. 5: Chemistry and Chemical Technology; pt. 3: Spagyrical Discovery and Invention: Historical Survey from Cinnabar Elixirs to Synthetic Insulin by Joseph Needham, et al. (review). Technology and Culture, 18(4):728–730, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892022/pdf.

Malone:1978:THE

[Mal78] Patrick M. Malone. Textile History ed. by K. G. Ponting, S. D. Chapman (review). Technology and Culture, 19(1):145–147, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892065/pdf.

Mantel:1972:GCT

[Man72] Samuel J. Mantel, Jr. Government Contracting and Technological Change by Clarence H. Danhof (review). Technology and Culture, 13(1):96–98, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893451/pdf.

Mantel:1973:RIM

[Man73] Samuel J. Mantel, Jr. Research and Innovation in the Modern Corporation by Edwin Mansfield, et al. (review). Technology and Culture, 14(4):663–665, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892900/pdf.

Maquet:1970:RCI

[Maq70] Jacques Maquet. The Raw and the Cooked: Introduction to a Science of Mythology by Claude Lévi–Strauss (review). Technology and Culture, 11(4):613–615, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894049/pdf.

Marcson:1972:ASE

[Mar72a] Simon Marcson. Automation and Society ed. by Ellis L. Scott, Roger W. Bolz (review). Technology and Culture, 13(1):106–107, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893455/pdf.

Marcson:1972:EIC

[Mar72b] Simon Marcson. The Engineer in the Industrial Corporation by R. R. Ritti (review). Technology and Culture, 13(1):104–106, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893454/pdf.

Marinoni:1972:SLV

[Mar72c] Augusto Marinoni. The Sublimations of Leonardo da Vinci with a Translation of the Codex Trivulzianus by Raymond S. Stites (review). Technology and Culture, 13(2):301–309, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893469/pdf.

Marling:1976:TIA

[Mar76a] Karal Ann Marling. The Tasteful Interlude: American Interiors through the Camera’s Eye, 1860–1917 by William Seale (review). Technology and Culture, 17(4):773–776, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891838/pdf.

Marling:1976:WPE

[Mar76b] Karal Ann Marling. Winterthur Portfolio 10 ed. by Ian M. C. Quimby (review). Technology and Culture, 17(3):606–610, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891807/pdf.

Marks:1977:RDB

[Mar77a] Leonard H. Marks. Regulation of Direct Broadcasting from Satellites: The UN Involvement by Benno Signitzer (review). Technology and Culture, 18(3):561–563, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891974/pdf.

Marling:1977:APS

[Mar77b] Karal Ann Marling. The Art of Paul Sawyier by Arthur F. Jones (review). Technology and Culture, 18(3):565–567, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891976/pdf.

Marsh:1977:SIC

[Mar77c] Barbara Marsh. Social Issues in Computing by C. C. Gotlieb, A. Borodin (review). Technology and Culture, 18(1):133–135, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891897/pdf.

Marling:1979:GWP

[Mar79a] Karal Ann Marling. The German Werkbund: The Politics of Reform in the Applied Arts by Joan Campbell (review). Technology and Culture, 20(2):392–394, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890516/pdf.

Marvin:1979:PPA

[Mar79b] Carolyn Marvin. The Printing Press as an Agent of Change: Communications and Cultural Transformations in Early-Modern Europe by Elizabeth Eisenstein (review). Technology and Culture, 20 (4):793–797, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890578/pdf.

Massouh:1973:CCA

[Mas73] Michael Massouh. The Cable Car in America by George W. Hilton (review). Technology and Culture, 14(4):640–641, October 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892886/pdf.

Massouh:1977:SEH

[Mas77a] Michael Massouh. A Social and Economic History of the Near East in the Middle Ages by Eliahu Ashtor (review). Technology and Culture, 18(3):522–524, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891953/pdf.

Massouh:1977:ISR

[Mas77b] Michael Massouh. Innovations in street railways before electric traction: Tom L. Johnson’s contributions. Technology and Culture, 18(2):202–217, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891910/pdf.

Mattoon:1974:BLP

[Mat74] Robert H. Mattoon, Jr. Brazil: The Land and People by Rollie E. Poppino (review). Technology and Culture, 15(3):510–511, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892759/pdf.

Mattoon:1975:SIP

[Mat75] Robert H. Mattoon, Jr. The Sugar Industry in Pernambuco: Modernization without Change, 1840–1910 by Peter L. Eisenberg (review). Technology and Culture, 16(2):318–320, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891550/pdf.

Mathes:1976:TCS

[Mat76] J. C. Mathes. The Technological Catch and Society ed. by Iraj Zandi (review). Technology and Culture, 17(3):589–591, July 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891798/pdf.

Mattoon:1978:BBS

[Mat78] Robert H. Mattoon, Jr. Beginnings of Brazilian Science: Oswaldo Cruz, Medical Research, and Policy, 1890–1920 by Nancy Stepan (review). Technology and Culture, 19(3):546–547, July 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892137/pdf.

Mayr:1970:HAR

[May70] Otto Mayr. History of Automation in Russia before 1917 by A. V. Khramoi (review). Technology and Culture, 11(2):324–326, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894174/pdf.

Mayr:1971:ASC

[May71a] Otto Mayr. Adam Smith and the concept of the feedback system: Economic thought and technology in 18th-Century Britain. Technology and Culture, 12(1):1–22, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893997/pdf.

Mayr:1971:OGN

[May71b] Otto Mayr. Otto von Guerickes neue (sogenannte) Magdeburger Versuche über den leeren Raum ed. by Hans Schimank (review). Technology and Culture, 12(3):501–504, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893605/pdf.

Mayr:1975:JBC

[May75a] Otto Mayr. Joseph Bramah: a Century of Invention, 1743–1841 by Ian McNeil (review). Technology and Culture, 16(3):488–489, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891579/pdf.

Mayr:1975:YPE

[May75b] Otto Mayr. Yankee practice and engineering theory: Charles T. Porter and the dynamics of the high-speed steam engine. Technology and Culture, 16(4):570–602, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891597/pdf.

Mayr:1976:STR

[May76] Otto Mayr. The science-technology relationship as a historiographic problem. Technology and Culture, 17(4):663–673, October 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891817/pdf.

Mayr:1977:MTE

[May77] Otto Mayr. Moderne Technikgeschichte ed. by Karin Hausen, Reinhard Rürup (review). Technology and Culture, 18(4):736–738, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892026/pdf.

Mayo:1978:BCM

[May78a] Edward L. Mayo. The Business Career of Moses Taylor: Merchant, Finance Capitalist and Industrialist by Daniel P. Hodas (review). Technology and Culture, 19(1):123–125, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892054/pdf.

Mayr:1978:KTZ

[May78b] Otto Mayr. Kultur und Technik: Zeitschrift des Deutschen Museums München (review). Technology and Culture, 19(4):724–725, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892175/pdf.

Mayr:1971:CMD

[MBB^{+}71] Otto Mayr, John G. Burke, John J. Beer, James Lee Cate, and Thomas R. De Gregori. The Chicago Meeting, December 26–30, 1970. Technology and Culture, 12(2):238–268, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893961/pdf.

Mark:1974:SAH

[MCA74] Robert Mark, James K. Chiu, and John F. Abel. Stress analysis of historic structures: Maillart’s warehouse at Chiasso. Technology and Culture, 15(1):49–63, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892824/pdf.

McCutcheon:1970:NWC

[McC70a] W. A. McCutcheon. Navigable Waterways by L. T. C. Rolt (review). Technology and Culture, 11(3):435–437, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894109/pdf.

McCutcheon:1970:EIN

[McC70b] W. A. McCutcheon. The Engineering Industry of the North of Ireland by W. E. Coe (review). Technology and Culture, 11(3):437– 439, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894110/pdf.

McClellan:1973:SPS

[McC73] Woodford McClellan. Science and Philosophy in the Soviet Union by Loren R. Graham (review). Technology and Culture, 14(4):676–677, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892907/pdf.

McGinn:1976:WIB

[McG76] Robert E. McGinn. Workshop on the interrelations between science and technology, and ethics and values: Reston, Virginia, April 10–12, 1975. Technology and Culture, 17(2):249–255, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891715/pdf.

McKeon:1973:CEP

[McK73] Robert M. McKeon. Coulomb and the Evolution of Physics and Engineering in Eighteenth-Century France by C. Stewart Gillmor (review). Technology and Culture, 14(2 (part I)):294–296, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893268/pdf.

McKinnon:1979:BHN

[McK79] Sarah M. McKinnon. Building a House in New France by Peter N. Moogk (review). Technology and Culture, 20(1):202–203, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890462/pdf.

McLaughlin:1970:CRC

[McL70] Andrew McLaughlin. Challenge to Reason by C. West Churchman (review). Technology and Culture, 11(1):127–129, January 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894208/pdf.

McLuhan:1975:MLM

[McL75] Marshall McLuhan. McLuhan’s laws of the media. Technology and Culture, 16(1):74–78, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891500/pdf.

McLuhan:1976:MML

[McL76] Marshall McLuhan. Misunderstanding the media’s laws. Technology and Culture, 17(2):263, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891717/pdf.

McLuhan:1978:TTC

[McL78] Marshall McLuhan. Television: Technology and Cultural Form by Raymond Williams (review). Technology and Culture, 19(2):259– 261, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892101/pdf.

McLuhan:1979:DIT

[McL79] Marshall McLuhan. The Dialectic of Ideology and Technology: The Origins of Grammar and Future of Ideology by A. W. Gouldner (review). Technology and Culture, 20(1):244–246, January 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890486/pdf.

McMath:1975:PEA

[McM75] Robert C. McMath, Jr. The Political Economy of Agrarian Change: an Essay on the Green Revolution by Keith Griffin (review). Technology and Culture, 16(4):676–678, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891629/pdf.

McMahon:1976:AWA

[McM76a] A. Michal McMahon. America’s Wooden Age: Aspects of Its Early Technology ed. by Brooke Hindle (review). Technology and Culture, 17(4):748–749, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891825/pdf.

McMath:1976:IES

[McM76b] Robert C. McMath, Jr. The Improbable Era: The South since World War II by Charles P. Roland (review). Technology and Culture, 17 (4):768–769, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891835/pdf.

McMath:1977:CMS

[McM77] Robert C. McMath, Jr. Coal-mining Safety in the Progressive Period: The Political Economy of Reform by William Graebner (review). Technology and Culture, 18(1):125–127, January 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891893/pdf.

McMahon:1978:PPP

[McM78] A. Michal McMahon. Physics, Patents and Politics: a Biography of Charles Grafton Page by Robert C. Post (review). Technology and Culture, 19(2):221–222, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892081/pdf.

McNeill:1978:PCD

[McN78] W. H. McNeill. The Promise of the Coming Dark Age by L. S. Stavrianos (review). Technology and Culture, 19(2):263–265, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892103/pdf.

Meadows:1971:TTD

[Mea71] A. J. Meadows. Tracing technological development by linguistic evidence. Technology and Culture, 12(3):464–468, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893596/pdf.

Meeker:1978:IAN

[Mee78] Joseph W. Meeker. The imminent alliance: New connections among art, science, and technology. Technology and Culture, 19(2):187– 198, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892070/pdf.

Meier:1970:ITD

[Mei70] Richard L. Meier. Industrial Technologies for Developing Economies by Jack Baranson (review). Technology and Culture, 11(2):329–330, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894177/pdf.

Meighan:1971:APG

[Mei71a] Clement W. Meighan. Aspects of Prehistory by Grahame Clark (review). Technology and Culture, 12(4):628–629, October 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893561/pdf.

Meighan:1971:FTS

[Mei71b] Clement W. Meighan. Frozen Tombs of Siberia: The Pazyryk Burials of Iron-Age Horsemen by Sergei I. Rudenko (review). Technology and Culture, 12(2):332–334, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893965/pdf.

Meighan:1972:HMP

[Mei72] Clement W. Meighan. From the Hand of Man: Primitive and Preindustrial Technologies by Robert F. G. Spier (review). Technology and Culture, 13(1):66, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893433/pdf.

Meier:1974:TID

[Mei74] Hugo A. Meier. Technological innovation and the decorative arts: 19th annual winterthur conference, March 1973. Technology and Culture, 15(2):228–241, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892793/pdf.

Merritt:1976:LSV

[Mer76a] Raymond H. Merritt. Lex et Scientia. Vol. 10 (October–December 1974) (review). Technology and Culture, 17(3):591–594, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891799/pdf.

Merritt:1976:TRS

[Mer76b] Raymond H. Merritt. Technology and Reform: Street Railways and the Growth of Milwaukee, 1887–1900 by Clay McShane (review). Technology and Culture, 17(3):551–552, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891777/pdf.

Meursinge:1971:PET

[Meu71] John H. Meursinge. Practical experience in the transfer of technology. Technology and Culture, 12(3):469–470, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893597/pdf.

Meursinge:1977:SVL

[Meu77] John H. Meursinge. Short- versus long-term patterns. Technology and Culture, 18(3):508, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891947/pdf.

Mileti:1978:TOM

[MGM78] Dennis S. Mileti, David F. Gillespie, and Elizabeth Morrissey. Technology and organizations: Methodological deficiencies and lacunae. Technology and Culture, 19(1):83–92, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892041/pdf.

Michalos:1970:TEG

[Mic70a] Alex C. Michalos. Technology and Empire by George Grant (review). Technology and Culture, 11(2):340–341, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894183/pdf.

Michalos:1970:SAH

[Mic70b] Alex C. Michalos. The Sciences of the Artificial by Herbert A. Simon (review). Technology and Culture, 11(1):118–120, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894205/pdf.

Middleton:1978:HMH

[Mid78] W. E. K. Middleton. The History of Meteorology to 1800 by H. Howard Frisinger (review). Technology and Culture, 19(3):522–523, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892125/pdf.

Milch:1976:NPN

[Mil76] Jerome Milch. National Politics and Nuclear Technology: Nuclear Reactor Development in Western Europe by Henry Nau (review). Technology and Culture, 17(1):180–182, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891693/pdf.

Mingay:1977:FML

[Min77a] G. E. Mingay. First Majority–Last Minority: The Transforming of Rural Life in America by John L. Shover (review). Technology and Culture, 18(4):710–712, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892011/pdf.

Mingay:1977:VLL

[Min77b] G. E. Mingay. Village Life and Labour ed. by Raphael Samuel (review). Technology and Culture, 18(4):708–710, October 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892010/pdf.

Minai:1978:CWR

[Min78] Asghar Talaye Minai. The Camel and the Wheel by Richard W. Bulliet (review). Technology and Culture, 19(2):204–207, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892072/pdf.

Minchinton:1979:ETM

[Min79a] W. E. Minchinton. Early tide mills: Some problems. Technology and Culture, 20(4):777–786, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890574/pdf.

Minchinton:1979:JTM

[Min79b] Walter E. Minchinton. John Taylor, Mining Entrepreneur and Engineer, 1779–1863 by Roger Burt (review). Technology and Culture, 20(3):637–638, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890541/pdf.

Mitcham:1973:I

[MM73] Carl Mitcham and Robert Mackly. Introduction. Technology and Culture, 14(2 (part II)):vii–xv, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893256/pdf.

Major:1974:PCP

[MM74] John S. Major and David C. Major. The Pattern of the Chinese Past by Mark Elvin (review). Technology and Culture, 15(3):511– 514, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892760/pdf.

McLuhan:1973:ACE

[MN73] Marshall McLuhan and Barrington Nevitt. The argument: Causality in the electric world. Technology and Culture, 14(1):1–18, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893349/pdf.

Moeller:1971:TMN

[Moe71] Walter O. Moeller. Textile Manufacture in the Northern Roman Provinces by J. P. Wild (review). Technology and Culture, 12(2):334– 335, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893966/pdf.

Mokyr:1977:FTM

[Mok77] Joel Mokyr. Failed Transitions to Modern Industrial Society: Renaissance Italy and Seventeenth Century Holland ed. by Paul M. Hohenberg, Frederick Krantz (review). Technology and Culture, 18(1):94–95, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891873/pdf.

Molella:1978:SIT

[Mol78] Arthur Molella. The Social Impact of the Telephone ed. by Ithiel de Sola Pool (review). Technology and Culture, 19(4):762–764, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892196/pdf.

Molgaard:1979:JHW

[Mol79] J. Molgaard. Joule, heat, and work. Technology and Culture, 20(3):686–687, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890566/pdf.

Montgomery:1971:UAM

[Mon71] Douglas G. Montgomery. Urban Analysis by Maurice D. Kilbridge, et al. (review). Technology and Culture, 12(4):677–679, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893585/pdf.

Montgomery:1972:MEP

[Mon72] Douglas Montgomery. The Metropolitan Economy: The Process of Employment Expansion by Thomas M. Stanback, Jr., Richard V. Knight (review). Technology and Culture, 13(2):337–338, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893490/pdf.

Montgomery:1975:TCL

[Mon75] Douglas G. Montgomery. Technology and Civic Life: Making and Implementing Development Decisions by John D. Montgomery (review). Technology and Culture, 16(4):689–691, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891636/pdf.

Morris:1970:SUE

[Mor70] Bertram Morris. The Search for Understanding ed. by Caryl P. Haskins (review). Technology and Culture, 11(4):653–655, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894071/pdf.

Morgan:1971:ACB

[Mor71] William Morgan. The Architecture of Charles Bulfinch by Harold Kirker (review). Technology and Culture, 12(2):347–349, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893973/pdf.

Moravcsik:1972:SGC

[Mor72a] Michael J. Moravcsik. Science: Growth and Change by Henry W. Menard (review). Technology and Culture, 13(4):673–674, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893407/pdf.

Morris:1972:CNE

[Mor72b] Bertram Morris. Chance and Necessity: an Essay on the Natural Philosophy of Modern Biology by Jacques Monod (review). Technology and Culture, 13(4):662–663, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893401/pdf.

Morris:1974:JDC

[Mor74] Robert J. Morris, Jr. John Dalton, Critical Assessments of His Life and Science by Arnold Thackray (review). Technology and Culture, 15(1):100–103, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892838/pdf.

Morrell:1977:SAB

[Mor77a] J. B. Morrell. Science and the Ante-Bellum American College by Stanley M. Guralnick (review). Technology and Culture, 18(3):547– 548, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891966/pdf.

Morris:1977:CT

[Mor77b] Bertram Morris. The context of technology. Technology and Culture, 18(3):395–418, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891937/pdf.

Morris:1977:MYE

[Mor77c] Stephanie Morris. Museums Yearbook: 1976 ed. by Valerie Shepard (review). Technology and Culture, 18(1):150–151, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891905/pdf.

Morgan:1978:UET

[Mor78] Robert P. Morgan. University education for technology and public policy: St. Louis, December 1976. Technology and Culture, 19 (1):106–108, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892045/pdf.

Morris:1979:RPT

[Mor79] Bertram Morris. Research in Philosophy and Technology: an Annual Compilation of Research ed. by Paul T. Durbin (review). Technology and Culture, 20(3):680–681, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890563/pdf.

Moss:1975:HRS

[Mos75a] Roger W. Moss, Jr. The Houses of the Royal Society of Arts: a History and a Guide by D. G. C. Allan (review). Technology and Culture, 16(4):650–651, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891616/pdf.

Moszynski:1975:HMZ

[Mos75b] J. R. Moszynski. Historia magnetycznego zapisu obrazów [The history of magnetic video recording] by Roman Wajdowicz (review). Technology and Culture, 16(1):106–107, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891515/pdf.

Moszynski:1978:SMZ

[Mos78] J. R. Moszynski. Studia i Materiaiy z dziejów nauki polskiej (review). Technology and Culture, 19(4):774–775, October 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892203/pdf.

Moszynski:1979:DEM

[Mos79] J. R. Moszynski. The Development of Excavating Machinery from Historical Perspective [Rozwoj Koparek w Przekroju Historycznym] by Alfred Wislicki (review). Technology and Culture, 20(3):630, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890536/pdf.

Muendel:1974:HMM

[Mue74] John Muendel. The horizontal mills of Medieval Pistoia. Technology and Culture, 15(2):194–225, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892791/pdf.

Multhauf:1971:FCP

[Mul71] Robert P. Multhauf. The French crash program for saltpeter production, 1776–94. Technology and Culture, 12(2):163–181, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893953/pdf.

Multhauf:1972:JBH

[Mul72a] Robert P. Multhauf. Jac. Berzelius: His Life and Work by J. Erik Jorpes (review). Technology and Culture, 13(1):73–74, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893438/pdf.

Multhauf:1972:MMP

[Mul72b] Robert P. Multhauf. The Myth of the Machine: The Pentagon of Power by Lewis Mumford (review). Technology and Culture, 13(2):295–299, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893467/pdf.

Muller:1973:CWS

[Mul73a] Herbert J. Muller. Can We Survive Our Future? ed. by G. R. Urban (review). Technology and Culture, 14(1):100–108, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893370/pdf.

Muller:1973:ECH

[Mul73b] Herbert J. Muller. Empire and Communications by Harold A. Innis (review). Technology and Culture, 14(3):482–485, July 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893220/pdf.

Muller:1974:TPS

[Mul74a] Herbert J. Muller. Technology, Power, and Social Change ed. by Charles A. Thrall, Jerold M. Starr, and: Tradition for the Future by Mirrit Boutros Ghali (review). Technology and Culture, 15(1):82–86, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892829/pdf.

Muller:1974:PIP

[Mul74b] Herbert J. Muller. The Post-industrial Prophets: Interpretations of Technology by William Kuhns (review). Technology and Culture, 15(2):357–360, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892814/pdf.

Multhauf:1974:CAC

[Mul74c] Robert P. Multhauf. Chemical Apparatus by C. R. Hill (review). Technology and Culture, 15(2):365–367, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892817/pdf.

Multhauf:1974:MHQ

[Mul74d] Robert P. Multhauf. Mercury: a History of Quicksilver by Leonard J. Goldwater (review). Technology and Culture, 15(1):91–92, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892832/pdf.

Multhauf:1974:HTS

[Mul74e] Robert P. Multhauf. The historiography of technology: Some observations on the state of the history of technology. Technology and Culture, 15(1):1–12, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892820/pdf. See comment [De 74e].

Multhauf:1976:GCP

[Mul76] Robert P. Multhauf. Geology, chemistry, and the production of common salt. Technology and Culture, 17(4):634–645, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891815/pdf.

Multhauf:1978:CCR

[Mul78] Robert P. Multhauf. A Century of Chemistry: The Role of Chemists and the American Chemical Society ed. by American Chemical Society (review). Technology and Culture, 19(1):127–129, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892057/pdf.

Multhauf:1979:MQS

[Mul79a] Robert P. Multhauf. Miners, Quarrymen and Saltworkers ed. by Raphael Samuel (review). Technology and Culture, 20(4):802–804, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890582/pdf.

Multhauf:1979:CPP

[Mul79b] Robert P. Multhauf. The Chemical Philosophy: Paracelsian Science and Medicine in the Sixteenth and Seventeenth Centuries by Allen G. Debus (review). Technology and Culture, 20(1):213–214, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890469/pdf.

Murphy:1979:ISG

[Mur79] Richard C. Murphy. Iron and Steed in the German Inflation, 1916– 1923 by Gerald D. Feldman (review). Technology and Culture, 20(2):382–384, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890511/pdf.

Needler:1970:EHA

[Nee70] Winifred Needler. The Excavations at Helwan: Art and Civilization in the First and Second Egyptian Dynasties by Zaki Y. Saad (review). Technology and Culture, 11(4):615–617, October 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894050/pdf.

Needham:1979:AME

[Nee79] Joseph Needham. Asia in the Making of Europe. Vol. 2: a Century of Wonder. Book 2: The Literary Arts. Book 3: The Scholarly Disciplines by Donald F. Lach (review). Technology and Culture, 20(3):622–624, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890533/pdf.

Nelson:1972:SWC

[Nel72] Alan H. Nelson. Six-wheeled carts: an underview. Technology and Culture, 13(3):391–416, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893923/pdf.

Nelson:1979:MDI

[Nel79] Daniel Nelson. A Matter of Dignity: Inquiries into the Humanization of Work ed. by W. J. Heisler, John W. Houck (review). Technology and Culture, 20(3):670–672, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890558/pdf.

Nesbitt:1977:YAG

[Nes77] Alexander Nesbitt. 200 Years of American Graphic Art: a Retrospective Survey of the Printing Arts and Advertising since the Colonial Period by Clarence P. Hornung, Fridolf Johnson (review). Technology and Culture, 18(2):247–248, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891920/pdf.

Nevitt:1979:CIE

[Nev79] Barrington Nevitt. The Changing Information Environment by John McHale (review). Technology and Culture, 20(3):677–679, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890562/pdf.

Nichols:1971:FBP

[Nic71] David A. Nichols. Federal Budget Priorities for Research and Development by Michael S. March (review). Technology and Culture, 12(2):374–375, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893987/pdf.

Nieburg:1970:NPW

[Nie70a] H. L. Nieburg. Nuclear Proliferation by Walter B. Wentz (review). Technology and Culture, 11(2):341–343, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894184/pdf.

Nieburg:1970:PWI

[Nie70b] H. L. Nieburg. The Politics of Weapons Innovation: The Thor– Jupiter Controversy by Michael H. Armacost (review). Technology and Culture, 11(3):470–471, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894127/pdf.

Noble:1971:ATS

[Nob71] Joseph Veach Noble. Art and Technology: a Symposium on Classical Bronzes ed. by Suzannah Doeringer, et al. (review). Technology and Culture, 12(4):629–630, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893562/pdf.

Noether:1977:III

[Noe77] Emiliana P. Noether. Industrial Imperialism in Italy, 1908–1915 by Richard A. Webster (review). Technology and Culture, 18(1): 144–146, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891902/pdf.

Norberg:1975:BFN

[Nor75] Arthur L. Norberg. Benjamin Franklin: New World Physicist by Raymond J. Seeger (review). Technology and Culture, 16(4):653–655, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891618/pdf.

Norberg:1977:SSO

[Nor77a] Arthur L. Norberg. Syntony and Spark: The Origins of Radio by Hugh G. J. Aitken (review). Technology and Culture, 18(3):551– 553, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891968/pdf.

North:1977:GRS

[Nor77b] J. D. North. Greek and Roman Sundials by Sharon L. Gibbs (review). Technology and Culture, 18(2):238–239, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891915/pdf.

Norberg:1978:TPA

[Nor78] Arthur L. Norberg. Turning Points in American Electrical History ed. by James E. Brittain (review). Technology and Culture, 19(3):564–567, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892149/pdf.

Nuttall:1977:EM

[Nut77] R. H. Nuttall. Edinburgh and medicine. Technology and Culture, 18(4):666–669, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891996/pdf.

Nuttall:1978:BMA

[Nut78] R. H. Nuttall. Benjamin Martin, Author, Instrument-Maker, and “Country Showman” by John R. Millburn (review). Technology and Culture, 19(1):114–116, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892049/pdf.

Nuttall:1979:APC

[Nut79a] R. H. Nuttall. Andrew Pritchard’s contribution to metallurgical microscopy. Technology and Culture, 20(3):569–577, July 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890525/pdf.

Nuttall:1979:BWM

[Nut79b] R. H. Nuttall. British Weights and Measures: a History from Antiquity to the Seventeenth Century by Ronald Edward Zupko (review). Technology and Culture, 20(1):210–211, January 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890467/pdf.

Nuttall:1979:CPI

[Nut79c] R. H. Nuttall. The Chamber of Physics: Instruments in the History of Sciences Collections of the Royal Swedish Academy of Sciences, Stockholm by Gunnar Pipping (review). Technology and Culture, 20(3):631–633, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890538/pdf.

OBrien:1971:SCA

[O’B71] Kenneth O’Brien. Symposium on 19th-Century American Science: Northwestern University, March 30–April 2, 1970. Technology and Culture, 12(1):56–63, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894001/pdf.

Okada:1978:JIS

[Oka78] Kazuhide Okada. Japanese Industrialization and Its Social Consequences ed. by Hugh Patrick (review). Technology and Culture, 19(2):244–246, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892093/pdf.

Olson:1974:TBR

[Ols74] Lawrence Olson. Three Bowls of Rice by Kusum Nair (review). Technology and Culture, 15(4):650–651, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892719/pdf.

Olson:1976:EFA

[Ols76] Laura Katz Olson. Energy: Future Alternatives and Risks by Academy Forum, National Academy of Sciences, and: A Time to Choose: America’s Energy Future by Energy Policy Project of the Ford Foundation (review). Technology and Culture, 17(1):177–180, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891692/pdf.

Olson:1977:FKS

[Ols77] Richard G. Olson. The Force of Knowledge: The Scientific Dimension of Society by John Ziman (review). Technology and Culture, 18(2):264–267, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891931/pdf.

Olszewski:1973:PGO

[OO73] Eugeniusz Olszewski and Bołeslaw Orłowski. Polish Galician oil. Technology and Culture, 14(2 (part I)):173–174, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893261/pdf.

Ordway:1970:HAS

[Ord70] Frederick I. Ordway. History of astronautics symposium: Mar Del Plata, Argentina, October 1969. Technology and Culture, 11(3):407– 416, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894105/pdf.

Ordway:1978:MVP

[Ord78] Frederick I. Ordway. Max Valier: a Pioneer of Space Travel by I. Essers (review). Technology and Culture, 19(2):251–253, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892097/pdf.

Orlowski:1973:PWS

[Orł73] Bolesław Orłowski. Polish works and serials in the history of technology. Technology and Culture, 14(3):461–473, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893216/pdf.

Ospovat:1979:STK

[Osp79] Alexander M. Ospovat. Der Seigerhüttenprozesz: Die Technologie des Kupferseigerns nach dem frühen metallurgischen Schrifttum by Lothar Suhling (review). Technology and Culture, 20(1):196–197, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890459/pdf.

Overton:1971:HSJ

[Ove71] Richard C. Overton. The Hiawatha Story by Jim Scribbins (review). Technology and Culture, 12(2):354–355, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893976/pdf.

Overfield:1975:CBI

[Ove75] Richard A. Overfield. Charles E. Bessey: The impact of the “New” botany on American agriculture, 1880–1910. Technology and Culture, 16(2):162–181, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891529/pdf.

Owens:1973:CEP

[Owe73] Joseph Owens. Comment: Effects precede causes. Technology and Culture, 14(1):19–21, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893350/pdf.

Packer:1974:VLR

[Pac74] James E. Packer. Vitruvius and Later Roman Building Manuals by Hugh Plommer (review). Technology and Culture, 15(1):89–91, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892831/pdf.

Packer:1976:CGR

[Pac76] James Packer. Craftsmen in Greek and Roman Society by Alison Burford (review). Technology and Culture, 17(3):537–541, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891770/pdf.

Packer:1978:AGA

[Pac78] James Packer. Ancient Greek Architects at Work by J. J. Coulton (review). Technology and Culture, 19(3):528–530, July 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892128/pdf.

Palliser:1979:EPP

[Pal79] David M. Palliser. Economic Policy and Projects: The Development of a Consumer Society in Early Modern England by Joan Thirsk (review). Technology and Culture, 20(3):635–636, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890540/pdf.

Parry:1970:KMP

[Par70] Albert Parry. K. Marks i problemy tekhniki [Karl Marx and the Problems of Technology] by A. A. Kuzin (review). Technology and Culture, 11(2):326–327, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894175/pdf.

Parascandola:1971:HPP

[Par71a] John Parascandola. The History of Penicillin Production ed. by Albert E. Elder (review). Technology and Culture, 12(2):355–357, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893977/pdf.

Parry:1971:AKN

[Par71b] Albert Parry. Andrei Konstantinovich Nartov by F. N. Zagorsky (review). Technology and Culture, 12(1):93–94, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894008/pdf.

Parr:1972:NFM

[Par72a] J. Gordon Parr. Non-ferrous Metals Casting, History and Forecast by Jerzy W. Meier (review). Technology and Culture, 13(2): 312, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893472/pdf.

Parr:1972:SMR

[Par72b] J. Gordon Parr. The sinking of the Ma Robert: an excursion into mid-19th-Century steelmaking. Technology and Culture, 13(2):209– 225, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893465/pdf.

Parkhurst:1974:LWC

[Par74] E. H. Parkhurst, Jr. The Lancashire Watch Company, 1889–1910 by Alan Smith, Henry G. Abbott (review). Technology and Culture, 15(3):505–506, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892756/pdf.

Parr:1978:HMR

[Par78] J. Gordon Parr. A History of Metallurgy by R. F. Tylecote (review). Technology and Culture, 19(1):111–112, January 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892047/pdf.

Parrott:1979:SIU

[Par79] Bruce Parrott. The Scientific Intelligentsia in the USSR: Structure and Dynamics of Personnel ed. by D. M. Gvishiani, et al. (review). Technology and Culture, 20(1):246–248, January 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890487/pdf.

Paterson:1975:OSS

[Pat75] Alan M. Paterson. Oranges, soot, and science: The development of frost protection in California. Technology and Culture, 16(3):360– 376, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891561/pdf.

Paul:1971:PHA

[Pau71] Rodman W. Paul. A Pictorial History of American Mining: The Adventure and Drama of Finding and Extracting Nature’s Wealth from the Earth, from Pre-Columbian Times to the Present by Howard N. Sloane, Lucille L. Sloane (review). Technology and Culture, 12 (4):650–652, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893572/pdf.

Peirce:1972:TIT

[Pei72] William S. Peirce. Technology and Information Transfer: a Survey of Practice in Industrial Organizations by Richard S. Rosenbloom, Francis W. Wolek (review). Technology and Culture, 13(3):519– 520, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893947/pdf.

Penn:1975:PSL

[Pen75] Theodore Z. Penn. Peter Stubs and the Lancashire Hand Tool Industry by E. Surrey Dane (review). Technology and Culture, 16(2):299– 301, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891541/pdf.

Penick:1977:AEH

[Pen77] James Penick, Jr. American Environmental History: The Exploitation and Conservation of Natural Resources by Joseph M. Petulla (review). Technology and Culture, 18(4):712, October 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892012/pdf.

Penn:1979:MPM

[Pen79] Theodore Z. Penn. The Mill at Philipsburg Manor Upper Mills and a Brief History of Milling by Charles Howell, Allan Keller (review). Technology and Culture, 20(3):639–640, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890543/pdf.

Perrot:1970:AEG

[Per70a] Paul N. Perrot. Ancient Egyptian Glass and Glazes in the Brooklyn Museum by Elizabeth Riefstahl (review). Technology and Culture, 11(2):299–301, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894159/pdf.

Perry:1970:DLR

[Per70b] Robert Perry. Development of the Lifesaving Rocket: a Study in 19th Century Technological Fallout by Mitchell R. Sharpe (review). Technology and Culture, 11(2):316–318, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/894168/pdf.

Perrot:1972:SGI

[Per72] Paul N. Perrot. The “Stourbridge Glass” Industry in the Nineteenth Century: a Study of the Glass Industry in Stourbridge, Brierley Hill, and Dudley by H. Jack Haden (review). Technology and Culture, 13(3):501–502, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893936/pdf.

Perrin:1973:HAB

[Per73] Eugene V. Perrin. Human Aspects of Biomedical Innovation by Everett Mendelsohn, et al. (review). Technology and Culture, 14(2 (part I)):340–342, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893288/pdf.

Perry:1974:MPP

[Per74] Robert Perry. Most Probable Position: a History of Aerial Navigation to 1941 by Monte Duane Wright (review). Technology and Culture, 15(1):116–119, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892846/pdf.

Perkins:1978:RTW

[Per78a] John H. Perkins. Reshaping technology in wartime: The effect of military goals on entomological research and insect-control practices. Technology and Culture, 19(2):169–186, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892069/pdf.

Pershey:1978:LPM

[Per78b] Edward J. Pershey. The Lower Peninsula of Michigan: an Inventory of Historic Engineering and Industrial Sites by Charles K. Hyde (review). Technology and Culture, 19(2):265–266, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892104/pdf.

Peterson:1975:GHJ

[Pet75] Charles E. Peterson. The Glass House by John Hix (review). Technology and Culture, 16(3):498–500, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891584/pdf.

Piaskowski:1970:TSW

[Pia70] J. Piaskowski. Termodynamika stopów w rozwoju history cznym [History of the thermodynamics of alloys] by Eugeniusz Tyrkiel (review). Technology and Culture, 11(3):459–460, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894122/pdf.

Piaskowski:1971:WZW

[Pia71] J. Piaskowski. Wykres Żelazo-Wegiel w rozwoju historycznym [History of the iron-carbon diagram] by E. Tyrkiel (review). Technology and Culture, 12(4):652–653, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893573/pdf.

Pickett:1977:ATP

[Pic77] William B. Pickett. American technology — past, present, and future: Rose–Hulman Institute, Terre Haute, Indiana, April 1–3, 1976. Technology and Culture, 18(4):651–654, October 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891993/pdf.

Ponting:1979:TH

[Pon79] K. G. Ponting. Textile history. Technology and Culture, 20(3): 692, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890570/pdf.

Porter:1974:AGC

[Por74] Glenn Porter. The Age of Giant Corporations: a Microeconomic History of American Business, 1914–1970 by Robert Sobel (review). Technology and Culture, 15(1):119–121, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892847/pdf.

Porter:1976:ISE

[Por76] Alan L. Porter. The Impacts of Snow Enhancement by Leo W. Weisbecker (review). Technology and Culture, 17(3):594–596, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891800/pdf.

Post:1972:PLF

[Pos72] Robert C. Post. The page locomotive: Federal sponsorship of invention in mid-19th-Century America. Technology and Culture, 13(2):140–169, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893463/pdf.

Post:1975:BPS

[Pos75a] Robert C. Post. British Paddle Steamers by Geoffrey Body, and: The British Internal-Combustion Locomotive: 1894–1940 by Brian Webb, and: The Fairlie Locomotive by Rowland A. S. Abbott (review). Technology and Culture, 16(2):301–304, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891542/pdf.

Post:1975:GAT

[Pos75b] Robert C. Post. The Golden Age of Tramways by Charles Klapper (review). Technology and Culture, 16(1):98–101, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891511/pdf.

Post:1976:HCC

[Pos76a] Robert C. Post. Horsecars, Cable Cars and Omnibuses: All 107 Photographs from the John Stephenson Company Album, 1888 by John H. White, Jr. (review). Technology and Culture, 17(4):765, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891833/pdf.

Post:1976:RUM

[Pos76b] Robert C. Post. Rails under the Mighty Hudson: The Story of the Hudson Tubes, the Pennsy Tunnels and Manhattan Transfer by Brian J. Cudahy (review). Technology and Culture, 17(4):763–765, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891832/pdf.

Post:1976:TST

[Pos76c] Robert C. Post. Transit’s Stepchild: The Trolley Coach by Mac Sebree, Paul Ward (review). Technology and Culture, 17(1):163–165, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891685/pdf.

Post:1976:LVS

[Pos76d] Robert C. Post. “Liberalizers” versus “Scientific Men” in the Antebellum Patent Office. Technology and Culture, 17(1):24–54, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891650/pdf.

Post:1978:TTR

[Pos78] Robert C. Post. Tramways and Trolleys: The Rise of Urban Mass Transit in Europe by John P. McKay (review). Technology and Culture, 19(4):746–747, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892187/pdf.

Post:1979:PRT

[Pos79] Robert C. Post. Personal Rapid Transit II: Progress, Problems, and Potential in a Promising New Form of Public Transportation Reported at the 1973 International Conference on Personal Rapid Transit ed. by J. Edward Anderson (review). Technology and Culture, 20(2):375–376, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890507/pdf.

Porter:1977:PMD

[PPH^{+}77] Glen Porter, Robert C. Post, David A. Hounshell, Brooke Hindle, James E. Brittain, Raymond H. Merritt, Carroll Pursell, Stephanie A. Morris, and Martha Moore Trescott. The Philadelphia Meeting, December 28–30, 1976. Technology and Culture, 18(3):479– 497, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891943/pdf.

Price:1975:VSU

[Pri75] David H. Price. Vietnam Studies, U.S. Army Engineers, 1965– 1970 by Robert R. Ploger (review). Technology and Culture, 16 (4):666–667, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891624/pdf.

Pi-Sunyer:1971:LDL

[PS71a] Oriol Pi-Sunyer. Labor and Development in Latin America by Joseph R. Ramos (review). Technology and Culture, 12(4):679–681, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893586/pdf.

Pi-Sunyer:1971:VET

[PS71b] Oriol Pi-Sunyer. Vunamami, Economic Transformation in a Traditional Society by Richard F. Salisbury (review). Technology and Culture, 12(3):525–528, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893618/pdf.

Pi-Sunyer:1979:CCT

[PS79] Oriol Pi-Sunyer. Contemporary Change in Traditional Communities of Mexico and Peru by Charles J. Erasmus, et al. (review). Technology and Culture, 20(3):655–657, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890550/pdf.

Pursell:1970:AVR

[Pur70] Carroll Pursell. America the Vanishing: Rural Life and the Price of Progress ed. by Samuel R. Ogden (review). Technology and Culture, 11(4):641–642, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894064/pdf.

Pursell:1971:JWS

[Pur71] Carroll W. Pursell, Jr. James Watt and the Steam Revolution. A Documentary History by Eric Robinson and A. E. Musson (review). Technology and Culture, 12(1):99, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894011/pdf.

Pursell:1974:DCI

[Pur74] Carroll Pursell. Destroy to Create: Interaction with the Natural Environment in the Building of America ed. by Duke Frederick, et al. (review). Technology and Culture, 15(1):106, January 1974. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892840/pdf.

Pursell:1975:TCT

[Pur75] Carroll W. Pursell, Jr. A trade catalog on the transmission of power by wire rope. Technology and Culture, 16(1):70–73, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891499/pdf.

Pursell:1976:TSP

[Pur76a] Carroll W. Pursell, Jr. The Technical Society of the Pacific Coast, 1884–1914. Technology and Culture, 17(4):702–717, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891820/pdf.

Pursell:1976:TCA

[Pur76b] Carroll W. Pursell, Jr. Testing a carriage: The “American industry” series of scientific American. Technology and Culture, 17(1):82–92, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891652/pdf.

Pursell:1978:RFJ

[Pur78] Carroll Pursell. Robert Fulton by John S. Morgan (review). Technology and Culture, 19(4):740–741, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892183/pdf.

Pursell:1979:GTG

[Pur79a] Carroll W. Pursell, Jr. Government and technology in the Great Depression. Technology and Culture, 20(1):162–174, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890457/pdf.

Pursell:1979:I

[Pur79b] Carroll W. Pursell, Jr. Introduction. Technology and Culture, 20(1):1–2, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890449/pdf.

Quinn:1971:TFI

[Qui71] James Brian Quinn. Technological Forecasting for Industry and Government, Methods and Applications ed. by James R. Bright (review). Technology and Culture, 12(2):376–378, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893989/pdf.

Quimby:1974:DAW

[Qui74] Ian M. G. Quimby. Design and Aesthetics in Wood ed. by Eric A. Anderson, George F. Earle (review). Technology and Culture, 15 (1):132–135, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892853/pdf.

Quinn:1977:FCM

[Qui77] David B. Quinn. Five Centuries of Map Printing ed. by David Woodward (review). Technology and Culture, 18(1):91–92, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891871/pdf.

Quimby:1978:AAG

[Qui78] Maureen O’Brien Quimby. The Arts and Architecture of German Settlements in Missouri: a Survey of a Vanishing Culture by Charles Van Ravenswaay (review). Technology and Culture, 19(4):731–733, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892180/pdf.

Rae:1971:FHF

[Rae71a] John B. Rae. The First Henry Ford: a Study in Personality and Business Leadership by Anne Jardim (review). Technology and Culture, 12(3):516–517, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893612/pdf.

Rae:1971:TDM

[Rae71b] John B. Rae. The Technical Development of Modern Aviation by Ronald Miller, David Sawers (review). Technology and Culture, 12(2):358–360, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893978/pdf.

Rae:1973:HFI

[Rae73a] John B. Rae. Harry Ferguson, Inventor and Pioneer by Colin Fraser (review). Technology and Culture, 14(3):511–513, July 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893235/pdf.

Rae:1973:MPA

[Rae73b] John B. Rae. Master Planning the Aviation Environment ed. by Angelo J. Cerchione, et al. (review). Technology and Culture, 14 (4):656–657, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892896/pdf.

Rae:1973:TRS

[Rae73c] John B. Rae. The Turnpike Road System in England, 1663–1840 by William Albert (review). Technology and Culture, 14(2 (part I)):307– 308, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893275/pdf.

Rae:1974:ESE

[Rae74] John B. Rae. The Engineer and Society ed. by E. G. Semmler (review). Technology and Culture, 15(3):523–524, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892766/pdf.

Rae:1975:EP

[Rae75] John B. Rae. Engineers are people. Technology and Culture, 16(3):404–418, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891563/pdf.

Rae:1976:IAG

[Rae76] John B. Rae. The Industrial Archaeology of Glasgow by John R. Hume (review). Technology and Culture, 17(4):798–800, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891851/pdf.

Rae:1979:BMI

[Rae79a] John B. Rae. The British Motor Industry, 1896–1939 by Kenneth Richardson (review). Technology and Culture, 20(1):233–235, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890480/pdf.

Rae:1979:WAI

[Rae79b] John B. Rae. The World Automotive Industry by Gerald Bloomfield (review). Technology and Culture, 20(2):376–379, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890508/pdf.

Rahman:1977:SAT

[Rah77] A. Rahman. Seminar on alternative technology: Simla, India, September 1975. Technology and Culture, 18(1):62–63, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891865/pdf.

Rahman:1979:STR

[Rah79] A. Rahman. Scientific-Technological Revolution: Social Aspects by Ralf Dahrendorf, et al. (review). Technology and Culture, 20 (4):837–839, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890597/pdf.

Ramberg:1970:EHE

[Ram70] E. G. Ramberg. Early History of the Electron Microscope by L. Marton (review). Technology and Culture, 11(2):320–321, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894171/pdf.

Rand:1972:HIM

[Ran72] Walter Rand. History and the Idea of Mankind ed. by W. Warren Wagar (review). Technology and Culture, 13(4):682–684, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893412/pdf.

Randall:1976:CSI

[Ran76] A. J. Randall. Class Struggle and the Industrial Revolution: Early Industrial Capitalism in Three English Towns by John Foster (review). Technology and Culture, 17(2):375–377, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891726/pdf.

Rapp:1975:GCN

[Rap75] Richard T. Rapp. Guerra e commercio nell’evoluzione della marina genovese tra XV e XVII secolo by Centro per la storia della tecnica in Italia del Consiglio nazionale delle ricerche (review). Technology and Culture, 16(2):289–291, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891536/pdf.

Rapp:1977:TPI

[Rap77] Friedrich Rapp. Das technologische Patt: Innovationen überwinden die Depression by Gerhard Mensch (review). Technology and Culture, 18(1):103–105, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891881/pdf.

Rasmussen:1971:AMP

[Ras71] Wayne D. Rasmussen. Agroclimatological Methods: Proceedings of the Reading Symposium by UNESCO (review). Technology and Culture, 12(1):150–151, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894035/pdf.

Rasmussen:1974:ULE

[Ras74] Wayne D. Rasmussen. The Use of the Land: Essays on the History of American Agriculture by John T. Schlebecker (review). Technology and Culture, 15(4):647–648, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892717/pdf.

Rasmussen:1975:ASW

[Ras75] Wayne D. Rasmussen. The Agricultural Systems of the World: an Evolutionary Approach by D. B. Grigg (review). Technology and Culture, 16(4):616–617, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891601/pdf.

Rasmussen:1976:WWT

[Ras76] Wayne D. Rasmussen. Whereby We Thrive: a History of American Farming, 1607–1972 by John T. Schlebecker (review). Technology and Culture, 17(3):543–544, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891772/pdf.

Ravetz:1970:MED

[Rav70] J. R. Ravetz. Machina ex Deo. Essays in the Dynamism of Western Culture by Lynn White, Jr. (review). Technology and Culture, 11(1):96–97, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894195/pdf.

Ravetz:1972:SHJ

[Rav72] J. R. Ravetz. Science in History by J. D. Bernal (review). Technology and Culture, 13(4):664–666, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893402/pdf.

Ray:1977:CRL

[Ray77] Dale C. Ray. The Conservation Response by Lloyd J. Dumas (review). Technology and Culture, 18(3):578–580, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891984/pdf.

Reader:1978:LFH

[Rea78] W. J. Reader. Lucas: The First Hundred Years. Vol. 1: The King of the Road by Harold Nockolds (review). Technology and Culture, 19(2):230–232, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892086/pdf.

Reingold:1970:ADB

[Rei70] Nathan Reingold. Alexander Dallas Bache: Science and technology in the American idiom. Technology and Culture, 11(2):163–177, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894147/pdf.

Reingold:1971:DRS

[Rei71] Nathan Reingold. Dollars for Research: Science and Its Patrons in Nineteenth-Century America by Howard S. Miller (review). Technology and Culture, 12(4):668–669, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893580/pdf.

Reti:1970:DAP

[Ret70] Ladislao Reti. The double-acting principle in East and West. Technology and Culture, 11(2):178–200, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/894148/pdf.

Reti:1972:LR

[Ret72] Ladislao Reti. Leonardo and ramelli. Technology and Culture, 13 (4):577–605, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893378/pdf.

Reynolds:1974:SSB

[Rey74] Clark G. Reynolds. Steam and Sail in Britain and North America by P. W. Brock, Basil Greenhill (review). Technology and Culture, 15(4):644–645, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892715/pdf.

Reynolds:1975:NNR

[Rey75] Clark G. Reynolds. The Nuclear Navy, 1946–1962 by Richard G. Hewlett, Francis Duncan (review). Technology and Culture, 16 (4):662–665, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891622/pdf.

Reynolds:1979:SWL

[Rey79a] Clark G. Reynolds. Supplying War: Logistics from Wallenstein to Patton by Martin van Creveld (review). Technology and Culture, 20(1):198–199, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890460/pdf.

Reynolds:1979:SIT

[Rey79b] Terry S. Reynolds. Scientific influences on technology: The case of the overshot waterwheel, 1752–1754. Technology and Culture, 20(2):270–295, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890492/pdf.

Rezneck:1970:EEA

[Rez70] Samuel Rezneck. The European education of an American chemist and its influence in 19th-Century America: Eben Norton Horsford. Technology and Culture, 11(3):366–388, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894101/pdf.

Ruttan:1973:TTA

[RH73] V. W. Ruttan and Yujiro Hayami. Technology transfer and agricultural development. Technology and Culture, 14(2 (part I)):119– 151, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893257/pdf.

Rae:1973:WMD

[RHTH73] John B. Rae, Richard S. Hartenberg, Arnold Thackray, and Thomas Parke Hughes. The Washington Meeting, December 27–29, 1972. Technology and Culture, 14(3):429–460, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893215/pdf.

Riepe:1970:CPM

[Rie70] Dale Riepe. Changing Perspectives on Man ed. by Ben Rothblatt (review). Technology and Culture, 11(3):466–469, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894126/pdf.

Risse:1978:SHM

[Ris78] Guenter B. Risse. A Social History of Medicine by Frederick F. Cartwright (review). Technology and Culture, 19(4):764–765, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892197/pdf.

Reingold:1976:I

[RM76] Nathan Reingold and Arthur Molella. Introduction. Technology and Culture, 17(4):624–633, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891814/pdf.

Robinson:1970:VEC

[Rob70] Eric Robinson. Victorian Engineering by L. T. C. Rolt (review). Technology and Culture, 11(4):631–632, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894058/pdf.

Robinson:1971:CMV

[Rob71] Eric Robinson. Capital et machine à vapeur au XVIII^{e} siècle: Les frères Perier et l’introduction en France de la machine à vapeur de Watt by J. Payen (review). Technology and Culture, 12(4):635–640, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893565/pdf.

Robinson:1972:JWL

[Rob72] Eric Robinson. James Watt and the law of patents. Technology and Culture, 13(2):115–139, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893462/pdf.

Robinson:1975:KHF

[Rob75a] Eric Robinson. Kenricks in Hardware, a Family Business: 1791– 1966 by R. A. Church (review). Technology and Culture, 16(3):489– 490, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891580/pdf.

Robinson:1975:WCW

[Rob75b] Eric H. Robinson. The Worshipful Company of Wheelwrights of the City of London, 1670–1970 by Eric Bennett (review). Technology and Culture, 16(1):94, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891508/pdf.

Robinson:1976:DWS

[Rob76a] Eric Robinson. The Development of the West of Scotland, 1750–1960 by Anthony Slaven (review). Technology and Culture, 17(4):800–802, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891852/pdf.

Robinson:1976:PIS

[Rob76b] Eric Robinson. The Paper Industry in Scotland: 1590–1861 by A. G. Thomson (review). Technology and Culture, 17(1):125–127, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891663/pdf.

Robinson:1976:AIW

[Rob76c] Eric H. Robinson. America’s Impact on the World by William Woodruff (review). Technology and Culture, 17(3):564–566, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891786/
pdf.

Robinson:1976:TBZ

[Rob76d] Eric H. Robinson. Technologische Beziehungen zwischen Deutschland und England von der Mitte 17. bis zum Ausgang des 18. Jahrhunderts by Hans-Joachim Braun (review). Technology and Culture, 17(3):554–555, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891779/pdf.

Rogers:1975:WSH

[Rog75] Jean Li Rogers. Woman in Science by H. J. Mozans, and: Women in Science and Technology: a Report on an M.I.T. Workshop ed. by Edith Ruina (review). Technology and Culture, 16(2):324–328, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891554/pdf.

Rogers:1977:SFS

[Rog77] Jean Li Rogers. The Social Function of Social Science by Duncan MacRae, Jr. (review). Technology and Culture, 18(3):588–590, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891988/pdf.

Rogers:1978:TDC

[Rog78] Everett M. Rogers. Technological Diffusion and the Computer Revolution: The UK Experience by Paul Stoneman (review). Technology and Culture, 19(1):140–142, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892063/pdf.

Rolt:1970:IAD

[Rol70a] L. T. C. Rolt. Industrial Archaeology of Derbyshire by Frank Nixon (review). Technology and Culture, 11(3):475–477, July 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894130/pdf.

Rolt:1970:MMF

[Rol70b] L. T. C. Rolt. Mines, Mills and Furnaces: Industrial Archaeology in Wales by D. Morgan Rees, and: The Industrial Archaeology of the Bristol Region by R. A. Buchanan, Neil Cossons, and: The Industrial Archaeology of Bath by R. A. Buchanan (review). Technology and Culture, 11(4):633–635, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894059/pdf.

Rolt:1972:BDH

[Rol72] L. T. C. Rolt. The Birth and Death of a Highland Railway by Duncan Kennedy (review). Technology and Culture, 13(1):77–78, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893441/pdf.

Rolt:1973:BHF

[Rol73] L. T. C. Rolt. Bicycling, a History by Frederick Alderson (review). Technology and Culture, 14(4):644–645, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892889/pdf.

Rolt:1974:GSE

[Rol74] L. T. C. Rolt. George Stephenson: The Engineer and His Letters by W. O. Skeat (review). Technology and Culture, 15(4):639–641, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892713/pdf.

Roland:1977:AJF

[Rol77a] Alex Roland. The Autobiography of John Fitch ed. by Frank D. Prager (review). Technology and Culture, 18(3):538–540, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891961/pdf.

Roland:1977:BSA

[Rol77b] Alex Roland. Bushnell’s submarine: American original or European import? Technology and Culture, 18(2):157–174, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891908/pdf.

Roland:1977:PD

[Rol77c] Alex Roland. Problems of definition. Technology and Culture, 18(3):509–511, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891948/pdf.

Roland:1978:HBM

[Rol78] Alex Roland. How the Battleship “Maine” Was Destroyed by H. G. Rickover (review). Technology and Culture, 19(2):222–224, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892082/pdf.

Ron:1978:SME

[Ron78] Moshe Ron. The Sidney M. Edelstein Collection in the History of Chemistry and Chemical Technology. Technology and Culture, 19(3):491–494, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892117/pdf.

Ropp:1972:GSB

[Rop72] Theodore Ropp. A Guide to the Sources of British Military History ed. by Robin Higham (review). Technology and Culture, 13 (4):691–692, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893417/pdf.

Ropp:1976:GSU

[Rop76a] Theodore Ropp. A Guide to the Sources of United States Military History ed. by Robin Higham (review). Technology and Culture, 17(3):613, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891810/pdf.

Ropp:1976:HTC

[Rop76b] Theodore Ropp. A History of Tactical Communication Techniques by David L. Woods (review). Technology and Culture, 17(1):151, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891677/pdf.

Ropp:1976:RRS

[Rop76c] Theodore Ropp. Railroads and Rifles: Soldiers, Technology, and the Unification of Germany by Dennis E. Showalter (review). Technology and Culture, 17(3):555–557, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891780/pdf.

Ropp:1976:DHP

[Rop76d] Theodore Ropp. The District: a History of the Philadelphia District U.S. Army Corps of Engineers, 1866–1971 by Frank E. Snyder, Brian H. Guss (review). Technology and Culture, 17(1):150–151, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891676/pdf.

Rosenberg:1970:EDT

[Ros70] Nathan Rosenberg. Economic development and the transfer of technology: Some historical perspectives. Technology and Culture, 11 (4):550–575, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894043/pdf.

Rosenberg:1971:SCA

[Ros71a] Nathan Rosenberg. Structural Change in the American Economy by Anne P. Carter (review). Technology and Culture, 12(3):511– 513, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893609/pdf.

Rosenberg:1971:TEE

[Ros71b] Nathan Rosenberg. Technology and the environment: an economic exploration. Technology and Culture, 12(4):543–561, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893555/pdf.

Rossiter:1972:AIN

[Ros72] Margaret W. Rossiter. The American Ideology of National Science, 1919–1930 by Ronald C. Tobey (review). Technology and Culture, 13(4):672–673, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893406/pdf.

Rosegger:1973:TCR

[Ros73a] Gerhard Rosegger. Technological Change in Regulated Industries ed. by William M. Capron (review). Technology and Culture, 14(2 (part I)):329–330, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893284/pdf.

Rossini:1973:RGS

[Ros73b] Frederick A. Rossini. The Relevance of General Systems Theory ed. by Ervin Laszlo (review). Technology and Culture, 14(4):674–676, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892906/pdf.

Rosen:1974:PLC

[Ros74] Edward Rosen. Poland, the Land of Copernicus ed. by Bogdan Suchodolski (review). Technology and Culture, 15(1):94–95, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892834/pdf.

Rosegger:1975:SPB

[Ros75a] Gerhard Rosegger. Science Policy and Business: The Changing Relation of Europe and, the United States ed. by David W. Ewing (review). Technology and Culture, 16(4):681–682, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891631/pdf.

Rossini:1975:SFE

[Ros75b] Frederick A. Rossini. A Strategy for the Future by Erwin Laszlo (review). Technology and Culture, 16(4):699–700, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891640/pdf.

Rossini:1975:GKS

[Ros75c] Frederick A. Rossini. Gabriel Kron and Systems Theory ed. by H. H. Happ (review). Technology and Culture, 16(1):117–118, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891521/pdf.

Rossini:1975:TIR

[Ros75d] Frederick A. Rossini. Technology as Institutionally Related to Human Values ed. by Philip C. Ritterbush (review). Technology and Culture, 16(2):331–333, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891557/pdf.

Rosenbloom:1976:TCE

[Ros76a] Richard S. Rosenbloom. Technological Change: Economics, Management and Environment ed. by Bela Gold (review). Technology and Culture, 17(4):782–785, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891842/pdf.

Rossiter:1976:FAS

[Ros76b] Margaret W. Rossiter. The Formation of the American Scientific Community: The American Association for the Advancement of Science, 1848–1860 by Sally Gregory Kohlstedt (review). Technology and Culture, 17(4):758–761, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891830/pdf.

Rossiter:1977:CFC

[Ros77] Margaret W. Rossiter. The Charles F. Chandler Collection. Technology and Culture, 18(2):222–230, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891912/pdf.

Rose:1979:SCT

[Ros79a] Mark H. Rose. Small Cities in Transition: The Dynamics of Growth and Decline ed. by Herrington J. Bryce (review). Technology and Culture, 20(3):665–667, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890555/pdf.

Rosenberg:1979:TIA

[Ros79b] Nathan Rosenberg. Technological interdependence in the American economy. Technology and Culture, 20(1):25–50, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890451/pdf.

Rossini:1979:TFH

[Ros79c] Frederick A. Rossini. Technology’s Future: The Hague Congress on Technology Assessment by Thomas J. Knight (review). Technology and Culture, 20(2):397–399, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890519/pdf.

Rossiter:1979:WMS

[Ros79d] Margaret W. Rossiter. Women and Mathematics, Science and Engineering: a Partially Annotated Bibliography with Emphasis on Mathematics and with References on Related Topics by Else Høyrup (review). Technology and Culture, 20(3):633–634, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890539/pdf.

Rotsch:1972:DLD

[Rot72] Melvin M. Rotsch. Design on the Land: The Development of Landscape Architecture by Norman T. Newton (review). Technology and Culture, 13(4):653–656, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893397/pdf.

Rouse:1971:RJB

[Rou71] Hunter Rouse. The Reminiscences of John B. Jervis, Engineer of the Old Croton ed. by Neal FitzSimons (review). Technology and Culture, 12(4):649–650, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893571/pdf.

Rouse:1972:HPF

[Rou72] Hunter Rouse. A History and Philosophy of Fluidmechanics by G. A. Tokaty (review). Technology and Culture, 13(3):498–499, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893934/pdf.

Rouse:1975:RKM

[Rou75] Hunter Rouse. The River Killers by Martin Heuvelmans (review). Technology and Culture, 16(1):103–104, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891513/pdf.

Rowlett:1972:EAE

[Row72a] Ralph M. Rowlett. Explanation in Archaeology: an Explicitly Scientific Approach by Patty Jo Watson, et al. (review). Technology and Culture, 13(4):688–690, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893416/pdf.

Rowlett:1972:BAC

[Row72b] Ralph M. Rowlett. The Bronze Age Cultures in Central and Eastern Europe by Marija Gimbutas (review). Technology and Culture, 13(2):299–301, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893468/pdf.

Rowlett:1973:OHS

[Row73] Ralph M. Rowlett. The Origin of Homo sapiens/Origine de l’homme moderne ed. by F. Bordes (review). Technology and Culture, 14 (4):614–615, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892872/pdf.

Rowlett:1976:PEH

[Row76a] Ralph M. Rowlett. Palaeoeconomy ed. by E. S. Higgs (review). Technology and Culture, 17(1):110–112, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891656/pdf.

Rowlett:1976:TCW

[Row76b] Ralph M. Rowlett. The Threshold of Civilization by Walter A. Fairservis (review). Technology and Culture, 17(2):365–368, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891721/
pdf.

Rowe:1977:ATP

[Row77] John Howland Rowe. Archaeological Theory and Practice ed. by D. E. Strong (review). Technology and Culture, 18(2):237–238, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891914/pdf.

Rowlett:1979:ESS

[Row79] Ralph M. Rowlett. Evolution of Social Systems ed. by J. Friedman, M. J. Rowlands (review). Technology and Culture, 20(3):620– 622, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890532/pdf.

Rurup:1974:HMT

[Rür74] Reinhard Rürup. Historians and modern technology: Reflections on the development and current problems of the history of technology. Technology and Culture, 15(2):161–193, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892790/pdf.

Schallenberg:1977:RMS

[SA77] Richard H. Schallenberg and David A. Ault. Raw materials supply and technological change in the American charcoal iron industry. Technology and Culture, 18(3):436–466, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891939/pdf.

Saloutos:1970:PMN

[Sal70] Theodore Saloutos. Populism: Its Meaning and National Characteristics by Ghita Ionescu, Ernest Gellner (review). Technology and Culture, 11(2):327–329, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894176/pdf.

Saloutos:1971:CAA

[Sal71] Theodore Saloutos. The Conquest of Arid America by William E. Smythe (review). Technology and Culture, 12(3):514–515, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893610/pdf.

Salsbury:1973:HSO

[Sal73] Stephen Salsbury. History of Standard Oil Company (New Jersey), 1927–1950: New Horizons by Henrietta M. Larson, et al. (review). Technology and Culture, 14(4):660–663, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892899/pdf.

Salsbury:1977:ACJ

[Sal77] Stephen Salsbury. The Age of Capital, 1848–1875 by E. J. Hobsbawm (review). Technology and Culture, 18(2):256–257, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891925/pdf.

Sandler:1970:EMC

[San70] Stanley Sandler. The emergence of the modern capital ship. Technology and Culture, 11(4):576–595, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/894044/pdf.

Sanders:1972:MAG

[San72] Ralph Sanders. Man and Atom by Glenn T. Seaborg, William R. Corliss (review). Technology and Culture, 13(4):680–682, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893411/pdf.

Sande:1973:IAB

[San73a] Ted Sande. Industrial Archaeology in Britain by R. A. Buchanan (review). Technology and Culture, 14(4):618–620, October 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892874/pdf.

Sande:1973:NPS

[San73b] Ted Sande. The National Park Service and the history of technology: The New England Textile Mill Survey. Technology and Culture, 14(3):404–414, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893210/pdf.

Sanford:1973:DIS

[San73c] Charles L. Sanford. Designing the. Industrial State: The Intellectual Pursuit of Collectivisim in America, 1880–1940 by James Gilbert, and: Good Tidings: The Belief in Progress from Darwin to Marcuse by W. Warren Wagar (review). Technology and Culture, 14(2 (part I)):333–337, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893286/pdf.

Sanders:1974:ASH

[San74] Ralph Sanders. Atomic Shield: a History of the United States Atomic Energy Commission, Vol. 2 (1947–52) by Richard G. Hewlett, Francis Duncan (review). Technology and Culture, 15(3):506–508, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892757/pdf.

Sanders:1977:AOT

[San77] Ralph Sanders. The Advisors: Oppenheimer, Teller and the Superbomb by Herbert York (review). Technology and Culture, 18 (1):129–131, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891895/pdf.

Sanders:1978:NPN

[San78] Ralph Sanders. Nuclear Proliferation and the Near-Nuclear Countries ed. by Onkar Marwah, Ann Schulz (review). Technology and Culture, 19(2):246–247, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892094/pdf.

Saunders:1972:PPS

[Sau72] Richard Saunders. Paternalism and Protest: Southern Cotton Mill Workers and Organized Labor, 1875–1905 by Melton Alonza McLaurin (review). Technology and Culture, 13(3):505–507, July 1972. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893940/pdf.

Stock:1972:RBA

[SB72] John T. Stock and D. J. Bryden. A Robinson balance by adie & son of Edinburgh. Technology and Culture, 13(1):44–54, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893430/pdf.

Schmandt-Besserat:1978:ETA

[SB78] Denise Schmandt-Besserat. Early technologies: Austin, Texas, October 1976. Technology and Culture, 19(1):104–105, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892044/pdf.

Scamehorn:1972:GLC

[Sca72] Lee Scamehorn. The Giant Leap: a Chronology of Ohio Aerospace Events and Personalities, 1815–1969 by Tom D. Crouch (review). Technology and Culture, 13(3):511–512, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893943/pdf.

Schillinger:1970:FFJ

[Sch70] A. George Schillinger. The Future of the Future by John McHale (review). Technology and Culture, 11(3):473–475, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894129/pdf.

Schioeler:1971:DSB

[Sch71a] Thorkild Schioeler. Dynamische und statische Betrachtungsweise in der antiken Mechanik by Fritz Krafft (review). Technology and Culture, 12(4):630–632, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893563/pdf.

Schmandt:1971:CFP

[Sch71b] Jurgen Schmandt. The Children of Frankenstein: a Primer on Modern Technology and Human Values by Herbert J. Muller (review). Technology and Culture, 12(1):137–139, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894029/pdf.

Schindler:1974:PPH

[Sch74] Charles R. Schindler. Printing Presses: History and Development from the Fifteenth Century to Modern Times by James Moran (review). Technology and Culture, 15(1):92–94, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892833/pdf.

Schofield:1976:REM

[Sch76] Robert E. Schofield. Reason, Experiment, and Mysticism in the Scientific Revolution ed. by M. L. Righini Bonelli, William R. Shea (review). Technology and Culture, 17(4):793–795, October 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891848/pdf.

Schallenberg:1977:TEP

[Sch77a] Richard H. Schallenberg. Thomas Edison: Professional Inventor by Thomas P. Hughes (review). Technology and Culture, 18(2):255, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891924/pdf.

Schofield:1977:PCS

[Sch77b] Robert E. Schofield. Philosophical Chemistry in the Scottish Enlightenment: The Doctrines and Discoveries of William Cullen and Joseph Black by A. L. Donovan (review). Technology and Culture, 18(3):528–529, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891956/pdf.

Schafer:1978:MIJ

[Sch78a] William J. Schafer. The Musical Instruments of Joseph Haydn: an Introduction by Helen Rice Hollis (review). Technology and Culture, 19(3):534–536, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892132/pdf.

Schallenberg:1978:ISA

[Sch78b] Richard H. Schallenberg. Iron and Steel in America by W. David Lewis (review). Technology and Culture, 19(1):118–119, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/892051/
pdf.

Schallenberg:1978:LJF

[Sch78c] Richard H. Schallenberg. The Lonaconing Journals: The Founding of a Coal and Iron Community, 1837–1840 ed. by Katherine A. Harvey (review). Technology and Culture, 19(3):558–559, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892145/pdf.

Seeger:1978:BFE

[See78] Raymond Seeger. Benjamin Franklin, Electrician by Bern Dibner (review). Technology and Culture, 19(2):216–217, April 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892078/pdf.

Segal:1977:CWC

[Seg77a] Howard P. Segal. Chicago’s White City of 1893 by David F. Burg, and: Civilizing the Machine: Technology and Republican Values in America, 1776–1900 by John F. Kasson (review). Technology and Culture, 18(2):250–253, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891922/pdf.

Segal:1977:TAD

[Seg77b] Howard P. Segal. Technocracy and the American Dream: The Technocrat Movement, 1900–1941 by William E. Akin (review). Technology and Culture, 18(4):714–715, October 1977. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892014/pdf.

Segal:1977:TPS

[Seg77c] Howard P. Segal. Toward a Planned Society: From Roosevelt to Nixon by Otis L. Graham, Jr. (review). Technology and Culture, 18(1):127–128, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891894/pdf.

Segal:1978:WSA

[Seg78] Howard P. Segal. Work without Salvation: America’s Intellectuals and Industrial Alienation, 1880–1910 by James B. Gilbert (review). Technology and Culture, 19(4):753–756, October 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892192/pdf.

Segal:1979:UUT

[Seg79a] Howard P. Segal. Urban Utopias in the Twentieth Century: Ebenezer Howard, Frank Lloyd Wright, and Le Corbusier by Robert Fishman (review). Technology and Culture, 20(4):815–817, October 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890588/pdf.

Segal:1979:UIT

[Seg79b] Howard P. Segal. Utopia by Ian Tod, Michael Wheeler (review). Technology and Culture, 20(3):667–668, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890556/pdf.

Selk:1977:THC

[Sel77] Eugene E. Selk. Technology and the Human Condition by Bernard Gendron (review). Technology and Culture, 18(3):518–520, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891951/
pdf.

Seller:1978:PPT

[Sel78] Anne E. Seller. Progress and Its Problems: Towards a Theory of Scientific Growth by Larry Laudan (review). Technology and Culture, 19(4):791–793, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892211/pdf.

Sewell:1972:SBT

[Sew72] Granville H. Sewell. The Struggle to Bring Technology to Cities by the Urban Institute (review). Technology and Culture, 13(4):687–688, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893415/pdf.

Sewell:1975:PPC

[Sew75] Granville H. Sewell. People or Penguins: The Case for Optimal Pollution by William F. Baxter (review). Technology and Culture, 16(4):687–689, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891635/pdf.

Schroeder-Gudehus:1975:SOS

[SG75] Brigitte Schroeder-Gudehus. Science Observed: Science as a Social and Intellectual Activity by F. R. Jevons (review). Technology and Culture, 16(1):120–122, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891523/pdf.

Sharrer:1971:IBS

[Sha71] G. Terry Sharrer. The indigo bonanza in South Carolina, 1740– 90. Technology and Culture, 12(3):447–455, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893594/pdf.

Sharlin:1972:HEM

[Sha72a] Harold I. Sharlin. A History of Electricity and Magnetism by Herbert W. Meyer (review). Technology and Culture, 13(4):630–631, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893384/pdf.

Sharpe:1972:HNW

[Sha72b] Mitchell R. Sharpe. History of the Naval Weapons Center, China Lake, California. Vol. 1: Sailors, Scientists, and Rockets: Origins of the Navy Rocket Program and the Naval Ordnance Test Station, Inyokern by Albert B. Christman (review). Technology and Culture, 13(4):677–678, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893409/pdf.

Sharrer:1972:DMH

[Sha72c] G. Terry Sharrer. The Depletion Myth: a History of Railroad Use of Timber by Sherry H. Olson (review). Technology and Culture, 13(2): 319, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893479/pdf.

Sharpe:1973:IIH

[Sha73] Mitchell R. Sharpe. Ignition! An Informal History of Liquid Rocket Propellants by John D. Clark (review). Technology and Culture, 14(3):514–515, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893236/pdf.

Sharlin:1974:BAG

[Sha74] Harold Issadore Sharlin. Bell: Alexander Graham Bell and the Conquest of Solitude by Robert V. Bruce (review). Technology and Culture, 15(2):334–336, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892801/pdf.

Sharpe:1975:MPH

[Sha75] Mitchell R. Sharpe. Mr. Piper and His Cubs by Devon Francis (review). Technology and Culture, 16(4):665–666, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891623/pdf.

Shapin:1976:AGE

[Sha76a] Steven Shapin. Artisan to Graduate: Essays to Commemorate the Foundation in 1824 of the Manchester Mechanics’ Institution, Now in 1974 the University of Manchester Institute of Science and Technology ed. by D. S. L. Cardwell (review). Technology and Culture, 17(1):143–145, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891672/pdf.

Sharlin:1976:WTR

[Sha76b] Harold Issadore Sharlin. Wireless Telegraphy: The Royal Institution Library of Science (Being the Friday Evening Discourses in Physical Sciences Held at the Royal Institution: 1851–1939) ed. by Sir Eric Eastwood (review). Technology and Culture, 17(1):145–148, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891673/pdf.

Sharpe:1976:MAT

[Sha76c] Mitchell R. Sharpe. Men against Tanks: a History of Anti-Tank Warfare by John Weeks (review). Technology and Culture, 17 (3):558–559, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891781/pdf.

Sharpe:1978:EES

[Sha78a] Mitchell R. Sharpe. The End of an Era in Space Exploration: From International Rivalry to International Cooperation by B. C. D. Blaine, and: Two Hundred Years of Flight in America, a Bicentennial Survey ed. by Eugene M. Emme, and: The Eagle Has Returned. Pt. 1 ed. by Dr. Ernst A. Steinhoff, and: The Eagle Has Returned. Pt. 2 ed. by Dr. Ernst A. Steinhoff (review). Technology and Culture, 19(3):575–580, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892153/pdf.

Sharpe:1978:OIE

[Sha78b] Mitchell R. Sharpe. The Origins and International Economics of Space Exploration by Sir Bernard Lovell (review). Technology and Culture, 19(2):249–251, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892096/pdf.

Sharpe:1978:RRG

[Sha78c] Mitchell R. Sharpe. The Rockets’ Red Glare: an Illustrated History of Rocketry through the Ages by Wernher von Braun, Frederick I. Ordway III (review). Technology and Culture, 19(2):247–248, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892095/pdf.

Sharpe:1979:STH

[Sha79] Mitchell R. Sharpe. On the Shoulders of Titans: a History of Project Gemini by Barton C. Hacker, James M. Grimwood (review). Technology and Culture, 20(1):238–240, January 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890483/pdf.

Sherwood:1974:HMC

[She74] Morgan Sherwood. Hiram Martin Chittenden: His Public Career by Gordon B. Dodds (review). Technology and Culture, 15(2):341– 342, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892805/pdf.

Shelby:1975:MTH

[She75a] Lon R. Shelby. Mariano taccola and his books on engines and machines. Technology and Culture, 16(3):466–475, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891572/pdf.

Sherwood:1975:INM

[She75b] Morgan Sherwood. Inventions Necessity Is Not the Mother Of: Patents Ridiculous and Sublime by Stacy V. Jones (review). Technology and Culture, 16(2):333–334, April 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891558/pdf.

Sherwood:1977:NII

[She77] Morgan Sherwood. New and Improved: Inventors and Inventions That Have Changed the Modern World by R. Baker (review). Technology and Culture, 18(4):697–698, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892004/pdf.

Shelby:1978:MFB

[She78a] Lon R. Shelby. Medieval French Bridges: a History by Marjorie Nice Boyer (review). Technology and Culture, 19(2):212–214, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892075/pdf.

Sherwood:1978:RCA

[She78b] Morgan Sherwood. Railroad in the Clouds: The Alaska Railroad in the Age of Steam, 1914–1945 by William H. Wilson (review). Technology and Culture, 19(4):744–746, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892186/pdf.

Shriver:1972:MHM

[Shr72] Donald W. Shriver, Jr. Man and his machines: Four angles of vision. Technology and Culture, 13(4):531–555, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893376/pdf.

Shriver:1973:CRH

[Shr73] Donald W. Shriver, Jr. Counterrevolution and Revolt by Herbert Marcuse (review). Technology and Culture, 14(2 (part I)):337–340, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893287/pdf.

Simonson:1971:TMS

[Sim71] G. R. Simonson. Technology and Market Structure: a Study of the Aircraft Industry by Almarin Phillips (review). Technology and Culture, 12(4):664–665, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893578/pdf.

Simms:1977:ABM

[Sim77] D. L. Simms. Archimedes and the burning mirrors of Syracuse. Technology and Culture, 18(1):1–24, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891861/pdf.

Sinclair:1970:CEJ

[Sin70a] Bruce Sinclair. Charles Ellet, Jr.: The Engineer as Individualist, 1810–1862 by Gene D. Lewis (review). Technology and Culture, 11(3):446–447, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894115/pdf.

Sinclair:1970:PAH

[Sin70b] Bruce Sinclair. The Papers of Alexander Hamilton ed. by Harold C. Syrett, Jacob E. Cooke (review). Technology and Culture, 11(3):444– 445, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894114/pdf.

Sinclair:1972:ASM

[Sin72] Bruce Sinclair. The American System of Manufactures: The Report of the Committee on Machinery of the United States, 1855, and the Special Reports of George Wallis and Joseph Whitworth, 1854 ed. by Nathan Rosenberg (review). Technology and Culture, 13(1):87–91, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893447/pdf.

Sinclair:1977:CPE

[Sin77] George Sinclair. A call for a philosophy of engineering. Technology and Culture, 18(4):685–689, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891998/pdf.

Sinclair:1979:CTB

[Sin79a] Bruce Sinclair. Canadian technology: British traditions and American influences. Technology and Culture, 20(1):108–123, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890455/
pdf.

Singer:1979:EGD

[Sin79b] S. Fred Singer. The Economic Growth Debate: an Assessment by E. J. Mishan (review). Technology and Culture, 20(3):663–665, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890554/pdf.

Sivin:1972:CWI

[Siv72] N. Sivin. China at Work: an Illustrated Record of the Primitive Industries of China’s Masses, Whose Life Is Toil, and Thus an Account of Chinese Civilization by Rudolf P. Hommel (review). Technology and Culture, 13(1):69–71, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893436/pdf.

Sivin:1975:STD

[Siv75] N. Sivin. Science and Technology in the Development of Modern China: an Annotated Bibliography ed. by Genevieve C. Dean, and: Bibliography on Technology and Social Change in China and Japan ed. by Roberta Abraham (review). Technology and Culture, 16 (4):707–710, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891646/pdf.

Sivin:1977:SRO

[Siv77] N. Sivin. Seiyû roku: On Oil Manufacturing by Ôkura Nagatsune (review). Technology and Culture, 18(3):524–526, July 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891954/pdf.

Skolimowski:1975:AEP

[Sko75] Henryk Skolimowski. Alfred Espinos, precurseur de la praxeologie, ses antecedents et ses successeurs by Jan Ostrowski, et al. (review). Technology and Culture, 16(4):682–683, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891632/pdf.

Skolnikoff:1976:TIP

[Sko76] Eugene B. Skolnikoff. Technology and International Politics: The Crisis of Wishing by William R. Kintner, Harvey Sicherman (review). Technology and Culture, 17(3):562–564, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891785/pdf.

Skydsgaard:1970:MWA

[Sky70] Jens Erik Skydsgaard. The Muses at Work: Arts, Crafts, and Professions in Ancient Greece and Rome ed. by Carl Roebuck (review). Technology and Culture, 11(4):617–620, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894051/pdf.

Skydsgaard:1973:RLS

[Sky73] Jens Erik Skydsgaard. The Roman Land-Surveyors: an Introduction to the Agrimensores by O. A. W. Dilke (review). Technology and Culture, 14(1):81–82, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893356/pdf.

Skydsgaard:1975:AEM

[Sky75] Jens Erik Skydsgaard. The Ancient Economy by M. I. Finley (review). Technology and Culture, 16(4):617–619, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891602/pdf.

Sleeswyk:1978:CRR

[Sle78] Andre Wegener Sleeswyk. The celestial river: a reconstruction. Technology and Culture, 19(3):423–449, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892111/pdf.

Sloan:1970:NAH

[Slo70] James J. Sloan. Northeast Aero Historians: Rhinebeck, New York, October 1969. Technology and Culture, 11(3):405–406, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894104/pdf.

Sloan:1972:CSW

[Slo72] Edward W. Sloan III. Confederate Shipbuilding by William N. Still, Jr. (review). Technology and Culture, 13(3):502, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893937/pdf.

Smallwood:1978:CCH

[Sma78] J. B. Smallwood, Jr. Custodians of the Coast: History of the United States Army Engineers at Galveston by Lynn M. Alperin (review). Technology and Culture, 19(4):741–743, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892184/pdf.

Smith:1970:ATS

[Smi70a] Cyril Stanley Smith. Art, technology, and science: Notes on their historical interaction. Technology and Culture, 11(4):493–549, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894042/pdf.

Smith:1970:RDS

[Smi70b] Norman A. F. Smith. The Roman dams of Subiaco. Technology and Culture, 11(1):58–68, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894190/pdf.

Smith:1970:MCH

[Smi70c] Peter H. Smith. Material Culture: Historical Agencies and the Historian ed. by Lucius F. Ellsworth, Maureen A. O’Brien (review). Technology and Culture, 11(4):661–664, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894074/pdf.

Smith:1971:PTA

[Smi71a] Cyril Stanley Smith. Philosophy, Technology, and the Arts in the Early Modern Era by Paolo Rossi (review). Technology and Culture, 12(2):344–345, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893971/pdf.

Smith:1971:CCS

[Smi71b] Peter H. Smith. Country Cabinetwork and Simple City Furniture ed. by John D. Morse (review). Technology and Culture, 12(1):103–104, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894014/pdf.

Smith:1972:ASH

[Smi72a] Cyril Stanley Smith. Architectural shapes of hot-rolled iron, 1753. Technology and Culture, 13(1):59–65, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893432/pdf.

Smith:1972:SAS

[Smi72b] Merritt Roe Smith. Science in American Society: a Social History by George H. Daniels (review). Technology and Culture, 13(4):666–669, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893403/pdf.

Smith:1972:CJI

[Smi72c] Merritt Roe Smith. The Contributions of Joseph Ives to Connecticut Clock Technology, 1810–1862 by Kenneth D. Roberts (review). Technology and Culture, 13(1):91–93, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893448/pdf.

Smith:1973:JHH

[Smi73a] Merritt Roe Smith. John H. Hall, Simeon North, and the milling machine: The nature of innovation among Antebellum arms makers. Technology and Culture, 14(4):573–591, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892869/pdf.

Smith:1973:HDE

[Smi73b] Cecil O. Smith, Jr. History and Development of Engineering by Malcolm S. Gregory (review). Technology and Culture, 14(2 (part I)):288–289, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893264/pdf.

Smith:1974:APM

[Smi74a] Merritt Roe Smith. The American Precision Museum by Robert R. Rhodehamel, Edwin A. Battison (review). Technology and Culture, 15(3):413–437, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892737/pdf.

Smith:1974:HRP

[Smi74b] Merritt Roe Smith. Hall’s rifle patent “Number”. Technology and Culture, 15(3):540, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892774/pdf.

Smith:1974:TSE

[Smi74c] Merritt Roe Smith. Toward a standard of excellence: The second installment. Technology and Culture, 15(1):76–79, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892827/pdf.

Smith:1975:MLI

[Smi75a] Cyril Stanley Smith. Mining Lore: an Illustrated Composition and Documentary Compilation with Emphasis on the Spirit and History of Mining by Wolfgang Paul (review). Technology and Culture, 16(1):85–86, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891504/pdf.

Smith:1975:TSC

[Smi75b] Merritt Roe Smith. Technology and Social Change in America ed. by Edwin T. Layton, Jr. (review). Technology and Culture, 16 (1):110–112, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891518/pdf.

Smith:1976:EIR

[Smi76a] Cecil O. Smith. Education and the Industrial Revolution by E. G. West (review). Technology and Culture, 17(2):377–378, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891727/pdf.

Smith:1976:ECG

[Smi76b] Cyril Stanley Smith. Eighteenth Century Gunfounding. The Verbruggens at the Royal Brass Foundry. A Chapter in the History of Technology by Melvin J. Jackson, Carel de Beer (review). Technology and Culture, 17(1):136–138, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891669/pdf.

Smith:1976:SDA

[Smi76c] Cyril Stanley Smith. O stali Damasceńskiej [About Damascus steel] by Jerzy Piaskowski (review). Technology and Culture, 17(2):371– 373, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891724/pdf.

Smith:1976:CCI

[Smi76d] Thomas M. Smith. The Concorde Conspiracy: The International Race for the SST by John Costello, Terry Hughes, and: The Superwarriors by James Canan (review). Technology and Culture, 17 (4):771–773, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891837/pdf.

Smith:1976:PWU

[Smi76e] Thomas M. Smith. Project Whirlwind: an unorthodox development project. Technology and Culture, 17(3):447–464, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891752/pdf.

Smith:1977:MRA

[Smi77a] Cyril Stanley Smith. Metallurgical Remains of Ancient China by Noel Bernard, Tamotsu Satō, and: The Cradle of the East: an Enquiry into the Indigenous Origins of Techniques and Ideas of Neolithic and Early Historic China, 5000–1000 B.C. by Ping-Ti Ho (review). Technology and Culture, 18(1):80–86, January 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891868/pdf.

Smith:1977:PAN

[Smi77b] David C. Smith. Papermaking in America by Norman B. Wilkinson (review). Technology and Culture, 18(3):540–541, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891962/pdf.

Smith:1977:MSU

[Smi77c] Merritt Roe Smith. Museums: In Search of a Usable Future by Alma S. Wittlin (review). Technology and Culture, 18(1):148–150, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891904/pdf.

Smith:1977:DWW

[Smi77d] Wilfred J. Smith. The Dragon’s Wings by William M. Leary, Jr. (review). Technology and Culture, 18(1):120–122, January 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891890/pdf.

Sokal:1975:IPD

[Sok75] Michael M. Sokal. The Intelligence of a People by Daniel Calhoun (review). Technology and Culture, 16(3):510–511, July 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891590/pdf.

Solo:1976:SDJ

[Sol76a] Robert Solo. Science for Development by Jacques Spaey, et al., and: Science Development by Michael J. Moravcsik (review). Technology and Culture, 17(3):566–568, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891787/pdf.

Solo:1976:EEG

[Sol76b] Robert Solo. The Economics of Efficiency and Growth: Lessons from Israel and the West Bank by Abba Lerner, Hiam Ben-Shahar (review). Technology and Culture, 17(1):183–186, January 1976. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891695/pdf.

Solo:1976:TTD

[Sol76c] Robert A. Solo. Transfer of technology and developing countries: Belgrade, September 9–11, 1975. Technology and Culture, 17(3):518– 520, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891761/pdf.

Sonnemann:1978:BWE

[Son78] Rolf Sonnemann. Berlins Weg zur Elektropolis. Technik- und Industriegeschichte an der Spree by Sigfrid von Weiher (review). Technology and Culture, 19(1):131–133, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892059/pdf.

Sparks:1970:RDG

[Spa70] Morgan Sparks. The R & D Game: Technical Men, Technical Managers, and Research Productivity ed. by David Allison (review). Technology and Culture, 11(4):648–649, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894068/pdf.

Spier:1974:ARA

[Spi74] Robert F. G. Spier. Annual Review of Anthropology ed. by Bernard J. Siegel, et al. (review). Technology and Culture, 15(1):156, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892863/pdf.

Spier:1975:AEJ

[Spi75a] Robert F. G. Spier. Archaeology by Experiment by John Coles (review). Technology and Culture, 16(4):703–704, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891644/pdf.

Spier:1975:PTA

[Spi75b] Robert F. G. Spier. Primitive technology and art: Calgary, November 15–17, 1974. Technology and Culture, 16(4):603–606, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891598/pdf.

Spier:1978:TCW

[Spi78] Robert F. G. Spier. Tomol: Chumash Watercraft as Described in the Ethnographic Notes of John P. Harrington ed. by Travis Hudson, et al. (review). Technology and Culture, 19(4):730–731, October 1978.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892179/pdf.

Spong:1972:BLT

[Spo72] Raymond A. Spong. The boyers’ locomotive toy. Technology and Culture, 13(3):373–390, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893922/pdf.

Sporn:1973:ESC

[Spo73] Philip Sporn. Economic and Social Consequences of Nuclear Energy ed. by Lord Sherfield (review). Technology and Culture, 14(1):97–98, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893368/pdf.

Springer:1976:APU

[Spr76] Lynn E. Springer. Art Pottery of the United States: an Encyclopedia of Producers and Their Marks by Paul F. Evans (review). Technology and Culture, 17(2):413–414, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891746/pdf.

Spiegel-Rosing:1973:CSC

[SR73] I. S. Spiegel-Rösing. Civilization and Science: In Conflict or Collaboration (review). Technology and Culture, 14(4):678–681, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892908/pdf.

Stapleton:1977:HPW

[Sta77a] Darwin H. Stapleton. History of Public Works in the United States, 1776–1976 ed. by Ellis L. Armstrong (review). Technology and Culture, 18(2):248–249, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891921/pdf.

Staudenmaier:1977:KPL

[Sta77b] John M. Staudenmaier. Knowledge and Power: Lectures by F. R. Jevons ed. by C. E. West (review). Technology and Culture, 18 (4):723–725, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892019/pdf.

Stapleton:1978:TCB

[Sta78a] Darwin H. Stapleton. Technological Change and the British Iron Industry, 1700–1870 by Charles K. Hyde (review). Technology and Culture, 19(3):552–553, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892142/pdf.

Stapleton:1978:IYM

[Sta78b] Darwin H. Stapleton. The Ingenious Yankees: The Men, Ideas, and Machines That Transformed a Nation, 1776–1876 by Joseph, Frances Gies (review). Technology and Culture, 19(1):121–122, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892053/pdf.

Starbuck:1978:IAE

[Sta78c] David R. Starbuck. Industrial archaeology at the Eli Whitney armory: a reply to Woodbury. Technology and Culture, 19(3):609– 611, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892166/pdf.

Staudenmaier:1978:ZAM

[Sta78d] John M. Staudenmaier. Zen and the Art of Motorcycle Maintenance by Robert M. Pirsig, and: Truck: On Rebuilding a Worn-out Pickup, and Other Post-technological Adventures by John Jerome (review). Technology and Culture, 19(2):257–259, April 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892100/pdf.

Stevens:1970:LMR

[Ste70] G. R. Stevens. The Liverpool and Manchester Railway Project, 1821– 1831 by Robert E. Carlson (review). Technology and Culture, 11 (4):628–630, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894056/pdf.

Stearns:1971:MCB

[Ste71] Peter N. Stearns. Men of Conscience by Brian Inglis (review). Technology and Culture, 12(4):646–648, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893569/pdf.

Stearns:1973:MVB

[Ste73] Peter N. Stearns. Mid-Victorian Britain, 1851–75 by Geoffrey Best (review). Technology and Culture, 14(3):503–505, July 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893231/pdf.

Stein:1975:WIR

[Ste75a] Margot B. Stein. Workers in the Industrial Revolution ed. by Peter N. Stearns, Daniel J. Walkowitz (review). Technology and Culture, 16(4):630–635, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891607/pdf.

Steinberg:1975:CTD

[Ste75b] Arthur Steinberg. Copper and Tin: The Distribution of Mineral Resources and the Nature of the Metals Trade in the Bronze Age by James David Muhly (review). Technology and Culture, 16(3):476– 478, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891573/pdf.

Stearns:1976:GCF

[Ste76a] Peter N. Stearns. The Glassworkers of Carmaux: French Craftsmen and Political Action in a Nineteenth-Century City by Joan Wallach Scott (review). Technology and Culture, 17(4):802–804, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891853/pdf.

Stearns:1976:WCI

[Ste76b] Peter N. Stearns. Working-Class Images of Society ed. by Martin Bulmer (review). Technology and Culture, 17(3):585–586, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891796/pdf.

Steen:1976:MMH

[Ste76c] Harold K. Steen. Mills and Markets: a History of the Pacific Coast Lumber Industry to 1900 by Thomas R. Cox (review). Technology and Culture, 17(2):387–389, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891732/pdf.

Steinberg:1976:AME

[Ste76d] Arthur Steinberg. Aegean Metalwork of the Early and Middle Bronze Age by Keith Branigan (review). Technology and Culture, 17(3):531– 535, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891768/pdf.

Steinberg:1976:PMS

[Ste76e] Arthur Steinberg. Le Problème des métaux dans la science antique by Robert Halleux (review). Technology and Culture, 17(3):535– 537, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891769/pdf.

Steinitz:1976:LRU

[Ste76f] Kate T. Steinitz. Ladislao Reti and “The Unknown Leonardo”. Technology and Culture, 17(2):264–270, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891718/pdf.

Stillman:1977:ONA

[Sti77] Peter G. Stillman. The Obsolete Necessity: America in Utopian Writings, 1888–1900 by Kenneth M. Roemer (review). Technology and Culture, 18(1):139–140, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891899/pdf.

Stokesberry:1970:WAR

[Sto70] James J. Stokesberry. Weapons of the American Revolution and Accoutrements by Warren Moore (review). Technology and Culture, 11(3):480–482, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894134/pdf.

Stover:1971:ABE

[Sto71] Carl F. Stover. Approaching the Benign Environment: The Franklin Lectures in the Sciences and Humanities by R. Buckminster Fuller, et al. (review). Technology and Culture, 12(1):140–141, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894030/pdf.

Stover:1973:EDO

[Sto73] John F. Stover. Enterprise Denied: Origins of the Decline of American Railroads, 1897–1917 by Albro Martin (review). Technology and Culture, 14(1):89–90, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893362/pdf.

Storer:1974:ICD

[Sto74] Norman W. Storer. Invisible Colleges: Diffusion of Knowledge in Scientific Communities by Diana Crane (review). Technology and Culture, 15(1):139–142, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892856/pdf.

Strauss:1971:MWC

[Str71a] Felix F. Strauss. “Mills without wheels” in the 16th-Century Alps. Technology and Culture, 12(1):23–42, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893998/pdf.

Strayer:1971:ISM

[Str71b] Joseph R. Strayer. Irrigation and Society in Medieval Valencia by Thomas F. Glick (review). Technology and Culture, 12(2):335–336, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893967/pdf.

Strauss:1972:TGT

[Str72] Felix F. Strauss. Triebwerke in Getreidemühlen – eine technischgeschichtliche Studie by Hermann Gleisberg (review). Technology and Culture, 13(4):626–627, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893381/pdf.

Strassmann:1973:PHI

[Str73a] W. Paul Strassmann. The Politics of Housing Innovation: The Fate of the Civilian Industrial Technology Program by Dorothy Nelkin (review). Technology and Culture, 14(1):90–91, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893363/pdf.

Struik:1973:LBB

[Str73b] Dirk J. Struik. The Life of Benjamin Banneker by Silvio A. Bedini (review). Technology and Culture, 14(1):84–86, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893359/pdf.

Strath:1975:MRN

[Str75] Frederick S. Strath. Man’s Responsibility for Nature: Ecological Problems and Western Traditions by John Passmore (review). Technology and Culture, 16(4):685–687, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891634/pdf.

Strassmann:1976:TCI

[Str76] W. Paul Strassmann. Technical Choice, Innovation and Economic Growth: Essays on American and British Experience in the Nineteenth Century by Paul A. David (review). Technology and Culture, 17(2):380–382, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891729/pdf.

Strayer:1978:PMT

[Str78] Joseph R. Strayer. On Pre-Modern Technology and Science: a Volume of Studies in Honor of Lynn White, jr ed. by Bert S. Hall, Delmo C. West (review). Technology and Culture, 19(1):109–111, January 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892046/pdf.

Sumner:1975:WWS

[Sum75] Philip L. Sumner. Wheels and Wheeling: The Smithsonian Cycle Collection by Smith Hempstone Oliver, Donald H. Berkebile (review). Technology and Culture, 16(3):514, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891592/pdf.

Susskind:1970:EWF

[Süs70] Charles Süsskind. Electronics in the West: The First Fifty Years by Jane Morgan (review). Technology and Culture, 11(2):320, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894170/pdf.

Susskind:1971:TBW

[Süs71] Charles Süsskind. Think: a Biography of the Watsons and IBM by William Rodgers (review). Technology and Culture, 12(1):125–127, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894023/pdf.

Susskind:1976:WWN

[Süs76a] Charles Süsskind. The World of Walther Nernst: The Rise and Fall of German Science, 1864–1941 by K. Mendelssohn (review). Technology and Culture, 17(1):151–153, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891678/pdf.

Susskind:1976:C

[Süs76b] Charles Süsskind. Commentary. Technology and Culture, 17(3):479– 481, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891754/pdf.

Sviedrys:1977:IPC

[Svi77] Romualdas Sviedrys. The Ideal of the Practical: Colombia’s Struggle to Form a Technical Elite by Frank Safford (review). Technology and Culture, 18(1):142–144, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891901/pdf.

Swanson:1970:MKM

[Swa70] Rowena Swanson. Moshava, Kibbutz, and Moshav by D. Weintraub, et al. (review). Technology and Culture, 11(4):645–648, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894067/pdf.

Swenson:1971:AAS

[Swe71] Loyd S. Swenson, Jr. American advancement into space: American Institute of Aeronautics and Astronautics Houston, Texas, October 19, 1970. Technology and Culture, 12(2):228–230, April 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893957/pdf.

Swenson:1972:GMC

[Swe72] Loyd S. Swenson, Jr. Goals and Means in the Conquest of Space by R. G. Perel’man (review). Technology and Culture, 13(2):338– 340, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893491/pdf.

Swenson:1973:IFC

[Swe73] Loyd S. Swenson, Jr. Interplanetary Flight and Communication by N. A. Rynin (review). Technology and Culture, 14(2 (part I)):317– 323, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893281/pdf.

Swenson:1975:HDP

[Swe75] Loyd S. Swenson, Jr. The Hugh L. Dryden Papers, 1898–1965: a Preliminary Catalogue of the Basic Collection ed. by Richard K. Smith (review). Technology and Culture, 16(3):514–515, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891593/pdf.

Swenson:1978:LIH

[Swe78a] Loyd S. Swenson, Jr. Lunar Impact: a History of Project Ranger by R. Cargill Hall (review). Technology and Culture, 19(4):777–779, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892205/pdf.

Swenson:1978:SRS

[Swe78b] Loyd S. Swenson, Jr. The Spaceflight Revolution: a Sociological Study by William Sims Bainbridge (review). Technology and Culture, 19(4):784–786, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892208/pdf.

Swierenga:1978:HHP

[Swi78] Robert P. Swierenga. Hired Hands and Plowboys: Farm Labor in the Midwest, 1815–60 by David E. Schob (review). Technology and Culture, 19(2):220–221, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892080/pdf.

Sylla:1972:BIA

[Syl72] Richard Sylla. British Investment in American Railways, 1834–1898 by Dorothy R. Adler (review). Technology and Culture, 13(1):78–79, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893442/pdf.

Tager:1972:UEJ

[Tag72] Jack Tager. The Urbanization of the Earth by Jorge Arango (review). Technology and Culture, 13(3):518–519, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893946/pdf.

Tann:1978:WSW

[Tan78] Jennifer Tann. Wiltshire and Somerset Woollen Mills by Kenneth Rogers (review). Technology and Culture, 19(2):226–228, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892084/pdf.

Tarpley:1974:PIE

[Tar74] Fred A. Tarpley, Jr. Patents, Invention, and Economic Change: Data and Selected Essays by Jacob Schmookler (review). Technology and Culture, 15(1):138–139, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892855/pdf.

Tarr:1977:RTA

[Tar77] Joel A. Tarr. Retrospective technology assessment: Carnegie– Mellon University, December 2–4, 1976. Technology and Culture, 18(4):655–658, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891994/pdf.

Tarr:1979:TPD

[Tar79] Joel A. Tarr. The Toilet Papers: Designs to Recycle Human Waste and Water: Dry Toilets, Greywater Systems, and Urban Sewage by Sim Van der Ryn (review). Technology and Culture, 20(1):231–233, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890479/pdf.

Taviss:1972:SFM

[Tav72a] Irene Taviss. Science and the Future of Man ed. by Robert L. Carovillano, James W. Skehan, S.J. (review). Technology and Culture, 13(2):331–332, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893486/pdf.

Taviss:1972:SPA

[Tav72b] Irene Taviss. A survey of popular attitudes toward technology. Technology and Culture, 13(4):606–621, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893379/pdf.

Taviss:1974:SSI

[Tav74] Irene Taviss. Supplement to the Sociology of Invention by S. Colum Gilfillan (review). Technology and Culture, 15(1):136–138, January 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892854/pdf.

Taylor:1975:OCE

[Tay75] James C. Taylor. Organizational Change: an Exploratory Study and Case History by Michael Tushman (review). Technology and Culture, 16(3):508–510, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891589/pdf.

Tebo:1974:ETS

[Teb74] Julian D. Tebo. The Electric Telegraph: a Social and Economic History by Jeffrey L. Keive (review). Technology and Culture, 15(2):332– 334, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892800/pdf.

Thackray:1972:CSS

[Tha72] Arnold Thackray. The Common Scientist in the Seventeenth Century: a Study of the Dublin Philosophical Society, 1683–1708 by K. Theodore Hoppen (review). Technology and Culture, 13(1):71–73, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893437/pdf.

Thackray:1973:PHS

[Tha73a] Arnold Thackray. Perspectives in the History of Science and Technology ed. by Duane H. D. Roller (review). Technology and Culture, 14(2 (part I)):287–288, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893263/pdf.

Thackray:1973:ARI

[Tha73b] Arnold Thackray. The Archives of the Royal Institution of Great Britain. Minutes of Managers’ Meetings, 1799–1900 (review). Technology and Culture, 14(3):531–532, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893244/pdf.

Thackray:1976:PPM

[Tha76] Arnold Thackray. Philadelphia’s Philosopher Mechanics: a History of the Franklin Institute, 1824–1865 by Bruce Sinclair (review). Technology and Culture, 17(2):382–385, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891730/pdf.

SHT:1974:DJG

[The74] The Society for the History of Technology. Dedication to Joseph Gray Jackson. Technology and Culture, 15(3):iv, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892734/pdf.

Thoday:1975:BSN

[Tho75a] A. G. Thoday. The Birth of Scientific Navigation: The Solving in the Eighteenth Century of the Problem of Finding Longitude at Sea by Eric G. Forbes (review). Technology and Culture, 16(1):86–88, January 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891505/pdf.

Thomas:1975:GTP

[Tho75b] Donald E. Thomas, Jr. German Trade Promotion in the First Half of the Nineteenth Century by Werner Plum (review). Technology and Culture, 16(4):635–636, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891608/pdf.

Thomas:1976:GGC

[Tho76] Phillip Drennon Thomas. Gunpowder and Galleys: Changing Technology and Mediterranean Warfare at Sea in the Sixteenth Century by John Francis Guilmartin, Jr. (review). Technology and Culture, 17(2):373–375, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891725/pdf.

Thomas:1978:DFS

[Tho78] Donald E. Thomas, Jr. Diesel, father and son: Social philosophies of technology. Technology and Culture, 19(3):376–393, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892109/pdf.

Tobey:1976:TST

[Tob76] Ronald Tobey. Theoretical science and technology in American ecology. Technology and Culture, 17(4):718–728, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891821/pdf.

Toffler:1972:PFP

[Tof72] Alvin Toffler. Prognostics by Fred L. Polak (review). Technology and Culture, 13(3):520–523, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893948/pdf.

Topper:1972:EAH

[Top72] David Topper. The Ethereal Aether: a History of the Michelson– Morley–Miller Aether-Drift Experiments, 1880–1930 by Loyd S. Swenson, Jr. (review). Technology and Culture, 13(4):669–670, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893404/pdf.

Topper:1974:MLB

[Top74] David R. Topper. The Master of Light: a Biography of Albert A. Michelson by Dorothy Michelson Livingston (review). Technology and Culture, 15(4):655–657, October 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892721/pdf.

Topper:1979:ETS

[Top79a] David Topper. The Essential Tension: Selected Studies in Scientific Tradition and Change by Thomas S. Kuhn (review). Technology and Culture, 20(1):218–220, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890472/pdf.

Topper:1979:SIC

[Top79b] David Topper. The Scientific Imagination: Case Studies by Gerald Holton (review). Technology and Culture, 20(1):216–218, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890471/pdf.

Tribe:1972:TAU

[Tri72] Laurence H. Tribe. Technology Assessment: Understanding the Social Consequences of Technological Applications ed. by Raphael G. Kasper (review). Technology and Culture, 13(4):686–687, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893414/pdf.

Trilling:1977:WBE

[Tri77] Leon Trilling. Wernher von Braun by Erik Bergaust (review). Technology and Culture, 18(4):717–718, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892016/pdf.

Troitzsch:1972:BTT

[Tro72] Ulrich Troitzsch. Beiträge zur Technikgeschichte Tirols ed. by Ernst Attlmayr (review). Technology and Culture, 13(2):344–345, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893494/pdf.

Tropp:1975:CPN

[Tro75] Henry S. Tropp. The Computer from Pascal to von Neumann by Herman H. Goldstine (review). Technology and Culture, 16(4):660–662, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891621/pdf.

Truitt:1971:AAA

[Tru71] Willis H. Truitt. Automation, Alienation, and Anomie ed. by Simon Marcson (review). Technology and Culture, 12(1):134–137, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894028/pdf.

Tuchman:1973:WWE

[Tuc73] Gaye Tuchman. Women at Work ed. by William L. O’Neill, and: Occupation Housewife by Helena Z. Lopata (review). Technology and Culture, 14(2 (part I)):326–328, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893283/pdf.

Tucker:1978:FVR

[Tuc78] D. Gordon Tucker. François van Rysselberghe: Pioneer of longdistance telephony. Technology and Culture, 19(4):650–674, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892170/pdf.

Turnbull:1975:CVN

[Tur75] Colin M. Turnbull. Cultures. Vol. 1, no. 1: Music and Society by UNESCO (review). Technology and Culture, 16(4):700–701, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891641/pdf.

Turner:1976:HSP

[Tur76] R. Steven Turner. Historical Studies in the Physical Sciences ed. by Russell McCormmach (review). Technology and Culture, 17(3):604– 606, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891806/pdf.

Tylecote:1977:OFP

[Tyl77] R. F. Tylecote. “Officina Ferraria”: a Polish Poem of 1612 Describing the Noble Craft of ironwork by Walenty Roździeński (review). Technology and Culture, 18(4):690–691, October 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891999/pdf.

Sun:1973:ACP

[tZS73a] E tu Zen Sun. Agricultural Change and the Peasant Economy in South China by Evelyn Sakakida Rawski (review). Technology and Culture, 14(3):495–497, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893226/pdf.

Sun:1973:SCC

[tZS73b] E tu Zen Sun. Science and Civilisation in China. Vol. 4: Physics and Physical Technology: Part III. Civil Engineering and Nautics by Joseph Needham (review). Technology and Culture, 14(2 (part I)):289–292, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893265/pdf.

zenSun:1975:CWC

[tzS75] E tu zen Sun. The Chinese Worker by Charles Hoffmann (review). Technology and Culture, 16(4):672–673, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891627/pdf.

Sun:1976:CME

[tZS76] E tu Zen Sun. China’s Modern Economy in Historical Perspective ed. by Dwight H. Perkins (review). Technology and Culture, 17(4):804–807, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891854/pdf.

Sun:1978:STE

[tZS78] E tu Zen Sun. Science and Technology in East Asia ed. by Nathan Sivin (review). Technology and Culture, 19(3):544–546, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892136/pdf.

Uchida:1975:JTC

[Uch75] Hoshimi Uchida. Japan’s Technological Challenge to the West, 1950– 1974: Motivation and Accomplishment by Terutomo Ozawa (review). Technology and Culture, 16(4):674–676, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891628/pdf.

Unger:1972:WHS

[Ung72] Richard W. Unger. Workshop on the history of science and technology: University of British Columbia, October 1971. Technology and Culture, 13(4):622–625, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893380/pdf.

Unger:1978:ABB

[Ung78] Richard W. Unger. Archaeology of the Boat by Basil Greenhill (review). Technology and Culture, 19(2):210–212, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892074/pdf.

Uselding:1974:EKR

[Use74] Paul Uselding. Elisha K. Root, forging, and the “American System”. Technology and Culture, 15(4):543–568, October 1974. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892701/pdf.

Uselding:1978:TAE

[Use78] Paul Uselding. Technology and American Economic Growth by Nathan Rosenberg (review). Technology and Culture, 19(2):214– 216, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892077/pdf.

Vaczek:1970:BMS

[Vac70] Louis Vaczek. Beyond Modern Sculpture: The Effects of Science and Technology on the Sculpture of This Century by Jack Burnham (review). Technology and Culture, 11(4):655–659, October 1970. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894072/pdf.

Vaczek:1971:AMN

[Vac71] Louis Vaczek. The Architecture Machine by Nicholas Negreponte, and: Behind Appearance: a Study of the Relations between Painting and the Natural Sciences in This Century by C. H. Waddington (review). Technology and Culture, 12(2):380–384, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893991/pdf.

Vaczek:1972:ESH

[Vac72a] Louis Vaczek. Eyewitness to Space by Hereward Lester Cooke (review). Technology and Culture, 13(3):512–515, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893944/pdf.

Vaczek:1972:MAS

[Vac72b] Louis Vaczek. Modern Art and Scientific Thought by John Adkins Richardson (review). Technology and Culture, 13(4):659–660, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893399/pdf.

Vaczek:1972:ODT

[Vac72c] Louis Vaczek. Overskill: The Decline of Technology in Modern Civilization by Eugene S. Schwartz (review). Technology and Culture, 13(4):685–686, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893413/pdf.

Vaczek:1974:AFD

[Vac74] Louis Vaczek. Art and the Future by Douglas Davis (review). Technology and Culture, 15(2):360–362, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892815/pdf.

Vaczek:1977:ACE

[Vac77a] Louis Vaczek. Artist and Computer ed. by Ruth Leavitt (review). Technology and Culture, 18(3):569–571, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891978/pdf.

Vaczek:1977:SMP

[Vac77b] Louis Vaczek. Science and the Media by Peter Farago (review). Technology and Culture, 18(2):268–269, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891932/pdf.

Vaczek:1978:PSS

[Vac78] Louis Vaczek. On Photography by Susan Sontag (review). Technology and Culture, 19(4):760–762, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892195/pdf.

VanDyke:1971:DGM

[Van71] Vernon Van Dyke. The Decision to Go to the Moon: Project Apollo and the National Interest by John M. Logsdon (review). Technology and Culture, 12(2):360–361, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893979/pdf.

Vanek:1978:FGE

[Van78a] Joann Vanek. The Factory Girls ed. by Philip S. Foner (review). Technology and Culture, 19(4):752–753, October 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892191/pdf.

Vanek:1978:WPL

[Van78b] Joann Vanek. Worlds of Pain: Life in the Working-Class Family by Lillian Breslow Rubin (review). Technology and Culture, 19(2):234– 236, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892088/pdf.

Vanek:1978:HTS

[Van78c] Joann Vanek. Household technology and social status: Rising living standards and status and residence differences in housework. Technology and Culture, 19(3):361–375, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/892108/pdf.

vanDeurs:1975:FAU

[vD75] George van Deurs. First Across: The U.S. Navy’s Transatlantic flight of 1919 by Richard K. Smith (review). Technology and Culture, 16(3):502–504, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891586/pdf.

Venable:1976:FML

[Ven76] Wm. Henry Venable. Flaws in McLuhan’s laws. Technology and Culture, 17(2):256–262, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891716/pdf.

Vincenti:1979:APT

[Vin79] Walter G. Vincenti. The air-propeller tests of W. F. Durand and E. P. Lesley: a case study in technological methodology. Technology and Culture, 20(4):712–751, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890572/pdf.

vanOss:1978:LGL

[vO78] Carel J. van Oss. Light on gas lighting. Technology and Culture, 19(3):602–603, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892162/pdf.

Volti:1978:RIC

[Vol78] Rudi Volti. Rural Industrialization in China by Jon Sigurdson, and: Rural Small-Scale Industry in the People’s Republic of China by the American Rural Small-Scale Industry Delegation (review). Technology and Culture, 19(3):541–543, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892135/pdf.

vanRiemsdijk:1978:CSB

[vR78] J. T. van Riemsdijk. Changing styles in “Best practice museology”. Technology and Culture, 19(3):606–607, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892164/pdf.

vanSpronsen:1972:DBJ

[vS72] Jan W. van Spronsen. Dr. B. J. Tideman, 1934–1883 by J. M. Dirkzwager (review). Technology and Culture, 13(3):503, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893938/pdf.

Wagar:1973:DNW

[Wag73] W. Warren Wagar. The Domination of Nature by William Leiss (review). Technology and Culture, 14(3):480–482, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893219/pdf.

Walker:1971:PPM

[Wal71a] Michael Walker. Pioneer Pottery by Michael Cardew (review). Technology and Culture, 12(2):331–332, April 1971. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893964/pdf.

Walker:1971:TAW

[Wal71b] Michael Walker. Technology in the Ancient World by Henry Hodges (review). Technology and Culture, 12(1):89–90, January 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894005/pdf.

Walker:1973:SAE

[Wal73a] Michael Walker. Science and Archaeology ed. by Robert H. Brill (review). Technology and Culture, 14(1):77–78, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893354/pdf.

Walker:1973:ECC

[Wal73b] Michael Walker. The Emergence of Civilisation: The Cyclades and the Aegean in the Third Millennium B.C. by Colin Renfrew, and: Lerna, a Preclassical Site in the Argolid: Results of Excavations Conducted by the American School of Classical Studies. Vol. 2: The People by J. Lawrence Angel (review). Technology and Culture, 14(1):78–81, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893355/pdf.

Walker:1974:PEP

[Wal74] Michael Walker. Papers in Economic Prehistory ed. by Eric S. Higgs (review). Technology and Culture, 15(1):86–89, January 1974. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892830/pdf.

Walker:1976:ECC

[Wal76a] Michael Walker. The Explanation of Culture Change: Models in Prehistory ed. by Colin Renfrew (review). Technology and Culture, 17(1):190–191, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891698/pdf.

Wallis:1976:TAL

[Wal76b] George W. Wallis. Tradition and Adaptation: Life in a Modern Yucatan Maya Village by Irwin Press (review). Technology and Culture, 17(4):807–808, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891855/pdf.

Warner:1970:SPH

[War70] Deborah Jean Warner. Star Performance by Harriett Pratt Lattin (review). Technology and Culture, 11(3):449, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894117/pdf.

Warner:1971:LMR

[War71] Deborah Jean Warner. Lewis M. Rutherfurd: Pioneer astronomical photographer and spectroscopist. Technology and Culture, 12 (2):190–216, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893955/pdf.

Warner:1973:CIG

[War73] Sam Bass Warner, Jr. Cities and Immigrants, a Geography of Change in Nineteenth-Century America by David Ward (review). Technology and Culture, 14(1):94, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893365/pdf.

Warner:1976:FPE

[War76] Deborah Jean Warner. Francis Place and the Early History of the Greenwich Observatory by Derek Howse (review). Technology and Culture, 17(2):414–415, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891747/pdf.

Watson:1972:SBI

[Wat72a] Sara Ruth Watson. A Span of Bridges: an Illustrated History by H. J. Hopkins (review). Technology and Culture, 13(1):81–83, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893444/pdf.

Watson:1972:FSK

[Wat72b] Sara Ruth Watson. The Fashionable Stone by Kenneth Hudson (review). Technology and Culture, 13(1):79–81, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893443/pdf.

Watson:1973:GBD

[Wat73] Sara Ruth Watson. The Great Bridge by David McCullough (review). Technology and Culture, 14(3):506–509, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893233/pdf.

Watson:1976:EPE

[Wat76a] Sara Ruth Watson. The Existential Pleasures of Engineering by Samuel C. Florman (review). Technology and Culture, 17(3):528– 529, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891766/pdf.

Watson:1976:TTJ

[Wat76b] Sara Ruth Watson. The Tallest Tower by Joseph Harriss (review). Technology and Culture, 17(3):552–554, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891778/pdf.

Watson:1977:OSP

[Wat77] Sara Ruth Watson. The Orchestra at San Petronio in the Baroque Era by Eugene Enrico (review). Technology and Culture, 18(2):243– 244, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891918/pdf.

Watson:1978:SIR

[Wat78a] Sara Ruth Watson. Studies in Italian Renaissance Architecture by Wolfgang Lotz (review). Technology and Culture, 19(3):531–533, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892130/pdf.

Watson:1978:PTM

[Wat78b] Sara Ruth Watson. The Palazzo del Te in Mantua: Images of Love and Politics by Egon Verheyen (review). Technology and Culture, 19(3):530–531, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892129/pdf.

Watson:1979:CIH

[Wat79a] Sara Ruth Watson. Cleveland: an Inventory of Historic Engineering and Industrial Sites ed. by Daniel M. Bluestone (review). Technology and Culture, 20(2):369–370, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890504/pdf.

Watson:1979:ECJ

[Wat79b] Sara Ruth Watson. Engineering Classics of James Kip Finch ed. by Neal FitzSimons (review). Technology and Culture, 20(4):787–788, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890575/pdf.

Webster:1974:EGW

[Web74] T. B. L. Webster. Early Greek Warfare: Horsemen and Chariots in the Homeric and Archaic Ages by P. A. L. Greenhalgh (review). Technology and Culture, 15(2):325–326, April 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892796/pdf.

Webb:1976:PSA

[Web76] Steven B. Webb. The Problem of Slavery in the Age of Revolution, 1770–1823 by David Brion Davis (review). Technology and Culture, 17(1):132–134, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891667/pdf.

Weber:1979:TIS

[Web79] Wolfhard Weber. Technology and its impact on society: Stockholm, August 1977. Technology and Culture, 20(3):614–617, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890530/pdf.

Weinberg:1971:IST

[Wei71] Alvin M. Weinberg. The Interaction of Science and Technology ed. by W. Dale Compton (review). Technology and Culture, 12(2):372– 374, April 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893986/pdf.

Weiss:1972:WSC

[Wei72] John H. Weiss. Work, Society, and Culture by Yves Simon (review). Technology and Culture, 13(1):98–100, January 1972. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893452/pdf.

Weinberg:1973:STA

[Wei73] Robert Weinberg. Science and Technology in Art Today by Jonathan Benthall (review). Technology and Culture, 14(4):667–669, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892902/pdf.

Weintraub:1975:PSP

[Wei75] Andrew Weintraub. The Productive Society by Patrick Haggerty (review). Technology and Culture, 16(4):683–684, October 1975. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891633/pdf.

Weintraub:1976:TCA

[Wei76] Andrew Weintraub. Two Cheers for the Affluent Society by Wilfred Beckerman (review). Technology and Culture, 17(1):188–189, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891697/pdf.

Weinstein:1978:MCE

[Wei78a] Jay Weinstein. Managing the Commons ed. by Garrett Hardin, John Baden (review). Technology and Culture, 19(3):589–592, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892157/pdf.

Weinstein:1978:TBN

[Wei78b] Jay Weinstein. Technologies for Basic Needs by Hans Singer, and: Science Policy Studies in India: a Status Report by A. Rahman (review). Technology and Culture, 19(4):797–800, October 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892215/pdf.

Weinstein:1979:FFM

[Wei79] Jay Weinstein. From Field to Factory by Morton Klass, and: The Net and the Quest by Ralph Glasser (review). Technology and Culture, 20(4):839–847, October 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/890598/pdf.

Welsh:1970:PMF

[Wel70] Peter C. Welsh. Pattern in the Material Folk Culture of the Eastern United States by Henry Glassie (review). Technology and Culture, 11(1):108–109, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894201/pdf.

Welsh:1971:POE

[Wel71] Peter C. Welsh. Planemakers and Other Edge Tool Enterprises in New York State in the Nineteenth Century by Kenneth D. Roberts, Jane W. Roberts (review). Technology and Culture, 12(4):648–649, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893570/pdf.

Welsh:1973:AAS

[Wel73a] Peter C. Welsh. American Axes: a Survey of Their Development and Their Makers by Henry J. Kauffman (review). Technology and Culture, 14(1):83–84, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893358/pdf.

Welsh:1973:WPE

[Wel73b] Peter C. Welsh. Winterthur Portfolio 7 ed. by Ian M. G. Quimby (review). Technology and Culture, 14(3):527–530, July 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893243/pdf.

Welsh:1977:WPC

[Wel77] Peter C. Welsh. Wooden Planes in 19th-Century America by Kenneth D. Roberts (review). Technology and Culture, 18(3):541–543, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891963/pdf.

Wertime:1970:ISW

[Wer70] Theodore A. Wertime. Iron and Steel by W. K. V. Gale (review). Technology and Culture, 11(4):630–631, October 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/894057/pdf.

Wertime:1978:SAS

[Wer78a] Theodore A. Wertime. Sir Aurel Stein, Archaeological Explorer by Jeannette Mirsky (review). Technology and Culture, 19(3):512–514, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892120/pdf.

Wertime:1978:ESP

[Wer78b] Theodore A. Wertime. The Emergence of Society: a Prehistory of the Establishment by John E. Pfeiffer (review). Technology and Culture, 19(3):509–512, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892119/pdf.

Wertime:1979:ENS

[Wer79] Theodore A. Wertime. Excavations at Nichoria in Southwest Greece. Vol. 1: Site, Environs, and Techniques ed. by George Rapp, Jr., S. E. Aschenbrenner (review). Technology and Culture, 20(3):618– 620, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890531/pdf.

Westbrook:1976:GCV

[Wes76] J. H. Westbrook. The Gunsmith in Colonial Virginia by H. B. Gill, Jr. (review). Technology and Culture, 17(4):752–753, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891827/pdf.

Westbrook:1977:WCA

[Wes77] J. H. Westbrook. Wind-Catchers: American Windmills of Yesterday and Tomorrow by Volta Torrey (review). Technology and Culture, 18(3):548–551, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891967/pdf.

West:1978:PCH

[Wes78] William Kyer West. Problems in the cultural history of the ellipse. Technology and Culture, 19(4):709–712, October 1978. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892173/pdf.

Whelan:1973:CSP

[Whe73] Robert K. Whelan. Culture Storm: Politics and the Ritual Order by H. L. Nieburg (review). Technology and Culture, 14(4):681–683, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892909/pdf.

Wheaton:1978:HSP

[Whe78] Bruce R. Wheaton. Historical Studies in the Physical Sciences ed. by Russell McCormmach, Lewis Pyenson (review). Technology and Culture, 19(3):596–599, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892160/pdf.

White:1970:BOT

[Whi70] John H. White, Jr. Baltimore and Ohio Transportation Museum by William E. Pyne (review). Technology and Culture, 11(1):70–84, January 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894192/pdf.

White:1971:AVP

[Whi71a] Dana F. White. The apocalyptic vision of Paolo Soleri. Technology and Culture, 12(1):75–88, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894004/pdf.

White:1971:ITR

[Whi71b] Dana F. White. Inside the Third Reich: Memoirs by Albert Speer tr. by Richard Winston, Clara Winston (review). Technology and Culture, 12(3):529–532, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893620/pdf.

White:1971:ALL

[Whi71c] John H. White. Articulated Locomotives by Lionel Wiener (review). Technology and Culture, 12(4):656–658, October 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893575/pdf.

White:1972:SLT

[Whi72a] John H. White. The Shay Locomotive, Titan of the Timber by Michael Koch (review). Technology and Culture, 13(2):316–318, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893477/pdf.

White:1972:CBD

[Whi72b] John H. White, Jr. The Car Builder’s Dictionary … 1879 by Matthias N. Forney, and: The Car Builder’s Dictionary … 1888 by Matthias N. Forney, and: The Car Builder’s Dictionary … 1906 by Rodney Hitt (review). Technology and Culture, 13(4):636–637, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893387/pdf.

White:1972:LBM

[Whi72c] John H. White, Jr. The Locomotives Built by Machinefabriek “Breda” by A. D. de Pater (review). Technology and Culture, 13(2):315–316, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/893476/pdf.

White:1973:ERD

[Whi73a] John H. White. Electric Railway Dictionary by Rodney Hitt (review). Technology and Culture, 14(2 (part I)):306–307, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893274/pdf.

White:1973:LDG

[Whi73b] John H. White. Locomotive Dictionary by George L. Fowler (review). Technology and Culture, 14(2 (part I)):304–305, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893273/pdf.

White:1973:IHF

[Whi73c] John H. White, Jr. An Illustrated Handbook of Fire Apparatus by George Anne Daly, John J. Robrecht, and: Fighting Fire with Fire by John M. Peckham (review). Technology and Culture, 14(4):630–632, October 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892881/pdf.

White:1973:RLR

[Whi73d] John H. White, Jr. Recent Locomotives by the Railroad Gazette (review). Technology and Culture, 14(4):635–636, October 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892883/pdf.

White:1973:OSF

[Whi73e] John H. White, Jr. Origins of the steam fire engine. Technology and Culture, 14(2 (part I)):166–169, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893259/pdf.

White:1973:RMP

[Whi73f] John H. White, Jr. The railway museum: Past, present, and future. Technology and Culture, 14(4):599–613, October 1973. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892871/pdf.

White:1974:RJF

[Whi74] Lynn White, Jr. Robert James Forbes (1900–1973). Technology and Culture, 15(3):438–439, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/892738/pdf.

White:1975:TRW

[Whi75a] K. D. White. The Transformation of the Roman World: Gibbon’s Problem after Two Centuries ed. by Lynn White, jr (review). Technology and Culture, 16(3):481–483, July 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891575/pdf.

White:1975:SMT

[Whi75b] Lynn White, Jr. The study of medieval technology, 1924–1974: Personal reflections. Technology and Culture, 16(4):519–530, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891595/pdf.

White:1976:CCJ

[Whi76a] Lawrence J. White. The Car Culture by James J. Flink (review). Technology and Culture, 17(2):400–401, April 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891739/pdf.

White:1976:RV

[Whi76b] John H. White, Jr. The railroad velocipede. Technology and Culture, 17(3):515–517, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891760/pdf.

White:1977:PLL

[Whi77a] John H. White, Jr. “A perfect light is a luxury”: Pintsch gas car lighting. Technology and Culture, 18(1):64–69, January 1977. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891866/pdf.

White:1977:HTH

[Whi77b] John H. White, Jr. A History of Tramways from Horse to Rapid Transit by R. J. Buckley (review). Technology and Culture, 18 (1):102–103, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891880/pdf.

White:1978:ML

[Whi78] John H. White, Jr. More light. Technology and Culture, 19(3):604– 605, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892163/pdf.

Wik:1970:HFE

[Wik70] Reynold M. Wik. Henry Ford ed. by John B. Rae (review). Technology and Culture, 11(2):311–313, April 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/894166/pdf.

Wik:1972:TFR

[Wik72] Reynold M. Wik. The Troubled Farmer: Rural Adjustment to Industrialism, 1850–1900 by Earl V. Hayter (review). Technology and Culture, 13(1):93–94, January 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893449/pdf.

Wik:1973:RCA

[Wik73] Reynold M. Wik. The Road and the Car in American Life by John B. Rae (review). Technology and Culture, 14(2 (part I)):308–311, April 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893276/pdf.

Wik:1977:NER

[Wik77] Reynold M. Wik. News for Everyman: Radio and Foreign Affairs in Thirties America by David Holbrook Culbert (review). Technology and Culture, 18(4):713, October 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/892013/pdf.

Wik:1979:BHI

[Wik79] Reynold M. Wik. Benjamin Holt and the invention of the tracktype tractor. Technology and Culture, 20(1):90–107, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890454/pdf.

Williams:1970:STD

[Wil70a] Wesley C. Williams. Scientific and Technological Documentation: Archival Evaluation and Processing of University Records Relating to Science and Technology by Maynard J. Brichford (review). Technology and Culture, 11(3):487–488, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/894139/pdf.

Wilson:1970:HSR

[Wil70b] David B. Wilson. Heralds of Science, as Represented by Two Hundred Epochal Books and Pamphlets Selected from the Burndy Library by Bern Dibner (review). Technology and Culture, 11(3):482– 483, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
894135/pdf.

Williams:1971:OCF

[Wil71a] Wesley C. Williams. The Ohio Canals by Frank N. Wilcox (review). Technology and Culture, 12(1):107–109, January 1971. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/894017/pdf.

Wilson:1971:NCS

[Wil71b] David B. Wilson. National Catalog of Sources for History of Physics; Report No. 1: a Selection of Manuscript Collections at American Repositories by Joan Nelson Warnow (review). Technology and Culture, 12(1):161–162, January 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/894041/pdf.

Wilkinson:1972:SPD

[Wil72] Kenneth P. Wilkinson. Sociological Perspectives of Domestic Development ed. by George M. Beal, et al. (review). Technology and Culture, 13(4):644–645, October 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893392/pdf.

Wilhelmsen:1973:CTR

[Wil73a] Frederick D. Wilhelmsen. Comment: Through a rearview mirror — darkly. Technology and Culture, 14(1):22–27, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893351/pdf.

Williamson:1973:TAE

[Wil73b] Harold F. Williamson. Technology and American Economic Growth by Nathan Rosenberg (review). Technology and Culture, 14(1):83, January 1973. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893357/pdf.

Wilson:1976:NNF

[Wil76] Richard Wilson. Non-nuclear Futures–the Case for an Ethical Energy Strategy by Amory B. Lovins, John H. Price (review). Technology and Culture, 17(4):811–813, October 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891858/pdf.

Williamson:1977:OPC

[Wil77] Harold F. Williamson. Oil Power by Carl Solberg (review). Technology and Culture, 18(1):113, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891885/pdf.

Wilson:1978:FPH

[Wil78] H. Clyde Wilson. Farming in Prehistory: From Hunter-Gatherer to Food-Producer by Barbara Bender, and: The Food Crisis in Prehistory: Overpopulation and the Origin of Agriculture by Mark Nathan Cohen (review). Technology and Culture, 19(2):207–210, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/892073/pdf.

Winters:1970:NBB

[Win70] Stanley B. Winters. Needles, Burrs, and Bibliographies: Study Resources: Technological Change, Human Values, and the Humanities ed. by Maxwell H. Goldberg (review). Technology and Culture, 11(3):484–485, July 1970. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/894137/pdf.

Winters:1971:DVT

[Win71] Stanley B. Winters. Dějiny věd a techniky [History of the sciences and technology] ed. by Luboš Nový, et al. (review). Technology and Culture, 12(3):487–490, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893600/pdf.

Winpenny:1978:AAL

[Win78] Thomas R. Winpenny. Automobiles in American Life ed. by Charles L. Sanford (review). Technology and Culture, 19(3):562– 564, July 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
892148/pdf.

Wise:1979:TFE

[Wis79] George Wise. Twenty-five Engineers and Inventors by Charles Susskind (review). Technology and Culture, 20(1):230–231, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890478/pdf.

Wachsmann:1971:IMI

[WK71] Klaus P. Wachsmann and Russell Kay. The interrelations of musical instruments, musical forms, and cultural systems in Africa. Technology and Culture, 12(3):399–413, July 1971. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893592/pdf.

Wolf:1978:NTO

[Wol78] Gerrit Wolf. New Technologies in Organization Development ed. by W. Warner Burke, John Adams (review). Technology and Culture, 19(2):255–257, April 1978. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/892099/pdf.

Woodbury:1972:FMH

[Woo72a] Robert S. Woodbury. Feedback Mechanisms in the Historical Collections of the National Museum of History and Technology by Otto Mayr (review). Technology and Culture, 13(3):523–524, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893949/pdf.

Woodbury:1972:HMM

[Woo72b] Robert S. Woodbury. Henry Maudslay, Machine Builder by K. R. Gilbert (review). Technology and Culture, 13(2):314, April 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893474/pdf.

Woodbury:1977:EWA

[Woo77a] Robert S. Woodbury. Eli Whitney Armory Survey (CT-2) by the Historic American Engineering Record (review). Technology and Culture, 18(1):146–148, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891903/pdf.

Woodcock:1977:NYS

[Woo77b] John Woodcock. The Next 200 Years: a Scenario for America and the World by Herman Kahn, et al. (review). Technology and Culture, 18(3):585–588, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891987/pdf.

Woods:1977:FM

[Woo77c] David L. Woods. “A fascinating mishmash”. Technology and Culture, 18(2):391–392, April 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891935/pdf.

Wood:1979:HBP

[Woo79a] Fred B. Wood. Hermes Bound: The Policy and Technology of Telecommunications by Clare D. McGillem, William P. McLaughlin (review). Technology and Culture, 20(3):676–677, July 1979. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890561/pdf.

Woods:1979:AFA

[Woo79b] David L. Woods. American Forts: Architectural Form and Function by Willard B. Robinson (review). Technology and Culture, 20(3):644– 645, July 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890546/pdf.

Woods:1979:UWA

[Woo79c] David L. Woods. Underwater Warfare in the Age of Sail by Alex Roland (review). Technology and Culture, 20(1):199–201, January 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890461/pdf.

Wright:1976:UCA

[Wri76] Monte D. Wright. The UFO Controversy in America by David Michael Jacobs (review). Technology and Culture, 17(3):596–598, July 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891801/pdf.

Young:1972:MEA

[You72] Otis E. Young, Jr. Mining Engineers and the American West: The Lace-Boot Brigade by Clark C. Spence (review). Technology and Culture, 13(3):507–509, July 1972. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893941/pdf.

Znachko-Iavorskii:1977:NMS

[ZIL77] Igor L. Znachko-Iavorskii and Henry M. Leicester. New methods for the study and contemporary aspects of the history of cementing materials. Technology and Culture, 18(1):25–42, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891862/pdf.

Zinberg:1979:TDD

[Zin79] Dorothy S. Zinberg. Technological Decisions and Democracy: European Experiments in Public Participation by Dorothy Nelkin, and: Basic Human Needs: a Framework for Action by John McHale, Magda Cordell McHale (review). Technology and Culture, 20(2):399– 402, April 1979. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890520/pdf.

Zupko:1977:RMS

[Zup77] Ronald Edward Zupko. The role of measurement standards in human civilization: Budapest, April 27–30, 1976. Technology and Culture, 18(3):498–502, July 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891944/pdf.